### v2

Abstract
This paper proposes a correlated random coefficient linear panel data model, where regressors can be correlated with time-varying and individual-specific random coefficients through both a fixed effect and a time-varying random shock. I develop a new panel data-based identification method to identify the average partial effect and the local average response function. The identification strategy employs a sufficient statistic to control for the fixed effect and a conditional control variable for the random shock. Conditional on these two controls, the residual variation in the regressors is driven solely by the exogenous instrumental variables, and thus can be exploited to identify the parameters of interest. The constructive identification analysis leads to three-step series estimators, for which I establish rates of convergence and asymptotic normality. To illustrate the method, I estimate a heterogeneous Cobb-Douglas production function for manufacturing firms in China, finding substantial variations in output elasticities across firms.
Keywords: Correlated random coefficients, panel data, time-varying endogeneity, index exclusion, conditional control variable, average partial effect, local average response function.
1 Introduction
Correlated random coefficient (CRC) linear panel models have proven useful due to their ability to accommodate complex forms of unobserved heterogeneity that are empirically relevant (Wooldridge (2010), Hsiao (2022)). A crucial consideration in these models is the correlation between random coefficients and regressors. Classical methods typically address this issue by allowing a time-invariant, individual-specific fixed effect—either in the form of an additive intercept or an individual-specific coefficient—to be correlated with the regressors (e.g., Hausman and Taylor (1981), Hsiao and Pesaran (2008)). While convenient, this approach may not fully capture agents’ optimization behavior. For instance, it is possible that a firm with a high fixed effect (e.g., strong management skills) endogenously selects lower input levels during certain periods. Such behavior could arise if the firm has information about its time-varying output efficiency shock, in addition to its time-invariant management skill, when optimally determining input values.
In this paper, I aim to address this gap by proposing a new time-varying endogenous random coefficient (TERC) linear panel model, where regressors can be correlated with time-varying and individual-specific random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock. Specifically, the baseline TERC model [^1] In the baseline model, I focus exclusively on endogenous regressors to address the issue of time-varying endogeneity through the random coefficients. The analysis remains largely unchanged when exogenous regressors are included in (1.1). I specify the necessary changes at the end of the proof of Theorem 1 in the Appendix. consists of two equations:
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (1.1)
$X_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right).$ (1.2)
In (1.1), I model the vector of random coefficients $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)$ as a vector of unknown and possibly nonlinear functions $\beta\left(\cdot\right)$ of a fixed effect $A_{i}$ and a time-varying random shock $\varepsilon_{it}$. $Y_{it}$ is then determined by the inner product between $X_{it}$ and $\beta_{it}$. In (1.2), I model the vector of regressors $X_{it}$ as a vector-valued, unknown and possibly nonlinear function $g\left(\cdot\right)$ of the vector of exogenous instrumental variables (IV) $Z_{it}$, fixed effect $A_{i}$, and per-period information $\eta_{it}$ about $\varepsilon_{it}$. The motivation for (1.2) is that the agent $i$ in period $t$ optimally chooses values of her regressors $X_{it}$ by solving an optimization problem (e.g., firm profit maximization), with $Z_{it},\ A_{i},\text{ and }\eta_{it}$ in her information set, leading to (1.2). All of $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$ can be scalar- or vector-valued [^2] In the baseline model, I assume in Assumption 1 that each coordinate of $g\left(\cdot\right)$ depends monotonically on a single coordinate of $\eta_{it}$ to invoke the control function argument. As an extension, I discuss how to relax this assumption to allow the entire vector $\eta_{it}$ to enter certain coordinates of $g\left(\cdot\right)$ at the end of Section 3., and any pair of the three variables can be correlated. The functional forms are potentially different across different coordinates of $\beta\left(\cdot\right)$ (and also $g\left(\cdot\right)$). As an analyst, I observe $\left\{X_{it},Y_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify the average partial effect $\mathbb{E}\beta_{it}$ (APE, see Graham and Powell (2012)) and the local average response function $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR, see Altonji and Matzkin (2005)).
A key feature of the TERC model is that $X_{it}$ can be correlated with $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$, potentially in a complicated manner. I refer to this feature as “time-varying endogeneity through the random coefficients.” Allowing such correlation is important to applications when agent $i$ in period $t$ has information about $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$ when optimally deciding $X_{it}$. I present three empirical examples in Section 2 that demonstrate the validity of such correlation in real-life scenarios. The associated technical challenge is that I need to control for the time-varying endogeneity through the random coefficients when both unknown (to the analyst) $A_{i}$ and $\eta_{it}$ appear in $g\left(\cdot\right)$ in a nonlinear and nonseparable way in (1.2). Furthermore, I do not specify the distributions of $A_{i}$, $\varepsilon_{it}$, or $\eta_{it}$, nor the functional form of $\beta\left(\cdot\right)$ or $g\left(\cdot\right)$. In this sense, the TERC model is a fixed effect panel data model (Laage (2024)). It has many important applications in economics; e.g., heterogeneous Cobb-Douglas production function estimation, return to schooling estimation, labor supply estimation, Engel curve analysis, and demand analysis, among others (Blundell, MaCurdy, and Meghir (2007b, a), Wooldridge (2009), Chernozhukov, Hausman, and Newey (2019), Li and Sasaki (2024), Keiller, de Paula, and Van Reenen (2024)).
[…]
The constructive identification analysis leads to easy-to-implement series estimators for the APE and LAR. I derive the convergence rates and establish the asymptotic normality of the proposed estimators. The inference results presented in this paper build on the literature on the asymptotic analysis of multi-step series estimators (Andrews (1991), Newey (1997), Imbens and Newey (2009)). The key departures from the existing literature include the fact that the object of interest is a partial mean process (Newey (1994)) of the derivative of the second-step estimator, and that the final step of the three-step estimator is an unknown but estimable functional of a conditional expectation function. Consequently, the estimation errors from each step must be properly accounted for to ensure accurate large sample properties.
As an empirical illustration, I estimate a Cobb-Douglas production function with heterogeneous output elasticities for Chinese manufacturing firms. My estimated APEs are broadly consistent with those derived from applying the classical methods with constant coefficients (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)) to the same data. I also estimate the LARs evaluated at the realized input levels for each firm and find significant variation across firms. [^4] Note that, with constant coefficient models, the LAR is equivalent to the APE. To support the empirical results, I conduct simulations motivated by production function estimation applications to evaluate the finite-sample performance of the estimators proposed in this paper. The results demonstrate that the proposed method performs well.
Next, I discuss the contributions and limitations of this paper. The first contribution concerns the features of the TERC model in (1.1)–(1.2). Allowing $X_{it}$ to be correlated with $\beta_{it}$ not only through $A_{i}$ but, more importantly, through $\eta_{it}$ significantly enhances the model’s flexibility and carries important empirical implications, as demonstrated by the examples discussed so far. Another novel aspect of the TERC model is the inclusion of $A_{i}$ in both (1.1) and (1.2). In particular, I argue that including $A_{i}$ in the first-step function $g\left(\cdot\right)$ better reflects the reality that $A_{i}$ is typically in the information set of agent $i$ when she optimally chooses the value of $X_{it}$, for instance, through repeated learning. Moreover, $A_{i}$ enters both (1.1) and (1.2) in a nonlinear and nonseparable way, which naturally arises from agent optimization behavior. Consequently, enabling these features enhances the model’s flexibility and broadens its applicability.
The second contribution is the proposal of a new panel data-based method to identify the APE and LAR of the TERC model. I construct a sufficient statistic $W_{i}$ to control for the time-invariant $A_{i}$ and a per-period conditional control variable $V_{it}$ to control for the time-varying shock $\eta_{it}$. I prove that, conditional on $V_{it}\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely attributable to the exogenous $Z_{it}$, thereby enabling the identification of the APE and LAR. The identification method proposed in this paper also contributes to the literature on control functions with multiple unobservables in nonseparable regression models (e.g., Blundell, Kristensen, and Matzkin (2013), Kasy (2014)).
The final contribution involves implementing the proposed method and characterizing its statistical properties. I develop three-step series estimators for the APE and LAR, establishing their convergence rates and asymptotic normality. The effectiveness of the method is demonstrated using a production dataset of Chinese manufacturing firms, revealing significant heterogeneity in capital and labor elasticities across firms. To further support the empirical findings, I conduct a simulation study inspired by applications in production function estimation, which produces satisfactory results.
I discuss two limitations of this paper. First, although both the unknown $A_{i}$ and $\eta_{it}$ determine the value of $X_{it}$, the baseline model imposes a restriction such that each coordinate of $g\left(\cdot\right)$ depends on a single coordinate of $\eta_{it}$ (as well as $Z_{it}\text{ and }A_{i}$). As noted by Imbens and Newey (2009), this restriction excludes nonseparable supply and demand models with one time-varying disturbance per equation. To relax this restriction, I propose two approaches under additional empirically motivated assumptions at the end of Section 3. Second, my method requires IVs. I provide a few examples of IVs for the applications discussed in this paper. However, depending on the specific application, identifying suitable IVs can still be challenging.
1.1 Related Literature
This paper contributes to the literature on CRC linear panel models. I highlight a few studies that are particularly relevant to this work. Chamberlain (1992) studies the regular identification of the first-order moments of the random coefficients when $T>d_{X}$, where $T$ is the number of periods and $d_{X}$ is the dimension of regressors. Wooldridge (2005) derives conditions for consistency of the fixed-effects estimators of the APE for a class of linear and nonlinear panel data models with time-invariant random coefficients. The condition requires the random coefficients to be mean independent of the residuals from regressing the covariates on the aggregate time variables. Graham and Powell (2012) prove irregular identification when $T=d_{X}$ and allow for more persistent processes in the regressors. They leverage different identifying content of the subpopulations of “movers” and “stayers” to identify the first-order moments of the time-varying random coefficients. Arellano and Bonhomme (2012) identify the first-order moments of the time-invariant random coefficients by using the within-group and between-group variation in the regressors. They rely on a mean independence assumption between the residuals and both the regressors and the random coefficients. They further identify the second-order moments and distribution functions of the random coefficients by exploiting assumptions on the time-dependence structure of the residuals. Graham, Hahn, Poirier, and Powell (2018) show how to identify quantiles of the random coefficients in a general linear quantile regression model. Masten (2018) considers a classical linear simultaneous equations model with random coefficients on the endogenous variables. By extending the Chamberlain-Mundlak approach for balanced panels, Wooldridge (2019) develops estimation strategies for CRC linear panel models that account for sample selection in unbalanced panels. Laage (2024) studies a time-invariant CRC linear panel data model with additive scalar-valued fixed effects. She proposes a two-step identification approach, where in the first step she identifies nonparametrically the conditional expectation of the disturbances given the regressors and the control variables, and then in the second step she uses between-group variation to identify the APE.
The key difference between my paper and the CRC literature discussed above is that my model, (1.1)–(1.2), features time-varying random coefficients that can be correlated with the regressors through both fixed effects and time-varying shocks. None of the papers, except Graham and Powell (2012), explicitly considers such correlation. Graham and Powell (2012) impose a time-stationarity assumption on the time-varying random shocks given the regressors for all periods and the fixed effect, and exploit the different identifying contents of the subpopulations for identification. The novel method proposed by Graham and Powell (2012) covers a wide range of empirical applications. However, their main time-stationarity assumption effectively rules out the time-varying endogeneity through the random coefficients, wherein $X_{it}$ and $\beta_{it}$ are correlated through both $A_{i}$ and $\eta_{it}$ [^5] To be precise, the time-stationarity assumption imposed by Graham and Powell (2012) is $\left.\eta_{it}\sim_{d}\eta_{is}\right|{\bf X}_{i},A_{i},\text{ for }t\neq s.$ To see why it rules out time-varying endogeneity through the random coefficients, consider a simplified example in which the number of periods $T=2$ and the data generating process (DGP) of a scalar-valued $X_{it}$ is $X_{it}=Z_{it}^{{}^{\prime}}\lambda+A_{i}+\eta_{it}$. Suppose one observes $X_{i2}>X_{i1}$ and $Z_{i2}=Z_{i1}$ in the data, which by the DGP of $X_{it}$ implies $\eta_{i2}>\eta_{i1}$ deterministically given ${\bf X}_{i}$ and $A_{i}$. Consequently, the time-stationarity condition is violated when the time-varying endogeneity through the random coefficients is present., which is the focus of this paper. Instead, I tackle this issue by proposing two controls for $A_{i}$ and $\eta_{it}$ and, subsequently, using the residual variation in $X_{it}$ to identify the APE and LAR. It should be clarified that my work is not necessarily more general than that of Graham and Powell (2012), as the set of assumptions used by each is not nested within the other’s. Instead, I consider this paper an initial step toward addressing the important issue of time-varying endogeneity through the random coefficients, which existing methods have not yet fully resolved.
In addition to the aforementioned papers, my work is also related to the literature on: (i) time-varying parameter panel data models (e.g., Li, Chen, and Gao (2011, 2016)), (ii) latent structure panel data models (e.g., Su, Shi, and Phillips (2016, 2019), Wang, Phillips, and Su (2024)), and (iii) linear panel data model with interactive fixed effects, which can be treated as imposing a specific structure on the random coefficient associated with the constant term (e.g., Pesaran (2006), Bai (2009), Moon and Weidner (2015)). The modeling technique and identification method of my paper are different from these seminal papers.
The second line of research concerns the techniques used in this paper. My method requires the existence of a sufficient statistic $W_{i}$ for the time-invariant $A_{i}$ and a feasible control variable $V_{it}$ for the time-varying $\eta_{it}$. To justify the sufficiency of $W_{i}$ for $A_{i}$, I present several techniques from the literature (e.g., Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)). See Liu, Poirier, and Shiu (2024) for a detailed discussion on the index sufficiency condition for the fixed effects in nonlinear semiparametric panel data models. It may be worth pointing out that the exchangeability condition used by Altonji and Matzkin (2005) concerns the conditional probability density function (PDF) of $A_{i}\text{ and }\eta_{it}$ given ${\bf X}_{i}$. I show that it is not directly applicable to the TERC model in which $X_{it}$ depends on $\eta_{it}$ for each $t$. Instead, I impose a lower-level exchangeability condition involving only the unobservable $A_{i}$ and $\eta_{it}$ and use it to establish a sufficiency result for $A_{i}$ in Proposition 1.
To find the feasible control variable for $\eta_{it}$ in (1.2), I generalize the technique of Imbens and Newey (2009) in two nontrivial ways. First, because the unknown $A_{i}$ enters equation (1.2) together with $\eta_{it}$, the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ proposed by Imbens and Newey (2009) does not uniquely determine $\eta_{it}$ even if it is a scalar. Instead, I use the index exclusion condition on $A_{i}$ to construct a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ given $A_{i}$ and $W_{i}$. Second, I allow $\eta_{it}$ to be vector-valued, enabling the model to capture the rich heterogeneity commonly observed in empirical applications. In the baseline model, I assume that each coordinate of $X_{it}$ depends upon a single coordinate of $\eta_{it}$. As an extension, I discuss how to relax this restriction when additional empirically motivated assumptions are imposed. Moreover, constructing $V_{it}$ requires IVs that satisfy specific exogeneity conditions. Murtazashvili and Wooldridge (2008) provide a set of conditions sufficient for the consistency of a general class of fixed effects IV estimators in a time-invariant individual-specific CRC panel data model. The model and method of this paper are different from theirs.
The asymptotic analysis in this paper builds on a series of foundational works. Andrews (1991) investigates the asymptotic properties of series estimators for nonparametric and semiparametric regression models. His results are applicable to a wide variety of estimands, including derivatives and integrals of the regression function. I use his results to cover vector-valued functionals of regression functions. Newey (1997) also studies series estimators and provides conditions for proving convergence rates and asymptotic normality for the estimators of conditional expectations. Newey, Powell, and Vella (1999) propose a two-step nonparametric estimator for a triangular simultaneous equation model with a separable first-step equation. They derive asymptotic normality for their two-step estimator with the first-step estimator plugged in. Imbens and Newey (2009) also analyze a triangular simultaneous equation model, but with a nonseparable first-step equation. They obtain mean-squared convergence rates for the first-step estimator and prove asymptotic normality (in their 2002 working paper version) for known functionals of the conditional expectation of the outcome variable given the regressors and control variables. I extend their analysis to establish asymptotic normality for unknown but estimable functionals of conditional expectation functions.
The rest of the paper is organized as follows. Section 2 formally introduces the TERC model and provides several empirical examples that fit within its framework. Section 3 outlines the identification idea and presents the assumptions, key identification theorem, and three extensions. Section 4 presents the series estimators for the APE and LAR and establishes their asymptotic properties. Section 5 provides an empirical illustration. Finally, Section 6 concludes. Proofs of all theorems, Proposition 1, and a simulation study are included in Appendix A, B, and C, respectively.
2 Model
Let $i\in\left\{1,...,n\right\}$ index $n$ agents and $t\in\left\{1,...,T\right\}$ index finite $T\geq 2$ periods. Let $d_{X}$ be the dimension of $X_{it}$ and similarly for $Z_{it},$ $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$. Consider the following baseline TERC model
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (2.1)
$X_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right),$ (2.2)
where:
  • $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{d_{X}}$, the central object of interest, is a vector of random coefficients modeled as a vector-valued, unknown, and potentially nonlinear function $\beta\left(\cdot\right)$ of $A_{i}\text{ and }\varepsilon_{it}$. Here, $A_{i}$ represents a fixed effect and $\varepsilon_{it}$ drives the time-varying behavior of $\beta_{it}$, both of which can be either scalar- or vector-valued. Furthermore, $A_{i}$ and $\varepsilon_{it}$ can be arbitrarily correlated.
  • $\eta_{it}\in\mathbb{R}^{d_{\eta}}$ is a continuously distributed and time-varying random vector that influences the value of $X_{it}$. $\eta_{it}$ is allowed to be correlated with both $A_{i}$ and $\varepsilon_{it}$. Additionally, different coordinates of $\eta_{it}$ can also be correlated.
  • $X_{it}\in\mathbb{R}^{d_{X}}$ is a vector of choice variables for agent $i$ in period $t$ [^6] For clarity of argument, in the baseline model I assume all coordinates of $X_{it}$ to be endogenous in the sense that they can be correlated with $\beta_{it}$ through both $A_{i}\text{ and }\eta_{it}$. In fact, the proposed method in this paper can benefit from additional exogenous regressors in (2.1). I discuss it in relation to Assumption 4 and note the changes needed when additional exogenous regressors are included in (2.1) at the end of the proof of Theorem 1. $Y_{it}\in\mathbb{R}$ denotes the scalar-valued outcome variable, and $Z_{it}\in\mathbb{R}^{d_{Z}}$ represents a vector of IVs.
  • $g\left(\cdot\right)$ is an unknown vector-valued function of $Z_{it},\ A_{i},\ \text{and }\eta_{it}$ that determines the value of each coordinate of $X_{it}$.
[…]
To clarify the model’s information structure, first nature draws $\left(A_{i},\varepsilon_{it}\right)$ for each agent and period in the economy. Then, agent $i$ in period $t$ chooses her endogenous $X_{it}$ by solving an optimization problem (e.g., profit maximization) after observing $A_{i}$, some signal $\eta_{it}$ about $\varepsilon_{it}$, and exogenously determined IVs $Z_{it}$, leading to (2.2). Finally, agent $i$ obtains its outcome $Y_{it}$ via (2.1). As an analyst, I observe $\left\{Y_{it},X_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify $\mathbb{E}\beta_{it}$ (APE) and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR). I illustrate the information structure in Figure 1.
Nature: draw ($A_{i}$,$\varepsilon_{it}$)Agent: choose $X_{it}$ given ($Z_{it}$,$A_{i}$,$\eta_{it}$)Agent: realize $Y_{it}$ via (2.1)Analyst: observe ($Y_{it}$,$X_{it}$,$Z_{it}$)
Figure 1: Information Structure
Model (2.1)–(2.2) appears in various economic applications. I describe three of them here.
Example 1 (Heterogeneous Production Function Estimation).
Suppose firm $i$ in year $t$ has a heterogeneous Cobb-Douglas production function in the form of (2.1). The capital and labor elasticities $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{2}$ are modeled as a two-dimensional function of the firm fixed effect $A_{i}$ (e.g., manager ability) and a time-varying random shock $\varepsilon_{it}$ (e.g., realized per-period technology shock). When deciding its input values $X_{it}\in\mathbb{R}^{2}$ for capital and labor, assume firm $i$ knows $A_{i}$ and $\eta_{it}$ (e.g., expected per-period technology shock). Clearly, $\eta_{it}$ is correlated with $A_{i}$ and $\varepsilon_{it}$. Assume firm $i$ also observes some exogenously determined input prices $Z_{it}$, such as interest rates and wages, and that the cost function is given by $c\left(x,z\right)$. Then, firm $i$ chooses input values $X_{it}$ by maximizing its expected profit with the knowledge of $\left(Z_{it},A_{i},\eta_{it}\right)$, i.e.,
$X_{it}=\arg\max_{x\in\mathbb{R}^{2}}\left[\mathbb{E}\left[\left.x^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right)\right|Z_{it},A_{i},\eta_{it}\right]-c\left(x,Z_{it}\right)\right],$
leading to (2.2). Subsequently, firm $i$ obtains its output $Y_{it}$ via (2.1).
Example 2 (Labor Supply Estimation).
Suppose agent $i$ has a linear labor supply function of the form (2.1), where $Y_{it}$ represents the number of hours worked and $X_{it}$ includes the endogenous hourly wage along with other exogenous demographic variables. The coordinate of $\beta_{it}$ corresponding to the wage is the key object of interest, as it quantifies how labor supply responds to variations in wage rate over time. Given some exogenous IVs $Z_{it}$ (e.g., county minimum wage and non-labor income), individual capability $A_{i}$, and expected health shocks $\eta_{it}$ (e.g., the likelihood of contracting Covid-19 at work), agent $i$ chooses a job with the wage that maximizes her expected utility, resulting in (2.2).
Example 3 (Engel Curve Estimation).
Suppose the budget share of gasoline $Y_{it}$ for household $i$ at time $t$ is a function of gas price and total expenditure represented in (2.1). Here, $\beta_{it}\in\mathbb{R}^{2}$ is modeled as a vector of unknown functions of the household fixed effect $A_{i}$ and a wealth shock $\varepsilon_{it}$, and captures the elasticity of gasoline demand with respect to total expenditure and gas price, respectively. Given $A_{i}$, a signal $\eta_{it}$ about the wealth shock $\varepsilon_{it}$, and an IV $Z_{it}$ (e.g., head of household’s gross income), household $i$ searches for a gas price and plans its total expenditure budget by maximizing its expected utility, leading to (2.2).
In these examples, the time-varying correlation between $X_{it}$ and $\beta_{it}$ highlights the significance and importance of the time-varying endogeneity through the random coefficients, which naturally arises due to agent optimization behavior. Numerous classic papers (Manski (1987), Altonji and Matzkin (2005), Graham and Powell (2012), Chernozhukov, Fernández-Val, Hahn, and Newey (2013)) have addressed such endogeneity by allowing for the correlation between $X_{it}$ and $A_{i}$. For instance, in production function estimation, a firm with a good management team is typically associated with larger input levels to leverage its managerial skill. This type of correlation has been addressed in the aforementioned papers. However, it carries important empirical implication to allow $X_{it}$ and $\beta_{it}$ to be correlated through factors beyond $A_{i}$. For instance, firms with a good management team may strategically select low levels of inputs for certain periods, which can not be explained only by the correlation between $X_{it}$ and $A_{i}$. To account for such behavior, it is useful to introduce an additional time-varying shock $\eta_{it}$ that contains important information about $\beta_{it}$ and affects the value of $X_{it}$.
Allowing $A_{i}$ to enter both the first-step equation (2.2) and the outcome equation (2.1) is another novel feature of this paper, distinguishing it from classic triangular simultaneous equations models (e.g., Newey, Powell, and Vella (1999), Imbens and Newey (2009)). In these models, the assumption is typically made on the first-step equation that determines the values of regressors that there is only one unobservable scalar-valued random variable $\eta_{it}$ that is correlated with both $A_{i}$ and $\varepsilon_{it}$, i.e., $X_{it}=g\left(Z_{it},\eta_{it}\right)$ instead of $X_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$. Under regularity conditions, they propose to control for $\eta_{it}$ via either the regression residual from regressing $X_{it}$ on $Z_{it}$ (e.g., Newey, Powell, and Vella (1999)) or the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ which equals $F_{\eta_{it}}\left(\eta_{it}\right)$ (e.g., Imbens and Newey (2009)). However, having $A_{i}$ appear in (2.2) renders the residual and CDF-based methods inapplicable because the one-to-one relationship between the control variable and $\eta_{it}$ no longer holds. Additionally, my model features a nonseparable first-step equation (2.2) in which $A_{i}$ enters $g\left(\cdot\right)$ potentially nonlinearly. The nonseparability and nonlinearity arise naturally from agent optimization behavior as illustrated in the empirical examples. The nonlinearity precludes the use of common demeaning or first-differencing techniques to eliminate $A_{i}$ (e.g., Laage (2024)). Instead, I rely on an index exclusion condition to deal with the fixed effects (e.g., Liu, Poirier, and Shiu (2024)).
3 Identification
3.1 Identification Idea
The main challenge to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ comes from the correlation between $X_{it}$ and $\beta_{it}$ in model (2.1)–(2.2) through their dependence on the pairwise correlated random variables $A_{i},\ \varepsilon_{it},$and $\eta_{it}$. Suppose I follow the standard identification argument for linear models by taking the conditional expectation of both sides of (2.1) given $X_{it}$:
$\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right].$ (3.1)
Then, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, the textbook methods to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ fail.
To see why, first I cannot identify $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by multiplying both sides of (3.1) by $X_{it}$, taking outer expectation with respect to $X_{it}$, and finally inverting $\mathbb{E}\left[X_{it}X_{it}^{{}^{\prime}}\right]$:
$\mathbb{E}\left[X_{it}Y_{it}\right]=\mathbb{E}\left\{X_{it}X_{it}^{{}^{\prime}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}}\mathinner{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left[\left.\beta_{it}\right|X_{it}\right]}\right\},$ (3.2)
since $X_{it}X_{it}^{{}^{\prime}}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$, a function of $X_{it}$, are not separable inside the expectation on the right hand side (RHS) of (3.2).
Second, the perturbation-based identification method in linear models does not go through either. This method perturbs $X_{it}$ on both sides of (3.1):
$\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]/\partial X_{it}=\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{{}^{\prime}}\partial\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]/\partial X_{it}}.$ (3.3)
Again, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, by the chain rule there is an additional term color-coded in red showing up on the RHS of (3.3), invalidating the perturbation-based method.
The key identification idea of this paper is to address the time-varying correlation between $X_{it}$ and $\beta_{it}$ using two control variables, $W_{i}$ and $V_{it}$. Conditional on $W_{i}$ and $V_{it}$, the residual variation in $X_{it}$ is determined solely by the exogenous $Z_{it}$. First, since the APE and LAR are generally not identified if the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted (Liu, Poirier, and Shiu (2024)), I exploit an index exclusion condition, $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$, to control for $A_{i}$, where the existence and construction of $W_{i}$ are supported by the established results in the literature (Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)). Then, for $\eta_{it}$ I propose a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and prove that $V_{it}=F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$, the RHS of which is a one-to-one function of $\eta_{it}$ given $A_{i}$ and $W_{i}$. Since $A_{i}$ is unknown, I apply the LIE to $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ by further conditioning on the unknown $A_{i}$ in the inner conditional expectation to fix $\eta_{it}$ via $V_{it}$, such that conditional on $A_{i},\ V_{it},\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely driven by the exogenous $Z_{it}$. This allows me to exclude $X_{it}$ from the inner conditional expectation. Next, I exclude $X_{it}$ from the conditioning set of the outer conditional expectation using the index exclusion condition of $W_{i}$ for $A_{i}$, i.e., $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$. Hence, I have
$\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|X_{it},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A_{i}},V_{it},W_{i}\right]\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right],$ (3.4)
which by (2.1) implies
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right].$ (3.5)
Given sufficient variation in $X_{it}$ conditional on $V_{it}$ and $W_{i}$, I identify $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ by either
$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right]$ (3.6)
or
$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}.$ (3.7)
Consequently, I use the LIE to identify
$\mathbb{E}\beta_{it}=\mathbb{E}\left[\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right]\text{ and }\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right|X_{it}\right].$ (3.8)
3.2 Assumptions and Main Identification Result
Next, I provide a list of assumptions on the model primitives required for the identification analysis and discuss them in relation to model (2.1)–(2.2).
Assumption 1 (Model).
Each coordinate of $g\left(Z_{it},A_{i},\eta_{it}\right)$ depends on one coordinate of $\eta_{it}$ strictly monotonically.
[…]
Assumption 1 requires justification when $d_{X}>1$, $d_{\eta}>1$, and each endogenous coordinate of $X_{it}$ admits a different coordinate of $\eta_{it}$. For example, it is possible that firms have different departments making hiring and investment decisions separately. Such decisions are based on different signals from the labor market and financial market. Therefore, it requires different coordinates of $\eta_{it}$ to enter the corresponding coordinate of $X_{it}$. Additionally, it is natural to allow firms to take advantage of positive shocks from labor market and financial market by hiring more employees and making more investments, leading to coordinate-wise strict monotonicity of $g\left(\cdot\right)$ in $\eta_{it}$. As a result, Assumption 1 is satisfied in this context.
It is worth highlighting that Assumption 1 does not restrict how different coordinates of $\eta_{it}$ are correlated. Thus, without loss of generality, I can redefine $\eta_{it}$ such that $d_{\eta}=d_{X}$ and each coordinate of $g\left(\cdot\right)$ depends on the corresponding coordinate of $\eta_{it}$. Moreover, Assumption 1 does not restrict how $A_{i}$ and $\eta_{it}$ are correlated, which could be important to empirical applications. Still, Assumption 1 restricts how the vector of $\eta_{it}$ enters the $g\left(\cdot\right)$ function that determines the value of $X_{it}$, which is a limitation of the paper. In Section 3.3, I discuss two other specifications to allow multiple coordinates of $\eta_{it}$ to enter one coordinate of $g\left(\cdot\right)$ when additional timing assumptions are considered plausible by practitioners and show how to adapt my method to these alternative specifications.
Assumption 2 (Index Exclusion).
Let $W_{i}\coloneqq W\left({\bf X}_{i},{\bf Z}_{i}\right)$, where $W:\mathbb{R}^{T\times\left(d_{X}+d_{Z}\right)}\mapsto\mathbb{R}^{d_{W}}$ is known. Suppose $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$ for any fixed $t$.
Assumption 2 is an index exclusion restriction that requires the conditional distribution of $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)$ to depend only on $W_{i}$ which is a known function of ${\bf X}_{i}$ and ${\bf Z}_{i}$. It is similar to Assumption 2.1 of Altonji and Matzkin (2005) and Assumption A2 of Liu, Poirier, and Shiu (2024). A similar condition is also imposed by Wooldridge (2019). Liu, Poirier, and Shiu (2024) prove a non-identification result when the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted in a nonlinear panel data model and suggest that restrictions on $A_{i}$ are generally required to identify the APE and LAR. Given $W_{i}$, the distribution of $A_{i}$ no longer depends on $X_{it}$ and $Z_{it}$ for any fixed $t$. In other words, $W_{i}$ is sufficient for $A_{i}$. Assumption 2 is used to (i) construct the conditional control variable for $\eta_{it}$, which I discuss in detail in relation to Assumption 3; and (ii) to control for $A_{i}$ when using the LIE argument to exclude $X_{it}$ from the conditioning set of $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ in (3.4), which I present in the proof of Theorem 1.
As pointed out by Liu, Poirier, and Shiu (2024), there are several ways to justify Assumption 2 from the literature. I follow their discussion to provide three justifications of Assumption 2. The first two methods are essentially parametric, resulting in a simpler expression for $W_{i}$, while the last one is nonparametric, leading to a more complex expression for $W_{i}$. This distinction highlights the trade-off between simplicity and generality. Moreover, I provide lower-level conditions for a nonparametric exchangeability condition and show that it is sufficient for Assumption 2 in Proposition 1.
First, Assumption 2 can be justified by generalizing equation (2.4) of Mundlak (1978) to be $A_{i}=h\left(T^{-1}\sum_{t=1}^{T}X_{it},\nu_{i}\right)$, where $\nu_{i}\perp\left(X_{it},Z_{it}\right)$ for any $t$. Then, Assumption 2 is satisfied by taking $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$. This is true because conditioning on $W_{i}$, any $t$-specific $X_{it}$ and $Z_{it}$ does not affect the distribution of $A_{i}$ as the residual randomness in $A_{i}$ is only driven by $\nu_{i}$. Notice that there is no need to specify the functional form of $h\left(\cdot\right)$ nor the distribution of $\nu_{i}$.
Second, insights from treatment assignment models in panel data (e.g., Arkhangelsky and Imbens (2022)) can be exploited to find candidate $W_{i}$. For example, if the random vector $\left(X_{it},Z_{it}\right)^{{}^{\prime}}\in\mathbb{R}^{2}$ conditional on $A_{i}$ is a two-dimensional normal random vector i.i.d. through time with known covariance [^7] The analysis can be extended to cover the case when the covariance matrix is unknown. The required change is to expand $W_{i}$ to include the averages of the second-order polynomials of $X_{it}$ and $Z_{it}$ through time. See Chapter 7 of Hogg, McKean, and Craig (2019) for more details. and ${\bf Z}_{i}\perp A_{i}$, then $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2. To see this, notice that the density of $\left.\left(X_{it},Z_{it}\right)\right|A_{i}$ depend on $A_{i}$ only through the mean of $X_{it}$. Hence, by the sufficient statistic for bi-variate normal random vectors with known covariance matrix (see Chapter 7 of Hogg, McKean, and Craig (2019)), $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ suffices for Assumption 2. Furthermore, sufficient conditions for the analysis above are (i) $\left.\eta_{i}\right|A_{i}$ is a normal random vector i.i.d. through time with known variance, (ii) ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$, and (iii) $g\left(\cdot\right)$ is linear in its arguments. Then, $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2.888Specifically, suppose $\left.\eta_{it}\right|A_{i}$ is i.i.d. normal across $t$ with mean $\mu\left(A_{i}\right)$ and known variance. Assume $X_{it}=Z_{it}^{{}^{\prime}}\gamma+A_{i}+\eta_{it}$ and ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$. Let $T=2$ for simplicity. Then, define $h_{t}=x_{t}-z_{t}^{{}^{\prime}}\gamma-a$ for $t=1\text{ and }2$ and by a change of variable argument, $f_{\left.\left(X_{i1},Z_{i1},X_{i2},Z_{i2}\right)\right|A_{i}}\left(\left.x_{1},z_{1},x_{2},z_{2}\right|a\right)$ $=f_{\left.\eta_{i1},\eta_{i2}\right|A}\left(\left.h_{1},h_{2}\right|a\right)f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right)$ $=\left[\Pi_{t=1}^{2}f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)\right]f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right).$ Using the formula of density of normal distribution for $f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)$ and applying the Fisher-Neyman factorization theorem yield the desired result.
Third, the nonparametric exchangeability condition (Assumption 2.3) from Altonji and Matzkin (2005) can also be adapted to justify Assumption 2. A similar idea is also used by Wooldridge (2019). Notice that the original exchangeability condition of Altonji and Matzkin (2005) is incompatible with the time-varying endogeneity through the random coefficients since it concerns the density of $\eta_{it}$ given ${\bf X}_{i}$, and I present how to adapt it to the TERC model below. To grasp the idea, suppose (i) $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ [^9] To be precise, by $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ I mean $f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{i1},...,h_{iT}\right|a_{i}\right)=f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{it_{1}},...,h_{it_{T}}\right|a_{i}\right)$ for any $\left(h_{i1},...,h_{iT}\right)$ belonging to the conditional support of $\eta_{i}$ given $A_{i}=a_{i}$, where $\left(t_{1},...,t_{T}\right)$ is any permutation of $\left(1,...,T\right)$. The definition is similar for $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$. where $\widetilde{\eta}_{i}$ is any permutation of the vector of $\eta_{i}$ in time, (ii) $Z_{it}\perp\left(\eta_{it},A_{i}\right)$, and (iii) $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}$ is continuous in $\left({\bf X}_{i},{\bf Z}_{i}\right)$. Then, combined with Assumption 1, it is straightforward to prove that $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$ where $\widetilde{{\bf X}}_{i}$ and $\widetilde{{\bf Z}}_{i}$ are permutations of ${\bf X}_{i}$ and ${\bf Z}_{i}$ in time, respectively. Then, following Altonji and Matzkin (2005), I use the Weierstrass approximation theorem and the fundamental theorem of symmetric functions to obtain that the vector $W_{i}$ of elementary symmetric functions of the paired elements of $\left\{\left(X_{i1},Z_{i1}\right),...,\left(X_{iT},Z_{iT}\right)\right\}$ satisfies Assumption 2. I provide the details of the above analysis in the proof of Proposition 1. To see an example of $W_{i}$ under the nonparametric exchangeability condition, suppose $T=2$ and both $X_{it}$ and $Z_{it}$ are scalars, and define $D_{it}\coloneqq\left(X_{it},Z_{it}\right)^{{}^{\prime}}$. Then, $W_{i}=\left(\sum_{t=1}^{T}D_{it},\sum_{t=1}^{T}D_{it}D_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ satisfies Assumption 2.101010In the concrete example discussed above, the maximum order of the polynomials in the construction of $W_{i}$ depends on $T$. In general, let $\varphi_{1}\left(u\right)=\sum_{t=1}^{T}u_{t}$, $\varphi_{2}\left(u,v\right)=\sum_{t\neq s}^{T}u_{t}v_{s}$, …, $\varphi_{T}\left(u,v,...,w\right)=\sum_{t\neq s\neq...\neq k}^{T}u_{t}v_{s}...w_{k}$, where $u,v,...,w$ are generic $T\times 1$ vectors. Then, I substitute each column vector of ${\bf D}_{i}=\left({\bf X}_{i},{\bf Z}_{i}\right)$ for each of the arguments $u,v,...,w$ (repetitions included) to construct $W_{i}$. Such a $W_{i}$ satisfies Assumption 2. See Chapter II.3 of Weyl (1939) for details of the construction and its proof. Notice that parametric assumptions on the distribution of $A_{i}$ are not needed for this result.
Once the fixed effect $A_{i}$ is controlled for by $W_{i}$, I still need to control for the time-varying $\eta_{it}$ that enters the $g\left(\cdot\right)$ function and is correlated with $\beta_{it}$ via its correlation with both $A_{i}$ and $\varepsilon_{it}$. When $A_{i}$ is not present in (2.2), I may follow the method of Imbens and Newey (2009) to use the conditional CDF $V_{it}^{\text{in09}}\coloneqq F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ as a control for $\eta_{it}$. This is true because Theorem 1 of Imbens and Newey (2009) proves that $V_{it}^{\text{in09}}=F_{\eta_{it}}\left(\eta_{it}\right)$, the RHS of which is a one-to-one mapping of $\eta_{it}$ when $d_{X}=d_{\eta}=1$ provided (i) $Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)$, (ii) $g\left(\cdot\right)$ is strictly monotonic in $\eta_{it}$, and (iii) $F_{\eta_{it}}\left(\cdot\right)$ is a strictly increasing function. Therefore, conditional on $V_{it}^{\text{in09}}$, $X_{it}$ and $\varepsilon_{it}$ are independent. The argument above can be generalized to the setting where $d_{X}>1$ by letting
$V_{it}^{\text{in09}}$ $=\left(F_{\left.X_{it}^{\left(1\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it}\right),...,F_{\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}\right)\right)^{{}^{\prime}}$
$=\left(F_{\eta_{it}^{\left(1\right)}}\left(\eta_{it}^{\left(1\right)}\right),...,F_{\eta_{it}^{\left(d_{X}\right)}}\left(\eta_{it}^{\left(d_{X}\right)}\right)\right)^{{}^{\prime}}.$
to control for the whole vector of $\eta_{it}$ coordinate by coordinate.
The challenge in applying their method to the TERC model arises from the existence of $A_{i}$ in (2.2), which is justified in many applications as agents usually know their $A_{i}$ (e.g., via repeated learning) when choosing the value of $X_{it}$. When $A_{i}$ is present in (2.2) and $d_{X}=1$, one may want to use $F_{\left.X_{it}\right|Z_{it},A_{i}}\left(\left.X_{it}\right|Z_{it},A_{i}\right)$ as the control variable for $\eta_{it}$, following a similar argument as in Imbens and Newey (2009). However, this approach is infeasible because $A_{i}$ is unknown. The idea of this paper is to replace $A_{i}$ with $W_{i}$ in the construction of $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ under a conditional independence assumption on $Z_{it}$ and a monotonicity condition on the conditional CDF $F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$ in $\eta_{it}$. I summarize them in the next assumption.
Assumption 3 (Control for $\boldsymbol{\eta_{it}}$).
Suppose the following conditions are satisfied:
  1. (a)
    $\left.Z_{it}\perp\left(\eta_{it},\varepsilon_{it}\right)\right|\left(A_{i},W_{i}\right)$.
  2. (b)
    $F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right)$ is strictly increasing in $\eta_{it}^{\left(l\right)}$ for every realization of $\left(A_{i},W_{i}\right)$ on its support and each $l=1,...,d_{X}$.
Assumption 3(a) concerns the vector of exogenous IVs $Z_{it}$. Since $W_{i}$ can be viewed as summarizing all the time-invariant information about $A_{i}$ in the data, this assumption, loosely speaking, requires $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|A_{i}$, which is already implied by the standard exogeneity assumption of $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$. For instance, when $d_{X}=1$ and suppose one follows Mundlak (1978) to assume $A_{i}=\left(\sum_{t=1}^{T}X_{it}\right)\alpha+\nu_{i}=W_{i}\alpha+\nu_{i}$, where $\nu_{i}$ is independent of everything else and $\alpha\neq 0$. Then, if $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it},\nu_{i}\right)$, I have $\left.Z_{it}\perp\left(\eta_{it},\varepsilon_{it}\right)\right|\left(A_{i},W_{i}\right)$ by the law of conditional independence because $W_{i}$ is a function of $A_{i}$ and $\nu_{i}$. [^11] I provide another parametric example where $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$ implies $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|\left(A_{i},W_{i}\right)$. For illustration purpose, suppress subscript $i$ and suppose $T=2$, $d_{X}=d_{Z}=1$, and $W=\left(X_{1}+X_{2},Z_{1}+Z_{2}\right)^{{}^{\prime}}$ satisfies Assumption 2 (Liu, Poirier, and Shiu (2024)). Furthermore, suppose $X_{t}=Z_{t}\gamma+A+\eta_{t}$ for $t=1,2$, and $\left.\left(Z_{1},Z_{2},\varepsilon_{1},\varepsilon_{2},\eta_{1},\eta_{2}\right)\right|A\sim N\left(\mu,\Sigma\right)$ where $\Sigma=\left[\begin{array}[]{cc}\Sigma_{1}&0_{2\times 4}\\ 0_{4\times 2}&\Sigma_{2}\end{array}\right]$ and $\Sigma_{1}$ and $\Sigma_{2}$ are arbitrary positive definite $2\times 2$ and $4\times 4$ matrices, respectively. Then, formula for the conditional density function of normally distributed random vectors shows that $\left.\left(Z_{t},\varepsilon_{t},\eta_{t}\right)\right|\left(A,W\right)$ is also normally distributed with diagonal covariance matrix for $t=1,2$, satisfying Assumption 3(a). Assumption 3(a) is used to ensure that $Z_{it}$ can be excluded from the conditioning set of $F_{\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}\right)$ for each $l=1,...,d_{X}$.
Assumption 3(b) is similar to condition (ii) of Theorem 1 in Imbens and Newey (2009). It is mild since it concerns a conditional CDF, which is typically strictly increasing for continuous random variables. Assumption 3(b) guarantees that fixing $\left(A_{i},V_{it},W_{i}\right)$ uniquely determines the value of $\eta_{it}$.
Assumption 3 allows me to construct a feasible control variable $V_{it}$ for $\eta_{it}$ given $A_{i}$ and $W_{i}$. To present it clearly, suppose $d_{X}=1$ and recall that $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$. Then,
$V_{it}$ $=F_{\left.X_{it}\right|Z_{it},A_{i},W_{i}}\left(\left.X_{it}\right|Z_{it},A_{i},W_{i}\right)=F_{\left.\eta_{it}\right|Z_{it},A_{i},W_{i}}\left(\left.\eta_{it}\right|Z_{it},A_{i},W_{i}\right)$
$=F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right),$ (3.9)
where the first equality holds by Assumption 2 which allows me to add $A_{i}$ to the conditioning set, the second equality holds by Assumption 1 and a change of variable argument, and the last equality holds by Assumption 3(a). Thus, $V_{it}$ uniquely determines $\eta_{it}$ given $A_{i}$ and $W_{i}$ by Assumption 3(b). Notice that $V_{it}$ is a one-to-one mapping of $\eta_{it}$ only when $A_{i}$ is also conditioned upon. In the proof of Theorem 1, I use the LIE on $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ where I first condition on $A_{i}$ to control for $\eta_{it}$ via $V_{it}$ in the inner expectation step. Then, I deal with the unknown $A_{i}$ with $W_{i}$ in the outer expectation step to prove (3.4). Finally, I use the residual variation in $X_{it}$, guaranteed by the next assumption, to identify the APE and LAR via (3.6) or (3.7).
The last assumption concerns the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$.
Assumption 4 (Residual Variation in $\boldsymbol{X_{it}}$).
There are at least $d_{X}$ linearly independent points in the support of $X_{it}$ a.s. given $V_{it}$ and $W_{i}$.
Assumption 4 guarantees that $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$ in (3.6) is invertible [^12] This is true because for any random vector $X\in\mathbb{R}^{d_{X}}$, $\mathbb{E}\left[XX^{{}^{\prime}}\right]$ is singular if and only if there exists a non-random vector $a\in\mathbb{R}^{d_{X}}\backslash\left\{0\right\}$ such that $a^{{}^{\prime}}X=0$ almost surely., thus the APE and LAR can be identified by (3.6) and the LIE. As discussed by Imbens and Newey (2009), requiring residual variation in $X_{it}$ conditioning on $V_{it}$ is generally not restrictive. Essentially, it requires $Z_{it}$ to move $X_{it}$ sufficiently well such that the set $\left\{X_{it}:V_{it}=v\text{ for }v\in\text{Supp}\left(V_{it}\right)\right\}$ is not a singleton. Therefore, it is similar to the relevance condition in the classical IV literature. [^13] I thank an anonymous referee for pointing this out. The restriction on the support of $X_{it}$ mainly comes from conditioning on $W_{i}$. When parametric justifications for Assumption 2 are used (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) so that $W_{i}$ only includes averages of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is not restrictive. However, when nonparametric justifications for Assumption 2 (e.g., the exchangeability condition of Altonji and Matzkin (2005)) are involved, Assumption 4 can be restrictive and generally requires that $T$ is larger than $d_{X}$. [^14] One way to justify Assumption 4 when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the $T^{\text{th}}$ order (Altonji and Matzkin (2005)) is to use the symmetry through time in the solution to $\left\{{\bf X}_{i}:V_{it}=v,W_{i}=w\right\}$. To see a concrete example, suppose $T=3$, $X_{it}\in\mathbb{R}^{2}$, $Z_{it}\in\mathbb{R}^{2}$, and $D_{it}\coloneqq\left(X_{it}^{{}^{\prime}},Z_{it}^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathbb{R}^{4}$. If $D_{i}=\left(D_{i1},D_{i2},D_{i3}\right)=\left(a,b,c\right)$ satisfies $V_{it}=v$ for $t=1,2$ and $W_{i}=w$ where all $a,b,c\in\mathbb{R}^{4}$ and the first two coordinates of $a$ and $b$ are linearly independent, by the symmetry of $W_{i}$ in $\left(X_{it},Z_{it}\right)$ through $t$, $\left(b,a,c\right)$ must also satisfy $V_{it}=v$ and $W_{i}=w$ for $t=1,2$. Therefore, given $V_{i1}=v$ and $W_{i}=w$ there are two linearly independent vectors of $\left(a^{\left(1\right)},a^{\left(2\right)}\right)$ and $\left(b^{\left(1\right)},b^{\left(2\right)}\right)$ that $X_{i1}$ can take, satisfying Assumption 4. It is worth pointing out that I only need to use the endogenous coordinates of $X_{it}$ to construct $V_{it}$ and $W_{i}$. Thus, when some coordinates of $X_{it}$ are exogenous in the sense that they are independent of $\left(A_{i},\eta_{it},\varepsilon_{it}\right)$, the unconditional variation in the exogenous coordinates of $X_{it}$ can be exploited to satisfy Assumption 4. See footnote 16 below for more details on this point.
Next, I present another assumption on the residual variation of $X_{it}$ given $V_{it}$ and $W_{i}$ to facilitate the derivative-based argument (3.7).
Assumption 4 (More Variation in $\boldsymbol{X_{it}}$).
The support of $X_{it}$ given $V_{it}$ and $W_{i}$ contains some ball of positive radius a.s. with respect to $\left(V_{it},W_{i}\right)$.
Assumption 4 is similar to Assumption 2.2 of Altonji and Matzkin (2005) and is stronger than Assumption 4. When parametric justifications (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) are used for Assumption 2 such that $W_{i}$ contains only mean of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is generally not restrictive. However, when nonparametric justifications (e.g., Altonji and Matzkin (2005)) are used for Assumption 2, Assumption 4 typically requires ex-ante information about $W_{i}$ (e.g., the existence of exogenous regressors in (2.1)). This is true because in this case $W_{i}$ includes averages of each endogenous coordinate of $X_{it}$, $Z_{it}$, and their interaction terms up to the $T^{\text{th}}$ order through time, conditioning on which there may not be enough residual variation in $X_{it}$. To address this issue, when nonparametric justifications for Assumption 2 are used, it is advisable to follow the recommendations by Altonji and Matzkin (2005) to ensure continuous residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$. [^15] Specifically, Altonji and Matzkin (2005) suggest to (i) leverage unconditional variation in the exogenous regressors that do not enter $W_{i}$, which I provide details in footnote 16; (ii) restrict the conditional distribution of $A_{i}$ given $W_{i}$ to depend on not all but only a subset of coordinates of $W_{i}$, which is testable by comparing the fit of $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ with $W_{i}$ replaced by a proper subset of it; (iii) assume $f_{\left.A_{i}\right|W_{i}}$ to depend on $W_{i}$ only through a linear combination of the elements of $W_{i}$, i.e., $f_{\left.A_{i}\right|W_{i}}=f_{\left.A_{i}\right|\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}}$ where $W_{i}=\left(W_{i}^{\left(1\right)},...,W_{i}^{\left(d_{W}\right)}\right)^{{}^{\prime}}$ and the $c_{k}$’s are unknown parameters, which implies that $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}\right]$ and the estimation of $c_{k}$’s can proceed by using the method of Ichimura and Lee (1991); (iv) impose a priori restrictions on model (2.1)–(2.2) such that $W_{i}$ enters $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ in a parametric way (e.g., $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}+\sum_{l=1}^{d_{V}}d_{l}V_{it}^{\left(l\right)}$ when $d_{X}=1$); and, (v) directly impose restrictions on $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$. [^16] To see a concrete example how the unconditional variations in the exogenous regressors help to satisfy Assumption 4, suppose $X_{it}^{\left(1\right)}$ is endogenous while $X_{it}^{\left(-1\right)}$, all other coordinates of $X_{it}$ except $X_{it}^{\left(1\right)}$, are exogenous (independent of $X_{it}^{\left(1\right)}$ and $\beta_{it}$). Suppose the conditional support of $X_{it}^{\left(1\right)}$ given $\left(V_{it},W_{i}\right)$ only contains a non-zero singleton, while the unconditional support of $X_{it}^{\left(-1\right)}$ contains some small ball of positive radius. Denote $b_{1}\left(V_{it},W_{i}\right)=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$. Then, I can identify the first-order moments of $\beta_{it}$ by taking partial derivative with respect to $X_{it}^{\left(-1\right)}$ on both sides of $\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]=X_{it}^{\left(1\right)}b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)+X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$ to identify $b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$. Finally, identification of $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)$ can be obtained as $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]-X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)\right)/X_{it}^{\left(1\right)}.$
Assumption 4 allows the straightforward perturbation-based method (3.7) to identify the APE and LAR without requiring the computation of the inverse of $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$. When residual variation is not a concern—such as when $W_{i}$ contains only the mean of $X_{it}$ through time or when exogenous regressors are included in (2.1)—Assumption 4 is preferred for simpler analysis. Accordingly, I impose it for the estimation and inference results in the next section.
I now state my main identification theorem.
Theorem 1 (Identification).
If Assumptions 13 and either Assumption 4 or 4 are satisfied, then $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ are identified.
3.3 Extensions
Vector-Valued $\boldsymbol{\eta_{it}}$
In my baseline model, I assume that each coordinate of $X_{it}$ depends on only one coordinate of $\eta_{it}$ (as well as $Z_{it}$ and $A_{i}$). Here, I discuss two ways to include vector-valued $\eta_{it}$ into coordinates of $X_{it}$ under additional assumptions motivated by empirical research. To clearly present the idea, I consider the case where $d_{X}=d_{\eta}=2$.
First, timing assumptions on the choice of regressors can be leveraged to allow vector-valued $\eta_{it}$ to enter certain coordinates of $X_{it}$. For instance, in production function applications, choices of later inputs such as capital typically do not depend on random shocks to earlier chosen inputs such as labor and/or material after conditioning on the level of the earlier chose inputs. [^17] Appendix 1 of the 2006 working paper version of Ackerberg, Caves, and Frazer (2015) discusses a similar idea. I thank an anonymous referee for this suggestion. If this timing assumption is imposed, I modify model (2.1)–(2.2) to be
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
$X_{it}^{\left(1\right)}$ $=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right),\text{ and }X_{it}^{\left(2\right)}=g^{\left(2\right)}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{\left(1\right)}},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right).$
Note that now $X_{it}^{\left(2\right)}$ depends on the whole vector of $\eta_{it}$ if $X_{it}^{\left(1\right)}$ is substituted into $g^{\left(2\right)}\left(\cdot\right)$. I propose to use $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}}\left(\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(2\right)}$ first. Then, I use $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(1\right)}$. The rest of the identification analysis follows as in this paper. [^18] Specifically, for this method to work, I maintain Assumptions 2, 3(a), and 4 (or 4) and change Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right)$ and $g^{\left(2\right)}\left(X_{it}^{\left(1\right)},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. I also modify Assumption 3(b) to be that $F_{\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
Second, the interaction between different coordinates of $X_{it}$ also provides useful information about certain coordinates of $\eta_{it}$. One way to exploit such information is to assume that while both coordinates of $\eta_{it}$ enter functions that determine each coordinate of $X_{it}$, the ratio of $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}$ may depend on only one coordinate of $\eta_{it}$, e.g., $\eta_{it}^{\left(2\right)}$ (and $Z_{it}$ and $A_{i}$). For example, in production function applications, the interpretation of $\eta_{it}$ could be that $\eta_{it}^{\left(1\right)}$ is the Hicks neutral productivity and $\eta_{it}^{\left(2\right)}$ is the labor-augmenting technology. Although both material and labor choices depend on both $\eta_{it}^{\left(1\right)}\text{ and }\eta_{it}^{\left(2\right)}$, it is plausible that material per worker depends only on the labor-augmenting technology $\eta_{it}^{\left(2\right)}$. See Proposition 2.1 of Demirer (2022) for more discussions about this assumption. Given this assumption, I can proceed to control for $\eta_{it}^{\left(2\right)}$ first by $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$, and then for $\eta_{it}^{\left(1\right)}$ by $V_{it}^{\left(1\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$. The rest of the identification analysis follows as in this paper. [^19] Specifically, suppose $X_{it}^{\left(1\right)}=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}=g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$. I maintain Assumptions 2, 3(a), and 4. I modify Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. I further change Assumption 3(b) to be $F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively. Then, I have $V_{it}^{\left(2\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right),$ and $V_{it}^{\left(1\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right),$ where the second equality holds because by Assumption (2), $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$ and that $V_{it}^{\left(2\right)}$ is a function of $\left(X_{it},Z_{it},W_{i}\right)$ and the last equality holds because $\left(V_{it}^{\left(2\right)},A_{i},W_{i}\right)$ uniquely determines $\eta_{it}^{\left(2\right)}$. Therefore, fixing $\left(X_{it},A_{i},V_{it},W_{i}\right)$ is equivalent to fixing $\left(X_{it},A_{i},\eta_{it},W_{i}\right)$. The rest of the argument follows as in the proof of Theorem 1.
Higher-Order Moments of $\boldsymbol{\beta_{it}}$
For clarity of exposition, suppose the regressors on the RHS of (2.1) include constant 1 and a scalar-valued $X_{it}$. It is straightforward to generalize the idea to cover $d_{X}$-dimensional $X_{it}$ for $d_{X}\geq 2$. With a slight abuse of notation, let $\left(\beta_{it},\omega_{it}\right)\in\mathbb{R}^{2}$ where $\beta_{it}$ is the random coefficient for $X_{it}$ and $\omega_{it}$ is for constant 1. Since the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$ is driven only by exogenous $Z_{it}$, (3.4) holds for any measurable function of $\beta_{it}$. Thus, I have
$\mathbb{E}\left[\left.\beta_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right],$ (3.10)
and similarly for $\mathbb{E}\left[\left.\omega_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ and $\mathbb{E}\left[\left.\omega_{it}\beta_{it}\right|X_{it},V_{it},W_{i}\right]$. Taking the conditional expectation of both sides of (2.1) squared given $\left(X_{it},V_{it},W_{i}\right)$ yields
$\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]+2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]+\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right].$ (3.11)
Then, I have
$\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]$ $=\left(\partial^{2}\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}^{2}\right)/2,$ (3.12)
$\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]$ $=\left(\partial\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}-2X_{it}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]\right)/2,\text{ and}$
$\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]-X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]-2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right],$
which identify the conditional (given $X_{it}$) and unconditional second-order moments of $\beta_{it}$ and $\omega_{it}$ by the LIE.
The above technique can also be used to identify the intertemporal correlations of the random coefficients. For example, $\mathbb{E}\left[\left.\beta_{it}\beta_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ can be identified from $\mathbb{E}\left[\left.Y_{it}Y_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ for any $t\neq s$ following an argument similar to (3.11)–(3.12).
Exogenous Shocks and Covariates
The identification argument of this paper goes through when I further include ex-post shocks $\upsilon_{it}$ to $\beta_{it}$ (i.e., $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$) and $\epsilon_{it}$ to $Y_{it}$ (i.e., $Y_{it}=X_{it}^{{}^{\prime}}\beta_{it}+\epsilon_{it}$), where both $\upsilon_{it}$ and $\epsilon_{it}$ are independent of all other variables with $\mathbb{E}\epsilon_{it}=0$. Note that due to the inclusion of $\upsilon_{it}$, $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ becomes a possibly time-varying function of $X_{it}$ as the distribution of $\upsilon_{it}$ may depend on time $t$. It does not affect my identification analysis because $\upsilon_{it}$ is independent of all other variables. I present the changes required to the proof due to the inclusion of $\upsilon_{it}$ into $\beta_{it}$ at the end of the proof of Theorem 1. I follow the more general specification $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$ in the estimation and examine its impact via simulations.
As for $\epsilon_{it}$, since it is independent of all other variables, has mean zero, and enters (2.1) in an additive way, equations (3.4)–(3.8) hold as before. Thus, the identification of $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is not affected. However, it has an impact on the identification of higher-order moments of $\beta_{it}$. For example, in (3.12) $\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ is not identified if $\epsilon_{it}$ is included with $\mathbb{E}\epsilon_{it}^{2}>0$. One way to handle this issue is by following Arellano and Bonhomme (2012) to impose and exploit additional assumptions on how $\epsilon_{it}$ evolves over time (e.g., $\epsilon_{it}$ follows an ARMA(p,q) process).
Finally, when exogenous covariates $Z_{1,it}$ are available and included in $X_{it}$ [^20] If $Z_{1,it}$ is included in $X_{it}$, I can write $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, $U_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$, and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$., my identification analysis actually benefits from it because Assumptions 13 concern only endogenous covariates $U_{it}$ of $X_{it}$. Thus, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 and in the construction of $V_{it}$. Furthermore, as discussed in footnote 16, the unconditional variation of $Z_{1,it}$ makes Assumption 4 and 4 easier to be satisfied. The details on how to adapt the current analysis to include $Z_{1,it}$ are provided at the end of the proof of Theorem 1.
4 Estimation and Large Sample Theory
In this section, I show how to estimate $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ via three-step series estimators based on the constructive identification argument. Next, I establish the convergence rates of the proposed estimators. Finally, I prove that the estimators are asymptotically normal and provide consistent estimators for their asymptotic covariance matrices.
4.1 Estimation
As discussed in Extension 3 of Section 3, I include the ex-post shock $\upsilon_{it}$ in $\beta_{it}$, resulting in $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]\eqqcolon b_{1t}\left(V_{it},W_{i}\right)$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]\eqqcolon b_{t}\left(X_{it}\right)$. The parameters of interest are
$\overline{b}\coloneqq\mathbb{E}\beta_{it}\ \text{and}\ b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right].$ (4.1)
$\overline{b}$ is the APE over the unconditional distribution of $\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$. $b_{t}\left(x\right)$ is the LAR function for a subpopulation characterized by $X_{it}=x$ in period $t$. Both objects are useful for answering policy-related questions. For example, plugging realizations $x_{it}$ of $X_{it}$ into $b_{t}\left(x\right)$ provides a fine approximation to $\beta_{it}$ for agent $i$ in time $t$. Therefore, distributional properties of $\beta_{it}$ may be inferred from the distribution of $b_{t}\left(x_{it}\right)$.
I propose to estimate the parameters in (4.1) with three-step series estimators. For clarity of exposition, I let $d_{X}=1$, and highlight the changes required for $d_{X}>1$ when necessary. To fix ideas, I follow Liu, Poirier, and Shiu (2024) to let $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$, which can be motivated by generalizing the method of Mundlak (1978).
First, for each $t$, I estimate $V_{t}\left(x,z,w\right)\coloneqq F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)$ by regressing $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ on the basis functions $q^{M_{1}}\left(\cdot\right)$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau\left(\cdot\right)$:
$\widehat{V}_{t}\left(x,z,w\right)$ $=\tau\left(\widehat{F}_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)\right)$
$=\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}n^{-1}\widehat{Q}_{t}^{-1}\sum_{j=1}^{n}q_{jt}\mathbf{\mathbbm{1}}\left\{x_{jt}\leq x\right\}\right)$
$\eqqcolon\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}\widehat{\gamma}_{t}^{M_{1}}\left(x\right)\right),$ (4.2)
[…]
Next, define $S\coloneqq\left(X,V,W\right)$ and let $\mathcal{X},\ \mathcal{Z},\ \mathcal{V},\ \mathcal{W},$ and $\mathcal{S}$ denote the supports of $X,\ Z,\ V,\ W,$ and $S$, respectively. Let $V_{it}\coloneqq V_{t}\left(X_{it},Z_{it},W_{i}\right),$$\widehat{V}_{it}\coloneqq\widehat{V}_{t}\left(X_{it},Z_{it},W_{i}\right)$, and $\widehat{v}_{it}\coloneqq\widehat{V}_{t}\left(x_{it},z_{it},w_{i}\right)$. For any $s=\left(x,v,w\right)\in\mathcal{S}$, I estimate $G_{t}\left(s\right)\coloneqq\mathbb{E}\left[\left.Y_{it}\right|S_{it}=s\right]$ by regressing $Y_{it}$ on the basis functions $p^{M_{2}}\left(\cdot\right)$ of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$:
$\widehat{G}_{t}\left(s\right)=p^{M_{2}}\left(s\right)^{{}^{\prime}}n^{-1}\widehat{P}_{t}^{-1}\widehat{p}_{t}^{{}^{\prime}}y_{t}\eqqcolon p^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$
where $\widehat{P}_{t}:=n^{-1}\sum_{i=1}^{n}\widehat{p}_{it}\widehat{p}_{it}^{{}^{\prime}}$, $\widehat{p}_{it}:=p^{M_{2}}\left(x_{it},\widehat{v}_{it},w_{i}\right)$, $\widehat{p}_{t}:=\left(\widehat{p}_{1t},...,\widehat{p}_{nt}\right)^{{}^{\prime}}$, and $y_{t}:=\left(y_{1t},...,y_{nt}\right)^{{}^{\prime}}$. Following Newey, Powell, and Vella (1999), I let $p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$ by exploiting the index structure of (2.1), which enables a faster convergence rate of $\widehat{G}_{t}\left(s\right)$ to $G_{t}\left(s\right)$.
Finally, I exploit the index structure of (2.1) again to estimate $b_{1t}\left(v,w\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|V_{it}=v,W_{i}=w\right]$. When (3.6) is used, I estimate $b_{1t}\left(v,w\right)$ by
$\widehat{b}_{1t}\left(v,w\right)=\left(\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it}=v,W_{i}=w\right]\right)^{-1}\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it}=v,W_{i}=w\right],$ (4.3)
where $\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it},W_{i}\right]$ is obtained by regressing each coordinate of $X_{it}Y_{it}$ on the basis functions of $\left(\widehat{V}_{it},W_{i}\right)$ and similarly for $\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it},W_{i}\right]$. When (3.7) is used, I estimate $b_{1t}\left(v,w\right)$ by
$\widehat{b}_{1t}\left(v,w\right)=\partial\widehat{G}_{t}\left(s\right)/\partial x=\left(I_{d_{X}}\otimes p^{m_{2}}\left(v,w\right)\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}}\eqqcolon\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$ (4.4)
where the second equality holds by the chain rule. I follow (4.4) in what follows for its simplicity in deriving the asymptotic properties.
To estimate $\overline{b}$ and $b_{t}\left(x\right)$, by the LIE I regress $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the basis function $r^{M_{3}}\left(\cdot\right)$ of constant one and $X_{it}$, respectively:
$\widehat{\overline{b}}$ $=\left(nT\right)^{-1}\sum_{t=1}^{T}\sum_{i=1}^{n}\widehat{b}_{1t}\left(\widehat{v}_{it},w_{i}\right),\text{ and}$
$\widehat{b}_{t}\left(x\right)$ $=r^{M_{3}}\left(x\right)^{{}^{\prime}}n^{-1}\widehat{R}_{t}^{-1}r_{t}^{{}^{\prime}}\widehat{B}_{t}\eqqcolon r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{\rho}_{t}^{M_{3}},$ (4.5)
where $\widehat{R}_{t}:=n^{-1}\sum_{i=1}^{n}r_{it}r_{it}^{{}^{\prime}}$, $r_{it}:=r^{M_{3}}\left(x_{it}\right)$, $r_{t}:=\left(r_{1t},...,r_{nt}\right)^{{}^{\prime}}$, and $\widehat{B}_{t}:=\left(\widehat{b}_{1t}\left(\widehat{v}_{1t},w_{1}\right),...,\widehat{b}_{1t}\left(\widehat{v}_{nt},w_{n}\right)\right)^{{}^{\prime}}$.
4.2 Convergence Rates and Asymptotic Normality
Since I let $n\rightarrow\infty$ for each fixed $t$ in the asymptotic analysis, the $t$-subscript is suppressed for brevity of exposition when there is no confusion. Let $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$ and $P\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}$. As Imbens and Newey (2009) have studied the convergence rates for the series estimators of conditional CDF (e.g., $F_{\left.X\right|Z}$) and conditional expectation functions (e.g., $\mathbb{E}\left[\left.Y\right|X,V\right]$), I adapt to their analysis to prove convergence rates for $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$. I impose the following assumption.
Assumption 5.
Suppose the following conditions hold:
  1. (a)
    $\left({\bf X}_{i},{\bf Z}_{i},A_{i},\varepsilon_{i},\eta_{i}\right)$ is i.i.d. across $i$.
  2. (b)
    $V\left(x,z,w\right)$ is continuously differentiable of order $d_{1}$ on the support with derivatives uniformly bounded in $\left(x,z,w\right)$ and $\mathcal{Z}\times\mathcal{W}\subset\mathbb{R}^{j_{1}}$.
  3. (c)
    $p^{m_{2}}\left(v,w\right)=p^{m_{2v}}\left(v\right)\otimes p^{m_{2w}}\left(w\right)$ and there exist constants $C,\theta>0$ such that $\lambda_{\min}\left(P\right)\geq C$ and
    $\inf_{w\in\mathcal{W}}f_{V,W}\left(v,w\right)\geq C\left[v\left(1-v\right)\right]^{\theta}.$
  4. (d)
    $G\left(s\right)$ is continuously differentiable of order $d_{2}>1$ on $\mathcal{S}\subset\mathbb{R}^{j_{2}}$.
  5. (e)
    $\mathbb{\mathbb{V}}\left(\left.Y\right|X,Z,W\right)$ is bounded.
Let $\zeta\left(m_{2}\right)\coloneqq m_{2v}^{\theta}m_{2}$ and $\zeta_{1}\left(m_{2}\right)\coloneqq m_{2v}^{\theta+2}m_{2}$. With Assumption 5 in position, the next lemma follows directly from Theorem 12 of Imbens and Newey (2009).
Lemma 1 (Convergence Rates of $\boldsymbol{\widehat{V}}$ and $\boldsymbol{\widehat{G}}$).
If the conditions of Theorem 1 and Assumption 5 are satisfied, and $m_{2}^{2}m_{2v}^{\theta+2}\left(n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}\right)\to 0$, then
$\mathbb{E}\left[n^{-1}\sum_{i}\left(\widehat{V}_{i}-V_{i}\right)^{2}\right]$ $=O\left(n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}\right)\eqqcolon O\left(\Delta_{1n}^{2}\right),$
[…]
$\sup_{s\in\mathcal{S}}\left|\widehat{G}\left(s\right)-G\left(s\right)\right|$ $=O_{p}\left(\zeta\left(m_{2}\right)\Delta_{2n}\right).$
Lemma 1 states that the mean-square convergence rate for $\widehat{G}\left(s\right)$ is the sum of the first-step rate $\Delta_{1n}^{2}$, the variance term $n^{-1}m_{2}$, and the squared bias term $m_{2}^{-2d_{2}/j_{2}}$. $d_{1}/j_{1}$ and $d_{2}/j_{2}$ are the uniform approximation rates that govern how well the unknown functions $V\left(x,z,w\right)$ and $G\left(s\right)$ can be approximated with $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$, respectively; see Assumption 3 and 5 of Imbens and Newey (2009) for the details. Note that even though the order of the basis function for the second-step estimation is $M_{2}$, by the index structure of (2.1) $M_{2}=d_{X}\times m_{2}$ and $d_{X}$ is finite. Thus, the effective order that matters for the convergence rate is $m_{2}$.
Let $\xi_{i}\coloneqq b_{1}\left(V_{i},W_{i}\right)-b\left(X_{i}\right)$ and $\xi\coloneqq\left(\xi_{1},...,\xi_{n}\right)^{{}^{\prime}}$. To derive the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, I impose the following assumption.
Assumption 6.
Suppose the following conditions hold:
  1. (a)
    $b\left(x\right)$ is continuously differentiable of order $d_{3}$ on $\mathcal{\mathcal{X}}\subset\mathbb{R}^{j_{3}}$.
  2. (b)
    There is a constant $C>0$ and $\zeta\left(M_{3}\right)$, such that for each $M_{3}$ there exists a non-singular constant matrix $N_{r}$ such that $\widetilde{r}^{M_{3}}\left(x\right)\coloneqq N_{r}r^{M_{3}}\left(x\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{r}^{M_{3}}\left(X_{i}\right)\widetilde{r}^{M_{3}}\left(X_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{x\in\mathcal{X}}\left\|\widetilde{r}^{M_{3}}\left(x\right)\right\|\leq C\zeta\left(M_{3}\right)$.
  3. (c)
    $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded.
Assumption 6 imposes conditions on the degree of smoothness of $b\left(x\right)$, the normalization of basis functions $r^{M_{3}}\left(x\right)$, and the boundedness of the second moment of $\xi_{i}$, similar to those in Assumption 5. Since $\widehat{b}_{1}\left(v,w\right)$ and $\widehat{G}\left(s\right)$ share the same series regression coefficient $\widehat{\alpha}^{M_{2}}$, the convergence rate of $\widehat{b}_{1}\left(v,w\right)$ to $b_{1}\left(v,w\right)$ is the same as that of $\widehat{G}\left(s\right)$ to $G\left(s\right)$. I use this result to prove the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, both of which are unknown but estimable functionals of $\widehat{G}\left(s\right)$.
Theorem 2 (Convergence Rates of $\boldsymbol{\widehat{\overline{b}}}$ and $\boldsymbol{\widehat{b}\left(x\right)}$).
If the conditions of Lemma 1 and Assumption 6 are satisfied, and $n^{-1}M_{3}\zeta\left(M_{3}\right)^{2}\to 0$, then
$\left\|\widehat{\overline{b}}-\overline{b}\right\|^{2}$ $=O_{p}\left(\Delta_{2n}^{2}\right),$
[…]
$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
The 2002 working paper version of Imbens and Newey (2009) (henceforth IN02) has obtained asymptotic normality for estimators of known and scalar-valued linear functionals of $G\left(s\right)$. I adapt to their results to analyze vector-valued functionals of $G\left(s\right)$ via a Cramér–Wold device and prove asymptotic normality for $\widehat{b}_{1}\left(v,w\right)$. With slight abuse of notation, I take $\widetilde{\overline{p}}^{M_{2}}\left(s\right)$ and $\widetilde{r}^{M_{3}}\left(x\right)$ that satisfy Assumptions 5 and 6 as $\overline{p}^{M_{2}}\left(s\right)$ and $r^{M_{3}}\left(x\right)$ in what follows.
Assumption 7.
Suppose the following conditions hold:
  1. (a)
    There is a constant $C>0$ and $\zeta\left(M_{1}\right)$, such that for each $M_{1}$ there exists a non-singular constant matrix $N_{q}$ such that $\widetilde{q}^{M_{1}}\left(z,w\right)\coloneqq N_{q}q^{M_{1}}\left(z,w\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{\left(z,w\right)\in\mathcal{Z\times\mathcal{W}}}\left\|\widetilde{q}^{M_{1}}\left(z,w\right)\right\|\leq C\zeta\left(M_{1}\right)$.
  2. (b)
    $G\left(s\right)$ is twice continuously differentiable with bounded first and second derivatives. For functional $a\left(\cdot\right)$ of $G$ and some constant $C>0$, it is true that $\left|a\left(G\right)\right|\leq C\sup_{s}\left|G\left(s\right)\right|$ and either (i) there is $\delta\left(s\right)$ and $\widetilde{\alpha}^{M_{2}}$ such that $\mathbb{E}\delta\left(S_{i}\right)^{2}<\infty$, $a\left(p_{m}^{M_{2}}\right)=\mathbb{E}\delta\left(S_{i}\right)p_{m}^{M_{2}}\left(S_{i}\right)$ for all $m=1,...,M_{2}$, $a\left(G\right)=\mathbb{E}\delta\left(S_{i}\right)G\left(S_{i}\right)$, and $\mathbb{E}\left(\delta\left(S_{i}\right)-p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)^{2}\rightarrow 0$; or (ii) for some $\widetilde{\alpha}^{M_{2}}$, $\mathbb{E}\left[p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right]^{2}\rightarrow 0$ and $a\left(p^{M_{2}}\left(\cdot\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)$ is bounded away from zero as $M_{2}\rightarrow\infty$.
  3. (c)
    $\mathbb{E}\left[\left.\left(Y-G\left(S\right)\right)^{4}\right|X,Z,W\right]$ is bounded and $\mathbb{\mathbb{V}}\left(\left.Y\right|X,Z,W\right)$ is bounded away from zero.
  4. (d)
    $nM_{1}^{1-2d_{1}/j_{1}}$, $nM_{2}^{-2d_{2}/j_{2}}$, $nM_{3}^{-2d_{3}/j_{3}}$, $n^{-1}M_{1}^{2}M_{2}\zeta_{1}\left(M_{2}\right)^{2}$, $n^{-1}M_{1}M_{3}\zeta_{1}\left(M_{2}\right)^{2}$, $n^{-1}M_{1}M_{3}\zeta\left(M_{3}\right)^{2}$, $n^{-1}M_{1}^{2}\zeta\left(M_{3}\right)^{2}$, $n^{-1}M_{2}\zeta\left(M_{3}\right)^{2}$, $n^{-1}M_{1}\zeta\left(M_{1}\right)^{4}\zeta\left(M_{2}\right)^{4}$, $n^{-1}M_{1}\zeta\left(M_{1}\right)^{4}\zeta\left(M_{3}\right)^{4}$, and $n^{-1}M_{1}^{4}\zeta\left(M_{2}\right)^{6}$ are all $o\left(1\right)$.
  5. (e)
    There exist $d_{4}$ and $\overline{\alpha}^{M_{2}}$ such that for each element $s_{j}$ of $s=\left(x,v,w\right)\in\mathcal{S}\subset\mathbb{R}^{j_{4}}$:
    $\sup\left\{\sup_{s\in\mathcal{S}}\left|G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right|,\sup_{s\in\mathcal{S}}\left|\partial\left(G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right)/\partial s_{j}\right|\right\}=O\left(M_{2}^{-d_{4}/j_{4}}\right).$
    Also, $nM_{2}^{-2d_{4}/j_{4}}$ and $M_{1}M_{2}^{-2d_{4}/j_{4}}\zeta_{1}\left(M_{2}\right)^{2}$ are $o\left(1\right)$.
  6. (f)
[…]
Assumptions 7(a)(e) are also imposed by IN02 and are regularity conditions required for the asymptotic normality of $\widehat{b}_{1}\left(v,w\right)$. Assumption 7(f) concerns the asymptotic covariance matrix of $\widehat{b}_{1}\left(v,w\right)$ and is used by Andrews (1991). It guarantees that the normality result of IN02 applies to vector-valued functionals of $G\left(s\right)$. Essentially, Assumption 7(f) requires that all the coordinates of $\widehat{b}_{1}\left(v,w\right)$ converge at the same speed, which is a mild assumption in my setting because ex-ante I do not distinguish any coordinate of $\beta_{it}$ from the others.
Lemma 2 (Asymptotic Normality of $\boldsymbol{\widehat{b}_{1}\left(v,w\right)}$).
If the conditions of Theorem 2 and Assumption 7 are satisfied, then,
[…]
where $\widehat{\Omega}_{1}$ is defined in (A.14).
Lemma 2 concerns $b_{1}\left(v,w\right)$, a known functional of $G\left(s\right)$. Therefore, the results of IN02 directly apply and I omit its proof in this paper. However, the results of IN02 do not directly apply to $\overline{b}$ and $b\left(x\right)$ because they are unknown functionals of $G\left(s\right)$. To explain, notice that by the LIE
$\overline{b}=\mathbb{E}\left[\partial G\left(S\right)/\partial X\right]\text{ and }b\left(x\right)=\mathbb{E}\left[\left.\partial G\left(S\right)/\partial X\right|X=x\right],$ (4.6)
both of which involve integrating $b_{1}\left(V,W\right)=\partial G\left(S\right)/\partial X$ with respect to the unknown but estimable distribution of $\left(V,W\right)$. Therefore, I need to estimate the unknown functionals in (4.6) and correctly account for the additional estimation bias in the asymptotic analysis.
Assumption 8.
Suppose the following conditions hold:
  1. (a)
    $\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}$ has full column rank.
  2. (b)
    $\mathbb{E}\left[\left.\left\|\xi\right\|^{4}\right|{\bf X}\right]$ is bounded and $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded away from zero.
  3. (c)
[…]
Assumption 8(a) is needed to show that the asymptotic covariance matrix $\Omega_{2}$ of $n^{1/2}\left(\widehat{b}\left(x\right)-b\left(x\right)\right)$ is positive definite. Assumption 8(b) is a regularity condition imposed for the Lindeberg–Feller central limit theorem (CLT). Assumption 8(c) is similar to Assumption 7(f) and is needed to prove that the asymptotic normality result holds for vector-valued functionals of $G\left(s\right)$.
Theorem 3 (Asymptotic Normality of $\boldsymbol{\widehat{\overline{b}}}$ and $\boldsymbol{\widehat{b}\left(x\right)}$).
If the conditions of Lemma 2 and Assumption 8 are satisfied, then
[…]
where $\widehat{\Omega}_{2}$ is defined in (A.45).
Furthermore, if $\mathbb{E}\left\|b_{1}\left(v,w\right)-\overline{b}\right\|^{4}<\infty$, then
[…]
where $\widehat{\Omega}_{3}$ is defined in (A.58).
5 Empirical Illustration
In this section, I apply my procedure to estimate a heterogeneous Cobb-Douglas production function for each of the five largest manufacturing sectors in China. I obtain unconditional means of the output elasticities and compare them with those derived using classic methods on the same data set. Furthermore, I estimate conditional means of the elasticities given the regressors. Results show that there are significant across-firm variations in the output elasticities. In Appendix C, I conduct a simulation study motivated by production function applications to support the findings of this illustration.
5.1 Data and Methodology
I use the China Annual Survey of Industrial Firms (CASIF), a longitudinal micro-level dataset collected by the National Bureau of Statistics of China that includes information on all state-owned industrial firms and non-state-owned firms with annual sales above 5 million RMB (~US\$770K). According to Brandt, Van Biesebroeck, Wang, and Zhang (2017), they account for 91% of the gross output, 71% of employment, 97% of exports, and 91% of total fixed assets in 2004, and thus are representative of industrial activities in China. Many papers on topics such as firm behavior, international trade, and growth theory have used the CASIF data (e.g., Hsieh and Klenow (2009), Brandt, Van Biesebroeck, Wang, and Zhang (2017)).
I focus on the five largest 2-digit sectors in terms of the number of firms between 2004 and 2007.222222There are 3140 textile firms, 2186 chemical firms, 1804 nonmetallic minerals firms, 2869 general equipment firms, and 1568 transportation equipment firms in the final dataset. I order the industries by their 2-digit Chinese Industry Classification codes. [^23] The CASIF dataset spans between 1998 and 2007. I choose year 2004 to 2007 to (i) ensure data consistency due to the change in the Chinese Industry Classification codes in 2003, (ii) avoid the financial crisis in the early 2000s, and (iii) use the most recent data. Following Brandt, Van Biesebroeck, and Zhang (2014), appropriate price deflators for inputs and outputs are applied separately. I preprocess the data so that firms with strictly positive amounts of capital, employment, value-added output, real wage expense, and real interest rate are used for estimation. The final dataset consists of a balanced panel of 11,567 firms over four years across five sectors. Summary statistics of the key variables are presented in Table 1.
Table 1: Summary Statistics
Variables N mean sd min max
[…]
Year 4 - - 2004 2007
Firm ID 11,567 - - - -
Industry Code 5 - - - -
The value-added production function I consider is
$y_{it}$ $=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
$\beta_{it}^{K}$ $=\beta^{K}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \beta_{it}^{L}=\beta^{L}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \omega_{it}=\omega\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),$
$k_{it}$ $=g^{K}\left(Z_{it},A_{i},\eta_{it}^{K}\right),\ l_{it}=g^{L}\left(Z_{it},A_{i},\eta_{it}^{L}\right).$ (5.1)
I highlight two features of model (5.1). First, output elasticities $\beta_{it}\coloneqq\left(\beta_{it}^{K},\beta_{it}^{L}\right)^{{}^{\prime}}$ are allowed to differ across firms and through time. Second, input choices $X_{it}\coloneqq\left(k_{it},l_{it}\right)^{{}^{\prime}}$ can be correlated with $\beta_{it}$ through their dependence on the pairwise correlated random variables $A_{i}$, $\eta_{it}$, and $\varepsilon_{it}$.
It is worth noting that the output is measured by the total revenue in dollars, rather than by the physical quantity produced in units, due to the absence of individual output prices in the data. When firms operate in distinct imperfectly competitive output markets, this may cause problems (Klette and Griliches (1996)). I use input prices $Z_{it}\coloneqq\left(r_{it},w_{it}\right)^{{}^{\prime}}$ as IVs for $X_{it}$. [^24] Real wages are very likely to be exogenous because the labor market in the manufacturing sectors of China is close to a perfectly competitive market in which firms face a relatively flat supply curve. Real interest rate may be considered as a valid IV because its fluctuation is mostly driven by monetary policy set by the central bank of China. For estimation, I take $W_{i}$ to be the mean through time of each coordinate of $X_{it}$. I estimate each coordinate of $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$ by regressing $\mathbf{\mathbbm{1}}\left(k\leq k_{it}\right)$ and $\mathbf{\mathbbm{1}}\left(l\leq l_{it}\right)$ on the second-degree polynomial spline basis of $\left(Z_{it},W_{i}\right)$ with its knot at the median, respectively. Next, I estimate $G_{it}\coloneqq\mathbb{E}\left[\left.y_{it}\right|X_{it},V_{it},W_{i}\right]$ by regressing $y_{it}$ on the same spline basis of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$. Estimation of $b_{1t}\left(V_{it},W_{i}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ is then obtained by taking the partial derivative of $\widehat{G}_{it}\left(X_{it},\widehat{V}_{it}W_{i}\right)$ with respect to $X_{it}$. Finally, I estimate $\overline{b}\coloneqq\mathbb{E}\beta_{it}$ by averaging $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ over $i$ and $t$ and $b_{t}\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by regressing $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the second-degree polynomial spline basis of $X_{it}$ with its knot at the median.
5.2 Results
First, I summarize the results for estimating the average output elasticities defined as $\widehat{\overline{b}}\coloneqq\left(\widehat{\overline{b}}^{K},\widehat{\overline{b}}^{L}\right)^{{}^{\prime}}$ in Table 2. I compare my estimated $\widehat{\overline{b}}$’s (TERC) with those derived by applying the methods of Olley and Pakes (1996) (OP), Levinsohn and Petrin (2003) (LP), and Ackerberg, Caves, and Frazer (2015) (ACF) to the same data set. Standard errors (se) and 95% confidence intervals (CI) are also included. [^25] I use the Stata commands prodest (Rovigatti and Mollisi (2018)) for OP and LP and acfest (Manjón and Manez (2016)) for ACF for point estimates of $\overline{b}$ and 95% CI. See Keiller, de Paula, and Van Reenen (2024) for a discussion of the implementation of the ACF method. For TERC, since $\widehat{\overline{b}}$ has slower than $n^{1/2}$ rate by Theorem 2, I follow Politis, Romano, and Wolf (1999) to randomly subsample $b=\lfloor 4n^{3/4}\rfloor$ firms from each sector without replacement for 1,000 times.
Table 2: Estimates of Average Output Elasticities
Textile OP LP ACF TERC
Capital Elasticity 0.359 0.252 0.300 0.387
se (0.030) (0.023) (0.037) (0.031)
95% CI [0.300, 0.417] [0.206, 0.297] [0.227, 0.373] [0.334, 0.455]
Labor Elasticity 0.470 0.175 0.567 0.375
se (0.014) (0.011) (0.065) (0.023)
95% CI [0.442, 0.498] [0.154, 0.196] [0.440, 0.695] [0.327, 0.418]
Chemical OP LP ACF TERC
Capital Elasticity 0.294 0.288 0.344 0.380
se (0.036) (0.031) (0.096) (0.037)
95% CI [0.223, 0.365] [0.228, 0.348] [0.155, 0.533] [0.306, 0.451]
Labor Elasticity 0.296 0.113 0.378 0.329
se (0.022) (0.015) (0.213) (0.026)
95% CI [0.253, 0.339] [0.084, 0.143] [-0.039, 0.795] [0.276, 0.377]
Nonmetallic Mineral OP LP ACF TERC
Capital Elasticity 0.697 0.311 0.236 0.379
se (0.101) (0.024) (0.101) (0.037)
95% CI [0.499, 0.895] [0.263, 0.358] [0.038, 0.434] [0.314, 0.451]
Labor Elasticity 0.353 0.071 0.601 0.459
se (0.021) (0.010) (0.384) (0.027)
95% CI [0.311, 0.394] [0.051, 0.091] [-0.151, 1.354] [0.404, 0.514]
General Equipment OP LP ACF TERC
Capital Elasticity 0.416 0.246 0.176 0.504
se (0.076) (0.023) (0.082) (0.033)
95% CI [0.267, 0.565] [0.202, 0.291] [0.016, 0.337] [0.451, 0.578]
Labor Elasticity 0.444 0.071 0.927 0.359
se (0.019) (0.009) (0.149) (0.020)
95% CI [0.406, 0.482] [0.053, 0.089] [0.635, 1.218] [0.320, 0.397]
Transportation Equipment OP LP ACF TERC
Capital Elasticity 0.523 0.281 0.217 0.383
se (0.064) (0.033) (0.084) (0.035)
95% CI [0.397, 0.649] [0.216, 0.346] [0.052, 0.381] [0.327, 0.469]
Labor Elasticity 0.523 0.137 1.042 0.460
se (0.025) (0.019) (0.138) (0.028)
95% CI [0.473, 0.573] [0.100, 0.173] [0.771, 1.313] [0.407, 0.517]
My estimated average capital elasticities are within the range of $\left[0.379,0.504\right]$ across the five sectors, which are consistent with the results obtained from applying OP, LP, and ACF to the same dataset. The estimated average labor elasticities present more discrepancies between methods. My estimates of $\widehat{\overline{b}}^{L}$’s lie between $\left[0.329,0.460\right]$ across the five sectors, similar to OP’s results. LP seem to derive small $\widehat{\overline{b}}^{L}$’s for all sectors, while ACF generate larger $\widehat{\overline{b}}^{L}$’s for the general equipment sector and transportation equipment sector. My 95% CI’s for $\widehat{\overline{b}}$’s are reasonably tight across all sectors. It is well documented in the literature that output elasticities in Cobb-Douglas production function estimation usually lie within $\left[0,1\right]$. Hsieh and Klenow (2009) also calculate that roughly half of output is distributed to capital, according to Chinese input-output tables and national accounts. My estimates are consistent with the empirical evidence in the literature.
Next, I investigate the distribution of the conditional means of the output elasticities given the regressors. Policymakers could use the information contained in the conditional means to formulate more accurate sector specific policies, as these conditional means provide a fine approximation to the average elasticities for each subgroup of firms with a certain level of capital and labor. Moreover, distributional properties about the true output elasticities can be inferred from the distribution of the conditional means of the output elasticities. For example, by the law of total variance, the variance of the conditional means of the output elasticities provides a lower bound for the variance of the true output elasticities. Demirer (2022) argues that the heterogeneity in output elasticities is largely explained by across-firm variation. To investigate into this explanation, for each firm I calculate the averages of $\widehat{b}_{t}^{K}\left(X_{it}\right)$ and $\widehat{b}_{t}^{L}\left(X_{it}\right)$ through time and denote them by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$, respectively. Then, I plot the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across firms for each of the five sectors.
Figure 2: Distribution of Conditional Means of Output Elasticities: Textile
In Figure 2, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the textile sector and indicate the corresponding $\widehat{\overline{b}}$’s labeled by “Mean” on the same graph. The left subplot of Figure 2 shows the histogram of $\widehat{\beta}_{i}^{K}$. All of its probability mass lies between zero and one, with its mode at 0.4. Over 90% of the probability mass of $\widehat{\beta}_{i}^{K}$ lies between 0.3 and 0.5. The right subplot of Figure 2 presents the histogram of $\widehat{\beta}_{i}^{L}$ across firms for the textile sector. Again, the majority of its probability mass lies between zero and one, with its mode at 0.38. The distributions of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ are concentrated around the mean, suggesting that firms in the textile industry tend to be homogeneous in terms of capital and labor efficiency.
Figure 3: Distribution of Conditional Means of Output Elasticities: Other Sectors
In Figure 3, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the other four sectors. First, the majority of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across all sectors lie between zero and one, which is consistent with empirical evidence. Second, I find significant across-firm variations in both capital and labor elasticities measured by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ within each of the five sectors. However, the degree of heterogeneity in these elasticities differs. For example, in the left subplot of Figure 3, there is a larger variance in $\widehat{\beta}_{i}^{K}$ in the nonmetallic mineral sector than in the chemical sector, which implies that the firms in the chemical industry may be more homogeneous in capital efficiency than those in the nonmetallic mineral sector.
6 Conclusion
In this paper, I propose a new TERC model in which the regressors are correlated with the random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock—a feature that aligns with agent optimization behavior in many empirical applications. I construct feasible control variables for both the fixed effect and the random shock, and subsequently utilize the residual variation in the regressors given the control variables to identify the APE and LAR. I provide three-step series estimators and establish their convergence rates and asymptotic normality. The empirical exercise of this paper reveals significant variation in the output elasticities across manufacturing firms in China.
I propose two directions for future research. First, beyond the first- and second-order moments analyzed in this paper, policymakers may also be interested in the density functions of the random coefficients. One possible, albeit laborious, approach is to identify the moments of all orders of the random coefficients by induction, which would uniquely determine their distribution (Stoyanov (2000)). Second, it remains an open question whether aspects of my method could be extended to dynamic linear or nonlinear panel data models, such as those discussed by Marx, Tamer, and Tang (2024) and Liu, Poirier, and Shiu (2024). Such an extension would likely require additional structure on how the lagged dependent variable correlates with the random coefficients.
Appendix A Proofs
Proof of Theorem 1.
I first prove that $V_{it}$ is a control for $\eta_{it}$ given $A_{i}$ and $W_{i}$. Then, I show $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ does not depend on $X_{it}$ via the LIE. Finally, I identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by leveraging the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$. I show how the inclusion of exogenous shocks $\upsilon_{it}$ and $\epsilon_{it}$ as well as exogenous coordinates in the regressors affects the analysis at the end of this proof.
Without loss of generality, I assume that $d_{\eta}=d_{X}$ and that each coordinate of $\eta_{it}$ enters the corresponding coordinate of $X_{it}$. [^26] This is without loss of generality because under Assumption 1, I can always redefine $\widetilde{\eta}$ to be the vector that collects the coordinates of $\eta$ that enters each coordinate of $g\left(\cdot\right)$ function. By Assumption 2, I have $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$, which implies $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$. Thus, for each $l\in\left\{1,...,d_{X}\right\}$ and any on-support $\left(x^{\left(l\right)},z,a,w\right)$, I have
$\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.x^{\left(l\right)}\right|z,w\right)$
$=$ $\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.x^{\left(l\right)}\right|z,a,w\right)$
$=$ $\ \mathbb{P}\left(\left.g^{\left(l\right)}\left(z,a,\eta_{it}^{\left(l\right)}\right)\leq x^{\left(l\right)}\right|Z_{it}=z,A_{i}=a,W_{i}=w\right)$
$=$ $\ \mathbb{P}\left(\left.\eta_{it}^{\left(l\right)}\leq g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|A_{i}=a,W_{i}=w\right)$
$=$ $\ F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|a,w\right),$ (A.1)
where the first equality holds by $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$, the second uses (2.2), the third holds by Assumptions 1 and 3(a), and the last holds by definition of the conditional CDF of $\eta_{it}^{\left(l\right)}$ given $\left(A_{i},W_{i}\right)$. By (2.2), the random variable $\eta_{it}^{\left(l\right)}=g^{\left(l\right)^{-1}}\left(X_{it}^{\left(l\right)},Z_{it},A_{i}\right)$ for each $l$, so that plugging in gives
$V_{it}^{\left(l\right)}\coloneqq F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}\right)=F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right),$ (A.2)
which establishes a one-to-one mapping to $\eta_{it}^{\left(l\right)}$ given $A_{i}$ and $W_{i}$ under Assumption 3(b). Let $V_{it}\coloneqq\left(V_{it}^{\left(1\right)},...,V_{it}^{\left(d_{X}\right)}\right)^{{}^{\prime}}$. By (A.2), $V_{it}$ uniquely determines the vector of $\eta_{it}$ given $A_{i}$ and $W_{i}$.
Next, I have
$\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it}\right)\right|g\left(Z_{it},A_{i},\eta_{it}\right),A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it}\right)\right|A_{i},V_{it},W_{i}\right]$
$\eqqcolon$ $\ b_{2}\left(A_{i},V_{it},W_{i}\right),$ (A.3)
where the second equality holds because conditioning on $\left(A_{i},V_{it},W_{i}\right)$ is equivalent to fixing $\left(A_{i},\eta_{it},W_{i}\right)$ by (A.2), and thus the residual variation in $X_{it}$ is driven solely by $Z_{it}$ which is independent of $\varepsilon_{it}$ given $\left(A_{i},\eta_{it},W_{i}\right)$ by Assumption 3(a). I exclude $X_{it}$ from the conditioning set of $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$ and write it out as $b_{2}\left(A_{i},V_{it},W_{i}\right)$.
By (A.3), I have
$\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.b_{2}\left(A_{i},V_{it},W_{i}\right)\right|X_{it},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\mathbb{E}\left[\left.b_{2}\left(A_{i},V_{it},W_{i}\right)\right|X_{it},Z_{it},V_{it},W_{i}\right]\right|X_{it},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\mathbb{E}\left[\left.b_{2}\left(A_{i},V_{it},W_{i}\right)\right|X_{it},Z_{it},W_{i}\right]\right|X_{it},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\int b_{2}\left(a,V_{it},W_{i}\right)f_{\left.A_{i}\right|X_{it},Z_{it},W_{i}}\left(\left.a\right|X_{it},Z_{it},W_{i}\right)\mu\left(da\right)\right|X_{it},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\int b_{2}\left(a,V_{it},W_{i}\right)f_{\left.A_{i}\right|W_{i}}\left(\left.a\right|W_{i}\right)\mu\left(da\right)\right|X_{it},V_{it},W_{i}\right]$
$\eqqcolon$ $\ b_{1}\left(V_{it},W_{i}\right),$ (A.4)
where the first and second equalities hold by the LIE, the third holds because $V_{it}$ is a measurable function of $X_{it},\ Z_{it},$ and $W_{i}$, all of which are also conditioned on, and the fifth equality holds by $\left.\left(X_{it},Z_{it}\right)\perp A_{i}\right|W_{i}$ by Assumption 2.
Finally, given (A.4), I have
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}b_{1}\left(V_{it},W_{i}\right).$ (A.5)
When Assumption 4 is satisfied, I pre-multiply both sides of (A.5) by $X_{it}$ and take conditional expectation of both sides conditioning on $V_{it}$ and $W_{i}$:
$\mathbb{E}\left[\left.\mathbb{E}\left[\left.X_{it}Y_{it}\right|X_{it},V_{it},W_{i}\right]\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]b_{1}\left(V_{it},W_{i}\right),$
which identifies $b_{1}\left(V_{it},W_{i}\right)$
$b_{1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]\right)^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right].$
When Assumption 4 is used, I take partial derivative of both sides of (A.5) with respect to $X_{it}$ and obtain
$b_{1}\left(V_{it},W_{i}\right)=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}.$
Given $b_{1}\left(V_{it},W_{i}\right)$, I use the LIE to identify $\overline{b}=\mathbb{E}\beta_{it}$ and $b\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by
$\overline{b}=\mathbb{E}\left[b_{1}\left(V_{it},W_{i}\right)\right]\text{ and }b\left(X_{it}\right)=\mathbb{E}\left[\left.b_{1}\left(V_{it},W_{i}\right)\right|X_{it}\right],$ (A.6)
respectively.
When ex-post shock $\epsilon_{it}$ is included additively in (2.1) and $\mathbb{E}\epsilon_{it}=0$, equations (A.4) and (A.5) still hold, so the identification result (A.6) holds without any changes. When ex-post shock $\upsilon_{it}$ is also included in $\beta_{it}\coloneqq\beta\left(A_{i},\eta_{it},\upsilon_{it}\right)$, I need to change $b_{2}\left(A_{i},V_{it},W_{i}\right)$ in (A.3) to be $b_{2t}\left(A_{i},V_{it},W_{i}\right)$ because although $\upsilon_{it}$ is independent of all other variables, the PDF of $\upsilon_{it}$ itself may depend on $t$, i.e.,
$\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|g\left(Z_{it},A_{i},\eta_{it}\right),A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|A_{i},V_{it},W_{i}\right]$
$\eqqcolon$ $\ b_{1{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t}}\left(A_{i},V_{it},W_{i}\right).$ (A.7)
As a result, I need to change $b\left(X_{it}\right)$ to be $b_{t}\left(X_{it}\right)$ in (A.6). The rest of the proof goes through as before.
When exogenous regressors are included in (2.1), I can define $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, and rewrite model (2.1)–(2.2) to be
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
$U_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right).$
Then, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 as well as in the definition of $V_{it}$, and the analysis goes through as before. Note that I keep $X_{it}$ in Assumption 4 or 4 because I need to perturb the whole $X_{it}$ vector instead of just endogenous coordinates $U_{it}$ in (A.5) to identify $b_{1}\left(V_{it},W_{i}\right)$. As discussed in footnote 16, the residual variation requirement of Assumption 4 or 4 is easier to be satisfied when exogenous $Z_{1,it}$ is included in $X_{it}$ as the support of $Z_{1,it}$ is unaffected by $V_{it}$ and $W_{i}$. ∎
Proof of Theorem 2.
I denote $\sum_{i=1}^{n}$ by $\sum_{i}$ and omit all $t$-subscripts. As the result for a finite-dimensional vector-valued $\beta$ can be established by proving for each of its coordinates and combining the results using the triangle inequality, I assume $\beta$ is a scalar in this proof. I focus on $\widehat{b}\left(x\right)$, since the result for $\widehat{\overline{b}}$ follows immediately by setting $r^{M_{3}}\left(\cdot\right)\equiv 1$. Let $p_{i}^{m_{2}}\coloneqq p^{m_{2}}\left(V_{i},W_{i}\right)$, $r_{i}\coloneqq r^{M_{3}}\left(X_{i}\right)$, and $\overline{p}_{i}=\overline{p}^{M_{2}}\left(S_{i}\right)$. Following Imbens and Newey (2009), I can normalize $\mathbb{E}p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}=I_{m_{2}}$ by Assumption 5 and $R\coloneqq\mathbb{E}r_{i}r_{i}^{{}^{\prime}}=I_{M_{3}}$ by Assumption 6, which imply $\overline{P}\coloneqq\mathbb{E}\overline{p}_{i}\overline{p}_{i}^{{}^{\prime}}=\mathbb{E}\left[I_{d_{X}}\otimes\left(p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}\right)\right]=I_{M_{2}}$. Furthermore, I have $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ with probability approaching one by Newey (1997). Let $\widetilde{B}\coloneqq\left(b_{1}\left(\widehat{v}_{1},w_{1}\right),...,b_{1}\left(\widehat{v}_{n},w_{n}\right)\right)^{{}^{\prime}}$.
By (4.5),
$\ \left\|n\widehat{R}^{1/2}\left(\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right)\right\|^{2}/4$
$\leq$ $\ \left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)+\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
$\ +\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)+\left(B^{X}-r\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-r\rho^{M_{3}}\right),$ (A.8)
where $B\coloneqq\left(b_{1}\left(v_{1},w_{1}\right),...,b_{1}\left(v_{n},w_{n}\right)\right)^{{}^{\prime}}$ and $B^{X}\coloneqq\left(b\left(x_{1}\right),...,b\left(x_{n}\right)\right)^{{}^{\prime}}$. Hence, $\xi=B-B^{X}$. I analyze the RHS of (A.8) term by term.
By Lemma S.5 of Imbens and Newey (2009), $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$. Then,
$\ n^{-2}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)\leq Cn^{-1}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
$=$ $\ Cn^{-1}\sum_{i}\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)\right)^{2}$
$\leq$ $\ C\left\|\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right\|^{2}+C\sup_{s\in\mathcal{S}}\left\|\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(v,w\right)\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}\right),$ (A.9)
where the first inequality holds because $n^{-1}r\widehat{R}^{-1}r^{{}^{\prime}}$ is idempotent, the last inequality holds by the Cauchy-Schwarz inequality (CS) and $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$, and the last equality uses Lemma 1.
Next,
$n^{-2}\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)\leq$ $\ Cn^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
$\leq$ $\ Cn^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}=O_{p}\left(\Delta_{1n}^{2}\right),$ (A.10)
where the last inequality holds by the mean value theorem and Assumption 5 and the equality holds by Lemma 1 and Markov’s inequality.
Finally, for the last two terms on the right-hand side of (A.8),
$\ n^{-2}\mathbb{E}\left[\left.\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)\right|{\bf X}\right]$
$=$ $\ n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\xi\right|{\bf X}\right]\right\}=n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\widehat{R}^{-1}r^{{}^{\prime}}\right\}$
$\leq$ $\ n^{-2}\text{tr}\left\{CIr\widehat{R}^{-1}r^{{}^{\prime}}\right\}=Cn^{-1}\text{tr}\left\{\widehat{R}^{-1}\widehat{R}\right\}=Cn^{-1}M_{3},$
and
$n^{-2}\left(B^{X}-R\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-R\rho^{M_{3}}\right)\leq$ $\ n^{-1}\left\|B^{X}-R\rho^{M_{3}}\right\|^{2}=O_{p}\left(M_{3}^{-2d_{3}/j_{3}}\right).$ (A.11)
By $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ and the result that $\mathbb{E}\left[\left.\left|Y_{n}\right|\right|Z_{n}\right]=O_{p}\left(r_{n}\right)$ implies $Y_{n}=O_{p}\left(r_{n}\right)$ (Conditional Markov’s inequality, henceforth denoted as “CM”),
$\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}+n^{-1}M_{3}+M_{3}^{-2d_{3}/j_{3}}\right)\eqqcolon O_{p}\left(\Delta_{3n}^{2}\right),$ (A.12)
which yields
[…]
$\leq$ $\ 2\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}+2\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|^{2}=O_{p}\left(\Delta_{3n}^{2}\right),$
and
$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $\leq\sup_{x\in\mathcal{X}}\left\|r^{M_{3}}\left(x\right)\right\|\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|+\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|$
$=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
Proof of Lemma 2.
Define
$\Omega_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}P^{-1}\left(\Sigma+\Sigma_{1}\right)P^{-1}\overline{p}^{M_{2}}\left(v,w\right),\ \Sigma\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2},$ (A.13)
$\Sigma_{1}$ $\coloneqq\mathbb{E}\overline{\mu}_{i}^{I}\overline{\mu}_{i}^{I^{\prime}},\ p_{i}\coloneqq p^{M_{2}}\left(S_{i}\right),\ q_{i}\coloneqq q^{M_{1}}\left(X_{i},Z_{i},W_{i}\right),$
$\overline{\mu}_{i}^{I}$ $\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau^{{}^{\prime}}\left(V_{j}\right)p_{j}q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\ u_{i}\coloneqq Y_{i}-G\left(S_{i}\right),$
$G_{V}\left(S_{j}\right)$ $\coloneqq\left.\partial G\left(s\right)/\partial v\right|_{s=S_{j}},\text{ and }v_{ji}\coloneqq\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-F\left(\left.x_{j}\right|z_{i},w_{i}\right).$
and
$\widehat{\Omega}_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}\widehat{P}^{-1}\left(\widehat{\Sigma}+\widehat{\Sigma}_{1}\right)\widehat{P}^{-1}\overline{p}^{M_{2}}\left(v,w\right),$ (A.14)
$\widehat{\Sigma}$ $\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\left(y_{i}-\widehat{G}\left(\widehat{s}_{i}\right)\right)^{2},\ \widehat{\Sigma}_{1}\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{\overline{\mu}}_{i}^{I}\widehat{\overline{\mu}}_{i}^{I^{\prime}},$
$\widehat{\overline{\mu}}_{i}^{I}$ $\coloneqq n^{-1}\sum_{j=1}^{n}\widehat{G}_{V}\left(\widehat{s}_{j}\right)\widehat{p}_{j}q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }\widehat{v}_{ji}\coloneqq\left(\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-\widehat{F}\left(\left.x_{j}\right|z_{i},w_{i}\right)\right).$
IN02 have proved asymptotic normality for known and scalar-valued functionals of $G\left(s\right)$. I apply their results to $c^{{}^{\prime}}\widehat{b}_{1}\left(v,w\right)$ for any constant vector $c^{{}^{\prime}}c=1$ and obtain
[…]
By (A.15) and Assumption 7(f),
[…]
which implies
[…]
Then,
$\ n^{1/2}\widehat{\Omega}_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
$=$ $\ \left(c_{1n}\widehat{\Omega}_{1}\right)^{-1/2}\left(c_{1n}\Omega_{1}\right)^{1/2}n^{1/2}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
[…]
where the convergence holds by (A.15), the Cramér–Wold device, (A.17), Assumption 7, and the continuous mapping theorem. ∎
Proof of Theorem 3.
Following the proof of Lemma 2, under regularity conditions one may use the Cramér–Wold device to extend the results from scalar-valued functionals to vector-valued functionals of $G\left(s\right)$. Therefore, I maintain that $b\left(x\right)$ is a scalar in this proof. First, I derive the influence functions for $\widehat{b}\left(x\right)$ that correctly account for the estimation errors from each step and prove its asymptotic normality. Then, I show consistency for the estimator of the variance of $\widehat{b}\left(x\right)$, which can be used to construct asymptotically valid confidence intervals. I write $r^{M_{3}}\left(x\right)$ as $r\left(x\right)$ when there is no confusion. The proof relies on the results from IN02, and I point out differences.
Define
$\Omega_{2}$ $\coloneqq\Omega_{21}+\Omega_{22},\ \Omega_{21}\coloneqq\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)\left(A_{1}P^{-1}p_{i}u_{i}\right)^{{}^{\prime}},$ (A.19)
$\Omega_{22}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{{}^{\prime}},$
$A_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}},\ \overline{\mu}_{i}^{II}\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\text{ and}$
$b_{1,V}\left(V_{j},W_{j}\right)$ $\coloneqq\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j},w=W_{j}},$
and $F\coloneqq\Omega_{2}^{-1/2}$, which is well-defined because $\Omega_{2}=\Omega_{21}+\Omega_{22}$ and
$\Omega_{21}$ $=A_{1}P^{-1}\left(\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2}\right)P^{-1}A_{1}^{{}^{\prime}}$
$=A_{1}P^{-1}\left(\mathbb{E}\left[p_{i}p_{i}^{{}^{\prime}}\mathbb{E}\left(\left.u_{i}^{2}\right|X_{i},Z_{i},W_{i}\right)\right]\right)P^{-1}A_{1}^{{}^{\prime}}$
$\geq CA_{1}P^{-1}A_{1}^{{}^{\prime}}=Cr\left(x\right)^{{}^{\prime}}\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)P^{-1}\left(\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}\right)r\left(x\right)>0,$ (A.20)
where the first inequality holds by Assumption 7(c) and the last inequality holds by Assumption 8(a).
Define functionals:
$\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.\widehat{b}_{1}\left(\widehat{V},W\right)\right|X=x\right]=\widehat{b}\left(x\right),$
$\widehat{a}\left(b_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.b_{1}\left(\widehat{V},W\right)\right|X=x\right],$
$\widehat{a}\left(b_{1},V\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.b_{1}\left(V,W\right)\right|X=x\right],\text{ and}$
$a\left(b_{1},V\right)$ $=\mathbb{E}\left[\left.b_{1}\left(V,W\right)\right|X=x\right]=b\left(x\right).$
I expand
$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)+\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)+\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.21)
and show that
$\psi_{1i}$ $=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right),\ \psi_{2i}=H_{2}\overline{\mu}_{i}^{II},\text{ and }\psi_{3i}=H_{2}r_{i}\xi_{i},$ (A.22)
[…]
Let $\widehat{H}_{1}\coloneqq F\widehat{A}_{1}\widehat{P}^{-1}$, $\widehat{A}_{1}\coloneqq r\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}n^{-1}\sum_{i=1}^{n}r_{i}\widehat{\overline{p}}_{i}$, $\widehat{R}\coloneqq n^{-1}\sum_{i=1}^{n}r_{i}r_{i}^{{}^{\prime}}$, $\widehat{\overline{p}}_{i}\coloneqq\overline{p}\left(\widehat{s}_{i}\right)$, $\widehat{H}_{2}\coloneqq FA_{2}\widehat{R}^{-1}$, $G\coloneqq\left(G\left(s_{1}\right),...,G\left(s_{n}\right)\right)^{{}^{\prime}}$, and $\widetilde{G}\coloneqq\left(G\left(\widehat{s}_{1}\right),...,G\left(\widehat{s}_{n}\right)\right)^{{}^{\prime}}$. First, for $\psi_{1i}$,
$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}Y-\widetilde{B}\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left[n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}\left(Y-G+G-\widetilde{G}+\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)+\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{1}\widehat{p}_{i}\left[u_{i}-\left(G\left(\widehat{s}_{i}\right)-G\left(s_{i}\right)\right)\right]+n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)$
$\ +n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)$ (A.23)
$\eqqcolon$ $\ D_{11}+D_{12}+D_{13}.$ (A.24)
I show $D_{11}=n^{-1/2}\sum_{i}\psi_{1i}+o_{p}\left(1\right)$, $D_{12}=o_{p}\left(1\right)$, and $D_{13}=o_{p}\left(1\right)$.
The proof of
$D_{11}=n^{-1/2}\sum_{i}\psi_{1i}+o_{p}\left(1\right)$ (A.25)
is analogous to that of Lemma B7 and B8 of IN02, except that I need to establish $\left\|\widehat{H}_{1}-H_{1}\right\|=o_{p}\left(1\right)$ and $\left\|\widehat{H}_{2}-H_{2}\right\|=o_{p}\left(1\right)$. To prove these two claims, notice that
$\left\|H_{1}\right\|=O\left(1\right)\text{ and }\left\|H_{2}\right\|=O\left(1\right),$ (A.26)
because $\left\|H_{1}\right\|^{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$ and $\left\|H_{2}\right\|^{2}=A_{2}A_{2}^{{}^{\prime}}/\Omega_{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$. In addition, $\left\|\widehat{P}-P\right\|=o_{p}\left(1\right)$, $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and $\left\|n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}-\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right\|=o_{p}\left(1\right)$ by Lemma A3 of IN02 and Newey (1997). I also know that $\lambda_{\min}\left(P\right)\geq C>0$ and $\left\|P\right\|=O\left(1\right)$ by Assumption 5. By Slutsky’s theorem, $\left\|\widehat{R}^{-1}-I\right\|=o_{p}\left(1\right)$ and $\left\|\widehat{P}^{-1}-P^{-1}\right\|=o_{p}\left(1\right)$. By the CS inequality and Lemma A3 of IN02,
$\left\|n^{-1}\sum_{i}r_{i}\left(\overline{p}_{i}-\widehat{\overline{p}}_{i}\right)^{{}^{\prime}}\right\|^{2}$ $\leq n^{-1}\sum_{i}\left\|r_{i}\right\|^{2}\times n^{-1}\sum_{i}\left\|\widehat{\overline{p}}_{i}-\overline{p}_{i}\right\|^{2}$
$=O_{p}\left(M_{3}\zeta_{1}\left(m_{2}\right)^{2}\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.27)
Therefore, by the triangle inequality with probability approaching one,
$\ \left\|\widehat{H}_{1}-H_{1}\right\|^{2}$
$=$ $\ \left\|F\widehat{A}_{1}\widehat{P}^{-1}-FA_{1}P^{-1}\right\|^{2}$
$\leq$ $\ 2\left\|F\left(\widehat{A}_{1}-A_{1}\right)\widehat{P}^{-1}\right\|^{2}+2\left\|FA_{1}\left(\widehat{P}^{-1}-P^{-1}\right)\right\|^{2}$
$=$ $\ 2\left\|F\left(r\left(x\right)^{{}^{\prime}}\left(I+o_{p}\left(1\right)\right)\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}+o_{p}\left(1\right)\right)-r\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)\widehat{P}^{-1}\right\|^{2}$
$\ +2\left\|FA_{1}P^{-1}\left(P-\widehat{P}\right)\widehat{P}^{-1}\right\|^{2}$
$\leq$ $\ \left\|H_{2}\right\|^{2}o_{p}\left(1\right)+\left\|H_{1}\right\|^{2}o_{p}\left(1\right)=o_{p}\left(1\right),$ (A.28)
and similarly $\left\|\widehat{H}_{2}-H_{2}\right\|=o_{p}\left(1\right)$, which establish (A.25).
Next, recall that by Assumption 7
$n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right),\text{ and}$
$n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right).$
Therefore, for $D_{12}$,
$\left|n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{1}\widehat{P}\widehat{H}_{1}^{{}^{\prime}}\right]\left[n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right]$
$\leq\left\|\widehat{H}_{1}\right\|^{2}O_{p}\left(nM_{2}^{-2d_{4}/j_{4}}\right)=o_{p}\left(1\right).$ (A.29)
For $D_{13}$, similarly to (A.29),
$\left|n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{2}\widehat{R}\widehat{H}_{2}\right]\left[n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
$\leq\left\|\widehat{H}_{2}\right\|^{2}O_{p}\left(nM_{2}^{-2d_{4}/j_{4}}\right)=o_{p}\left(1\right).$ (A.30)
Combining the results for $D_{11}$, $D_{12,}$ and $D_{13}$, I have
$\psi_{1i}=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right).$ (A.31)
To prove $\psi_{2i}=H_{2}\overline{\mu}_{i}^{II}$, first notice that by Taylor expansion,
$\ n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
$=$ $\ \widehat{H}_{2}n^{-1/2}\sum_{i}r_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)$
$=$ $\ \widehat{H}_{2}n^{-1/2}\sum_{i}r_{i}b_{1,V}\left(v_{i},w_{i}\right)\left(\widehat{v}_{i}-v_{i}\right)+\widehat{H}_{2}n^{-1/2}\sum_{i}r_{i}b_{1,VV}\left(\widetilde{v}_{i},w_{i}\right)\left(\widehat{v}_{i}-v_{i}\right)^{2}/2$
$\eqqcolon$ $\ D_{21}+D_{22},$ (A.32)
where $b_{1,VV}\left(v_{i},w_{i}\right)\coloneqq\left.\partial^{2}b_{1}\left(v,w\right)/\partial v^{2}\right|_{v=v_{i},w=w_{i}}$ and $\widetilde{v}_{i}$ lies between $v_{i}$ and $\widehat{v}_{i}$. I prove $D_{21}=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right)$ and $D_{22}=o_{p}\left(1\right)$.
For $D_{21}$,
$D_{21}=$ $\ \widehat{H}_{2}n^{-1/2}\sum_{i}r_{i}b_{1,V}\left(v_{i},w_{i}\right)\left(\widehat{v}_{i}-v_{i}\right)$
$=$ $\ H_{2}n^{-1/2}\sum_{i}r_{i}b_{1,V}\left(v_{i},w_{i}\right)\Delta_{i}^{I}+\left(\widehat{H}_{2}-H_{2}\right)n^{-1/2}\sum_{i}r_{i}b_{1,V}\left(v_{i},w_{i}\right)\left(\widehat{v}_{i}-v_{i}\right)$
$+H_{2}n^{-1/2}\sum_{i}r_{i}b_{1,V}\left(v_{i},w_{i}\right)\left(\Delta_{i}^{II}+\Delta_{i}^{III}\right)$
$\eqqcolon$ $\ D_{211}+D_{212}+D_{213},$ (A.33)
where
$\delta_{ij}=F_{\left.X_{i}\right|Z_{j},W_{j}}\left(\left.x_{i}\right|z_{j},w_{j}\right)-q_{j}^{{}^{\prime}}\gamma^{M_{1}}\left(x_{i}\right),\ \Delta_{i}^{I}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}v_{ij},$
$\Delta_{i}^{II}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}\delta_{ij},\text{ and }\Delta_{i}^{III}=-\delta_{ii}.$ (A.34)
Following an argument identical to the proof of Lemma B7 of IN02,
$D_{211}=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.35)
For $D_{212}$,
$\left|D_{212}\right|^{2}$ $\leq Cn\left[\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right]$
$=O_{p}\left\{n\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)\Delta_{1n}^{2}\right\}=o_{p}\left(1\right).$ (A.36)
For $D_{213}$,
$\left|D_{213}\right|^{2}\leq Cn\left[H_{2}\widehat{R}H_{2}^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\left(\Delta_{i}^{II}\right)^{2}+\left(\Delta_{i}^{III}\right)^{2}\right)\right]=O_{p}\left(nM_{1}^{1-2d_{1}/j_{1}}\right)=o_{p}\left(1\right),$ (A.37)
where the first equality is established in the proof of Theorem 4 of IN02.
Next, for $D_{22}$,
$\left|D_{22}\right|$ $\leq Cn^{1/2}\left\|\widehat{H}_{2}\right\|\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|\left|n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right|$
$=O_{p}\left(n^{1/2}\zeta\left(M_{3}\right)\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.38)
Combining the results for $D_{21}$ and $D_{22}$,
$n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.39)
To show $\psi_{3i}=H_{2}r_{i}\xi_{i}$, I expand
$\ n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}b_{1}\left(v_{i},w_{i}\right)-n^{1/2}Fb\left(x\right)$
$=$ $\ n^{-1/2}\sum_{i}H_{2}r_{i}\left(b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)\right)+n^{-1/2}\sum_{i}\left(\widehat{H}_{2}-H_{2}\right)r_{i}\left(b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)\right)$
$+n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)-n^{1/2}F\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)$
$\eqqcolon$ $\ D_{31}+D_{32}+D_{33}+D_{34},$
where I use the definition of $\widehat{H}_{2}$ and $\widehat{R}$ for the decomposition. Recall that $D_{31}=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}$ by the definition of $\xi_{i}$. Thus, I only need to show $D_{32}$, $D_{33}$, and $D_{34}$ are all $o_{p}\left(1\right)$.
For $D_{32}$,
$\mathbb{E}\left[\left.\left|D_{32}\right|^{2}\right|{\bf X}\right]$ $=n^{-1}\left(\widehat{H}_{2}-H_{2}\right)r^{{}^{\prime}}\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left\|\widehat{H}_{2}-H_{2}\right\|^{2}\left(1+\left\|\widehat{R}-I\right\|\right)$
$=O_{p}\left\{\left\|\widehat{H}_{2}-H_{2}\right\|^{2}\right\}$
$=O_{p}\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)=o_{p}\left(1\right),$
where the first inequality holds by Assumption 6(c) and the fact that $\widehat{H}_{2}$ and $r$ are functions of $X$ only, the second equality holds by $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and the third equality follows similarly as in equation (A.1) and (A.6) of Newey (1997). Therefore, $D_{32}=o_{p}\left(1\right)$ by the CM inequality.
For $D_{33}$, by the CS inequality,
$\left|D_{33}\right|^{2}$ $\leq n\left(\widehat{H}_{2}\widehat{R}\widehat{H}_{2}^{{}^{\prime}}\right)n^{-1}\sum_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)^{2}$
$=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right),$
where the first equality holds by Assumption 6(a).
Finally, for $D_{34}$,
$\left|D_{34}\right|^{2}=nF^{2}\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)^{2}=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right).$
Combining the results for $D_{31},$ $D_{32},$ $D_{33}$, and $D_{34}$, I have
$n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}+o_{p}\left(1\right).$ (A.40)
In sum, I have proved
$n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$ $=n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.41)
where
$\psi_{1i}=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right),\ \psi_{2i}=H_{2}\overline{\mu}_{i}^{II},\text{ and }\psi_{3i}=H_{2}r_{i}\xi_{i}.$ (A.42)
Furthermore, notice that
$H_{1}p_{i}u_{i}\perp\left(H_{1}\overline{\mu}_{i}^{I},H_{2}\overline{\mu}_{i}^{II},H_{2}r_{i}\xi_{i}\right)$ (A.43)
because $\mathbb{E}\left[\left.u_{i}\right|X_{i},V_{i},W_{i}\right]=0$.
Let $\Psi_{in}=n^{-1/2}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)$. It is clear that $\mathbb{E}\Psi_{in}=0$ and $\mathbb{\mathbb{V}}\left(\Psi_{in}\right)=n^{-1}$. For any $\epsilon>0$, under Assumptions 7 and 8,
$\ n\mathbb{E}\left[\mathbf{\mathbbm{1}}\left\{\left|\Psi_{in}\right|>\epsilon\right\}\Psi_{in}^{2}\right]$
$\leq$ $\ n\epsilon^{2}\mathbb{E}\left[\mathbf{\mathbbm{1}}\left\{\left|\Psi_{in}\right|>\epsilon\right\}\left(\Psi_{in}/\epsilon\right)^{4}\right]\leq n\epsilon^{-2}\mathbb{E}\Psi_{in}^{4}$
$\leq$ $\ Cn^{-1}\mathbb{E}\left[\left(H_{1}p_{i}u_{i}\right)^{4}+\left(H_{1}\overline{\mu}_{i}^{I}\right)^{4}+\left(H_{2}\overline{\mu}_{i}^{II}\right)^{4}+\left(H_{2}r_{i}\xi_{i}\right)^{4}\right]$
$\leq$ $\ Cn^{-1}\left(\zeta\left(M_{2}\right)^{2}M_{2}+\zeta\left(M_{2}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{2}M_{3}\right)\rightarrow 0,$
where the last inequality holds by Lemma B5 of IN02. Then, by the Lindeberg–Feller CLT,
[…]
where $\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)=\widehat{b}\left(x\right)$ and $a\left(b_{1},V\right)=b\left(x\right)$.
To construct a feasible confidence interval, one needs a consistent estimator of the covariance matrix $\Omega_{2}$. Define
$\widehat{\Omega}_{2}$ $\coloneqq\widehat{\Omega}_{21}+\widehat{\Omega}_{22},\ \widehat{\Omega}_{21}\coloneqq\widehat{A}_{1}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{1}^{{}^{\prime}},$ (A.45)
$\widehat{\Omega}_{22}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
$\widehat{A}_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}\left(n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}\right),\ \text{$\widehat{\overline{p}}_{i}$}\coloneqq\overline{p}\left(\widehat{s}_{i}\right),$
$\widehat{\overline{\mu}}_{i}^{II}$ $\coloneqq n^{-1}\sum_{j}r_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{and }\widehat{\xi}_{i}\coloneqq\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right).$
[…]
$\Omega_{2}=\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)^{2}+\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-A_{2}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{2}=\Omega_{21}+\Omega_{22}$ (A.46)
and
$\widehat{\Omega}_{2}=n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{p}_{i}\widehat{u}_{i}\right)^{2}+n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\widehat{A}_{2}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{2}\eqqcolon\widehat{\Omega}_{21}+\widehat{\Omega}_{22}.$ (A.47)
[…]
For $\widehat{\Omega}_{22}$, I need to show
$n^{-1}\sum_{i}\left(\widehat{H}_{1}\widehat{\overline{\mu}}_{i}^{I}-H_{1}\overline{\mu}_{i}^{I}\right)^{2}$ $=o_{p}\left(1\right),$
$n^{-1}\sum_{i}\left(\widehat{H}_{2}\widehat{\overline{\mu}}_{i}^{II}-H_{2}\overline{\mu}_{i}^{II}\right)^{2}$ $=o_{p}\left(1\right),\text{ and}$
$n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}$ $=o_{p}\left(1\right).$ (A.48)
The first two convergence results have been established in Lemma B9 of IN02. For the last one,
$\ \widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}$
$=$ $\ \widehat{H}_{2}r_{i}\left(\widehat{\xi}_{i}-\xi_{i}\right)+\left(\widehat{H}_{2}-H_{2}\right)r_{i}\xi_{i}$
$=$ $\ \widehat{H}_{2}r_{i}\left(\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right)-b_{1}\left(v_{i},w_{i}\right)+b\left(x_{i}\right)\right)+\left(\widehat{H}_{2}-H_{2}\right)r_{i}\xi_{i}$
$=$ $\ \widehat{H}_{2}r_{i}\left(\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)+\widehat{H}_{2}r_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)$
$+\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-\widehat{b}\left(x_{i}\right)\right)+\left(\widehat{H}_{2}-H_{2}\right)r_{i}\xi_{i}$
$\eqqcolon$ $\ D_{41i}+D_{42i}+D_{43i}+D_{44i}.$ (A.49)
For $D_{41}$,
$n^{-1}\sum_{i}D_{41i}^{2}$ $\leq\left\|\widehat{H}_{2}\right\|^{2}\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|^{2}n^{-1}\sum_{i}\left(\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}$
$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left[\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)\right)^{2}+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}\right]$
$=O_{p}\left(\zeta\left(M_{3}\right)^{2}\Delta_{2n}^{2}\right)=o_{p}\left(1\right),$ (A.50)
where the second inequality holds by $\left\|\widehat{H}_{2}\right\|=O_{p}\left(1\right)$ and Assumption 8(a) and the first equality holds by (A.9).
For $D_{42}$,
$n^{-1}\sum_{i}D_{42i}^{2}$ $\leq\left\|\widehat{H}_{2}\right\|^{2}\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|^{2}n^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}=O_{p}\left(\zeta\left(M_{3}\right)^{2}\Delta_{1n}^{2}\right)=o_{p}\left(1\right),$ (A.51)
where the first equality holds by Lemma 1.
The proof of $n^{-1}\sum_{i}D_{43i}^{2}=o_{p}\left(1\right)$ is completely analogous to (A.50) and is omitted.
For $D_{44}$,
$\mathbb{E}\left[\left.n^{-1}\sum_{i}D_{44i}^{2}\right|{\bf X}\right]$ $=\left(\widehat{H}_{2}-H_{2}\right)n^{-1}\sum_{i}r_{i}r_{i}^{{}^{\prime}}\mathbb{E}\left(\left.\xi_{i}^{2}\right|X_{i}\right)\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left\|\widehat{H}_{2}-H_{2}\right\|^{2}=o_{p}\left(1\right),$ (A.52)
where the first equality holds by the fact that both $\widehat{H}_{2}$ and $r_{i}$ are functions of ${\bf X}$, the first inequality holds by Assumption 6(c), and the last inequality uses $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$. Then, by the CM inequality,
$n^{-1}\sum_{i}D_{44i}^{2}=o_{p}\left(1\right).$ (A.53)
Combining the results for $D_{41}$, $D_{42}$, $D_{43}$, and $D_{44}$, I obtain
$n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$ (A.54)
Therefore, by (A.48),
$\ n^{-1}\sum_{i}\left(\left(\widehat{H}_{1}\widehat{\overline{\mu}}_{i}^{I}-\widehat{H}_{2}\widehat{\overline{\mu}}_{i}^{II}-\widehat{H}_{2}r_{i}\widehat{\xi}_{i}\right)-\left(H_{1}\overline{\mu}_{i}^{I}-H_{2}\overline{\mu}_{i}^{II}-H_{2}r_{i}\xi_{i}\right)\right)^{2}$
$\leq$ $\ 3n^{-1}\sum_{i}\left(\widehat{H}_{1}\widehat{\overline{\mu}}_{i}^{I}-H_{1}\overline{\mu}_{i}^{I}\right)^{2}+3n^{-1}\sum_{i}\left(\widehat{H}_{2}\widehat{\overline{\mu}}_{i}^{II}-H_{2}\overline{\mu}_{i}^{II}\right)^{2}$
$+3n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$
Since $\mathbb{E}\left(H_{1}\overline{\mu}_{i}^{I}-H_{2}\overline{\mu}_{i}^{II}-H_{2}r_{i}\xi_{i}\right)^{2}=\Omega_{22}/\Omega_{2}\leq 1$, by Markov’s inequality and Lemma B6 of IN02,
$\left|\widehat{\Omega}_{22}/\Omega_{2}-n^{-1}\sum_{i}\left(H_{1}\overline{\mu}_{i}^{I}-H_{2}\overline{\mu}_{i}^{II}-H_{2}r_{i}\xi_{i}\right)^{2}\right|=o_{p}\left(1\right).$ (A.55)
By the law of large numbers,
$\left|n^{-1}\sum_{i}\left(H_{1}\overline{\mu}_{i}^{I}-H_{2}\overline{\mu}_{i}^{II}-H_{2}r_{i}\xi_{i}\right)^{2}-\Omega_{22}/\Omega_{2}\right|=o_{p}\left(1\right).$ (A.56)
Therefore, by the triangle inequality,
$\widehat{\Omega}_{22}/\Omega_{2}-\Omega_{22}/\Omega_{2}=o_{p}\left(1\right).$ (A.57)
Combining the results for $\widehat{\Omega}_{21}$ and $\widehat{\Omega}_{22}$, I obtain
[…]
For the last result of Theorem 3, first define
$\widehat{\Omega}_{3}$ $\coloneqq\widehat{\Omega}_{31}+\widehat{\Omega}_{32},\ \widehat{\Omega}_{31}\coloneqq\widehat{A}_{3}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{3},$ (A.58)
$\widehat{\Omega}_{32}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
$\widehat{A}_{3}$ $\coloneqq n^{-1}\sum_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}},\text{ and}\ \widehat{\overline{\mu}}_{i}^{III}\coloneqq n^{-1}\sum_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji}.$
[…]
Appendix B Sufficient Conditions for Assumption 2
Proposition 1 (Sufficient Conditions for Assumption 2 based on Altonji and Matzkin (2005)).
Suppose the following conditions hold:
  1. (a)
    $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ for any permutation $\widetilde{\eta}_{i}$ of the $\eta_{i}$ vector in time,
  2. (b)
    $Z_{it}\perp\left(A_{i},\eta_{it}\right)$ and the support of $\left(X_{it},Z_{it}\right)$ is compact, and,
  3. (c)
    $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}$ is continuous in $\left({\bf X}_{i},{\bf Z}_{i}\right)$.
Then, Assumption 2 is satisfied.
Proof of Proposition 1.
To begin with, we use the first condition $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ to establish the exchangeability condition
$f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}},$ (B.1)
where $\left(\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}\right)$ is any permutation of $\left({\bf X}_{i},{\bf Z}_{i}\right)$ in time. It is worth emphasizing that (B.1), which will be proved in the following, is different from Assumption 2.3 of Altonji and Matzkin (2005). [^27] To be specific, Altonji and Matzkin (2005) assume an exchangeability condition involving both $A$ and $\eta_{t}$, i.e., $f_{\left.A_{i},\eta_{it}\right|{\bf X}_{i}}=f_{\left.A_{i},\eta_{it}\right|{\bf\widetilde{X}}_{i}},$ which effectively rules out time-varying endogeneity through the random coefficients since it requires that the density of $\eta_{it}$ given $X_{it}=x_{it}$ is the same as that given $X_{it}=x_{is}$ for any $s\neq t$. Without loss of generality, we prove (B.1) for $T=2$ since any ordering of $\left\{1,...,T\right\}$ for finite $T$ can be achieved via a finite number of pairwise permutations. Then, we prove that one can construct $W_{i}$ such that Assumption 2 holds. For simplicity of notation, we assume $X_{it}$ and $Z_{it}$ are both scalars and suppress $i$ subscripts in all variables. Thus, in this proof all subscripts of the variables denote the time period.
By condition (a) of Proposition 1,
$f_{A,\eta_{1},\eta_{2}}\left(a,h_{1},h_{2}\right)=f_{A,\eta_{1},\eta_{2}}\left(a,h_{2},h_{1}\right).$ (B.2)
Let $g^{-1}\left(X,Z,A\right)$ denote the inverse function of $g\left(Z,A,\eta\right)$ with respect to $\eta$, which exists by Assumption 1. Define $h_{1}=g^{-1}\left(x_{1},z_{1},a\right)$ and $h_{2}=g^{-1}\left(x_{2},z_{2},a\right)$. Calculate the determinants of the Jacobians as
$D_{1}\coloneqq$ $\left|\begin{array}[]{ccc}\dfrac{\partial A}{\partial X_{1}}&\dfrac{\partial A}{\partial X_{2}}&\dfrac{\partial A}{\partial A}\\ \dfrac{\partial g^{-1}\left(X_{1},Z_{1},A\right)}{\partial X_{1}}&\dfrac{\partial g^{-1}\left(X_{1},Z_{1},A\right)}{\partial X_{2}}&\dfrac{\partial g^{-1}\left(X_{1},Z_{1},A\right)}{\partial A}\\ \dfrac{\partial g^{-1}\left(X_{2},Z_{2},A\right)}{\partial X_{1}}&\dfrac{\partial g^{-1}\left(X_{2},Z_{2},A\right)}{\partial X_{2}}&\dfrac{\partial g^{-1}\left(X_{2},Z_{2},A\right)}{\partial A}\end{array}\right|_{\begin{array}[]{c}\left(X_{1},X_{2},Z_{1},Z_{2},A\right)\\ =\left(x_{1},x_{2},z_{1},z_{2},a\right)\end{array}}$
$=$ $\left|\begin{array}[]{ccc}0&0&1\\ \dfrac{\partial g^{-1}\left(X_{1},Z_{1},A\right)}{\partial X_{1}}&0&\dfrac{\partial g^{-1}\left(X_{1},Z_{1},A\right)}{\partial A}\\ 0&\dfrac{\partial g^{-1}\left(X_{2},Z_{2},A\right)}{\partial X_{2}}&\dfrac{\partial g^{-1}\left(X_{2},Z_{2},A\right)}{\partial A}\end{array}\right|_{\begin{array}[]{c}\left(X_{1},X_{2},Z_{1},Z_{2},A\right)\\ =\left(x_{1},x_{2},z_{1},z_{2},a\right)\end{array}}$
$=$ $\left.\partial g^{-1}\left(X,Z,A\right)/\partial X\right|_{\left(X,Z,A\right)=\left(x_{1},z_{1},a\right)}\times\left.\partial g^{-1}\left(X,Z,A\right)/\partial X\right|_{\left(X,Z,A\right)=\left(x_{2},z_{2},a\right)},$
and similarly
$D_{2}\coloneqq$ $\left|\begin{array}[]{ccc}\dfrac{\partial g\left(Z_{1},A,\eta_{1}\right)}{\partial A}&\dfrac{\partial g\left(Z_{1},A,\eta_{1}\right)}{\partial\eta_{1}}&\dfrac{\partial g\left(Z_{1},A,\eta_{1}\right)}{\partial\eta_{2}}\\ \dfrac{\partial g\left(Z_{2},A,\eta_{2}\right)}{\partial A}&\dfrac{\partial g\left(Z_{2},A,\eta_{2}\right)}{\partial\eta_{1}}&\dfrac{\partial g\left(Z_{2},A,\eta_{2}\right)}{\partial\eta_{2}}\\ \dfrac{\partial A}{\partial A}&\dfrac{\partial A}{\partial\eta_{1}}&\dfrac{\partial A}{\partial\eta_{2}}\end{array}\right|_{\begin{array}[]{c}\left(Z_{1},Z_{2},A,\eta_{1},\eta_{2}\right)\\ =\left(z_{2},z_{1},a,h_{2},h_{1}\right)\end{array}}$
$=$ $\left.\partial g\left(Z,A,\eta\right)/\partial\eta\right|_{\left(Z,A,\eta\right)=\left(z_{2},a,h_{2}\right)}\times\left.\partial g\left(Z,A,\eta\right)/\partial\eta\right|_{\left(Z,A,\eta\right)=\left(z_{1},a,h_{1}\right)}.$
Then,
$\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2},a\right|z_{1},z_{2}\right)$
$=$ $\ f_{\left.A,\eta_{1},\eta_{2}\right|Z_{1},Z_{2}}\left(\left.a,g^{-1}\left(x_{1},z_{1},a\right),g^{-1}\left(x_{2},z_{2},a\right)\right|z_{1},z_{2}\right)\left|D_{1}\right|$
$=$ $\ f_{\left.A,\eta_{1},\eta_{2}\right|Z_{1},Z_{2}}\left(\left.a,g^{-1}\left(x_{2},z_{2},a\right),g^{-1}\left(x_{1},z_{1},a\right)\right|z_{2},z_{1}\right)\left|D_{1}\right|$
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right)\left|D_{2}D_{1}\right|$
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right),$ (B.3)
where the first equality holds by the change of variables for $\eta_{1}$ and $\eta_{2}$, the second equality uses (B.2) and $Z\perp\left(A,\eta\right)$, the third equality holds by $X_{1}=g\left(z_{2},a,g^{-1}\left(x_{2},z_{2},a\right)\right)=x_{2}$ and $X_{2}=g\left(z_{1},a,g^{-1}\left(x_{1},z_{1},a\right)\right)=x_{1}$, and the last equality uses the fact that the product of derivatives of inverse functions equals one.
Given (B.3), I integrate $A$ out in (B.3) to obtain
$f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$ $=f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right),$ (B.4)
which implies
$\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2},a\right|z_{1},z_{2}\right)/f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right)/f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right)$
$=$ $\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right),$ (B.5)
where the second equality uses (B.3) and (B.4).
The rest of the proof follows the same argument as in Section 2.2 of Altonji and Matzkin (2005). Specifically, I show that for any on-support $a$, the conditional density $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a function of the form $f_{\left.A\right|W}\left(\left.a\right|w\right)$, where $W$ is a vector-valued function symmetric in the paired elements of $\left(X,Z\right)$. See footnote 10 for the detailed description of $W$. By condition (b) of Proposition 1, the supports of $X$ and $Z$ are compact. By condition (c) of Proposition 1, $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}$ is continuous in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$. Therefore, by the Stone-Weierstrass theorem, there exists a function $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}$ that is a polynomial in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$ over a compact set with the property that for any fixed $\delta$ that is arbitrarily close to 0,
$\max_{x_{1},x_{2}\in\mathcal{X},z_{1},z_{2}\in\mathcal{Z}}\left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|\leq\delta.$ (B.6)
Let
$\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$
$\coloneqq$ $\left[f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)+f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right)\right]/2!$
denote the simple average of $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ over all $T!$ (here $T=2$) unique permutations of $\left(x_{t},z_{t}\right)$, and similarly for $\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$. By (B.5),
$\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)=f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right).$
By (B.5), (B.6), and the triangle inequality,
$\ \left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|$
$=$ $\ \left|\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|$
$\leq$ $\ T!\times\left(\delta/T!\right)=\delta.$ (B.7)
Since $f^{w}$ can be chosen to make $\delta$ arbitrarily small, equation (B.7) implies that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a polynomial $\overline{f}^{w}$ that is symmetric in $\left(x_{t},z_{t}\right)$ pairs for $t=1,2$. Furthermore, by the fundamental theorem of symmetric functions, for any $a$ on its support, $\overline{f}^{w}$ can be written as a polynomial function of the elementary symmetric functions $W\left(\cdot\right)$ of $\left(\left(x_{1},z_{1}\right),\left(x_{2},z_{2}\right)\right)$. We denote this function by $f_{\left.A\right|W}\left(\left.a\right|w\right)$ and obtain that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by $f_{\left.A\right|W}\left(\left.a\right|W\right)$. Let $\delta\rightarrow 0$ in (B.7). [^28] Note that to let $\delta\to 0$ we need the order of polynomials in the approximating function $f^{w}$ to increase. However, it does not affect the order of the elements of $W$ since by the fundamental theorem of symmetric functions, the elements of $W$ has a fixed order of $T$. The key is we have two different orders of polynomials; one is the order of the approximating functions $f^{w}$, the other is the order of the arguments of $W$. For example, one may let the order $P_{1}$ and $P_{2}$ of a polynomial function $h\left(x,y\right)=\sum_{i=1}^{P_{1}}\left(x+y\right)^{i}+\sum_{i=1}^{P_{2}}\left(xy\right)^{i}$ to increase while keeping the order of its symmetric arguments (in this example, it is $x+y$ and $xy$) fixed. See footnote 9 and 10 of Altonji and Matzkin (2005) for a more detailed discussion. Then, for any $t\in\left\{1,...,T\right\}$ and on-support $\left(x_{t},z_{t},a,w\right)$
$f_{\left.A\right|X_{t},Z_{t},W}\left(\left.a\right|x_{t},z_{t},w\right)=f_{\left.A\right|W}\left(\left.a\right|w\right).$
Appendix C Simulation
In this section, I examine the finite-sample performance of the series estimators via a simulation study. A discussion of the data generating process (DGP) motivated by production function applications is first provided. Then, I present the baseline results of estimating the APE $\overline{b}\coloneqq\mathbb{E}\beta_{it}$. Next, I show how the method performs for estimating the LAR $b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$. Finally, as robustness checks, I (i) vary the number of firms and periods, (ii) change the order of basis functions used for series estimation, and (iii) include ex-post shocks $\epsilon_{it}$ and $\upsilon_{it}$ to the DGP.
C.1 DGP
The baseline revenue-based DGP is
$Y_{it}=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
[…]
$k_{it}$ $=\dfrac{r_{it}-\left(A_{i}+\eta_{it}^{K}\right)\left(r_{it}-w_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{K}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{K}/2+1\right)}{\left(A_{i}+\eta_{it}^{K}\right)/5-1},$
$l_{it}$ $=\dfrac{w_{it}-\left(A_{i}+\eta_{it}^{L}\right)\left(w_{it}-r_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{L}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{L}/2+1\right)}{\left(A_{i}+\eta_{it}^{L}\right)/5-1}.$
Let $X_{it}=\left(k_{it},l_{it}\right)^{{}^{\prime}}$. It is clear that $X_{it}$ is correlated with $\beta_{it}$ in each period via $\left(A_{i},\eta_{it}^{K},\eta_{it}^{L}\right)$ in a nonseparable way. I use $N$, $T$, and $M$ to denote the total number of firms, periods, and simulations, respectively. The observed data is $\left\{X_{it,m},Y_{it,m},Z_{it,m}\right\}$ for $i=1,...,n$, $t=1,...,T$, and $m=1,...,M$, which are used to construct $\widehat{\overline{b}}_{m}$ and $\widehat{b}_{t,m}\left(x\right)$ to estimate $\overline{b}$ and $b_{t}\left(x\right)$, respectively, via the estimation procedure outlined in Section 4.1. I evaluate the performance of the estimators by their biases, root-mean squared errors (rMSE), and mean normed deviations (MND), with the explicit mathematical definitions provided in the tables below.
C.2 Results
For the baseline results, I set $N=1,000$ and $T=2$, and use a second-order polynomial spline basis with its knot at the median for all series estimation steps. I run $M=1,000$ simulations. For each $i$, I construct $W_{i}$ as the mean over time of each coordinate of $X_{it}$. Following the theory, I set $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$. The performance of $\widehat{\overline{\omega}}$, $\widehat{\overline{b}}^{K},$ and $\widehat{\overline{b}}^{L}$ is summarized in Table 3. For notational simplicity, I use $\sum_{m}$ for $\sum_{m=1}^{M}$ and $\sum_{i,t,m}$ for $\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{m=1}^{M}$.
Table 3: Performance of $\widehat{\overline{b}}$
Formula $\widehat{\overline{\omega}}$ $\widehat{\overline{b}}^{K}$ $\widehat{\overline{b}}^{L}$
Bias $M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.22% 3.24%
rMSE $\sqrt{M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 2.89% 3.65% 3.62%
MND $M^{-1}\sum_{m}\left|\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.26% 3.26%
The first row of Table 3 reports the normalized bias of each coordinate of $\widehat{\overline{b}}$. The bias is reasonably small across all three coordinates, with a magnitude between 2.81% and 3.24% of the length of the corresponding coordinate of $\overline{b}$. The second row shows the normalized rMSE of each coordinate of $\widehat{\overline{b}}$. My method achieves low normalized rMSEs between 2.89% and 3.65% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$. The last row presents the normalized MNDs of each coordinate of $\widehat{\overline{b}}$. Again, the method performs well with an MND between 2.81% and 3.26% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$.
Next, I investigate the performance of $\widehat{b}_{t,m}\left(x\right)$, which is obtained by regressing the estimated $\widehat{b}_{1t,m}\left(\widehat{V}_{it,m},W_{i,m}\right)$ on the second-degree polynomial spline basis of $X_{it,m}$ with its knot at the median for each $\left(t,m\right)$ combination. For the true $b_{t}\left(x\right)$, due to the complex dependence structure of $X_{it}$ on $\beta_{it}$, there is no analytical solution available. Therefore, for each $t$ I pool the observations across all $M$ simulations ($N\times M$ observations for each $t$) and approximate $b_{t}\left(x\right)$ by regressing the true $\beta_{it,m}$ on the same spline basis functions of $X_{it,m}$.
Table 4 presents the results. Note that by definition the bias of $\widehat{b}_{t}\left(x\right)$ is the same as $\widehat{\overline{b}}$ so I omit it here. The normalized rMSE of $\widehat{b}_{t}\left(x\right)$ is bigger than that of $\widehat{\overline{b}}$, with a magnitude between 3.36% and 5.59% of the size of the corresponding $\overline{b}^{\left(d\right)}$. The normalized MND follows a similar pattern. The performance of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ is not as good as that of $\widehat{\overline{b}}$ for $\overline{b}$, which is expected because (i) $b_{t}\left(x\right)$ is a function rather than a finite-dimensional vector and (ii) there is approximation error in calculating the true $b_{t}\left(x\right)$ via simulations.
Table 4: Performance of $\widehat{b}_{t}\left(x\right)$
Formula $\widehat{\omega}_{t}\left(x\right)$ $\widehat{b}_{t}^{K}\left(x\right)$ $\widehat{b}_{t}^{L}\left(x\right)$
rMSE $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left(\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 3.36% 5.57% 5.59%
MND $\left(NTM\right)^{-1}\sum_{i,t,m}\left|\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.87% 4.36% 4.37%
To show how robust my method is in estimating $\overline{b}$ and $b_{t}\left(x\right)$, I conduct another set of exercises. I evaluate the performance of my estimator using rMSE defined as $\sqrt{M^{-1}\sum_{m}\left\|\widehat{\overline{b}}_{m}-\overline{b}\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $\overline{b}$, and $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left\|\widehat{b}_{t,m}\left(x_{it,m}\right)-b_{t}\left(x_{it,m}\right)\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $b_{t}\left(x\right)$. First, I vary $N$ and $T$, and present the results in Table 5. As expected, a larger $N$ is good for overall performance. However, the magnitude in the improvement of performance is mild, possibly caused by the fact that with more agents I need more data to control for the increasing degree of heterogeneity. On the other hand, I find that the method performs reasonably well even with a small sample size of $N=500$. Having a larger $T$ improves the performance of my method, which is again expected as I can exploit more information from repeated observations of the same individual to better control for the fixed effect $A_{i}$.
Table 5: Performance under Varying $N$ and $T$
$N$ ($T=2$) rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
$500$ 3.25% 4.23%
$1,000$ 2.99% 3.69%
$2,000$ 2.87% 3.43%
$T$ ($N=1,000$) rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
2 2.99% 3.69%
3 2.74% 3.38%
4 2.66% 3.27%
Second, I vary the degree of the spline basis functions used to construct the series estimators for all steps, and present the results in Table 6. The knot is placed at the median for all cases. I find that increasing the degree of basis functions from one to two improves estimation accuracy significantly. When I increase the degree of basis functions from two to three, the performance deteriorates. I attribute it to the fact that there are many more regressors in each step of regression when $d=3$, leading to a possible over-fitting problem. Based on this result, I use the second-degree splines in the empirical application. [^29] One may also use the AIC criterion to select the degree of basis functions.
Table 6: Performance under Varying Degree of Basis Functions
Degree of Basis Functions rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
1 5.27% 6.41%
2 2.99% 3.69%
3 4.82% 6.40%
[…]
Table 7: Performance with and without Ex Post Shocks
Add $\epsilon_{it}$ to $Y_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
No 2.99% 3.69%
Yes 3.64% 5.43%
Add $\upsilon_{it}$ to $\left(\beta_{it}^{K},\beta_{it}^{L}\right)$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
No 2.99% 3.69%
Yes 3.57% 5.19%

### v3

@@
 This paper proposes a correlated random coefficient linear panel data model, where regressors can be correlated with time-varying and individual-specific random coefficients through both a fixed effect and a time-varying random shock.
-I develop a new panel data-based identification method to identify the average partial effect and the local average response function.
-The identification strategy employs a sufficient statistic to control for the fixed effect and a conditional control variable for the random shock.
+I develop a new panel data-based method to identify the average partial effect and the local average response function.
+The identification strategy employs a sufficient statistic to control for the fixed effect and a control variable for the random shock.
 Conditional on these two controls, the residual variation in the regressors is driven solely by the exogenous instrumental variables, and thus can be exploited to identify the parameters of interest.
 The constructive identification analysis leads to three-step series estimators, for which I establish rates of convergence and asymptotic normality.
-To illustrate the method, I estimate a heterogeneous Cobb-Douglas production function for manufacturing firms in China, finding substantial variations in output elasticities across firms.
-Keywords: Correlated random coefficients, panel data, time-varying endogeneity, index exclusion, conditional control variable, average partial effect, local average response function.
+To illustrate the method, I estimate a heterogeneous Cobb-Douglas production function for manufacturing firms in China, finding substantial variations in output elasticities across firms that can be related to various firm characteristics.
+Keywords: Correlated random coefficients, panel data, time-varying endogeneity, semiparametric estimation.
 1 Introduction
@@
 While convenient, this approach may not fully capture agents’ optimization behavior.
-For instance, it is possible that a firm with a high fixed effect (e.g., strong management skills) endogenously selects lower input levels during certain periods.
-Such behavior could arise if the firm has information about its time-varying output efficiency shock, in addition to its time-invariant management skill, when optimally determining input values.
+For example, even high-capability firms may cut inputs in some periods when input prices are flat because input choices also respond to transitory, time-varying shocks to output efficiency or productivity.
+Throughout the sample, I assume capability is time-invariant.
+Thus, both persistent capability and transitory shocks can affect realized output efficiency and may be reflected in input choices, thereby generating endogeneity that is not fully absorbed by a fixed effect.
 In this paper, I aim to address this gap by proposing a new time-varying endogenous random coefficient (TERC) linear panel model, where regressors can be correlated with time-varying and individual-specific random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock.
-Specifically, the baseline TERC model [^1] In the baseline model, I focus exclusively on endogenous regressors to address the issue of time-varying endogeneity through the random coefficients.
-The analysis remains largely unchanged when exogenous regressors are included in (1.1).
-I specify the necessary changes at the end of the proof of Theorem 1 in the Appendix.
-consists of two equations:
-$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (1.1)
+Specifically, the baseline TERC model consists of two equations:
+$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$ (1.1)
 $X_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right).$ (1.2)
-In (1.1), I model the vector of random coefficients $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)$ as a vector of unknown and possibly nonlinear functions $\beta\left(\cdot\right)$ of a fixed effect $A_{i}$ and a time-varying random shock $\varepsilon_{it}$.
+The vector of random coefficients $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)$ is modeled as a vector of unknown functions $\beta$ of a fixed effect $A_{i}$ and a time-varying random shock $\varepsilon_{it}$.
 $Y_{it}$ is then determined by the inner product between $X_{it}$ and $\beta_{it}$.
-In (1.2), I model the vector of regressors $X_{it}$ as a vector-valued, unknown and possibly nonlinear function $g\left(\cdot\right)$ of the vector of exogenous instrumental variables (IV) $Z_{it}$, fixed effect $A_{i}$, and per-period information $\eta_{it}$ about $\varepsilon_{it}$.
+In (1.2), I model the vector of regressors $X_{it}$ as a vector-valued, time-varying function $g$ of the vector of exogenous instrumental variables (IV) $Z_{it}$, fixed effect $A_{i}$, and per-period information $\eta_{it}$ about $\beta_{it}$.
 The motivation for (1.2) is that the agent $i$ in period $t$ optimally chooses values of her regressors $X_{it}$ by solving an optimization problem (e.g., firm profit maximization), with $Z_{it},\ A_{i},\text{ and }\eta_{it}$ in her information set, leading to (1.2).
-All of $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$ can be scalar- or vector-valued [^2] In the baseline model, I assume in Assumption 1 that each coordinate of $g\left(\cdot\right)$ depends monotonically on a single coordinate of $\eta_{it}$ to invoke the control function argument.
-As an extension, I discuss how to relax this assumption to allow the entire vector $\eta_{it}$ to enter certain coordinates of $g\left(\cdot\right)$ at the end of Section 3., and any pair of the three variables can be correlated.
-The functional forms are potentially different across different coordinates of $\beta\left(\cdot\right)$ (and also $g\left(\cdot\right)$).
-As an analyst, I observe $\left\{X_{it},Y_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify the average partial effect $\mathbb{E}\beta_{it}$ (APE, see Graham and Powell (2012)) and the local average response function $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR, see Altonji and Matzkin (2005)).
+All of $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$ can be scalar- or vector-valued, and any pair of the three variables can be correlated.
+As an analyst, I observe $\left\{X_{it},Y_{it},Z_{it}\right\}$ for $i=1,\ldots,n$ and $t=1,\ldots,T$ and aim to identify the pooled average partial effect $T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$ (APE, see Graham and Powell (2012)) and the local average response function $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR, see Altonji and Matzkin (2005)).
+[^1] Note that the joint distribution of $(\varepsilon_{it},\eta_{it})$ may vary over $t$.
+Consequently, any functionals of their distribution (e.g., $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$) are in principle time-varying.
+I suppress the subscript $t$ whenever it is clear from context.
 A key feature of the TERC model is that $X_{it}$ can be correlated with $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$, potentially in a complicated manner.
 I refer to this feature as “time-varying endogeneity through the random coefficients.” Allowing such correlation is important to applications when agent $i$ in period $t$ has information about $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$ when optimally deciding $X_{it}$.
-I present three empirical examples in Section 2 that demonstrate the validity of such correlation in real-life scenarios.
-The associated technical challenge is that I need to control for the time-varying endogeneity through the random coefficients when both unknown (to the analyst) $A_{i}$ and $\eta_{it}$ appear in $g\left(\cdot\right)$ in a nonlinear and nonseparable way in (1.2).
-Furthermore, I do not specify the distributions of $A_{i}$, $\varepsilon_{it}$, or $\eta_{it}$, nor the functional form of $\beta\left(\cdot\right)$ or $g\left(\cdot\right)$.
-In this sense, the TERC model is a fixed effect panel data model (Laage (2024)).
-It has many important applications in economics; e.g., heterogeneous Cobb-Douglas production function estimation, return to schooling estimation, labor supply estimation, Engel curve analysis, and demand analysis, among others (Blundell, MaCurdy, and Meghir (2007b, a), Wooldridge (2009), Chernozhukov, Hausman, and Newey (2019), Li and Sasaki (2024), Keiller, de Paula, and Van Reenen (2024)).
-[…]
-The constructive identification analysis leads to easy-to-implement series estimators for the APE and LAR.
-I derive the convergence rates and establish the asymptotic normality of the proposed estimators.
-The inference results presented in this paper build on the literature on the asymptotic analysis of multi-step series estimators (Andrews (1991), Newey (1997), Imbens and Newey (2009)).
-The key departures from the existing literature include the fact that the object of interest is a partial mean process (Newey (1994)) of the derivative of the second-step estimator, and that the final step of the three-step estimator is an unknown but estimable functional of a conditional expectation function.
-Consequently, the estimation errors from each step must be properly accounted for to ensure accurate large sample properties.
-As an empirical illustration, I estimate a Cobb-Douglas production function with heterogeneous output elasticities for Chinese manufacturing firms.
-My estimated APEs are broadly consistent with those derived from applying the classical methods with constant coefficients (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)) to the same data.
-I also estimate the LARs evaluated at the realized input levels for each firm and find significant variation across firms.
-[^4] Note that, with constant coefficient models, the LAR is equivalent to the APE.
-To support the empirical results, I conduct simulations motivated by production function estimation applications to evaluate the finite-sample performance of the estimators proposed in this paper.
-The results demonstrate that the proposed method performs well.
-Next, I discuss the contributions and limitations of this paper.
-The first contribution concerns the features of the TERC model in (1.1)–(1.2).
-Allowing $X_{it}$ to be correlated with $\beta_{it}$ not only through $A_{i}$ but, more importantly, through $\eta_{it}$ significantly enhances the model’s flexibility and carries important empirical implications, as demonstrated by the examples discussed so far.
-Another novel aspect of the TERC model is the inclusion of $A_{i}$ in both (1.1) and (1.2).
-In particular, I argue that including $A_{i}$ in the first-step function $g\left(\cdot\right)$ better reflects the reality that $A_{i}$ is typically in the information set of agent $i$ when she optimally chooses the value of $X_{it}$, for instance, through repeated learning.
-Moreover, $A_{i}$ enters both (1.1) and (1.2) in a nonlinear and nonseparable way, which naturally arises from agent optimization behavior.
-Consequently, enabling these features enhances the model’s flexibility and broadens its applicability.
-The second contribution is the proposal of a new panel data-based method to identify the APE and LAR of the TERC model.
-I construct a sufficient statistic $W_{i}$ to control for the time-invariant $A_{i}$ and a per-period conditional control variable $V_{it}$ to control for the time-varying shock $\eta_{it}$.
-I prove that, conditional on $V_{it}\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely attributable to the exogenous $Z_{it}$, thereby enabling the identification of the APE and LAR.
-The identification method proposed in this paper also contributes to the literature on control functions with multiple unobservables in nonseparable regression models (e.g., Blundell, Kristensen, and Matzkin (2013), Kasy (2014)).
-The final contribution involves implementing the proposed method and characterizing its statistical properties.
-I develop three-step series estimators for the APE and LAR, establishing their convergence rates and asymptotic normality.
-The effectiveness of the method is demonstrated using a production dataset of Chinese manufacturing firms, revealing significant heterogeneity in capital and labor elasticities across firms.
-To further support the empirical findings, I conduct a simulation study inspired by applications in production function estimation, which produces satisfactory results.
-I discuss two limitations of this paper.
-First, although both the unknown $A_{i}$ and $\eta_{it}$ determine the value of $X_{it}$, the baseline model imposes a restriction such that each coordinate of $g\left(\cdot\right)$ depends on a single coordinate of $\eta_{it}$ (as well as $Z_{it}\text{ and }A_{i}$).
-As noted by Imbens and Newey (2009), this restriction excludes nonseparable supply and demand models with one time-varying disturbance per equation.
-To relax this restriction, I propose two approaches under additional empirically motivated assumptions at the end of Section 3.
-Second, my method requires IVs.
-I provide a few examples of IVs for the applications discussed in this paper.
-However, depending on the specific application, identifying suitable IVs can still be challenging.
+In Section 2, I present three empirical examples—heterogeneous production function estimation, a labor supply model, and demand estimation—that demonstrate the validity of this correlation (Deaton (1986), Blundell, MaCurdy, and Meghir (2007), Li and Sasaki (2024), Keiller, de Paula, and Van Reenen (2024)).
+The associated technical challenge is to control for the time-varying endogeneity through the random coefficients when both $A_{i}$ and $\eta_{it}$ enter $g$ in a nonlinear and nonseparable manner in (1.2).
+Moreover, I do not impose parametric assumptions on the distributions of $A_{i}$, $\varepsilon_{it}$, or $\eta_{it}$, nor do I specify the functional form of $\beta$ or $g$.
+In this sense, the TERC model is a fixed-effects panel data model, and addressing these challenges requires a new method.
+I propose a new panel data-based method for identifying the APE and LAR in the TERC model.
+The idea is to control for $A_{i}$ and $\eta_{it}$ via a sufficient statistic and a generalized residual, respectively, such that, conditional on these two controls, the residual variation in $X_{it}$ is driven solely by the exogenous IV $Z_{it}$.
+In other words, the residual variation in $X_{it}$ is causal and thus can be exploited to identify the parameters of interest.
+Specifically, I first impose an index exclusion assumption that supplies a sufficient statistic $W_{i}$ for $A_{i}$.
+I present justifications for this assumption from the literature.
+Next, I construct a feasible conditional cumulative distribution function (CDF) $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it}$.
+Given $(A_{i},W_{i})$, I establish a one-to-one mapping between $V_{it}$ and $\eta_{it}$.
+Finally, conditioning on these two controls—which effectively fixes both $A_{i}$ and $\eta_{it}$—(1.2) implies that the residual variation in $X_{it}$ is driven solely by $Z_{it}$.
+This residual, instrument-induced variation can therefore be exploited to identify the APE and LAR via perturbation arguments and iterated expectations.
+At the end of Section 3, I present three extensions: (i) allowing $g$ to depend on multiple components of $\eta_{it}$, (ii) identifying higher-order moments of $\beta_{it}$, and (iii) adding ex-post shocks to $\beta_{it}$ and $Y_{it}$, and including exogenous covariates into $X_{it}$.
+I estimate the parameters of interest using a three-step series procedure.
+In the first step, I estimate the conditional CDF $V_{it}$, which serves as a control for $\eta_{it}$.
+In the second step, I exploit the linear structure of (1.1) to estimate $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$.
+In the third step, I use these estimates to recover the APE and LAR.
+When working with large datasets, the main computational challenge comes from the first-step estimation of $V_{it}$ and the second-step estimation of $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$.
+I address these issues in Remark 3.
+Inference is less standard because these estimators are multi-step: they rely on intermediate nonparametric estimates, include derivatives of an earlier-step estimate, and then apply an additional unknown but estimable mapping to a conditional expectation.
+As a result, sampling error from the first and second steps can affect the limiting distribution of the final estimators, so valid standard errors must account for how uncertainty propagates across steps.
+I therefore derive convergence rates and asymptotic normality by explicitly tracking the contribution of each step, building on existing results for multi-step series estimators (Andrews (1991), Newey (1997), Imbens and Newey (2009)).
+A discussion of how the errors from each step aggregate into the final standard error is provided at the end of Subsection 4.2.
+To illustrate the method, I estimate a Cobb–Douglas production function with heterogeneous output elasticities using data on Chinese manufacturing firms.
+The instruments are inspired by Berry, Levinsohn, and Pakes (1995) and constructed as competitors’ leave-one-out weighted-average input prices within the same city and industry.
+The resulting APE estimates are broadly consistent with those obtained from classical constant-coefficient approaches applied to the same data (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)).
+I also estimate the LAR functions evaluated at each firm’s realized input choices, and I find substantial cross-firm variation.
+I then relate heterogeneity in output elasticities to firm characteristics and obtain economically intuitive patterns.
+Finally, I conduct extensive robustness checks in Appendix B and find that the results are stable.
+To complement the empirical analysis, Appendix C presents Monte Carlo simulations designed to mirror the empirical environment; the results show that the proposed estimator performs well and reinforce the empirical findings.
+In particular, in the baseline design ($n=1,000$, $T=2$, and 1,000 replications, using second-degree spline bases), the APE estimator exhibits small normalized bias (about 2.8–3.2% relative to the length of the true parameter) and low normalized rMSE (about 2.9–3.7%) across coefficients.
+Estimating the function-valued LAR is naturally less precise, but still yields normalized rMSEs on the order of 3.4–5.6% for all the coefficients.
+Performance improves with larger $n$ and additional time periods $T$.
+Robustness checks also indicate that second-degree splines outperform first- or third-degree bases, and that adding ex-post shocks increases rMSEs only slightly.
+Before turning to the related literature, I note two limitations of the paper.
+First, the baseline model assumes each coordinate of $g$ is monotone in a single coordinate of $\eta_{it}$, which (as in Imbens and Newey (2009)) rules out certain nonseparable supply-and-demand systems.
+Subsection 3.3 discusses remedies under additional, empirically motivated assumptions.
+Second, the approach requires IVs and a tractable way to summarize the fixed effect $A_{i}$ via a sufficient statistic $W_{i}$; while I provide examples for the applications considered, finding credible instruments (and justifying the required structure) may be challenging in some settings.
 1.1 Related Literature
 This paper contributes to the literature on CRC linear panel models.
-I highlight a few studies that are particularly relevant to this work.
-Chamberlain (1992) studies the regular identification of the first-order moments of the random coefficients when $T>d_{X}$, where $T$ is the number of periods and $d_{X}$ is the dimension of regressors.
-Wooldridge (2005) derives conditions for consistency of the fixed-effects estimators of the APE for a class of linear and nonlinear panel data models with time-invariant random coefficients.
-The condition requires the random coefficients to be mean independent of the residuals from regressing the covariates on the aggregate time variables.
-Graham and Powell (2012) prove irregular identification when $T=d_{X}$ and allow for more persistent processes in the regressors.
-They leverage different identifying content of the subpopulations of “movers” and “stayers” to identify the first-order moments of the time-varying random coefficients.
-Arellano and Bonhomme (2012) identify the first-order moments of the time-invariant random coefficients by using the within-group and between-group variation in the regressors.
-They rely on a mean independence assumption between the residuals and both the regressors and the random coefficients.
-They further identify the second-order moments and distribution functions of the random coefficients by exploiting assumptions on the time-dependence structure of the residuals.
-Graham, Hahn, Poirier, and Powell (2018) show how to identify quantiles of the random coefficients in a general linear quantile regression model.
-Masten (2018) considers a classical linear simultaneous equations model with random coefficients on the endogenous variables.
-By extending the Chamberlain-Mundlak approach for balanced panels, Wooldridge (2019) develops estimation strategies for CRC linear panel models that account for sample selection in unbalanced panels.
-Laage (2024) studies a time-invariant CRC linear panel data model with additive scalar-valued fixed effects.
-She proposes a two-step identification approach, where in the first step she identifies nonparametrically the conditional expectation of the disturbances given the regressors and the control variables, and then in the second step she uses between-group variation to identify the APE.
-The key difference between my paper and the CRC literature discussed above is that my model, (1.1)–(1.2), features time-varying random coefficients that can be correlated with the regressors through both fixed effects and time-varying shocks.
-None of the papers, except Graham and Powell (2012), explicitly considers such correlation.
-Graham and Powell (2012) impose a time-stationarity assumption on the time-varying random shocks given the regressors for all periods and the fixed effect, and exploit the different identifying contents of the subpopulations for identification.
-The novel method proposed by Graham and Powell (2012) covers a wide range of empirical applications.
-However, their main time-stationarity assumption effectively rules out the time-varying endogeneity through the random coefficients, wherein $X_{it}$ and $\beta_{it}$ are correlated through both $A_{i}$ and $\eta_{it}$ [^5] To be precise, the time-stationarity assumption imposed by Graham and Powell (2012) is $\left.\eta_{it}\sim_{d}\eta_{is}\right|{\bf X}_{i},A_{i},\text{ for }t\neq s.$ To see why it rules out time-varying endogeneity through the random coefficients, consider a simplified example in which the number of periods $T=2$ and the data generating process (DGP) of a scalar-valued $X_{it}$ is $X_{it}=Z_{it}^{{}^{\prime}}\lambda+A_{i}+\eta_{it}$.
-Suppose one observes $X_{i2}>X_{i1}$ and $Z_{i2}=Z_{i1}$ in the data, which by the DGP of $X_{it}$ implies $\eta_{i2}>\eta_{i1}$ deterministically given ${\bf X}_{i}$ and $A_{i}$.
-Consequently, the time-stationarity condition is violated when the time-varying endogeneity through the random coefficients is present., which is the focus of this paper.
-Instead, I tackle this issue by proposing two controls for $A_{i}$ and $\eta_{it}$ and, subsequently, using the residual variation in $X_{it}$ to identify the APE and LAR.
-It should be clarified that my work is not necessarily more general than that of Graham and Powell (2012), as the set of assumptions used by each is not nested within the other’s.
-Instead, I consider this paper an initial step toward addressing the important issue of time-varying endogeneity through the random coefficients, which existing methods have not yet fully resolved.
+Related work includes Chamberlain (1992) on regular identification of first moments when $T>d_{X}$; Graham and Powell (2012) on irregular identification when $T=d_{X}$ using movers/stayers; and Arellano and Bonhomme (2012) on identifying first (and, under additional assumptions, higher) moments and distributions via within–between variation and restrictions on residual dependence.
+Wooldridge (2005, 2009) develops estimation strategies for APEs—respectively under mean-independence conditions and with sample selection in unbalanced panels—while Murtazashvili and Wooldridge (2008) study consistency of fixed-effects IV estimators in time-invariant CRC models.
+Additional contributions include identification of quantiles (Graham, Hahn, Poirier, and Powell (2018)), random-coefficient simultaneous equations (Masten (2018)), and a two-step identification approach with additive fixed effects and time-invariant random coefficients (Laage (2024)).
+My contribution to the CRC literature is to study a model, (1.1) and (1.2), with time-varying random coefficients that can be correlated with regressors through both fixed effects and time-varying shocks.
+The closest paper to my work is Graham and Powell (2012), who identify moments by exploiting movers and stayers under a time-stationarity restriction.
+This time-stationarity restriction conditions on the regressors for all periods and thus effectively rules out the time-varying endogeneity central here.
+I instead introduce controls for $A_{i}$ and importantly $\eta_{it}$ to identify the APE and LAR.
+The two approaches rely on non-nested assumptions; I view this paper as a step toward addressing time-varying endogeneity through random coefficients.
 In addition to the aforementioned papers, my work is also related to the literature on: (i) time-varying parameter panel data models (e.g., Li, Chen, and Gao (2011, 2016)), (ii) latent structure panel data models (e.g., Su, Shi, and Phillips (2016, 2019), Wang, Phillips, and Su (2024)), and (iii) linear panel data model with interactive fixed effects, which can be treated as imposing a specific structure on the random coefficient associated with the constant term (e.g., Pesaran (2006), Bai (2009), Moon and Weidner (2015)).
@@
 The second line of research concerns the techniques used in this paper.
-My method requires the existence of a sufficient statistic $W_{i}$ for the time-invariant $A_{i}$ and a feasible control variable $V_{it}$ for the time-varying $\eta_{it}$.
-To justify the sufficiency of $W_{i}$ for $A_{i}$, I present several techniques from the literature (e.g., Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)).
-See Liu, Poirier, and Shiu (2024) for a detailed discussion on the index sufficiency condition for the fixed effects in nonlinear semiparametric panel data models.
-It may be worth pointing out that the exchangeability condition used by Altonji and Matzkin (2005) concerns the conditional probability density function (PDF) of $A_{i}\text{ and }\eta_{it}$ given ${\bf X}_{i}$.
-I show that it is not directly applicable to the TERC model in which $X_{it}$ depends on $\eta_{it}$ for each $t$.
-Instead, I impose a lower-level exchangeability condition involving only the unobservable $A_{i}$ and $\eta_{it}$ and use it to establish a sufficiency result for $A_{i}$ in Proposition 1.
-To find the feasible control variable for $\eta_{it}$ in (1.2), I generalize the technique of Imbens and Newey (2009) in two nontrivial ways.
-First, because the unknown $A_{i}$ enters equation (1.2) together with $\eta_{it}$, the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ proposed by Imbens and Newey (2009) does not uniquely determine $\eta_{it}$ even if it is a scalar.
-Instead, I use the index exclusion condition on $A_{i}$ to construct a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ given $A_{i}$ and $W_{i}$.
-Second, I allow $\eta_{it}$ to be vector-valued, enabling the model to capture the rich heterogeneity commonly observed in empirical applications.
-In the baseline model, I assume that each coordinate of $X_{it}$ depends upon a single coordinate of $\eta_{it}$.
-As an extension, I discuss how to relax this restriction when additional empirically motivated assumptions are imposed.
-Moreover, constructing $V_{it}$ requires IVs that satisfy specific exogeneity conditions.
-Murtazashvili and Wooldridge (2008) provide a set of conditions sufficient for the consistency of a general class of fixed effects IV estimators in a time-invariant individual-specific CRC panel data model.
-The model and method of this paper are different from theirs.
-The asymptotic analysis in this paper builds on a series of foundational works.
-Andrews (1991) investigates the asymptotic properties of series estimators for nonparametric and semiparametric regression models.
-His results are applicable to a wide variety of estimands, including derivatives and integrals of the regression function.
-I use his results to cover vector-valued functionals of regression functions.
-Newey (1997) also studies series estimators and provides conditions for proving convergence rates and asymptotic normality for the estimators of conditional expectations.
-Newey, Powell, and Vella (1999) propose a two-step nonparametric estimator for a triangular simultaneous equation model with a separable first-step equation.
-They derive asymptotic normality for their two-step estimator with the first-step estimator plugged in.
-Imbens and Newey (2009) also analyze a triangular simultaneous equation model, but with a nonseparable first-step equation.
-They obtain mean-squared convergence rates for the first-step estimator and prove asymptotic normality (in their 2002 working paper version) for known functionals of the conditional expectation of the outcome variable given the regressors and control variables.
-I extend their analysis to establish asymptotic normality for unknown but estimable functionals of conditional expectation functions.
+For nonseparable panel models with generalized fixed effects, restrictions on the fixed effect are generally needed to identify global objects such as the APE and LAR (Hoderlein and White (2012, p.2); see also Gao, Li, and Xu (2023) and Gao and Li (2026) on dealing with the issue of fixed effects in network formation and panel multinomial choice models, respectively).
+My method requires a sufficient statistic $W_{i}$ for $A_{i}$ and a control variable $V_{it}$ for $\eta_{it}$.
+To justify the sufficiency of $W_{i}$ for $A_{i}$, I adapt several techniques from the literature (e.g., Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)).
+See Liu, Poirier, and Shiu (2025) for a discussion on the index sufficiency condition in nonlinear semiparametric panel data models.
+To justify $V_{it}$ as a valid control for $\eta_{it}$, I generalize the technique of Imbens and Newey (2009) in two nontrivial ways and discuss them in relation to (3.15).
+More recently, Nagasawa (2024) studies treatment effect estimation with noisy conditioning variables using control variables from the joint distribution of noisy proxy measures.
+The asymptotic analysis builds on foundational results for series and control-function estimators.
+Andrews (1991) and Newey (1997) provide conditions for convergence rates and asymptotic normality of series estimators, which I use to handle vector-valued functionals of regression functions.
+Newey, Powell, and Vella (1999) establish asymptotic normality for two-step nonparametric estimators in triangular models with a separable first stage, while Imbens and Newey (2009) extend the triangular framework to nonseparable first stages and derive rates and asymptotic normality for known functionals of conditional expectations.
+I build on Imbens and Newey (2009) to obtain asymptotic normality for unknown but estimable functionals of conditional expectation functions.
+Organization.
 The rest of the paper is organized as follows.
@@
 Finally, Section 6 concludes.
-Proofs of all theorems, Proposition 1, and a simulation study are included in Appendix A, B, and C, respectively.
+Proofs, additional empirical results, and a simulation study are included in Appendix A, B, and C, respectively.
+Notation.
+Let $i\in\left\{1,\ldots,n\right\}$ index agents and $t\in\left\{1,\ldots,T\right\}$ index periods, with finite $T\geq 2$.
+I use boldface upper case letters for random matrices, regular upper case letters for random variables and random vectors, and lower case letters for their values.
+Let ${\bf X}_{i}=\left(X_{i1},...,X_{iT}\right)^{\prime}$, $Y_{i}=\left(Y_{i1},...,Y_{iT}\right)^{\prime}$, ${\bf Z}_{i}=\left(Z_{i1},...,Z_{iT}\right)^{\prime}$, $\varepsilon_{i}=\left(\varepsilon_{i1},...,\varepsilon_{iT}\right)^{\prime}$, and $\eta_{i}=\left(\eta_{i1},...,\eta_{iT}\right)^{\prime}$.
+I assume $\left({\bf X}_{i},{\bf Z}_{i},A_{i},\varepsilon_{i},\eta_{i}\right)$ are i.i.d.
+across $i$.
+Let $d_{X}$ be the dimension of $X_{it}$ and $X_{it,l}$ denote the $l^{\text{th}}$ coordinate of vector $X_{it}$.
+Similar notation is used for all other variables.
+I use $\coloneqq$ to define a new random variable, $\sim_{d}$ to indicate that two random variables are identically distributed, $F$ for CDF, $f$ for probability density function (PDF), $\mathbb{E}$ for expectation, $\mathbb{\mathbb{V}}$ for variance, $I_{d}$ for $d\times d$ identity matrix, $\overset{p}{\longrightarrow}$ for convergence in probability, and $\overset{d}{\longrightarrow}$ for convergence in distribution.
+All convergence results are stated under $n\to\infty$.
 2 Model
-Let $i\in\left\{1,...,n\right\}$ index $n$ agents and $t\in\left\{1,...,T\right\}$ index finite $T\geq 2$ periods.
-Let $d_{X}$ be the dimension of $X_{it}$ and similarly for $Z_{it},$ $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$.
 Consider the following baseline TERC model
-$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (2.1)
+$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$ (2.1)
 $X_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right),$ (2.2)
-where:
-  • $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{d_{X}}$, the central object of interest, is a vector of random coefficients modeled as a vector-valued, unknown, and potentially nonlinear function $\beta\left(\cdot\right)$ of $A_{i}\text{ and }\varepsilon_{it}$.
-Here, $A_{i}$ represents a fixed effect and $\varepsilon_{it}$ drives the time-varying behavior of $\beta_{it}$, both of which can be either scalar- or vector-valued.
-Furthermore, $A_{i}$ and $\varepsilon_{it}$ can be arbitrarily correlated.
-  • $\eta_{it}\in\mathbb{R}^{d_{\eta}}$ is a continuously distributed and time-varying random vector that influences the value of $X_{it}$.
-$\eta_{it}$ is allowed to be correlated with both $A_{i}$ and $\varepsilon_{it}$.
-Additionally, different coordinates of $\eta_{it}$ can also be correlated.
-  • $X_{it}\in\mathbb{R}^{d_{X}}$ is a vector of choice variables for agent $i$ in period $t$ [^6] For clarity of argument, in the baseline model I assume all coordinates of $X_{it}$ to be endogenous in the sense that they can be correlated with $\beta_{it}$ through both $A_{i}\text{ and }\eta_{it}$.
-In fact, the proposed method in this paper can benefit from additional exogenous regressors in (2.1).
-I discuss it in relation to Assumption 4 and note the changes needed when additional exogenous regressors are included in (2.1) at the end of the proof of Theorem 1.
-$Y_{it}\in\mathbb{R}$ denotes the scalar-valued outcome variable, and $Z_{it}\in\mathbb{R}^{d_{Z}}$ represents a vector of IVs.
-  • $g\left(\cdot\right)$ is an unknown vector-valued function of $Z_{it},\ A_{i},\ \text{and }\eta_{it}$ that determines the value of each coordinate of $X_{it}$.
-[…]
-To clarify the model’s information structure, first nature draws $\left(A_{i},\varepsilon_{it}\right)$ for each agent and period in the economy.
-Then, agent $i$ in period $t$ chooses her endogenous $X_{it}$ by solving an optimization problem (e.g., profit maximization) after observing $A_{i}$, some signal $\eta_{it}$ about $\varepsilon_{it}$, and exogenously determined IVs $Z_{it}$, leading to (2.2).
-Finally, agent $i$ obtains its outcome $Y_{it}$ via (2.1).
-As an analyst, I observe $\left\{Y_{it},X_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify $\mathbb{E}\beta_{it}$ (APE) and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR).
-I illustrate the information structure in Figure 1.
-Nature: draw ($A_{i}$,$\varepsilon_{it}$)Agent: choose $X_{it}$ given ($Z_{it}$,$A_{i}$,$\eta_{it}$)Agent: realize $Y_{it}$ via (2.1)Analyst: observe ($Y_{it}$,$X_{it}$,$Z_{it}$)
-Figure 1: Information Structure
-Model (2.1)–(2.2) appears in various economic applications. I describe three of them here.
-Example 1 (Heterogeneous Production Function Estimation).
+where $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{d_{X}}$, the central object of interest, is a vector of random coefficients modeled as an unknown vector-valued function $\beta$ of $(A_{i},\varepsilon_{it})$.
+Here, $A_{i}$ is a fixed effect and $\varepsilon_{it}$ governs the time-varying behavior of $\beta_{it}$; both may have arbitrary dimension.
+The mapping $\beta$ itself can be time-varying, because $\varepsilon_{it}$ captures time variation in the functional form.
+$\eta_{it}\in\mathbb{R}^{d_{\eta}}$ is a continuously distributed, time-varying random vector that may be correlated with $A_{i}$ and $\varepsilon_{it}$; its coordinates may also be mutually correlated.
+This correlation can be important in applications; see the examples below.
+Let $X_{it}\in\mathbb{R}^{d_{X}}$ denote choice variables (e.g., capital and labor), $Y_{it}\in\mathbb{R}$ the outcome, and $Z_{it}\in\mathbb{R}^{d_{Z}}$ instruments.
+Finally, $g$ is an unknown vector-valued function of $(Z_{it},A_{i},\eta_{it})$ that determines each coordinate of $X_{it}$.
+The TERC model is widely used in empirical studies.
+It extends the classic constant-coefficient linear panel model while preserving directly interpretable targets—such as the APE and LAR—that carry clear policy relevance.
+It also accommodates natural extensions, including higher-order moments of the APE and LAR.
+In what follows, I describe three applications of the TERC model.
+Example 1 (Production Function Estimation).
 Suppose firm $i$ in year $t$ has a heterogeneous Cobb-Douglas production function in the form of (2.1).
-The capital and labor elasticities $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{2}$ are modeled as a two-dimensional function of the firm fixed effect $A_{i}$ (e.g., manager ability) and a time-varying random shock $\varepsilon_{it}$ (e.g., realized per-period technology shock).
-When deciding its input values $X_{it}\in\mathbb{R}^{2}$ for capital and labor, assume firm $i$ knows $A_{i}$ and $\eta_{it}$ (e.g., expected per-period technology shock).
+The capital and labor elasticities $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{2}$ are modeled as a two-dimensional function of the firm fixed effect $A_{i}$ (e.g., manager ability that is assumed to be stable over time) and a time-varying random shock $\varepsilon_{it}$ (e.g., realized per-period technology shock).
+When deciding its $X_{it}\in\mathbb{R}^{2}$ of capital and labor, assume firm $i$ knows $A_{i}$ and $\eta_{it}$ (e.g., expected per-period technology shock).
 Clearly, $\eta_{it}$ is correlated with $A_{i}$ and $\varepsilon_{it}$.
-Assume firm $i$ also observes some exogenously determined input prices $Z_{it}$, such as interest rates and wages, and that the cost function is given by $c\left(x,z\right)$.
-Then, firm $i$ chooses input values $X_{it}$ by maximizing its expected profit with the knowledge of $\left(Z_{it},A_{i},\eta_{it}\right)$, i.e.,
-$X_{it}=\arg\max_{x\in\mathbb{R}^{2}}\left[\mathbb{E}\left[\left.x^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right)\right|Z_{it},A_{i},\eta_{it}\right]-c\left(x,Z_{it}\right)\right],$
+Assume firm $i$ also observes $Z_{it}$, which captures local input market dynamics that affect $X_{it}$—for example, competitors’ debt-weighted interest rates and employment-weighted wages.
+Let the cost function be given by $c\left(x,z\right)$.
+Then, firm $i$ selects $X_{it}$ by maximizing its expected profit with the knowledge of $\left(Z_{it},A_{i},\eta_{it}\right)$, i.e.,
+$X_{it}=\arg\max_{x\in\mathbb{R}^{2}}\left[\mathbb{E}\left[\left.x^{\prime}\beta\left(A_{i},\varepsilon_{it}\right)\right|Z_{it},A_{i},\eta_{it}\right]-c\left(x,Z_{it}\right)\right],$
 leading to (2.2). Subsequently, firm $i$ obtains its output $Y_{it}$ via (2.1).
-Example 2 (Labor Supply Estimation).
-Suppose agent $i$ has a linear labor supply function of the form (2.1), where $Y_{it}$ represents the number of hours worked and $X_{it}$ includes the endogenous hourly wage along with other exogenous demographic variables.
-The coordinate of $\beta_{it}$ corresponding to the wage is the key object of interest, as it quantifies how labor supply responds to variations in wage rate over time.
-Given some exogenous IVs $Z_{it}$ (e.g., county minimum wage and non-labor income), individual capability $A_{i}$, and expected health shocks $\eta_{it}$ (e.g., the likelihood of contracting Covid-19 at work), agent $i$ chooses a job with the wage that maximizes her expected utility, resulting in (2.2).
-Example 3 (Engel Curve Estimation).
-Suppose the budget share of gasoline $Y_{it}$ for household $i$ at time $t$ is a function of gas price and total expenditure represented in (2.1).
-Here, $\beta_{it}\in\mathbb{R}^{2}$ is modeled as a vector of unknown functions of the household fixed effect $A_{i}$ and a wealth shock $\varepsilon_{it}$, and captures the elasticity of gasoline demand with respect to total expenditure and gas price, respectively.
-Given $A_{i}$, a signal $\eta_{it}$ about the wealth shock $\varepsilon_{it}$, and an IV $Z_{it}$ (e.g., head of household’s gross income), household $i$ searches for a gas price and plans its total expenditure budget by maximizing its expected utility, leading to (2.2).
-In these examples, the time-varying correlation between $X_{it}$ and $\beta_{it}$ highlights the significance and importance of the time-varying endogeneity through the random coefficients, which naturally arises due to agent optimization behavior.
-Numerous classic papers (Manski (1987), Altonji and Matzkin (2005), Graham and Powell (2012), Chernozhukov, Fernández-Val, Hahn, and Newey (2013)) have addressed such endogeneity by allowing for the correlation between $X_{it}$ and $A_{i}$.
-For instance, in production function estimation, a firm with a good management team is typically associated with larger input levels to leverage its managerial skill.
-This type of correlation has been addressed in the aforementioned papers.
-However, it carries important empirical implication to allow $X_{it}$ and $\beta_{it}$ to be correlated through factors beyond $A_{i}$.
-For instance, firms with a good management team may strategically select low levels of inputs for certain periods, which can not be explained only by the correlation between $X_{it}$ and $A_{i}$.
-To account for such behavior, it is useful to introduce an additional time-varying shock $\eta_{it}$ that contains important information about $\beta_{it}$ and affects the value of $X_{it}$.
-Allowing $A_{i}$ to enter both the first-step equation (2.2) and the outcome equation (2.1) is another novel feature of this paper, distinguishing it from classic triangular simultaneous equations models (e.g., Newey, Powell, and Vella (1999), Imbens and Newey (2009)).
-In these models, the assumption is typically made on the first-step equation that determines the values of regressors that there is only one unobservable scalar-valued random variable $\eta_{it}$ that is correlated with both $A_{i}$ and $\varepsilon_{it}$, i.e., $X_{it}=g\left(Z_{it},\eta_{it}\right)$ instead of $X_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$.
-Under regularity conditions, they propose to control for $\eta_{it}$ via either the regression residual from regressing $X_{it}$ on $Z_{it}$ (e.g., Newey, Powell, and Vella (1999)) or the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ which equals $F_{\eta_{it}}\left(\eta_{it}\right)$ (e.g., Imbens and Newey (2009)).
-However, having $A_{i}$ appear in (2.2) renders the residual and CDF-based methods inapplicable because the one-to-one relationship between the control variable and $\eta_{it}$ no longer holds.
-Additionally, my model features a nonseparable first-step equation (2.2) in which $A_{i}$ enters $g\left(\cdot\right)$ potentially nonlinearly.
-The nonseparability and nonlinearity arise naturally from agent optimization behavior as illustrated in the empirical examples.
-The nonlinearity precludes the use of common demeaning or first-differencing techniques to eliminate $A_{i}$ (e.g., Laage (2024)).
-Instead, I rely on an index exclusion condition to deal with the fixed effects (e.g., Liu, Poirier, and Shiu (2024)).
+Example 2 (Labor Supply).
+Suppose worker $i$ has a linear labor supply function of the form (2.1), where $Y_{it}$ is the number of hours worked and $X_{it}$ includes the endogenous hourly wage ($X_{it,1}$) along with other exogenous demographic variables.
+The coordinate of $\beta_{it}$ corresponding to the wage is the key object of interest because it quantifies the elasticity of labor supply with respect to the wage rate.
+Given exogenous factors $Z_{it}$ (e.g., county minimum-wage changes or earned income tax credit expansions), individual ability $A_{i}$, and shock $\eta_{it}$ (e.g., her expectation of a near-term surge in labor demand in her sector, which increases the expected marginal product of labor and wage offers), she chooses a job with the wage that maximizes her expected utility, resulting in (2.2).
+Finally, worker $i$ commits $Y_{it}$ hours of working via (2.1).
+Example 3 (Almost Ideal Demand System (AIDS)).
+Suppose household $i$’s gasoline budget share $Y_{it}$ at time $t$ depends on the gasoline price $X_{it,1}$ and total expenditure $X_{it,2}$ as in (2.1).
+Let $\beta_{it}\in\mathbb{R}^{2}$ denote the possibly heterogeneous elasticities with respect to $X_{it}$, modeled as unknown functions of a household fixed effect $A_{i}$ (e.g., a persistent propensity to drive versus use transit) and a time-varying wealth shock $\varepsilon_{it}$.
+Given instruments $Z_{it}$ (e.g., head of household’s earned income and gas tax change), $A_{i}$, and a signal $\eta_{it}$ about $\beta_{it}$ (e.g., expected near-term income change such as upcoming bonus), the household searches for gasoline prices and sets its total expenditure budget by maximizing expected utility, which yields (2.2).
+The time-varying correlation between $X_{it}$ and $\beta_{it}$ illustrates a distinct source of endogeneity: optimization-induced selection into regressors through random coefficients.
+Classic work (Manski (1987), Altonji and Matzkin (2005), Graham and Powell (2012), Chernozhukov, Fernández-Val, Hahn, and Newey (2013)) allows $X_{it}$ to correlate with $A_{i}$ (e.g., better-managed firms choose higher inputs).
+But empirically, $X_{it}$ can also co-move with $\beta_{it}$ beyond $A_{i}$, because time variation in choices is rarely fully spanned by instrument variation $Z_{it}$.
+For example, even with little movement in input prices, a well-managed firm may temporarily scale inputs up or down in response to private information about returns.
+This motivates a time-varying shock $\eta_{it}$ that affects $X_{it}$ and is informative about $\beta_{it}$.
+A second departure from standard triangular models (Newey, Powell, and Vella (1999), Imbens and Newey (2009)) is that $A_{i}$ enters both the outcome equation and the first-step choice equation: $X_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$, not $X_{it}=g\left(Z_{it},\eta_{it}\right)$.
+Consequently, residual- or CDF-based control-function approaches that rely on a one-to-one mapping between the control and $\eta_{it}$ no longer apply.
+Moreover, $g$ is nonseparable and can depend on $A_{i}$ nonlinearly, as implied by optimization, so demeaning or first-differencing cannot eliminate $A_{i}$.
+I instead use an index exclusion condition to handle the fixed effects.
 3 Identification
 3.1 Identification Idea
-The main challenge to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ comes from the correlation between $X_{it}$ and $\beta_{it}$ in model (2.1)–(2.2) through their dependence on the pairwise correlated random variables $A_{i},\ \varepsilon_{it},$and $\eta_{it}$.
-Suppose I follow the standard identification argument for linear models by taking the conditional expectation of both sides of (2.1) given $X_{it}$:
-$\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right].$ (3.1)
-Then, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, the textbook methods to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ fail.
-To see why, first I cannot identify $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by multiplying both sides of (3.1) by $X_{it}$, taking outer expectation with respect to $X_{it}$, and finally inverting $\mathbb{E}\left[X_{it}X_{it}^{{}^{\prime}}\right]$:
-$\mathbb{E}\left[X_{it}Y_{it}\right]=\mathbb{E}\left\{X_{it}X_{it}^{{}^{\prime}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}}\mathinner{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left[\left.\beta_{it}\right|X_{it}\right]}\right\},$ (3.2)
-since $X_{it}X_{it}^{{}^{\prime}}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$, a function of $X_{it}$, are not separable inside the expectation on the right hand side (RHS) of (3.2).
-Second, the perturbation-based identification method in linear models does not go through either. This method perturbs $X_{it}$ on both sides of (3.1):
-$\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]/\partial X_{it}=\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{{}^{\prime}}\partial\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]/\partial X_{it}}.$ (3.3)
-Again, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, by the chain rule there is an additional term color-coded in red showing up on the RHS of (3.3), invalidating the perturbation-based method.
-The key identification idea of this paper is to address the time-varying correlation between $X_{it}$ and $\beta_{it}$ using two control variables, $W_{i}$ and $V_{it}$.
-Conditional on $W_{i}$ and $V_{it}$, the residual variation in $X_{it}$ is determined solely by the exogenous $Z_{it}$.
-First, since the APE and LAR are generally not identified if the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted (Liu, Poirier, and Shiu (2024)), I exploit an index exclusion condition, $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$, to control for $A_{i}$, where the existence and construction of $W_{i}$ are supported by the established results in the literature (Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)).
-Then, for $\eta_{it}$ I propose a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and prove that $V_{it}=F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$, the RHS of which is a one-to-one function of $\eta_{it}$ given $A_{i}$ and $W_{i}$.
-Since $A_{i}$ is unknown, I apply the LIE to $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ by further conditioning on the unknown $A_{i}$ in the inner conditional expectation to fix $\eta_{it}$ via $V_{it}$, such that conditional on $A_{i},\ V_{it},\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely driven by the exogenous $Z_{it}$.
-This allows me to exclude $X_{it}$ from the inner conditional expectation.
-Next, I exclude $X_{it}$ from the conditioning set of the outer conditional expectation using the index exclusion condition of $W_{i}$ for $A_{i}$, i.e., $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$.
-Hence, I have
-$\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|X_{it},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A_{i}},V_{it},W_{i}\right]\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right],$ (3.4)
-which by (2.1) implies
-$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right].$ (3.5)
-Given sufficient variation in $X_{it}$ conditional on $V_{it}$ and $W_{i}$, I identify $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ by either
-$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right]$ (3.6)
-or
-$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}.$ (3.7)
-Consequently, I use the LIE to identify
-$\mathbb{E}\beta_{it}=\mathbb{E}\left[\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right]\text{ and }\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right|X_{it}\right].$ (3.8)
-3.2 Assumptions and Main Identification Result
+Why Standard Linear Identification Fails
+The main obstacle to identifying the APE and the LAR function in the TERC model using standard linear identification is that the regressor is chosen by the agent using information that is also informative about the random coefficients.
+To see where the standard argument breaks down, take conditional expectations in (2.1) given $X_{it}$:
+$\mathbb{E}[\left.Y_{it}\right|X_{it}]=X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|X_{it}].$ (3.1)
+In a constant-coefficient model, the term $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$ would be a constant vector and could be recovered from moment conditions.
+Here, however, $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$ is generally a function of $X_{it}$.
+As a result, the familiar argument based on multiplying both sides of (3.1) by $X_{it}$, taking expectations, and inverting $\mathbb{E}[X_{it}X_{it}^{\prime}]$ does not identify $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$:
+$\mathbb{E}[X_{it}Y_{it}]=\mathbb{E}\left[X_{it}X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|X_{it}]\right],$ (3.2)
+because $X_{it}X_{it}^{\prime}$ and $\mathbb{E}[\beta_{it}\mid X_{it}]$ move together inside the expectation.
+Likewise, the “derivative/perturbation” argument for linear models does not isolate $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$, since differentiating (3.1) yields
+$\frac{\partial\mathbb{E}[\left.Y_{it}\right|X_{it}]}{\partial X_{it}}=\mathbb{E}[\left.\beta_{it}\right|X_{it}]+\left(\frac{\partial\mathbb{E}[\left.\beta_{it}\right|X_{it}]}{\partial X_{it}}\right)^{\prime}X_{it},$ (3.3)
+and the second term captures how selection into different $X_{it}$ changes the conditional distribution of $\beta_{it}$.
+Motivation and Interpretation of the Two Control Variables
+The identification strategy in this paper is to recreate, using observables, the thought experiment “vary $X_{it}$ while holding fixed everything the agent knows that is related to $\beta_{it}$.” Doing so requires two controls—one for the time-invariant heterogeneity $A_{i}$ and one for the time-varying information $\eta_{it}$.
+The first control is a statistic $W_{i}=W({\bf X}_{i},{\bf Z}_{i})$ constructed from the individual’s panel history.
+The maintained index exclusion condition implies that once I condition on $W_{i}$, the remaining variation in $(X_{it},Z_{it})$ for a fixed $t$ contains no additional information about $A_{i}$:
+$\left.A_{i}\right|(X_{it},Z_{it},W_{i})\sim_{d}\left.A_{i}\right|W_{i}.$ (3.4)
+Intuitively, $W_{i}$ is chosen to summarize the persistent, individual-specific component of behavior that is linked to $A_{i}$.
+For example, in many panel applications a time average of $X_{it}$ (possibly together with time averages of instruments) is a natural candidate: firms with higher managerial ability or individuals with higher baseline productivity tend to exhibit persistently different choices, and $W_{i}$ is designed to capture that persistent component.
+Even after accounting for $A_{i}$, endogeneity remains because agents may respond to time-varying information $\eta_{it}$ when choosing $X_{it}$.
+The second control is constructed as the conditional CDF (a “rank” or “generalized residual”):
+$V_{it}:=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right).$ (3.5)
+The interpretation is that $V_{it}$ records where the realized $X_{it}$ sits in the distribution of $X_{it}$ among agents with the same $(Z_{it},W_{i})$.
+Since $A_{i}$ and $Z_{it}$ appear in the first-step equation (1.2) and $W_{i}$ controls for $A_{i}$ by (3.4), the residual variation of $X_{it}$ given $(Z_{it},W_{i})$ is only driven by $\eta_{it}$ which is assumed to enter (1.2) strictly monotonically.
+For example, consider two firms with the same observed average inputs over all periods ($W_{i}$) and input prices in period $t$ ($Z_{it}$).
+If firm 1’s choice of capital ($X_{it}$) in this period is higher than that of firm 2, it can be inferred that firm 1 receives a larger productivity shock ($\eta_{it}$) to its capital choice function in period $t$ than firm 2 since (3.4) ensures that the residual variation in $A_{i}$ given $W_{i}$ is not informative about the choice of $X_{it}$.
+As a result, $V_{it}$ records where the realized $\eta_{it}$ sits in the distribution of $\eta_{it}$ among agents with the same $(Z_{it},A_{i},W_{i})$.
+By an exogeneity assumption between $Z_{it}$ and $\eta_{it}$, it drops out from the conditioning set and hence $V_{it}$ controls for $\eta_{it}$ given $(A_{i},W_{i})$.
+Thus, conditioning on $(V_{it},W_{i})$ is intended to hold fixed both the individual’s persistent type (through $W_{i}$) and the individual’s period-$t$ information that drives endogenous adjustment (through $V_{it}$).
+This two-control structure matches the economics of the motivating examples in Section 2.
+In the production function estimation example, $W_{i}$ captures persistent firm heterogeneity (e.g., management skill) that affects both average input choices and average elasticities, while $V_{it}$ captures time-$t$ private information or expectations about technology that shift input choices and are correlated with realized productivity shocks.
+In the labor supply application, $W_{i}$ summarizes persistent ability/tastes and $V_{it}$ captures time-varying job-match information that affects wage choice.
+In the AIDS application, $W_{i}$ captures persistent household heterogeneity and $V_{it}$ captures time-varying wealth or liquidity information that affects expenditure plans.
+Steps to Identify APE and LAR
+Given the two controls $V_{it}$ and $W_{i}$, the central step to identification of the APE and LAR is to show that once I condition on the two controls, $X_{it}$ no longer carries additional information about $\beta_{it}$.
+The argument proceeds in two layers.
+The first step tackles the time-varying part of the endogeneity by assuming $A_{i}$ is known and showing
+$\mathbb{E}[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}]=\mathbb{E}[\left.\beta_{it}\right|A_{i},V_{it},W_{i}].$ (3.6)
+Specifically, since $V_{it}$ records where the realized $\eta_{it}$ sits in the distribution of $\eta_{it}$ among agents with the same $(A_{i},W_{i})$, holding $(A_{i},V_{it},W_{i})$ fixed amounts to holding $(A_{i},\eta_{it})$ fixed.
+In this step, Assumption 1 (Componentwise Monotonicity) is used to invert $\eta_{it}$ out from function $g$, Assumption 2 (Index Exclusion) allows one to add $A_{i}$ into the conditioning set of the CDF $V_{it}$, and Assumption 3(a) (Exogeneity of IV) enables one to exclude $Z_{it}$ from the conditioning set of the CDF $V_{it}$.
+Assumption 3(b) (Monotonicity of the Conditional CDF) then implies that $V_{it}$ is a one-to-one transformation of $\eta_{it}$ given $(A_{i},W_{i})$, proving the claim that $(A_{i},V_{it},W_{i})$ uniquely pins down $(A_{i},\eta_{it},W_{i})$.
+Hence, equation (1.2) implies that the remaining variation in $X_{it}$ is generated only by $Z_{it}$.
+Under the maintained IV exogeneity condition, this residual variation is independent of $\beta_{it}$.
+This is the point at which the time-varying part of the endogeneity problem is resolved: within cells defined by $(A_{i},V_{it},W_{i})$, movements in $X_{it}$ are “as good as randomized” by $Z_{it}$.
+In the production function estimation example, if two firms have the same management ability ($A_{i}$) and expected per-period technology shock ($\eta_{it}$) but selects different capital or labor, such difference can only be explained by the variation in exogenous $Z_{it}$, which is causal and thus can be exploited to identify the APE and LAR.
+The second step tackles the time-invariant part of the endogeneity.
+Since $A_{i}$ is not observed, it is not feasible to condition on it in (3.6).
+I deal with unknown $A_{i}$ by integrating it out.
+This involves iterating expectations and the index exclusion restriction of $W_{i}$.
+Specifically, I show
+$\mathbb{E}[\left.\mathbb{E}[\left.\beta_{it}\right|A_{i},V_{it},W_{i}]\right|X_{it},V_{it},W_{i}]=\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.7)
+Notice that on the left-hand side of (3.7), $X_{it}$ affects the conditional expectation only through the conditional density $f_{\left.A_{i}\right|X_{it},V_{it},W_{i}}$.
+The index exclusion restriction (Assumption 2) implies that, conditional on $W_{i}$, $A_{i}$ is independent of $(X_{it},Z_{it})$ and thus of $V_{it}$.
+For instance, in the production function estimation example, $W_{i}$ captures the firm’s time-invariant management ability $A_{i}$, so once $W_{i}$ is held fixed, period-$t$ inputs $X_{it}$ and control for productivity shock $V_{it}$ only reflect transitory variation and add no extra information about $A_{i}$.
+Therefore, I have
+$f_{\left.A_{i}\right|X_{it},V_{it},W_{i}}(\left.A_{i}\right|X_{it},V_{it},W_{i})=f_{\left.A_{i}\right|W_{i}}(\left.A_{i}\right|W_{i}),$ (3.8)
+which ensures that (3.7) holds.
+Combining (3.6) and (3.7) gives
+$\mathbb{E}[\left.\beta_{it}\right|X_{it},V_{it},W_{i}]=\mathbb{E}\left[\left.\mathbb{E}[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}]\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.9)
+Equation (3.9) is the key identification result: after conditioning on $(V_{it},W_{i})$, the dependence of $\beta_{it}$ on $X_{it}$ is removed.
+Given (3.9), equation (2.1) implies a conditional linear representation
+$\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]=X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.10)
+With sufficient residual variation in $X_{it}$ conditional on $(V_{it},W_{i})$, $\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]$ can be identified either from conditional second moments
+$\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]=\mathbb{E}[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}]^{-1}\mathbb{E}[\left.X_{it}Y_{it}\right|V_{it},W_{i}],$ (3.11)
+or from the derivative of the conditional mean
+$\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]=\frac{\partial\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]}{\partial X_{it}}.$ (3.12)
+Finally, identification of the APE and LAR follows by the law of iterated expectations
+$\overline{b}=\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]\right],\quad\text{and}\quad b_{t}(X_{it})=\mathbb{E}\left[\left.\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]\right|X_{it}\right].$ (3.13)
+In the production function estimation example, $\overline{b}$ summarizes output elasticities by averaging across firms and time, so it plays the same role as the constant coefficient in standard linear panel regressions.
+$b_{t}(x)$ denotes the period-$t$ average elasticity conditional on operating at input level $x$, a policy-relevant local parameter when elasticities are systematically related to firms’ input choices.
+Analogous interpretations apply to labor supply (average and local wage elasticities) and demand estimation (average and local expenditure/price elasticities).
+The purpose of $(W_{i},V_{it})$ is precisely to make these interpretations credible by ensuring that the remaining variation used to identify (3.13) is driven by the exogenous instruments rather than by unobserved, coefficient-relevant information.
+3.2 Assumptions
 Next, I provide a list of assumptions on the model primitives required for the identification analysis and discuss them in relation to model (2.1)–(2.2).
-Assumption 1 (Model).
-Each coordinate of $g\left(Z_{it},A_{i},\eta_{it}\right)$ depends on one coordinate of $\eta_{it}$ strictly monotonically.
-[…]
-Assumption 1 requires justification when $d_{X}>1$, $d_{\eta}>1$, and each endogenous coordinate of $X_{it}$ admits a different coordinate of $\eta_{it}$.
-For example, it is possible that firms have different departments making hiring and investment decisions separately.
-Such decisions are based on different signals from the labor market and financial market.
-Therefore, it requires different coordinates of $\eta_{it}$ to enter the corresponding coordinate of $X_{it}$.
-Additionally, it is natural to allow firms to take advantage of positive shocks from labor market and financial market by hiring more employees and making more investments, leading to coordinate-wise strict monotonicity of $g\left(\cdot\right)$ in $\eta_{it}$.
-As a result, Assumption 1 is satisfied in this context.
-It is worth highlighting that Assumption 1 does not restrict how different coordinates of $\eta_{it}$ are correlated.
-Thus, without loss of generality, I can redefine $\eta_{it}$ such that $d_{\eta}=d_{X}$ and each coordinate of $g\left(\cdot\right)$ depends on the corresponding coordinate of $\eta_{it}$.
-Moreover, Assumption 1 does not restrict how $A_{i}$ and $\eta_{it}$ are correlated, which could be important to empirical applications.
-Still, Assumption 1 restricts how the vector of $\eta_{it}$ enters the $g\left(\cdot\right)$ function that determines the value of $X_{it}$, which is a limitation of the paper.
-In Section 3.3, I discuss two other specifications to allow multiple coordinates of $\eta_{it}$ to enter one coordinate of $g\left(\cdot\right)$ when additional timing assumptions are considered plausible by practitioners and show how to adapt my method to these alternative specifications.
+Assumption 1 (Componentwise Monotonicity).
+For every $(Z_{it},A_{i})$ and each $j\in\{1,\ldots,d_{X}\}$, $g_{j}\left(Z_{it},A_{i},\eta_{it}\right)$ depends on $\eta_{it}$ only through $\eta_{it,j}$ and is strictly monotone in $\eta_{it,j}$.
+The role of Assumption 1 is to deliver the invertibility/control-function step used in the identification argument.
+This assumption is mild when $\eta_{it}$ is scalar: it is automatically satisfied in additive triangular specifications (e.g., Newey, Powell, and Vella (1999)), and in the nonseparable scalar case it is essentially the same as the monotonicity condition in Imbens and Newey (2009, eq.
+(2)), except that I also allow for time-invariant heterogeneity $A_{i}$.
+When $d_{X}>1$, as in production-function estimation (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)), one can further weaken the requirement to monotonicity in only one endogenous input, mirroring the identifying restriction in those classic approaches.
+Assumption 1 requires justification when $\eta_{it}$ is vector-valued and each coordinate of $X_{it}$ admits one distinct coordinate of $\eta_{it}$.
+In the production function estimation example, this is plausible if firms make hiring and investment decisions in different units (e.g., HR and finance), each responding primarily to its own market-specific signal—labor-market conditions for hiring and financial conditions for investment—satisfying the “single-coordinate” dependence.
+Moreover, positive shocks in each market naturally raise the corresponding choice (more hiring when labor conditions improve, more investment when financing conditions improve), implying that $g_{j}$ is strictly monotone in $\eta_{it,j}$ for each coordinate $j$ and thus satisfying componentwise monotonicity.
+Importantly, Assumption 1 places no restrictions on the dependence structure across coordinates of $\eta_{it}$, nor on the dependence between $A_{i}$ and $\eta_{it}$.
+Accordingly, without loss of generality, one may normalize $\eta_{it}$ so that $d_{\eta}=d_{X}$ and each coordinate of $g$ depends on the corresponding coordinate of $\eta_{it}$.
+Remark 1.
+Assumption 1 excludes settings where multiple components of $\eta_{it}$ enter a single coordinate of $X_{it}$, as in simultaneous supply–demand systems.
+Subsection 3.3 relaxes this restriction via timing assumptions, global univalence conditions, or single-index structure.
 Assumption 2 (Index Exclusion).
 Let $W_{i}\coloneqq W\left({\bf X}_{i},{\bf Z}_{i}\right)$, where $W:\mathbb{R}^{T\times\left(d_{X}+d_{Z}\right)}\mapsto\mathbb{R}^{d_{W}}$ is known.
-Suppose $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$ for any fixed $t$.
-Assumption 2 is an index exclusion restriction that requires the conditional distribution of $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)$ to depend only on $W_{i}$ which is a known function of ${\bf X}_{i}$ and ${\bf Z}_{i}$.
-It is similar to Assumption 2.1 of Altonji and Matzkin (2005) and Assumption A2 of Liu, Poirier, and Shiu (2024).
-A similar condition is also imposed by Wooldridge (2019).
-Liu, Poirier, and Shiu (2024) prove a non-identification result when the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted in a nonlinear panel data model and suggest that restrictions on $A_{i}$ are generally required to identify the APE and LAR.
-Given $W_{i}$, the distribution of $A_{i}$ no longer depends on $X_{it}$ and $Z_{it}$ for any fixed $t$.
-In other words, $W_{i}$ is sufficient for $A_{i}$.
-Assumption 2 is used to (i) construct the conditional control variable for $\eta_{it}$, which I discuss in detail in relation to Assumption 3; and (ii) to control for $A_{i}$ when using the LIE argument to exclude $X_{it}$ from the conditioning set of $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ in (3.4), which I present in the proof of Theorem 1.
-As pointed out by Liu, Poirier, and Shiu (2024), there are several ways to justify Assumption 2 from the literature.
-I follow their discussion to provide three justifications of Assumption 2.
-The first two methods are essentially parametric, resulting in a simpler expression for $W_{i}$, while the last one is nonparametric, leading to a more complex expression for $W_{i}$.
-This distinction highlights the trade-off between simplicity and generality.
-Moreover, I provide lower-level conditions for a nonparametric exchangeability condition and show that it is sufficient for Assumption 2 in Proposition 1.
-First, Assumption 2 can be justified by generalizing equation (2.4) of Mundlak (1978) to be $A_{i}=h\left(T^{-1}\sum_{t=1}^{T}X_{it},\nu_{i}\right)$, where $\nu_{i}\perp\left(X_{it},Z_{it}\right)$ for any $t$.
-Then, Assumption 2 is satisfied by taking $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$.
+Suppose $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$ for each $t$.
+Assumption 2 serves two roles.
+First, it makes $W_{i}$ a sufficient statistic for $A_{i}$, which is used to establish (3.9).
+Second, it is used to construct $V_{it}$ and establish its connection with $\eta_{it}$ by holding $A_{i}$ fixed.
+Under this assumption, the channel of endogeneity operating through $A_{i}$ is handled entirely via $W_{i}$, while the remaining “time-varying endogeneity through the random coefficients” is driven by the relationship between $\eta_{it}$ and $\varepsilon_{it}$ and is addressed via $V_{it}$.
+Assumption 2 is similar to Assumption 2.1 of Altonji and Matzkin (2005) and Assumption A2 of Liu, Poirier, and Shiu (2025).
+Considering the heterogeneity and endogeneity afforded in the TERC model, some restriction that isolates $A_{i}$ is typically needed.
+To motivate $W_{i}$ in empirical settings, consider steady-state input policies with convex adjustment costs in the production application.
+In standard investment and hiring models with convex (e.g., quadratic) adjustment costs (Cooper and Haltiwanger (2006)), optimality implies convergence to a firm-specific long-run target $X_{i}^{\star}(A_{i})$ that is increasing in ability $A_{i}$.
+Observed inputs fluctuate around this target due to transitory shocks and gradual adjustment.
+Averaging $X_{it}$ over time filters out these high-frequency deviations and provides a consistent estimator of $X_{i}^{\star}(A_{i})$.
+Because $X_{i}^{\star}(\cdot)$ is monotone, conditioning on $\overline{X}_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$ induces a control for $A_{i}$, satisfying Assumption 2.
+Next, I provide three sufficient conditions for Assumption 2.
+First, Assumption 2 can be justified by nonparametrically generalizing equation (2.4) of Mundlak (1978) to be $A_{i}=h\left(\overline{X}_{i},\nu_{i}\right)$, where $\nu_{i}\perp\left(X_{it},Z_{it}\right)$ for all $t$.
+Here, the functional form of $h$ and distribution of $\nu_{i}$ can be unknown.
+Then, Assumption 2 is satisfied by taking $W_{i}$ to be $\overline{X}_{i}$.
 This is true because conditioning on $W_{i}$, any $t$-specific $X_{it}$ and $Z_{it}$ does not affect the distribution of $A_{i}$ as the residual randomness in $A_{i}$ is only driven by $\nu_{i}$.
-Notice that there is no need to specify the functional form of $h\left(\cdot\right)$ nor the distribution of $\nu_{i}$.
 Second, insights from treatment assignment models in panel data (e.g., Arkhangelsky and Imbens (2022)) can be exploited to find candidate $W_{i}$.
-For example, if the random vector $\left(X_{it},Z_{it}\right)^{{}^{\prime}}\in\mathbb{R}^{2}$ conditional on $A_{i}$ is a two-dimensional normal random vector i.i.d.
-through time with known covariance [^7] The analysis can be extended to cover the case when the covariance matrix is unknown.
-The required change is to expand $W_{i}$ to include the averages of the second-order polynomials of $X_{it}$ and $Z_{it}$ through time.
-See Chapter 7 of Hogg, McKean, and Craig (2019) for more details.
-and ${\bf Z}_{i}\perp A_{i}$, then $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2.
-To see this, notice that the density of $\left.\left(X_{it},Z_{it}\right)\right|A_{i}$ depend on $A_{i}$ only through the mean of $X_{it}$.
-Hence, by the sufficient statistic for bi-variate normal random vectors with known covariance matrix (see Chapter 7 of Hogg, McKean, and Craig (2019)), $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ suffices for Assumption 2.
-Furthermore, sufficient conditions for the analysis above are (i) $\left.\eta_{i}\right|A_{i}$ is a normal random vector i.i.d.
-through time with known variance, (ii) ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$, and (iii) $g\left(\cdot\right)$ is linear in its arguments.
-Then, $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2.888Specifically, suppose $\left.\eta_{it}\right|A_{i}$ is i.i.d.
-normal across $t$ with mean $\mu\left(A_{i}\right)$ and known variance.
-Assume $X_{it}=Z_{it}^{{}^{\prime}}\gamma+A_{i}+\eta_{it}$ and ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$.
-Let $T=2$ for simplicity.
-Then, define $h_{t}=x_{t}-z_{t}^{{}^{\prime}}\gamma-a$ for $t=1\text{ and }2$ and by a change of variable argument, $f_{\left.\left(X_{i1},Z_{i1},X_{i2},Z_{i2}\right)\right|A_{i}}\left(\left.x_{1},z_{1},x_{2},z_{2}\right|a\right)$ $=f_{\left.\eta_{i1},\eta_{i2}\right|A}\left(\left.h_{1},h_{2}\right|a\right)f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right)$ $=\left[\Pi_{t=1}^{2}f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)\right]f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right).$ Using the formula of density of normal distribution for $f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)$ and applying the Fisher-Neyman factorization theorem yield the desired result.
-Third, the nonparametric exchangeability condition (Assumption 2.3) from Altonji and Matzkin (2005) can also be adapted to justify Assumption 2.
-A similar idea is also used by Wooldridge (2019).
-Notice that the original exchangeability condition of Altonji and Matzkin (2005) is incompatible with the time-varying endogeneity through the random coefficients since it concerns the density of $\eta_{it}$ given ${\bf X}_{i}$, and I present how to adapt it to the TERC model below.
-To grasp the idea, suppose (i) $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ [^9] To be precise, by $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ I mean $f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{i1},...,h_{iT}\right|a_{i}\right)=f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{it_{1}},...,h_{it_{T}}\right|a_{i}\right)$ for any $\left(h_{i1},...,h_{iT}\right)$ belonging to the conditional support of $\eta_{i}$ given $A_{i}=a_{i}$, where $\left(t_{1},...,t_{T}\right)$ is any permutation of $\left(1,...,T\right)$.
-The definition is similar for $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$.
-where $\widetilde{\eta}_{i}$ is any permutation of the vector of $\eta_{i}$ in time, (ii) $Z_{it}\perp\left(\eta_{it},A_{i}\right)$, and (iii) $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}$ is continuous in $\left({\bf X}_{i},{\bf Z}_{i}\right)$.
-Then, combined with Assumption 1, it is straightforward to prove that $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$ where $\widetilde{{\bf X}}_{i}$ and $\widetilde{{\bf Z}}_{i}$ are permutations of ${\bf X}_{i}$ and ${\bf Z}_{i}$ in time, respectively.
-Then, following Altonji and Matzkin (2005), I use the Weierstrass approximation theorem and the fundamental theorem of symmetric functions to obtain that the vector $W_{i}$ of elementary symmetric functions of the paired elements of $\left\{\left(X_{i1},Z_{i1}\right),...,\left(X_{iT},Z_{iT}\right)\right\}$ satisfies Assumption 2.
-I provide the details of the above analysis in the proof of Proposition 1.
-To see an example of $W_{i}$ under the nonparametric exchangeability condition, suppose $T=2$ and both $X_{it}$ and $Z_{it}$ are scalars, and define $D_{it}\coloneqq\left(X_{it},Z_{it}\right)^{{}^{\prime}}$.
-Then, $W_{i}=\left(\sum_{t=1}^{T}D_{it},\sum_{t=1}^{T}D_{it}D_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ satisfies Assumption 2.101010In the concrete example discussed above, the maximum order of the polynomials in the construction of $W_{i}$ depends on $T$.
-In general, let $\varphi_{1}\left(u\right)=\sum_{t=1}^{T}u_{t}$, $\varphi_{2}\left(u,v\right)=\sum_{t\neq s}^{T}u_{t}v_{s}$, …, $\varphi_{T}\left(u,v,...,w\right)=\sum_{t\neq s\neq...\neq k}^{T}u_{t}v_{s}...w_{k}$, where $u,v,...,w$ are generic $T\times 1$ vectors.
-Then, I substitute each column vector of ${\bf D}_{i}=\left({\bf X}_{i},{\bf Z}_{i}\right)$ for each of the arguments $u,v,...,w$ (repetitions included) to construct $W_{i}$.
-Such a $W_{i}$ satisfies Assumption 2.
-See Chapter II.3 of Weyl (1939) for details of the construction and its proof.
-Notice that parametric assumptions on the distribution of $A_{i}$ are not needed for this result.
-Once the fixed effect $A_{i}$ is controlled for by $W_{i}$, I still need to control for the time-varying $\eta_{it}$ that enters the $g\left(\cdot\right)$ function and is correlated with $\beta_{it}$ via its correlation with both $A_{i}$ and $\varepsilon_{it}$.
-When $A_{i}$ is not present in (2.2), I may follow the method of Imbens and Newey (2009) to use the conditional CDF $V_{it}^{\text{in09}}\coloneqq F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ as a control for $\eta_{it}$.
-This is true because Theorem 1 of Imbens and Newey (2009) proves that $V_{it}^{\text{in09}}=F_{\eta_{it}}\left(\eta_{it}\right)$, the RHS of which is a one-to-one mapping of $\eta_{it}$ when $d_{X}=d_{\eta}=1$ provided (i) $Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)$, (ii) $g\left(\cdot\right)$ is strictly monotonic in $\eta_{it}$, and (iii) $F_{\eta_{it}}\left(\cdot\right)$ is a strictly increasing function.
-Therefore, conditional on $V_{it}^{\text{in09}}$, $X_{it}$ and $\varepsilon_{it}$ are independent.
-The argument above can be generalized to the setting where $d_{X}>1$ by letting
-$V_{it}^{\text{in09}}$ $=\left(F_{\left.X_{it}^{\left(1\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it}\right),...,F_{\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}\right)\right)^{{}^{\prime}}$
-$=\left(F_{\eta_{it}^{\left(1\right)}}\left(\eta_{it}^{\left(1\right)}\right),...,F_{\eta_{it}^{\left(d_{X}\right)}}\left(\eta_{it}^{\left(d_{X}\right)}\right)\right)^{{}^{\prime}}.$
-to control for the whole vector of $\eta_{it}$ coordinate by coordinate.
-The challenge in applying their method to the TERC model arises from the existence of $A_{i}$ in (2.2), which is justified in many applications as agents usually know their $A_{i}$ (e.g., via repeated learning) when choosing the value of $X_{it}$.
-When $A_{i}$ is present in (2.2) and $d_{X}=1$, one may want to use $F_{\left.X_{it}\right|Z_{it},A_{i}}\left(\left.X_{it}\right|Z_{it},A_{i}\right)$ as the control variable for $\eta_{it}$, following a similar argument as in Imbens and Newey (2009).
-However, this approach is infeasible because $A_{i}$ is unknown.
-The idea of this paper is to replace $A_{i}$ with $W_{i}$ in the construction of $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ under a conditional independence assumption on $Z_{it}$ and a monotonicity condition on the conditional CDF $F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$ in $\eta_{it}$.
-I summarize them in the next assumption.
+For example, if $(X_{it},Z_{it})$ conditional on $A_{i}$ is a two-dimensional normal random vector i.i.d.
+through time with known covariance and ${\bf Z}_{i}\perp A_{i}$, then the density of $(X_{it},Z_{it})$ given $A_{i}$ depends on $A_{i}$ only through $\overline{X}_{i}$.
+Hence, by the sufficient statistic for bi-variate normal random vectors with known covariance matrix (Hogg, McKean, and Craig (2019, ch.
+7)), $W_{i}=(\overline{X}_{i},\overline{Z}_{i})$ suffices for Assumption 2.
+Third, the nonparametric exchangeability condition (e.g., Altonji and Matzkin (2005)) can be adapted to justify Assumption 2.
+I present its details in Proposition 1 of Appendix A.2.
+The key idea is time exchangeability: conditional on $A_{i}$, the joint behavior of $(X_{it},Z_{it})$ is unchanged if I relabel periods, so the time order contains no information about $A_{i}$.
+Hence only order-invariant summaries of the panel matter for learning about $A_{i}$, and symmetric functions of the observed pairs $(X_{it},Z_{it})$ can be used as $W_{i}$ to satisfy Assumption 2.
+Next, I control for the time-varying $\eta_{it}$. When $A_{i}$ is not present in (2.2) and under Assumption 1, Imbens and Newey (2009) propose
+$V_{it}^{\text{IN09}}\coloneqq F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)\overset{\text{by IN09}}{=}F_{\eta_{it}}\left(\eta_{it}\right)$ (3.14)
+as a control for $\eta_{it}$.
+The intuition is that, holding input prices $Z_{it}$ fixed, a firm choosing a higher capital $X_{it}$ must be experiencing a higher productivity shock $\eta_{it}$; otherwise the higher input choice would not be optimal.
+However, this argument breaks down in the TERC model because the unobserved, time-invariant component $A_{i}$ also shifts input choices: firm 1’s higher $X_{it}$ could reflect greater managerial ability rather than a larger transitory shock $\eta_{it}$.
+Therefore $A_{i}$ must also be conditioned on, but this makes $F_{\left.X_{it}\right|A_{i},Z_{it}}\left(\left.X_{it}\right|A_{i},Z_{it}\right)$ infeasible since $A_{i}$ is unobserved.
+I proceed in two steps to generalize the argument of Imbens and Newey (2009), adapting it to the TERC model.
+First, to make the control for $\eta_{it}$ feasible, I replace $A_{i}$ with the observable proxy $W_{i}$ and define
+$V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right).$ (3.15)
+Then, I show in (A.1) that, under the next Assumption, $V_{it}$ controls for $\eta_{it}$ conditional on $A_{i}$ and $W_{i}$.
+Second, I further extend their analysis to allow coordinates of $g$ to depend on multiple elements of $\eta_{it}$ in Subsection 3.3.
 Assumption 3 (Control for $\boldsymbol{\eta_{it}}$).
-Suppose the following conditions are satisfied:
+Suppose the following conditions hold:
   1. (a)
@@
   2. (b)
-    $F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right)$ is strictly increasing in $\eta_{it}^{\left(l\right)}$ for every realization of $\left(A_{i},W_{i}\right)$ on its support and each $l=1,...,d_{X}$.
-Assumption 3(a) concerns the vector of exogenous IVs $Z_{it}$.
-Since $W_{i}$ can be viewed as summarizing all the time-invariant information about $A_{i}$ in the data, this assumption, loosely speaking, requires $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|A_{i}$, which is already implied by the standard exogeneity assumption of $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$.
-For instance, when $d_{X}=1$ and suppose one follows Mundlak (1978) to assume $A_{i}=\left(\sum_{t=1}^{T}X_{it}\right)\alpha+\nu_{i}=W_{i}\alpha+\nu_{i}$, where $\nu_{i}$ is independent of everything else and $\alpha\neq 0$.
-Then, if $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it},\nu_{i}\right)$, I have $\left.Z_{it}\perp\left(\eta_{it},\varepsilon_{it}\right)\right|\left(A_{i},W_{i}\right)$ by the law of conditional independence because $W_{i}$ is a function of $A_{i}$ and $\nu_{i}$.
-[^11] I provide another parametric example where $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$ implies $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|\left(A_{i},W_{i}\right)$.
-For illustration purpose, suppress subscript $i$ and suppose $T=2$, $d_{X}=d_{Z}=1$, and $W=\left(X_{1}+X_{2},Z_{1}+Z_{2}\right)^{{}^{\prime}}$ satisfies Assumption 2 (Liu, Poirier, and Shiu (2024)).
-Furthermore, suppose $X_{t}=Z_{t}\gamma+A+\eta_{t}$ for $t=1,2$, and $\left.\left(Z_{1},Z_{2},\varepsilon_{1},\varepsilon_{2},\eta_{1},\eta_{2}\right)\right|A\sim N\left(\mu,\Sigma\right)$ where $\Sigma=\left[\begin{array}[]{cc}\Sigma_{1}&0_{2\times 4}\\ 0_{4\times 2}&\Sigma_{2}\end{array}\right]$ and $\Sigma_{1}$ and $\Sigma_{2}$ are arbitrary positive definite $2\times 2$ and $4\times 4$ matrices, respectively.
-Then, formula for the conditional density function of normally distributed random vectors shows that $\left.\left(Z_{t},\varepsilon_{t},\eta_{t}\right)\right|\left(A,W\right)$ is also normally distributed with diagonal covariance matrix for $t=1,2$, satisfying Assumption 3(a).
-Assumption 3(a) is used to ensure that $Z_{it}$ can be excluded from the conditioning set of $F_{\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}\right)$ for each $l=1,...,d_{X}$.
-Assumption 3(b) is similar to condition (ii) of Theorem 1 in Imbens and Newey (2009).
-It is mild since it concerns a conditional CDF, which is typically strictly increasing for continuous random variables.
-Assumption 3(b) guarantees that fixing $\left(A_{i},V_{it},W_{i}\right)$ uniquely determines the value of $\eta_{it}$.
-Assumption 3 allows me to construct a feasible control variable $V_{it}$ for $\eta_{it}$ given $A_{i}$ and $W_{i}$.
+    For every $(A_{i},W_{i})$ and each $l=1,...,d_{X}$, $F_{\left.\eta_{it,l}\right|A_{i},W_{i}}(\left.\eta_{it,l}\right|A_{i},W_{i})$ is strictly increasing in $\eta_{it,l}$.
+Assumption 3 enables construction of a feasible control variable $V_{it}$ for $\eta_{it}$ given $A_{i}$ and $W_{i}$.
 To present it clearly, suppose $d_{X}=1$ and recall that $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$.
@@
 $=F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right),$ (3.16)
-where the first equality holds by Assumption 2 which allows me to add $A_{i}$ to the conditioning set, the second equality holds by Assumption 1 and a change of variable argument, and the last equality holds by Assumption 3(a).
+where the first equality holds by Assumption 2, the second equality holds by Assumption 1 and a change of variable argument, and the last equality holds by Assumption 3(a).
 Thus, $V_{it}$ uniquely determines $\eta_{it}$ given $A_{i}$ and $W_{i}$ by Assumption 3(b).
-Notice that $V_{it}$ is a one-to-one mapping of $\eta_{it}$ only when $A_{i}$ is also conditioned upon.
-In the proof of Theorem 1, I use the LIE on $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ where I first condition on $A_{i}$ to control for $\eta_{it}$ via $V_{it}$ in the inner expectation step.
-Then, I deal with the unknown $A_{i}$ with $W_{i}$ in the outer expectation step to prove (3.4).
-Finally, I use the residual variation in $X_{it}$, guaranteed by the next assumption, to identify the APE and LAR via (3.6) or (3.7).
-The last assumption concerns the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$.
+Note that (3.16) relies on Assumption 1; when multiple elements of $\eta_{it}$ appear in one coordinate of $X_{it}$, the one-to-one relationship does not hold in general, and additional structure or assumption is needed to recover (3.16).
+I discuss them in Subsection 3.3.
+Assumption 3(a) is similar to condition (i) of Theorem 1 in Imbens and Newey (2009), except that I further condition on $(A_{i},W_{i})$.
+Since $W_{i}$ can be viewed as summarizing all the time-invariant information about $A_{i}$ in the data, Assumption 3(a), loosely speaking, requires $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|A_{i}$, which is already implied by the standard exogeneity condition $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$.
+Note that $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$ is stronger than Assumption 3(a) as the former rules out the possibility that $Z_{it}$ and $A_{i}$ are correlated.
+[^2] For instance, Bartik instruments use national industry shocks, weighted by pre-existing local industry shares.
+The intuition is the national shocks are plausibly “as-good-as-random” conditional on the fixed effects and controls, so the conditional exogeneity is plausible.
+But the baseline shares can reflect deep local traits—so unconditional exogeneity is harder to defend.
+Nonetheless, the unconditional exogeneity condition remains a useful benchmark since it is widely used in applied work.
+Assumption 3(a) can be justified in many applications with the usual choice of IVs from the literature.
+For example, in labor-supply applications, $Z_{it}$ can be county minimum-wage changes or EITC expansions—policy-driven shocks that, conditional on time-invariant ability $A_{i}$ and its sufficient statistic $W_{i}$, vary independently of expected sectoral labor-demand shocks $\eta_{it}$ and labor-supply-elasticity shocks $\varepsilon_{it}$.
+In gasoline-demand estimation, $Z_{it}$ can be the head of household’s earned income or gas-tax changes, assumed (conditional on a household fixed effect $A_{i}$ and $W_{i}$) to be independent of anticipated near-term income changes $\eta_{it}$ (e.g., an upcoming bonus) and transitory wealth shocks $\varepsilon_{it}$.
+For the empirical application of Section 5, I construct the instruments toward BLP-style (Berry, Levinsohn, and Pakes (1995)) cost shifters: leave-one-out, competitor-weighted input-price (wages and interest rates) averages at the industry–city–year level.
+These instruments are relevant because firms compete locally for labor and capital, so competitors’ input prices shift firm $i$’s cost environment and input choices.
+They are plausibly exogenous to firm $i$’s idiosyncratic productivity shocks conditional on the fixed effect and its control variable because firm $i$’s own input prices are excluded from the construction, so identification comes from competitors’ input-price variation rather than firm $i$’s unobserved shocks.
+Robustness checks in Appendix B provide evidence that these IVs are plausibly valid and satisfy Assumption 3(a).
+Assumption 3(b) is analogous to condition (ii) of Theorem 1 in Imbens and Newey (2009).
+It is mild because it concerns a conditional CDF, which is typically strictly increasing for continuous random variables.
+The last assumption concerns the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$, which is used to identify the APE and LAR by (3.11).
 Assumption 4 (Residual Variation in $\boldsymbol{X_{it}}$).
-There are at least $d_{X}$ linearly independent points in the support of $X_{it}$ a.s. given $V_{it}$ and $W_{i}$.
-Assumption 4 guarantees that $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$ in (3.6) is invertible [^12] This is true because for any random vector $X\in\mathbb{R}^{d_{X}}$, $\mathbb{E}\left[XX^{{}^{\prime}}\right]$ is singular if and only if there exists a non-random vector $a\in\mathbb{R}^{d_{X}}\backslash\left\{0\right\}$ such that $a^{{}^{\prime}}X=0$ almost surely., thus the APE and LAR can be identified by (3.6) and the LIE.
-As discussed by Imbens and Newey (2009), requiring residual variation in $X_{it}$ conditioning on $V_{it}$ is generally not restrictive.
-Essentially, it requires $Z_{it}$ to move $X_{it}$ sufficiently well such that the set $\left\{X_{it}:V_{it}=v\text{ for }v\in\text{Supp}\left(V_{it}\right)\right\}$ is not a singleton.
-Therefore, it is similar to the relevance condition in the classical IV literature.
-[^13] I thank an anonymous referee for pointing this out.
-The restriction on the support of $X_{it}$ mainly comes from conditioning on $W_{i}$.
-When parametric justifications for Assumption 2 are used (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) so that $W_{i}$ only includes averages of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is not restrictive.
-However, when nonparametric justifications for Assumption 2 (e.g., the exchangeability condition of Altonji and Matzkin (2005)) are involved, Assumption 4 can be restrictive and generally requires that $T$ is larger than $d_{X}$.
-[^14] One way to justify Assumption 4 when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the $T^{\text{th}}$ order (Altonji and Matzkin (2005)) is to use the symmetry through time in the solution to $\left\{{\bf X}_{i}:V_{it}=v,W_{i}=w\right\}$.
-To see a concrete example, suppose $T=3$, $X_{it}\in\mathbb{R}^{2}$, $Z_{it}\in\mathbb{R}^{2}$, and $D_{it}\coloneqq\left(X_{it}^{{}^{\prime}},Z_{it}^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathbb{R}^{4}$.
-If $D_{i}=\left(D_{i1},D_{i2},D_{i3}\right)=\left(a,b,c\right)$ satisfies $V_{it}=v$ for $t=1,2$ and $W_{i}=w$ where all $a,b,c\in\mathbb{R}^{4}$ and the first two coordinates of $a$ and $b$ are linearly independent, by the symmetry of $W_{i}$ in $\left(X_{it},Z_{it}\right)$ through $t$, $\left(b,a,c\right)$ must also satisfy $V_{it}=v$ and $W_{i}=w$ for $t=1,2$.
-Therefore, given $V_{i1}=v$ and $W_{i}=w$ there are two linearly independent vectors of $\left(a^{\left(1\right)},a^{\left(2\right)}\right)$ and $\left(b^{\left(1\right)},b^{\left(2\right)}\right)$ that $X_{i1}$ can take, satisfying Assumption 4.
-It is worth pointing out that I only need to use the endogenous coordinates of $X_{it}$ to construct $V_{it}$ and $W_{i}$.
-Thus, when some coordinates of $X_{it}$ are exogenous in the sense that they are independent of $\left(A_{i},\eta_{it},\varepsilon_{it}\right)$, the unconditional variation in the exogenous coordinates of $X_{it}$ can be exploited to satisfy Assumption 4.
-See footnote 16 below for more details on this point.
-Next, I present another assumption on the residual variation of $X_{it}$ given $V_{it}$ and $W_{i}$ to facilitate the derivative-based argument (3.7).
+There are at least $d_{X}$ linearly independent points in the support of $X_{it}$ conditional on $(V_{it},W_{i})$ almost surely.
+Assumption 4 ensures that $\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]$ in (3.11) is invertible.
+Thus, the APE and LAR can be identified by (3.11) and the law of iterated expectations.
+Assumption 4 is best interpreted as a conditional support (non-degeneracy) requirement: for each $t$, $X_{it}$ must retain sufficient variation conditional on $(V_{it},W_{i})$ to identify the target objects.
+Requiring residual variation in $X_{it}$ conditional on $V_{it}$ is generally not restrictive (Imbens and Newey (2009)).
+Essentially, it requires $Z_{it}$ to induce enough variation in $X_{it}$ so that, for each $v\in\text{Supp}(V_{it})$, the set $\left\{X_{it}:V_{it}=v\right\}$ is not a singleton.
+Therefore, it corresponds to the classical IV relevance condition of $\partial g(z,a,\eta)/\partial z\neq 0$.
+The restriction on the support of $X_{it}$ in Assumption 4 mainly comes from conditioning on $W_{i}$, which also concerns Assumption 2.
+There is an inherent trade-off between Assumptions 2 and 4.
+To see this, suppose $W_{i}=(\mathbf{X}_{i},\mathbf{Z}_{i})$ which satisfies Assumption 2 trivially.
+However, because $X_{it}$ is included in $W_{i}$, it makes the support of $X_{it}$ given $W_{i}$ a singleton.
+Hence, Assumption 4 does not hold unless $X_{it}$ is a scalar and the point in the support of $X_{it}$ given $W_{i}$ is nonzero.
+More generally, including more elements in $W_{i}$ tends to make Assumption 2 easier to satisfy while making Assumption 4 harder to maintain.
+Because of this trade-off, I describe next how the conditions used to justify Assumption 2 affect the plausibility of Assumption 4, and give corresponding sufficient conditions.
+When justifications for Assumption 2 are used so that $W_{i}$ only includes averages of $X_{it}$ and/or $Z_{it}$ through time (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2025)), Assumption 4 is not restrictive.
+For instance, with $d_{X}=2$, $T=2$, and $W_{i}=\overline{X}_{i}$, Assumption 4 holds whenever $X_{i1}$ and $X_{i2}$ are not perfectly linearly dependent, so that $\mathrm{Supp}(X_{it}\mid\overline{X}_{i})$ remains non-degenerate.
+By (1.2) and the relevance condition above, it in turn requires that the serial dependence in $Z_{it}$ cannot be so high that the value of $Z_{it}$ (hence $X_{it}$) for any $t$ is uniquely determined once $W_{i}$ is fixed.
+By contrast, when a high-dimensional $W_{i}$ is used to justify Assumption 2 (e.g., via the exchangeability condition of Altonji and Matzkin (2005)), Assumption 4 becomes more restrictive.
+Intuitively, a high-dimensional $W_{i}$ imposes many constraints on the path $(X_{i1},\ldots,X_{iT})$, so one typically needs longer panel to leave enough feasible configurations and satisfy Assumption 4.
+In this case, one possibility is to explore the symmetry through time in the solution to $\left\{{\bf X}_{i}:V_{it}=v,W_{i}=w\right\}$.
+In particular, when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the $T^{\text{th}}$ order, Assumption 2 is satisfied under a nonparametric exchangeability condition on $f_{\left.\eta_{i}\right|A_{i}}$ by adapting the proof of Altonji and Matzkin (2005) in Proposition 1.
+Then, since $W_{i}$ is symmetric in $\left(X_{it},Z_{it}\right)$ through time, one may permute the order of $\left(X_{it},Z_{it}\right)$ in time without changing the value of $W_{i}$.
+This creates enough linearly independent points in the support of $X_{it}$ given $W_{i}$ when $T\geq d_{X}$.
+When such permutation also does not change the value of $V_{it}$ which is usually satisfied under the relevance condition, Assumption 4 is satisfied.
+I provide an example in Appendix A.3.
+Assumption 4 pertains to the inversion-based argument in (3.11).
+To facilitate the derivative-based argument in (3.12), I introduce the next assumption, which requires more residual variation of $X_{it}$ conditional on $V_{it}$ and $W_{i}$ than Assumption 4.
 Assumption 4 (More Variation in $\boldsymbol{X_{it}}$).
-The support of $X_{it}$ given $V_{it}$ and $W_{i}$ contains some ball of positive radius a.s. with respect to $\left(V_{it},W_{i}\right)$.
-Assumption 4 is similar to Assumption 2.2 of Altonji and Matzkin (2005) and is stronger than Assumption 4.
-When parametric justifications (e.g., Mundlak (1978), Arkhangelsky and Imbens (2022), Liu, Poirier, and Shiu (2024)) are used for Assumption 2 such that $W_{i}$ contains only mean of $X_{it}$ and/or $Z_{it}$ through time, Assumption 4 is generally not restrictive.
-However, when nonparametric justifications (e.g., Altonji and Matzkin (2005)) are used for Assumption 2, Assumption 4 typically requires ex-ante information about $W_{i}$ (e.g., the existence of exogenous regressors in (2.1)).
-This is true because in this case $W_{i}$ includes averages of each endogenous coordinate of $X_{it}$, $Z_{it}$, and their interaction terms up to the $T^{\text{th}}$ order through time, conditioning on which there may not be enough residual variation in $X_{it}$.
-To address this issue, when nonparametric justifications for Assumption 2 are used, it is advisable to follow the recommendations by Altonji and Matzkin (2005) to ensure continuous residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$.
-[^15] Specifically, Altonji and Matzkin (2005) suggest to (i) leverage unconditional variation in the exogenous regressors that do not enter $W_{i}$, which I provide details in footnote 16; (ii) restrict the conditional distribution of $A_{i}$ given $W_{i}$ to depend on not all but only a subset of coordinates of $W_{i}$, which is testable by comparing the fit of $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ with $W_{i}$ replaced by a proper subset of it; (iii) assume $f_{\left.A_{i}\right|W_{i}}$ to depend on $W_{i}$ only through a linear combination of the elements of $W_{i}$, i.e., $f_{\left.A_{i}\right|W_{i}}=f_{\left.A_{i}\right|\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}}$ where $W_{i}=\left(W_{i}^{\left(1\right)},...,W_{i}^{\left(d_{W}\right)}\right)^{{}^{\prime}}$ and the $c_{k}$’s are unknown parameters, which implies that $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}\right]$ and the estimation of $c_{k}$’s can proceed by using the method of Ichimura and Lee (1991); (iv) impose a priori restrictions on model (2.1)–(2.2) such that $W_{i}$ enters $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ in a parametric way (e.g., $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\sum_{k=1}^{d_{W}}c_{k}W_{i}^{\left(k\right)}+\sum_{l=1}^{d_{V}}d_{l}V_{it}^{\left(l\right)}$ when $d_{X}=1$); and, (v) directly impose restrictions on $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$.
-[^16] To see a concrete example how the unconditional variations in the exogenous regressors help to satisfy Assumption 4, suppose $X_{it}^{\left(1\right)}$ is endogenous while $X_{it}^{\left(-1\right)}$, all other coordinates of $X_{it}$ except $X_{it}^{\left(1\right)}$, are exogenous (independent of $X_{it}^{\left(1\right)}$ and $\beta_{it}$).
-Suppose the conditional support of $X_{it}^{\left(1\right)}$ given $\left(V_{it},W_{i}\right)$ only contains a non-zero singleton, while the unconditional support of $X_{it}^{\left(-1\right)}$ contains some small ball of positive radius.
-Denote $b_{1}\left(V_{it},W_{i}\right)=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$.
-Then, I can identify the first-order moments of $\beta_{it}$ by taking partial derivative with respect to $X_{it}^{\left(-1\right)}$ on both sides of $\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]=X_{it}^{\left(1\right)}b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)+X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$ to identify $b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)$.
-Finally, identification of $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)$ can be obtained as $b_{1}^{\left(1\right)}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.Y_{it}\right|X_{it}^{\left(1\right)},X_{it}^{\left(-1\right)},V_{it},W_{i}\right]-X_{it}^{\left(-1\right)^{\prime}}b_{1}^{\left(-1\right)}\left(V_{it},W_{i}\right)\right)/X_{it}^{\left(1\right)}.$
-Assumption 4 allows the straightforward perturbation-based method (3.7) to identify the APE and LAR without requiring the computation of the inverse of $\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]$.
+The support of $X_{it}$ conditional on $(V_{it},W_{i})$ contains an open ball of positive radius almost surely.
+Assumption 4 allows one to perturb $X_{it}$ in $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$ when $(V_{it},W_{i})$ is fixed so that conditional moments of $\beta_{it}$ can be recovered by (3.12).
+It is similar to Assumption 2.2 of Altonji and Matzkin (2005).
+The same trade-off between Assumptions 2 and 4 also applies to Assumption 4.
+When Assumption 2 is justified such that $W_{i}$ is low-dimensional (e.g., $\overline{X}_{i}$ as in Mundlak (1978)), Assumption 4 is generally not restrictive.
+In essence, $Z_{it}$ must vary enough to generate sufficient residual variation in $X_{it}$ conditional on $(V_{it},W_{i})$.
+However, when Assumption 2 is justified using nonparametric arguments (e.g., the exchangeability condition of Altonji and Matzkin (2005)), $d_{W}$ is large and Assumption 4 can be restrictive.
+To address this issue, I present two sets of solutions in Appendix A.3.
+First, the unconditional variation of exogenous regressors outside of $W_{i}$ can be leveraged.
+Second, following the suggestions of Altonji and Matzkin (2005), one may restrict how $W_{i}$ enters $f_{\left.A_{i}\right|W_{i}}$ so that the relevant conditional expectations only concern a subset of $W_{i}$ or a linear index of its components.
+Assumption 4 allows the straightforward perturbation-based method (3.12) to identify the APE and LAR without requiring the computation of the inverse of $\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]$.
 When residual variation is not a concern—such as when $W_{i}$ contains only the mean of $X_{it}$ through time or when exogenous regressors are included in (2.1)—Assumption 4 is preferred for simpler analysis.
-Accordingly, I impose it for the estimation and inference results in the next section.
-I now state my main identification theorem.
+I compare these two approaches in Remark 2 below.
+The next theorem summarizes the main identification result of this paper.
 Theorem 1 (Identification).
-If Assumptions 13 and either Assumption 4 or 4 are satisfied, then $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ are identified.
+If Assumptions 13 and either Assumption 4 or 4 are satisfied, then the APE $\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$ and the LAR function $b_{t}(x)=\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$ are both identified.
 3.3 Extensions
 Vector-Valued $\boldsymbol{\eta_{it}}$
-In my baseline model, I assume that each coordinate of $X_{it}$ depends on only one coordinate of $\eta_{it}$ (as well as $Z_{it}$ and $A_{i}$).
-Here, I discuss two ways to include vector-valued $\eta_{it}$ into coordinates of $X_{it}$ under additional assumptions motivated by empirical research.
-To clearly present the idea, I consider the case where $d_{X}=d_{\eta}=2$.
-First, timing assumptions on the choice of regressors can be leveraged to allow vector-valued $\eta_{it}$ to enter certain coordinates of $X_{it}$.
-For instance, in production function applications, choices of later inputs such as capital typically do not depend on random shocks to earlier chosen inputs such as labor and/or material after conditioning on the level of the earlier chose inputs.
-[^17] Appendix 1 of the 2006 working paper version of Ackerberg, Caves, and Frazer (2015) discusses a similar idea.
-I thank an anonymous referee for this suggestion.
-If this timing assumption is imposed, I modify model (2.1)–(2.2) to be
-$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
-$X_{it}^{\left(1\right)}$ $=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right),\text{ and }X_{it}^{\left(2\right)}=g^{\left(2\right)}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{\left(1\right)}},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right).$
-Note that now $X_{it}^{\left(2\right)}$ depends on the whole vector of $\eta_{it}$ if $X_{it}^{\left(1\right)}$ is substituted into $g^{\left(2\right)}\left(\cdot\right)$.
-I propose to use $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}}\left(\left.X_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(2\right)}$ first.
-Then, I use $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it}^{\left(1\right)}$.
-The rest of the identification analysis follows as in this paper.
-[^18] Specifically, for this method to work, I maintain Assumptions 2, 3(a), and 4 (or 4) and change Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)}\right)$ and $g^{\left(2\right)}\left(X_{it}^{\left(1\right)},Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
-I also modify Assumption 3(b) to be that $F_{\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|X_{it}^{\left(1\right)},A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
-Second, the interaction between different coordinates of $X_{it}$ also provides useful information about certain coordinates of $\eta_{it}$.
-One way to exploit such information is to assume that while both coordinates of $\eta_{it}$ enter functions that determine each coordinate of $X_{it}$, the ratio of $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}$ may depend on only one coordinate of $\eta_{it}$, e.g., $\eta_{it}^{\left(2\right)}$ (and $Z_{it}$ and $A_{i}$).
-For example, in production function applications, the interpretation of $\eta_{it}$ could be that $\eta_{it}^{\left(1\right)}$ is the Hicks neutral productivity and $\eta_{it}^{\left(2\right)}$ is the labor-augmenting technology.
-Although both material and labor choices depend on both $\eta_{it}^{\left(1\right)}\text{ and }\eta_{it}^{\left(2\right)}$, it is plausible that material per worker depends only on the labor-augmenting technology $\eta_{it}^{\left(2\right)}$.
+I present four methods for incorporating vector-valued $\eta_{it}$ into $X_{it}$ under additional assumptions. To illustrate the idea clearly, I assume $d_{X}=d_{\eta}=2$.
+First, timing assumptions on the choice of regressors can be exploited.
+For instance, in production function estimation, choices of later inputs such as capital typically do not depend on random shocks to earlier chosen inputs such as labor after conditioning on the level of the earlier chosen inputs; see Ackerberg, Caves, and Frazer (2015, (Appendix 1 of 2006 Working Paper)) for a similar idea.
+Under this timing assumption, I modify (2.1)–(2.2) to be:
+$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$
+$X_{it,1}$ $=g_{1}\left(Z_{it},A_{i},\eta_{it,1}\right),\text{ and }X_{it,2}=g_{2}\left({\color[rgb]{1,0,0}X_{it,1}},Z_{it},A_{i},\eta_{it,2}\right).$ (3.17)
+Note that now $X_{it,2}$ depends on the whole vector of $\eta_{it}$ if $X_{it,1}=g_{1}\left(Z_{it},A_{i},\eta_{it,1}\right)$ is substituted into $g_{2}$.
+I first use $V_{it,2}\coloneqq F_{\left.X_{it,2}\right|X_{it,1},Z_{it},W_{i}}\left(\left.X_{it,2}\right|X_{it,1},Z_{it},W_{i}\right)$ to control for $\eta_{it,2}$ first.
+Then, I use $F_{\left.X_{it,1}\right|Z_{it},W_{i}}\left(\left.X_{it,1}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it,1}$.
+Second, interaction across different coordinates of $X_{it}$ can provide useful information about $\eta_{it}$.
+One possibility is to assume that, although the full vector $\eta_{it}$ enters each coordinate of $X_{it}$, the ratio $X_{it,1}/X_{it,2}$ depends only on a single component of $\eta_{it}$.
+For example, in production function estimation, $\eta_{it,1}$ represents the Hicks neutral productivity and $\eta_{it,2}$ denotes the labor-augmenting technology.
+Although both capital and labor choices depend on vector $\eta_{it}$, it is plausible that capital per worker depends only on the labor-augmenting technology $\eta_{it,2}$.
 See Proposition 2.1 of Demirer (2025) for more discussions about this assumption.
-Given this assumption, I can proceed to control for $\eta_{it}^{\left(2\right)}$ first by $V_{it}^{\left(2\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$, and then for $\eta_{it}^{\left(1\right)}$ by $V_{it}^{\left(1\right)}\coloneqq F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$.
-The rest of the identification analysis follows as in this paper.
-[^19] Specifically, suppose $X_{it}^{\left(1\right)}=g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}=g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$.
-I maintain Assumptions 2, 3(a), and 4.
-I modify Assumption 1 to be that $g^{\left(1\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(1\right)},\eta_{it}^{\left(2\right)}\right)$ and $g^{\left(2\right)}\left(Z_{it},A_{i},\eta_{it}^{\left(2\right)}\right)$ are strictly monotonic in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
-I further change Assumption 3(b) to be $F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right)$ and $F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right)$ are strictly increasing in $\eta_{it}^{\left(1\right)}$ and $\eta_{it}^{\left(2\right)}$, respectively.
-Then, I have $V_{it}^{\left(2\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}/X_{it}^{\left(2\right)}\right|Z_{it},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(2\right)}\right|A_{i},W_{i}\right),$ and $V_{it}^{\left(1\right)}\coloneqq$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},W_{i}\right)$ $=$ $F_{\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it},V_{it}^{\left(2\right)},A_{i},W_{i}\right)=F_{\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(1\right)}\right|\eta_{it}^{\left(2\right)},A_{i},W_{i}\right),$ where the second equality holds because by Assumption (2), $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$ and that $V_{it}^{\left(2\right)}$ is a function of $\left(X_{it},Z_{it},W_{i}\right)$ and the last equality holds because $\left(V_{it}^{\left(2\right)},A_{i},W_{i}\right)$ uniquely determines $\eta_{it}^{\left(2\right)}$.
-Therefore, fixing $\left(X_{it},A_{i},V_{it},W_{i}\right)$ is equivalent to fixing $\left(X_{it},A_{i},\eta_{it},W_{i}\right)$.
-The rest of the argument follows as in the proof of Theorem 1.
+Given this assumption, I can proceed to control for $\eta_{it,2}$ first by
+$V_{it,2}\coloneqq F_{\left.X_{it,1}/X_{it,2}\right|Z_{it},W_{i}}\left(\left.X_{it,1}/X_{it,2}\right|Z_{it},W_{i}\right),$ (3.18)
+and then for $\eta_{it,1}$ by
+$V_{it,1}\coloneqq F_{\left.X_{it,1}\right|Z_{it},V_{it,2},W_{i}}\left(\left.X_{it,1}\right|Z_{it},V_{it,2},W_{i}\right).$ (3.19)
+Third, general “global univalence” results—such as Gale and Nikaido (1965) or Palais (1959), Hadamard (1906a, b)—suffice to establish the invertibility required here.
+In production function estimation, the inverse isotonicity approach in Berry, Gandhi, and Haile (2013) appears promising to obtain global injectivity of $g$ in $\eta_{it}$.
+In particular, suppose each input choice increases in its own component of $\eta_{it}$, while cross-partials are weakly negative due to substitutability.
+This happens when, for example, inputs are substitutes in the firm’s optimal input demand; a favorable shock to one input in the form of higher efficiency or lower effective cost leads the firm to substitute toward that input and away from others.
+Then, the Jacobian matrix of $g$ with respect to $\eta_{it}$, denoted by $J_{g}(\eta_{it})$, has positive diagonal entries and nonpositive off-diagonal entries, i.e., the $Z$-matrix sign pattern required for an $M$-matrix.
+If, in addition, the own effects are strong enough that the matrix is strictly diagonally dominant, then $J_{g}(\eta_{it})$ is an $M$-matrix.
+This, in turn, guarantees a unique monotone inverse mapping from observables back to the vector $\eta_{it}$.
+Finally, if $\eta_{it}$ enters the structural function $g$ through a scalar index $\eta_{it}^{\prime}\tau$ and $g$ is monotone in that index (Ichimura and Lee (1991)), then the effective unobservable is scalar, satisfying Assumption 1.
+This aligns with common practice in asset pricing, where a scalar market factor emerges as an index of aggregate shocks and affects stock returns monotonically.
+A similar structure arises in production function estimation, where scalar productivity summarizes richer underlying states and affects input choices strictly increasingly.
 Higher-Order Moments of $\boldsymbol{\beta_{it}}$
-For clarity of exposition, suppose the regressors on the RHS of (2.1) include constant 1 and a scalar-valued $X_{it}$.
-It is straightforward to generalize the idea to cover $d_{X}$-dimensional $X_{it}$ for $d_{X}\geq 2$.
-With a slight abuse of notation, let $\left(\beta_{it},\omega_{it}\right)\in\mathbb{R}^{2}$ where $\beta_{it}$ is the random coefficient for $X_{it}$ and $\omega_{it}$ is for constant 1.
-Since the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$ is driven only by exogenous $Z_{it}$, (3.4) holds for any measurable function of $\beta_{it}$.
-Thus, I have
-$\mathbb{E}\left[\left.\beta_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right],$ (3.10)
-and similarly for $\mathbb{E}\left[\left.\omega_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ and $\mathbb{E}\left[\left.\omega_{it}\beta_{it}\right|X_{it},V_{it},W_{i}\right]$.
-Taking the conditional expectation of both sides of (2.1) squared given $\left(X_{it},V_{it},W_{i}\right)$ yields
-$\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]+2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]+\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right].$ (3.11)
-Then, I have
-$\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]$ $=\left(\partial^{2}\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}^{2}\right)/2,$ (3.12)
-$\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]$ $=\left(\partial\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}-2X_{it}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]\right)/2,\text{ and}$
-$\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]-X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]-2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right],$
-which identify the conditional (given $X_{it}$) and unconditional second-order moments of $\beta_{it}$ and $\omega_{it}$ by the LIE.
-The above technique can also be used to identify the intertemporal correlations of the random coefficients.
-For example, $\mathbb{E}\left[\left.\beta_{it}\beta_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ can be identified from $\mathbb{E}\left[\left.Y_{it}Y_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ for any $t\neq s$ following an argument similar to (3.11)–(3.12).
+I identify the $P^{\text{th}}$-order moments of $\beta_{it}$.
+For clarity of illustration, suppose the regressor is $X_{it}=(K_{it},L_{it})^{\prime}$ and $\beta_{it}=(\beta_{it,K},\beta_{it,L})^{\prime}$.
+The extension to $d_{X}>2$ is straightforward.
+Since the residual variation in $X_{it}$ given $(V_{it},W_{i})$ is driven only by exogenous $Z_{it}$, (3.9) holds for any measurable function $h$ of $\beta_{it}$:
+$\mathbb{E}\left[h(\beta_{it})\mid X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[h(\beta_{it})\mid V_{it},W_{i}\right].$ (3.20)
+In particular, let $h=(h_{1},\ldots,h_{P})^{\prime}$ and $h_{q}(\beta_{it})=\beta_{it,K}^{q}\beta_{it,L}^{P-q}$ for $q=0,\ldots,P$.
+By (1.1), the conditional $P^{\text{th}}$-order moment of $Y_{it}$ given $(X_{it},V_{it},W_{i})$ is
+$\mathbb{E}\left[Y_{it}^{P}\mid X_{it},V_{it},W_{i}\right]=\sum_{q=0}^{P}\binom{P}{q}K_{it}^{q}L_{it}^{P-q}\mathbb{E}\left[\beta_{it,K}^{q}\beta_{it,L}^{P-q}\mid V_{it},W_{i}\right].$ (3.21)
+Under $P$-times differentiability of $\mathbb{E}\left[Y_{it}^{P}\mid X_{it},V_{it},W_{i}\right]$ in $X_{it}$, the coefficients on $K_{it}^{P}$ and $L_{it}^{P}$ are identified via differentiation, yielding
+$\mathbb{E}\left[\beta_{it,K}^{q}\beta_{it,L}^{P-q}\mid V_{it},W_{i}\right]$ $=\dfrac{1}{P!}\frac{\partial^{P}}{\partial k^{q}\partial l^{P-q}}\mathbb{E}\left[Y_{it}^{P}\mid K_{it}=k,L_{it}=l,V_{it},W_{i}\right],\text{ for }q=0,\ldots,P,$ (3.22)
+and hence the $P^{\text{th}}$-order moments of $\beta_{it}$ follow by the law of iterated expectations
+$\mathbb{E}[\beta_{it,K}^{q}\beta_{it,L}^{P-q}]$ $=\mathbb{E}\left[\dfrac{1}{P!}\frac{\partial^{P}}{\partial k^{q}\partial l^{P-q}}\mathbb{E}\left[Y_{it}^{P}\mid K_{it}=k,L_{it}=l,V_{it},W_{i}\right]\right],\text{ for }q=0,\ldots,P.$ (3.23)
+Repeating this argument for all $P\in\mathbb{N}$ identifies the entire moment sequence of $\beta_{it}$.
+Under standard moment-determinacy conditions (e.g., Stoyanov (2000)), this sequence uniquely determines the distribution of $\beta_{it}$.
 Exogenous Shocks and Covariates
-The identification argument of this paper goes through when I further include ex-post shocks $\upsilon_{it}$ to $\beta_{it}$ (i.e., $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$) and $\epsilon_{it}$ to $Y_{it}$ (i.e., $Y_{it}=X_{it}^{{}^{\prime}}\beta_{it}+\epsilon_{it}$), where both $\upsilon_{it}$ and $\epsilon_{it}$ are independent of all other variables with $\mathbb{E}\epsilon_{it}=0$.
-Note that due to the inclusion of $\upsilon_{it}$, $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ becomes a possibly time-varying function of $X_{it}$ as the distribution of $\upsilon_{it}$ may depend on time $t$.
-It does not affect my identification analysis because $\upsilon_{it}$ is independent of all other variables.
-I present the changes required to the proof due to the inclusion of $\upsilon_{it}$ into $\beta_{it}$ at the end of the proof of Theorem 1.
-I follow the more general specification $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$ in the estimation and examine its impact via simulations.
-As for $\epsilon_{it}$, since it is independent of all other variables, has mean zero, and enters (2.1) in an additive way, equations (3.4)–(3.8) hold as before.
-Thus, the identification of $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is not affected.
-However, it has an impact on the identification of higher-order moments of $\beta_{it}$.
-For example, in (3.12) $\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ is not identified if $\epsilon_{it}$ is included with $\mathbb{E}\epsilon_{it}^{2}>0$.
-One way to handle this issue is by following Arellano and Bonhomme (2012) to impose and exploit additional assumptions on how $\epsilon_{it}$ evolves over time (e.g., $\epsilon_{it}$ follows an ARMA(p,q) process).
-Finally, when exogenous covariates $Z_{1,it}$ are available and included in $X_{it}$ [^20] If $Z_{1,it}$ is included in $X_{it}$, I can write $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, $U_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$, and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$., my identification analysis actually benefits from it because Assumptions 13 concern only endogenous covariates $U_{it}$ of $X_{it}$.
-Thus, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 and in the construction of $V_{it}$.
-Furthermore, as discussed in footnote 16, the unconditional variation of $Z_{1,it}$ makes Assumption 4 and 4 easier to be satisfied.
-The details on how to adapt the current analysis to include $Z_{1,it}$ are provided at the end of the proof of Theorem 1.
+The identification argument goes through when I include ex-post shocks $\upsilon_{it}$ to $\beta_{it}$ (i.e., $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$) and $\widetilde{\epsilon}_{it}$ to $Y_{it}$ (i.e., $Y_{it}=X_{it}^{\prime}\beta_{it}+\widetilde{\epsilon}_{it}$ with $\mathbb{E}\widetilde{\epsilon}_{it}=0$), where $\upsilon_{it}$ and $\widetilde{\epsilon}_{it}$ are independent of all other variables.
+Note that $\beta(\cdot)$ may a priori be time-varying and the joint distribution of $(\varepsilon_{it},\eta_{it})$ may also vary with $t$.
+Introducing ex-post shocks $\upsilon_{it}$ provides an additional and economically distinct source of time variation (e.g., unexpected technology shock).
+While the identification result remains valid with the inclusion of $\upsilon_{it}$, I present the required changes to the proof at the end of the proof of Theorem 1.
+I also include $\upsilon_{it}$ into $\beta_{it}$ when estimating the APE and LAR and examine its impact via simulations.
+As for $\widetilde{\epsilon}_{it}$, since it is independent of all other variables and enters (2.1) in an additive way, equations (3.9)–(3.13) hold as before.
+Thus, the identification analysis is unaffected.
+The presence of exogenous covariates in $X_{it}$ can strengthen identification.
+The key point is that Assumptions 13 pertain only to the endogenous covariates of $X_{it}$.
+In particular, the control variables $V_{it}$ and $W_{i}$ depend only on the endogenous covariates of $X_{it}$, which alleviates concerns that the controls may be high-dimensional and makes the required residual variation in $X_{it}$ easier to satisfy.
+Moreover, as discussed in Appendix A.3, unconditional variation in exogenous covariates of $X_{it}$ can be directly exploited to make Assumptions 4 and 4 easier to satisfy.
+Details on how to adapt the analysis to incorporate exogenous covariates are provided at the end of the proof of Theorem 1.
 4 Estimation and Large Sample Theory
-In this section, I show how to estimate $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ via three-step series estimators based on the constructive identification argument.
-Next, I establish the convergence rates of the proposed estimators.
-Finally, I prove that the estimators are asymptotically normal and provide consistent estimators for their asymptotic covariance matrices.
+In this section, I first describe how to estimate the parameters using three-step series estimators.
+I then establish the convergence rates of the proposed estimators, and finally show that they are asymptotically normal.
 4.1 Estimation
-As discussed in Extension 3 of Section 3, I include the ex-post shock $\upsilon_{it}$ in $\beta_{it}$, resulting in $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]\eqqcolon b_{1t}\left(V_{it},W_{i}\right)$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]\eqqcolon b_{t}\left(X_{it}\right)$.
 The parameters of interest are
-$\overline{b}\coloneqq\mathbb{E}\beta_{it}\ \text{and}\ b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right].$ (4.1)
-$\overline{b}$ is the APE over the unconditional distribution of $\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$.
-$b_{t}\left(x\right)$ is the LAR function for a subpopulation characterized by $X_{it}=x$ in period $t$.
-Both objects are useful for answering policy-related questions.
-For example, plugging realizations $x_{it}$ of $X_{it}$ into $b_{t}\left(x\right)$ provides a fine approximation to $\beta_{it}$ for agent $i$ in time $t$.
-Therefore, distributional properties of $\beta_{it}$ may be inferred from the distribution of $b_{t}\left(x_{it}\right)$.
+$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}\quad\text{and}\quad b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right].$ (4.1)
+$\overline{b}$ is the pooled APE across all firms and periods.
+$b_{t}\left(x\right)$ is the LAR function for a subpopulation characterized by $X_{it}=x$ in period $t$ and is also useful for answering policy-related questions.
+For example, plugging realizations $x_{it}$ of $X_{it}$ into $b_{t}\left(x\right)$ provides a fine approximation to $\beta_{it}$.
 I propose to estimate the parameters in (4.1) with three-step series estimators.
-For clarity of exposition, I let $d_{X}=1$, and highlight the changes required for $d_{X}>1$ when necessary.
+For clarity of exposition, I set $d_{X}=1$ and highlight the modifications required for $d_{X}>1$ when needed.
 To fix ideas, I follow Liu, Poirier, and Shiu (2025) to let $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$, which can be motivated by generalizing the method of Mundlak (1978).
-First, for each $t$, I estimate $V_{t}\left(x,z,w\right)\coloneqq F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)$ by regressing $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ on the basis functions $q^{M_{1}}\left(\cdot\right)$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau\left(\cdot\right)$:
+First, for each $t$, I estimate $V_{t}\left(x,z,w\right)\coloneqq F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)$ by regressing $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ on the basis functions $q^{M_{1}}$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau$:
 $\widehat{V}_{t}\left(x,z,w\right)$ $=\tau\left(\widehat{F}_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)\right)$
-$=\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}n^{-1}\widehat{Q}_{t}^{-1}\sum_{j=1}^{n}q_{jt}\mathbf{\mathbbm{1}}\left\{x_{jt}\leq x\right\}\right)$
-$\eqqcolon\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}\widehat{\gamma}_{t}^{M_{1}}\left(x\right)\right),$ (4.2)
-[…]
-Next, define $S\coloneqq\left(X,V,W\right)$ and let $\mathcal{X},\ \mathcal{Z},\ \mathcal{V},\ \mathcal{W},$ and $\mathcal{S}$ denote the supports of $X,\ Z,\ V,\ W,$ and $S$, respectively.
+$=\tau\left(q^{M_{1}}\left(z,w\right)^{\prime}n^{-1}\widehat{Q}_{t}^{-1}\sum_{j=1}^{n}q_{jt}\mathbf{\mathbbm{1}}\left\{x_{jt}\leq x\right\}\right)$
+$\eqqcolon\tau\left(q^{M_{1}}\left(z,w\right)^{\prime}\widehat{\gamma}_{t}^{M_{1}}\left(x\right)\right),$ (4.2)
+where $\widehat{Q}_{t}:=n^{-1}\sum_{i=1}^{n}q_{it}q_{it}^{\prime}$ and $q_{it}:=q^{M_{1}}\left(z_{it},w_{i}\right)$.
+Examples of $q^{M_{1}}$ include power series and spline functions.
+When $d_{X}>1$, I regress $\mathbf{\mathbbm{1}}\left\{X_{it,l}\leq x_{l}\right\}$ on the basis functions $q^{M_{1}}$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau$ for each $l$ and obtain $\widehat{V}_{t}\left(x,z,w\right)=\left(\widehat{V}_{t,1}\left(x,z,w\right),\ldots,\widehat{V}_{t,d_{X}}\left(x,z,w\right)\right)^{\prime}$.
+I highlight two properties of $\widehat{V}_{t}\left(x,z,w\right)$.
+First, the regression coefficient $\widehat{\gamma}_{t}^{M_{1}}\left(x\right)$ in (4.2) depends on $x$ because the dependent variable $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ is a function of $x$.
+This fact causes the convergence rate of $\widehat{V}_{t}\left(x,z,w\right)$ to be slower than the standard rates for series estimators (Imbens and Newey (2009)).
+Second, a trimming function $\tau$ is needed since I am estimating a conditional CDF.
+An example of $\tau$ is $\tau\left(x\right)=\mathbf{\mathbbm{1}}\left\{x>0\right\}\cdot\min\left(x,1\right)$.
+Next, define $S\coloneqq\left(X,V,W\right)$ and let $\mathcal{X},\ \mathcal{Z},\ \mathcal{V},\ \mathcal{W},$ and $\mathcal{S}$ denote the supports of X,Z,V,W,X,\ Z,\ V,\ W, and $S$, respectively.
 Let $V_{it}\coloneqq V_{t}\left(X_{it},Z_{it},W_{i}\right),$$\widehat{V}_{it}\coloneqq\widehat{V}_{t}\left(X_{it},Z_{it},W_{i}\right)$, and $\widehat{v}_{it}\coloneqq\widehat{V}_{t}\left(x_{it},z_{it},w_{i}\right)$.
-For any $s=\left(x,v,w\right)\in\mathcal{S}$, I estimate $G_{t}\left(s\right)\coloneqq\mathbb{E}\left[\left.Y_{it}\right|S_{it}=s\right]$ by regressing $Y_{it}$ on the basis functions $p^{M_{2}}\left(\cdot\right)$ of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$:
-$\widehat{G}_{t}\left(s\right)=p^{M_{2}}\left(s\right)^{{}^{\prime}}n^{-1}\widehat{P}_{t}^{-1}\widehat{p}_{t}^{{}^{\prime}}y_{t}\eqqcolon p^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$
-where $\widehat{P}_{t}:=n^{-1}\sum_{i=1}^{n}\widehat{p}_{it}\widehat{p}_{it}^{{}^{\prime}}$, $\widehat{p}_{it}:=p^{M_{2}}\left(x_{it},\widehat{v}_{it},w_{i}\right)$, $\widehat{p}_{t}:=\left(\widehat{p}_{1t},...,\widehat{p}_{nt}\right)^{{}^{\prime}}$, and $y_{t}:=\left(y_{1t},...,y_{nt}\right)^{{}^{\prime}}$.
-Following Newey, Powell, and Vella (1999), I let $p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$ by exploiting the index structure of (2.1), which enables a faster convergence rate of $\widehat{G}_{t}\left(s\right)$ to $G_{t}\left(s\right)$.
+For any $s=\left(x,v,w\right)\in\mathcal{S}$, I estimate $G_{t}\left(s\right)\coloneqq\mathbb{E}\left[\left.Y_{it}\right|S_{it}=s\right]$ by regressing $Y_{it}$ on the basis functions $p^{M_{2}}$ of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$:
+$\widehat{G}_{t}\left(s\right)=p^{M_{2}}\left(s\right)^{\prime}n^{-1}\widehat{P}_{t}^{-1}\widehat{p}_{t}^{\prime}y_{t}\eqqcolon p^{M_{2}}\left(s\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}},$ (4.3)
+where $\widehat{P}_{t}:=n^{-1}\sum_{i=1}^{n}\widehat{p}_{it}\widehat{p}_{it}^{\prime}$, $\widehat{p}_{it}:=p^{M_{2}}\left(x_{it},\widehat{v}_{it},w_{i}\right)$, $\widehat{p}_{t}$ is an $n\times M_{2}$ matrix with $\widehat{p}_{it}^{\prime}$ being its $i^{\text{th}}$ row, and $y_{t}$ is the vector of $y_{it}$s.
+Following Newey, Powell, and Vella (1999), I let
+$p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$ (4.4)
+by exploiting the index structure of (2.1). Hence, the effective degree that matters for the convergence rate is $m_{2}$ since $M_{2}=d_{X}\times m_{2}$ and $d_{X}$ is finite.
 Finally, I exploit the index structure of (2.1) again to estimate $b_{1t}\left(v,w\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|V_{it}=v,W_{i}=w\right]$.
 When (3.11) is used, I estimate $b_{1t}\left(v,w\right)$ by
-$\widehat{b}_{1t}\left(v,w\right)=\left(\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it}=v,W_{i}=w\right]\right)^{-1}\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it}=v,W_{i}=w\right],$ (4.3)
-where $\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it},W_{i}\right]$ is obtained by regressing each coordinate of $X_{it}Y_{it}$ on the basis functions of $\left(\widehat{V}_{it},W_{i}\right)$ and similarly for $\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it},W_{i}\right]$.
+$\widetilde{b}_{1t}\left(v,w\right)=\left(\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{\prime}\right|\widehat{V}_{it}=v,W_{i}=w\right]\right)^{-1}\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it}=v,W_{i}=w\right],$ (4.5)
+where $\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it},W_{i}\right]$ is obtained by regressing each coordinate of $X_{it}Y_{it}$ on the basis functions of $\left(\widehat{V}_{it},W_{i}\right)$ and similarly for $\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{\prime}\right|\widehat{V}_{it},W_{i}\right]$.
 When (3.12) is used, I estimate $b_{1t}\left(v,w\right)$ by
-$\widehat{b}_{1t}\left(v,w\right)=\partial\widehat{G}_{t}\left(s\right)/\partial x=\left(I_{d_{X}}\otimes p^{m_{2}}\left(v,w\right)\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}}\eqqcolon\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$ (4.4)
-where the second equality holds by the chain rule. I follow (4.4) in what follows for its simplicity in deriving the asymptotic properties.
-To estimate $\overline{b}$ and $b_{t}\left(x\right)$, by the LIE I regress $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the basis function $r^{M_{3}}\left(\cdot\right)$ of constant one and $X_{it}$, respectively:
-$\widehat{\overline{b}}$ $=\left(nT\right)^{-1}\sum_{t=1}^{T}\sum_{i=1}^{n}\widehat{b}_{1t}\left(\widehat{v}_{it},w_{i}\right),\text{ and}$
-$\widehat{b}_{t}\left(x\right)$ $=r^{M_{3}}\left(x\right)^{{}^{\prime}}n^{-1}\widehat{R}_{t}^{-1}r_{t}^{{}^{\prime}}\widehat{B}_{t}\eqqcolon r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{\rho}_{t}^{M_{3}},$ (4.5)
-where $\widehat{R}_{t}:=n^{-1}\sum_{i=1}^{n}r_{it}r_{it}^{{}^{\prime}}$, $r_{it}:=r^{M_{3}}\left(x_{it}\right)$, $r_{t}:=\left(r_{1t},...,r_{nt}\right)^{{}^{\prime}}$, and $\widehat{B}_{t}:=\left(\widehat{b}_{1t}\left(\widehat{v}_{1t},w_{1}\right),...,\widehat{b}_{1t}\left(\widehat{v}_{nt},w_{n}\right)\right)^{{}^{\prime}}$.
+$\widehat{b}_{1t}\left(v,w\right)=\partial\widehat{G}_{t}\left(s\right)/\partial x=\left(I_{d_{X}}\otimes p^{m_{2}}\left(v,w\right)\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}}\eqqcolon\overline{p}^{M_{2}}\left(s\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}},$ (4.6)
+where the second equality holds by the chain rule. I follow (4.6) in what follows, as it simplifies both implementation and the derivation of asymptotic properties.
+Remark 2 (Compare $\boldsymbol{\widetilde{b}_{1t}\left(v,w\right)}$ with $\boldsymbol{\widehat{b}_{1t}\left(v,w\right)}$).
+While both approaches are valid, this paper recommends using the derivative-based estimator $\widehat{b}_{1t}(v,w)$ rather than the inversion-based estimator $\widetilde{b}_{1t}(v,w)$ in practice when Assumption 4 holds.
+The derivative-based route operates through the estimated one-dimensional regression function $\widehat{G}_{t}(x,v,w)$ and obtains $\widehat{b}_{1t}(v,w)$ via the analytic partial derivative with respect to $x$.
+This approach exploits the linear random-coefficient structure directly, avoids an additional “matrix estimation + inversion” step, and leads to a cleaner and more stable asymptotic argument in the paper’s framework.
+By contrast, $\widetilde{b}_{1t}(v,w)$ requires estimating and inverting the conditional second-moment matrix $\mathbb{E}[X_{it}X_{it}^{\prime}\mid V_{it}=v,W_{i}=w]$, whose sample analogue can be ill-conditioned in finite samples and hence numerically unstable (Carrasco, Florens, and Renault (2007)).
+This inversion also complicates the asymptotic analysis, because one must control how estimation error in the conditional second moment propagates through a matrix inverse.
+To estimate $\overline{b}$ and $b_{t}\left(x\right)$, by the law of iterated expectations I regress $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the basis function $r^{M_{3}}$ of constant one and $X_{it}$, respectively:
+$\widehat{\overline{b}}=\dfrac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\widehat{b}_{1t}\left(\widehat{v}_{it},w_{i}\right)\quad\text{and}\quad\widehat{b}_{t}\left(x\right)=r^{M_{3}}\left(x\right)^{\prime}n^{-1}\widehat{R}_{t}^{-1}r_{t}^{\prime}\widehat{B}_{t}\eqqcolon r^{M_{3}}\left(x\right)^{\prime}\widehat{\rho}_{t}^{M_{3}},$ (4.7)
+where $\widehat{R}_{t}:=n^{-1}\sum_{i=1}^{n}r_{it}r_{it}^{\prime}$, $r_{it}:=r^{M_{3}}(x_{it})$, $r_{t}$ is an $n\times M_{3}$ matrix with $r_{it}^{\prime}$ being its $i^{\text{th}}$ row, and $\widehat{B}_{t}$ is an $n\times d_{X}$ matrix with $\widehat{b}_{1t}(\widehat{v}_{it},w_{i})^{\prime}$ being its $i^{\text{th}}$ row.
+Remark 3 (Computational Considerations).
+While the three-step estimators are based on standard series regressions and are typically computationally manageable, two implementation aspects are worth noting for large datasets.
+First, when estimating $V_{t}\left(x,z,w\right)$, a key computational simplification is that the regressors—the basis functions $q^{M_{1}}(Z_{it},W_{i})$—are identical across all thresholds, with only the dependent variable $\mathbf{\mathbbm{1}}\{X_{it}\leq x\}$ changing in $x$.
+Hence, I can compute the inverse Gram matrix $\widehat{Q}_{t}^{-1}$ once and reuse it to obtain the coefficient vector $\widehat{\gamma}_{t}^{M_{1}}\left(x\right)$ for every threshold $x_{it}$.
+This reuse is substantially cheaper than running $n$ fully separate regressions.
+Second, when estimating $G_{t}\left(s\right)$, I exploit the linear structure in $X_{it}$ using the separable specification $p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$, which reduces the effective regressor dimensionality and alleviates the associated computational burden.
 4.2 Convergence Rates and Asymptotic Normality
-Since I let $n\rightarrow\infty$ for each fixed $t$ in the asymptotic analysis, the $t$-subscript is suppressed for brevity of exposition when there is no confusion.
-Let $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$ and $P\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}$.
-As Imbens and Newey (2009) have studied the convergence rates for the series estimators of conditional CDF (e.g., $F_{\left.X\right|Z}$) and conditional expectation functions (e.g., $\mathbb{E}\left[\left.Y\right|X,V\right]$), I adapt to their analysis to prove convergence rates for $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$.
-I impose the following assumption.
+Since I let $n\rightarrow\infty$ for each fixed $t$ in the asymptotic analysis, the $t$-subscript is suppressed when it is clear.
+Let $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$ and $P\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}$.
+In the analysis, I use results from Imbens and Newey (2009) to establish convergence rates for $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$.
+The following assumption is needed.
 Assumption 5.
@@
   1. (a)
-    $\left({\bf X}_{i},{\bf Z}_{i},A_{i},\varepsilon_{i},\eta_{i}\right)$ is i.i.d. across $i$.
+    $V\left(x,z,w\right)$ is continuously differentiable of order $d_{1}$ on the support with derivatives uniformly bounded in $\left(x,z,w\right)$ and $\mathcal{Z}\times\mathcal{W}\subset\mathbb{R}^{j_{1}}$.
   2. (b)
-    $V\left(x,z,w\right)$ is continuously differentiable of order $d_{1}$ on the support with derivatives uniformly bounded in $\left(x,z,w\right)$ and $\mathcal{Z}\times\mathcal{W}\subset\mathbb{R}^{j_{1}}$.
-  3. (c)
     $p^{m_{2}}\left(v,w\right)=p^{m_{2v}}\left(v\right)\otimes p^{m_{2w}}\left(w\right)$ and there exist constants $C,\theta>0$ such that $\lambda_{\min}\left(P\right)\geq C$ and
     $\inf_{w\in\mathcal{W}}f_{V,W}\left(v,w\right)\geq C\left[v\left(1-v\right)\right]^{\theta}.$
+  3. (c)
+    $G\left(s\right)$ is continuously differentiable of order $d_{2}>1$ on $\mathcal{S}\subset\mathbb{R}^{j_{2}}$.
   4. (d)
-    $G\left(s\right)$ is continuously differentiable of order $d_{2}>1$ on $\mathcal{S}\subset\mathbb{R}^{j_{2}}$.
-  5. (e)
     $\mathbb{\mathbb{V}}\left(\left.Y\right|X,Z,W\right)$ is bounded.
 Let $\zeta\left(m_{2}\right)\coloneqq m_{2v}^{\theta}m_{2}$ and $\zeta_{1}\left(m_{2}\right)\coloneqq m_{2v}^{\theta+2}m_{2}$.
-With Assumption 5 in position, the next lemma follows directly from Theorem 12 of Imbens and Newey (2009).
+With Assumption 5 in place, the next lemma follows directly from Theorem 12 of Imbens and Newey (2009).
+Let $\Delta_{1n}^{2}:=n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}$ and $\Delta_{2n}^{2}:=\Delta_{1n}^{2}+n^{-1}m_{2}+m_{2}^{-2d_{2}/j_{2}}$ in the next lemma.
 Lemma 1 (Convergence Rates of $\boldsymbol{\widehat{V}}$ and $\boldsymbol{\widehat{G}}$).
 If the conditions of Theorem 1 and Assumption 5 are satisfied, and $m_{2}^{2}m_{2v}^{\theta+2}\left(n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}\right)\to 0$, then
-$\mathbb{E}\left[n^{-1}\sum_{i}\left(\widehat{V}_{i}-V_{i}\right)^{2}\right]$ $=O\left(n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}\right)\eqqcolon O\left(\Delta_{1n}^{2}\right),$
-[…]
-$\sup_{s\in\mathcal{S}}\left|\widehat{G}\left(s\right)-G\left(s\right)\right|$ $=O_{p}\left(\zeta\left(m_{2}\right)\Delta_{2n}\right).$
+$\mathbb{E}\left[n^{-1}\sum_{i}\left(\widehat{V}_{i}-V_{i}\right)^{2}\right]$ $=O\left(\Delta_{1n}^{2}\right),$
+$\int\left[\widehat{G}\left(s\right)-G\left(s\right)\right]^{2}dF\left(s\right)$ $=O_{p}\left(\Delta_{2n}^{2}\right),\quad\text{and}\quad\sup_{s\in\mathcal{S}}\left|\widehat{G}\left(s\right)-G\left(s\right)\right|=O_{p}\left(\zeta\left(m_{2}\right)\Delta_{2n}\right).$
 Lemma 1 states that the mean-square convergence rate for $\widehat{G}\left(s\right)$ is the sum of the first-step rate $\Delta_{1n}^{2}$, the variance term $n^{-1}m_{2}$, and the squared bias term $m_{2}^{-2d_{2}/j_{2}}$.
-$d_{1}/j_{1}$ and $d_{2}/j_{2}$ are the uniform approximation rates that govern how well the unknown functions $V\left(x,z,w\right)$ and $G\left(s\right)$ can be approximated with $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$, respectively; see Assumption 3 and 5 of Imbens and Newey (2009) for the details.
-Note that even though the order of the basis function for the second-step estimation is $M_{2}$, by the index structure of (2.1) $M_{2}=d_{X}\times m_{2}$ and $d_{X}$ is finite.
-Thus, the effective order that matters for the convergence rate is $m_{2}$.
-Let $\xi_{i}\coloneqq b_{1}\left(V_{i},W_{i}\right)-b\left(X_{i}\right)$ and $\xi\coloneqq\left(\xi_{1},...,\xi_{n}\right)^{{}^{\prime}}$.
-To derive the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, I impose the following assumption.
+$d_{1}/j_{1}$ and $d_{2}/j_{2}$ are the uniform approximation rates that govern how well the unknown functions $V\left(x,z,w\right)$ and $G\left(s\right)$ can be approximated with $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$, respectively; see Assumptions 3 and 5 of Imbens and Newey (2009) for the details.
+Let $\xi_{i}\coloneqq b_{1}\left(V_{i},W_{i}\right)-b\left(X_{i}\right)$ and $\xi\coloneqq\left(\xi_{1},\ldots,\xi_{n}\right)^{\prime}$.
+To derive the convergence rates for the APE and LAR estimators, I impose the following assumption.
 Assumption 6.
@@
   2. (b)
-    There is a constant $C>0$ and $\zeta\left(M_{3}\right)$, such that for each $M_{3}$ there exists a non-singular constant matrix $N_{r}$ such that $\widetilde{r}^{M_{3}}\left(x\right)\coloneqq N_{r}r^{M_{3}}\left(x\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{r}^{M_{3}}\left(X_{i}\right)\widetilde{r}^{M_{3}}\left(X_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{x\in\mathcal{X}}\left\|\widetilde{r}^{M_{3}}\left(x\right)\right\|\leq C\zeta\left(M_{3}\right)$.
+    There is a constant $C>0$ and $\zeta\left(M_{3}\right)$, such that for each $M_{3}$ there exists a non-singular constant matrix $N_{r}$ such that $\widetilde{r}^{M_{3}}\left(x\right)\coloneqq N_{r}r^{M_{3}}\left(x\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{r}^{M_{3}}\left(X_{i}\right)\widetilde{r}^{M_{3}}\left(X_{i}\right)^{\prime}\right)\geq C$ and $\sup_{x\in\mathcal{X}}\left\|\widetilde{r}^{M_{3}}\left(x\right)\right\|\leq C\zeta\left(M_{3}\right)$.
   3. (c)
-    $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded.
+    $\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]$ is bounded.
 Assumption 6 imposes conditions on the degree of smoothness of $b\left(x\right)$, the normalization of basis functions $r^{M_{3}}\left(x\right)$, and the boundedness of the second moment of $\xi_{i}$, similar to those in Assumption 5.
@@
 I use this result to prove the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, both of which are unknown but estimable functionals of $\widehat{G}\left(s\right)$.
+Let $\Delta_{3n}^{2}:=n^{-1}M_{3}+M_{3}^{-2d_{3}/j_{3}}+\Delta_{2n}^{2}$ in the next theorem.
 Theorem 2 (Convergence Rates of $\boldsymbol{\widehat{\overline{b}}}$ and $\boldsymbol{\widehat{b}\left(x\right)}$).
@@
 $\left\|\widehat{\overline{b}}-\overline{b}\right\|^{2}$ $=O_{p}\left(\Delta_{2n}^{2}\right),$
-[…]
-$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
+$\int\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|^{2}dF(x)$ $=O_{p}\left(\Delta_{3n}^{2}\right),\quad\text{and}\quad\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
 The 2002 working paper version of Imbens and Newey (2009) (henceforth IN02) has obtained asymptotic normality for estimators of known and scalar-valued linear functionals of $G\left(s\right)$.
-I adapt to their results to analyze vector-valued functionals of $G\left(s\right)$ via a Cramér–Wold device and prove asymptotic normality for $\widehat{b}_{1}\left(v,w\right)$.
-With slight abuse of notation, I take $\widetilde{\overline{p}}^{M_{2}}\left(s\right)$ and $\widetilde{r}^{M_{3}}\left(x\right)$ that satisfy Assumptions 5 and 6 as $\overline{p}^{M_{2}}\left(s\right)$ and $r^{M_{3}}\left(x\right)$ in what follows.
+I build on their results to analyze vector-valued functionals of $G\left(s\right)$ via a Cramér–Wold device and prove asymptotic normality for $\widehat{b}_{1}\left(v,w\right)$.
+To simplify the notation, I take $\widetilde{\overline{p}}^{M_{2}}\left(s\right)$ and $\widetilde{r}^{M_{3}}\left(x\right)$ that satisfy Assumptions 5 and 6 as $\overline{p}^{M_{2}}\left(s\right)$ and $r^{M_{3}}\left(x\right)$ in what follows.
 Assumption 7.
@@
   1. (a)
-    There is a constant $C>0$ and $\zeta\left(M_{1}\right)$, such that for each $M_{1}$ there exists a non-singular constant matrix $N_{q}$ such that $\widetilde{q}^{M_{1}}\left(z,w\right)\coloneqq N_{q}q^{M_{1}}\left(z,w\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{\left(z,w\right)\in\mathcal{Z\times\mathcal{W}}}\left\|\widetilde{q}^{M_{1}}\left(z,w\right)\right\|\leq C\zeta\left(M_{1}\right)$.
+    There is a constant $C>0$ and $\zeta\left(M_{1}\right)$, such that for each $M_{1}$ there exists a non-singular constant matrix $N_{q}$ such that $\widetilde{q}^{M_{1}}\left(z,w\right)\coloneqq N_{q}q^{M_{1}}\left(z,w\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)^{\prime}\right)\geq C$ and $\sup_{\left(z,w\right)\in\mathcal{Z\times\mathcal{W}}}\left\|\widetilde{q}^{M_{1}}\left(z,w\right)\right\|\leq C\zeta\left(M_{1}\right)$.
   2. (b)
     $G\left(s\right)$ is twice continuously differentiable with bounded first and second derivatives.
-For functional $a\left(\cdot\right)$ of $G$ and some constant $C>0$, it is true that $\left|a\left(G\right)\right|\leq C\sup_{s}\left|G\left(s\right)\right|$ and either (i) there is $\delta\left(s\right)$ and $\widetilde{\alpha}^{M_{2}}$ such that $\mathbb{E}\delta\left(S_{i}\right)^{2}<\infty$, $a\left(p_{m}^{M_{2}}\right)=\mathbb{E}\delta\left(S_{i}\right)p_{m}^{M_{2}}\left(S_{i}\right)$ for all $m=1,...,M_{2}$, $a\left(G\right)=\mathbb{E}\delta\left(S_{i}\right)G\left(S_{i}\right)$, and $\mathbb{E}\left(\delta\left(S_{i}\right)-p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)^{2}\rightarrow 0$; or (ii) for some $\widetilde{\alpha}^{M_{2}}$, $\mathbb{E}\left[p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right]^{2}\rightarrow 0$ and $a\left(p^{M_{2}}\left(\cdot\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)$ is bounded away from zero as $M_{2}\rightarrow\infty$.
+For functional $a$ of $G$ and some constant $C>0$, it is true that $\left|a\left(G\right)\right|\leq C\sup_{s}\left|G\left(s\right)\right|$ and either (i) there is $\delta\left(s\right)$ and $\widetilde{\alpha}^{M_{2}}$ such that $\mathbb{E}\delta\left(S_{i}\right)^{2}<\infty$, $a\left(p_{m}^{M_{2}}\right)=\mathbb{E}\delta\left(S_{i}\right)p_{m}^{M_{2}}\left(S_{i}\right)$ for all $m=1,...,M_{2}$, $a\left(G\right)=\mathbb{E}\delta\left(S_{i}\right)G\left(S_{i}\right)$, and $\mathbb{E}\left(\delta\left(S_{i}\right)-p^{M_{2}}\left(S_{i}\right)^{\prime}\widetilde{\alpha}^{M_{2}}\right)^{2}\rightarrow 0$; or (ii) for some $\widetilde{\alpha}^{M_{2}}$, $\mathbb{E}\left[p^{M_{2}}\left(S_{i}\right)^{\prime}\widetilde{\alpha}^{M_{2}}\right]^{2}\rightarrow 0$ and $a\left(p^{M_{2}}{}^{\prime}\widetilde{\alpha}^{M_{2}}\right)$ is bounded away from zero as $M_{2}\rightarrow\infty$.
   3. (c)
@@
     There exist $d_{4}$ and $\overline{\alpha}^{M_{2}}$ such that for each element $s_{j}$ of $s=\left(x,v,w\right)\in\mathcal{S}\subset\mathbb{R}^{j_{4}}$:
-    $\sup\left\{\sup_{s\in\mathcal{S}}\left|G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right|,\sup_{s\in\mathcal{S}}\left|\partial\left(G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right)/\partial s_{j}\right|\right\}=O\left(M_{2}^{-d_{4}/j_{4}}\right).$
+    $\sup\left\{\sup_{s\in\mathcal{S}}\left|G\left(s\right)-p^{M_{2}}\left(s\right)^{\prime}\overline{\alpha}^{M_{2}}\right|,\sup_{s\in\mathcal{S}}\left|\partial\left(G\left(s\right)-p^{M_{2}}\left(s\right)^{\prime}\overline{\alpha}^{M_{2}}\right)/\partial s_{j}\right|\right\}=O\left(M_{2}^{-d_{4}/j_{4}}\right).$
     Also, $nM_{2}^{-2d_{4}/j_{4}}$ and $M_{1}M_{2}^{-2d_{4}/j_{4}}\zeta_{1}\left(M_{2}\right)^{2}$ are $o\left(1\right)$.
   6. (f)
-[…]
-Assumptions 7(a)(e) are also imposed by IN02 and are regularity conditions required for the asymptotic normality of $\widehat{b}_{1}\left(v,w\right)$.
+    (Assumption J(iii) of Andrews (1991)) For a bounded sequence of constants $\left\{c_{1n}:n\geq 1\right\}$ and constant pd matrix $\overline{\Omega}_{1}$, it is true that $c_{1n}\Omega_{1}\overset{p}{\longrightarrow}\overline{\Omega}_{1}$, where $\Omega_{1}$ is defined in (A.12).
+Assumption 7(a)(e) is also imposed by IN02 and is a regularity condition required for the asymptotic normality of $\widehat{b}_{1}\left(v,w\right)$.
 Assumption 7(f) concerns the asymptotic covariance matrix of $\widehat{b}_{1}\left(v,w\right)$ and is used by Andrews (1991).
 It guarantees that the normality result of IN02 applies to vector-valued functionals of $G\left(s\right)$.
-Essentially, Assumption 7(f) requires that all the coordinates of $\widehat{b}_{1}\left(v,w\right)$ converge at the same speed, which is a mild assumption in my setting because ex-ante I do not distinguish any coordinate of $\beta_{it}$ from the others.
+Essentially, Assumption 7(f) requires that all the coordinates of $\widehat{b}_{1}\left(v,w\right)$ converge at the same speed, which is mild because ex-ante I do not distinguish any coordinate of $\beta_{it}$ from the others.
 Lemma 2 (Asymptotic Normality of $\boldsymbol{\widehat{b}_{1}\left(v,w\right)}$).
-If the conditions of Theorem 2 and Assumption 7 are satisfied, then,
-[…]
+If the conditions of Theorem 2 and Assumption 7 are satisfied, then
+$\sqrt{n}\,\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)\overset{d}{\longrightarrow}N\left(0,I\right),$
+where $\Omega_{1}$ is defined in (A.12).
+Furthermore,
+$\widehat{\Omega}_{1}\Omega_{1}^{-1}\overset{p}{\longrightarrow}I,$
 where $\widehat{\Omega}_{1}$ is defined in (A.13).
@@
 Therefore, the results of IN02 directly apply and I omit its proof in this paper.
-However, the results of IN02 do not directly apply to $\overline{b}$ and $b\left(x\right)$ because they are unknown functionals of $G\left(s\right)$.
-To explain, notice that by the LIE
-$\overline{b}=\mathbb{E}\left[\partial G\left(S\right)/\partial X\right]\text{ and }b\left(x\right)=\mathbb{E}\left[\left.\partial G\left(S\right)/\partial X\right|X=x\right],$ (4.6)
+However, the results of IN02 do not directly apply to $\widehat{\overline{b}}$ and $\widehat{b}(x)$ because $\overline{b}$ and $b(x)$ are unknown functionals of $G\left(s\right)$.
+To explain, notice that by the law of iterated expectations
+$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\left[\partial G\left(S\right)/\partial X\right],\quad\text{and}\quad b\left(x\right)=\mathbb{E}\left[\left.\partial G\left(S\right)/\partial X\right|X=x\right],$ (4.8)
 both of which involve integrating $b_{1}\left(V,W\right)=\partial G\left(S\right)/\partial X$ with respect to the unknown but estimable distribution of $\left(V,W\right)$.
@@
   1. (a)
-    $\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}$ has full column rank.
+    $\mathbb{E}\overline{p}_{i}r_{i}^{\prime}$ has full column rank.
   2. (b)
-    $\mathbb{E}\left[\left.\left\|\xi\right\|^{4}\right|{\bf X}\right]$ is bounded and $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded away from zero.
+    $\mathbb{E}\left[\left.\left\|\xi\right\|^{4}\right|{\bf X}\right]$ is bounded and $\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]$ is bounded away from zero.
   3. (c)
-[…]
-Assumption 8(a) is needed to show that the asymptotic covariance matrix $\Omega_{2}$ of $n^{1/2}\left(\widehat{b}\left(x\right)-b\left(x\right)\right)$ is positive definite.
+    For a sequence of bounded constants $\left\{c_{2n},c_{3n}:n\geq 1\right\}$ and some constant pd matrix $\overline{\Omega}_{2}$ and $\overline{\Omega}_{3}$, $c_{2n}\Omega_{2}\overset{p}{\longrightarrow}\overline{\Omega}_{2}$ and $c_{3n}\Omega_{3}\overset{p}{\longrightarrow}\overline{\Omega}_{3}$, where $\Omega_{2}$ and $\Omega_{3}$ are defined in (A.18) and (A.64), respectively.
+  4. (d)
+    $\mathbb{E}\left\|b_{1}\left(v,w\right)-\overline{b}\right\|^{4}<\infty$.
+Assumption 8(a) is needed to show that the asymptotic covariance matrix $\Omega_{2}$ of $\sqrt{n}(\widehat{b}\left(x\right)-b\left(x\right))$ is positive definite and similarly for $\Omega_{3}$.
 Assumption 8(b) is a regularity condition imposed for the Lindeberg–Feller central limit theorem (CLT).
-Assumption 8(c) is similar to Assumption 7(f) and is needed to prove that the asymptotic normality result holds for vector-valued functionals of $G\left(s\right)$.
+Assumption 8(c) is similar to Assumption 7(f) and is needed to prove that the asymptotic normality result holds for vector-valued functionals of $G\left(s\right)$ by the Cramér–Wold device.
 Theorem 3 (Asymptotic Normality of $\boldsymbol{\widehat{\overline{b}}}$ and $\boldsymbol{\widehat{b}\left(x\right)}$).
 If the conditions of Lemma 2 and Assumption 8 are satisfied, then
+$\sqrt{n}\,\Omega_{2}^{-1/2}(\widehat{b}\left(x\right)-b\left(x\right))$ $\overset{d}{\longrightarrow}N\left(0,I\right)\quad\text{and}\quad\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})\overset{d}{\longrightarrow}N(0,I),$
+where $\Omega_{2}$ and $\Omega_{3}$ are defined in (A.18) and (A.64), respectively.
+Furthermore,
+$\widehat{\Omega}_{2}\Omega_{2}^{-1}\overset{p}{\longrightarrow}I\quad\text{and}\quad\widehat{\Omega}_{3}\Omega_{3}^{-1}\overset{p}{\longrightarrow}I,$
+where $\widehat{\Omega}_{2}$ and $\widehat{\Omega}_{3}$ are defined in (A.48) and (A.67), respectively.
+I present the main idea to the proof of Theorem 3, focusing on how the error propagates through the three-step analysis and the roles of $\widehat{\Omega}_{2}$ and $\widehat{\Omega}_{3}$.
+Define the functionals
+$a\left(b_{1},V\right)$ $\coloneqq\mathbb{E}\left[\left.b_{1}\left(V,W\right)\right|X=x\right]=b\left(x\right),$
+$\widehat{a}\left(b_{1},V\right)$ $:=\widehat{\mathbb{E}}\left[\left.b_{1}\left(V,W\right)\right|X=x\right],$
 […]
-where $\widehat{\Omega}_{2}$ is defined in (A.45).
-Furthermore, if $\mathbb{E}\left\|b_{1}\left(v,w\right)-\overline{b}\right\|^{4}<\infty$, then
 […]
-where $\widehat{\Omega}_{3}$ is defined in (A.58).
-5 Empirical Illustration
+Then
+[…]
+This decomposition makes clear that the asymptotic covariance of $\sqrt{n}\,(\widehat{b}(x)-b(x))$, denoted by $\Omega_{2}$ as in (A.18), must account for errors of all three sources.
+In the proof of Theorem 3, I show that
+$\sqrt{n}\,\Omega_{2}^{-1/2}\big(\widehat{b}(x)-b(x)\big)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\sum_{k=1}^{3}\psi_{ki}+o_{p}(1),$ (4.10)
+where the influence functions $\psi_{1i},\ \psi_{2i},\text{ and }\psi_{3i}$ correspond exactly to the three error terms on the right-hand side of (4.9) up to proper normalization.
+Once (4.10) is established, I verify the Lindeberg-Feller condition on $\Psi_{in}:=n^{-1/2}\sum_{k=1}^{3}\psi_{ki}$, which proves its convergence to the standard normal distribution and that $\Omega_{2}$ is indeed the appropriate asymptotic covariance associated with $\sqrt{n}\,(\widehat{b}(x)-b(x))$.
+For feasible inference, a consistent estimator for $\Omega_{2}$ is needed, and I show in the proof of Theorem 3 that $\widehat{\Omega}_{2}$ is consistent.
+Note that $\Omega_{2}$ may diverge because $\widehat{b}(x)$ converges more slowly than $n^{-1/2}$ as shown in Theorem 2.
+Hence, all convergence results are expressed in self-normalized way, for example,
+$\sqrt{n}\,\Omega_{2}^{-1/2}\big(\widehat{b}(x)-b(x)\big)\overset{d}{\longrightarrow}N(0,I)\quad\text{and}\quad\widehat{\Omega}_{2}\Omega_{2}^{-1}\overset{p}{\longrightarrow}I.$ (4.11)
+The same argument applies to $\widehat{\Omega}_{3}$ with the basis function $r^{M_{3}}\left(x\right)=1$ and the influence function capturing serial correlation of estimating the APE across periods.
+5 Empirical Application
 In this section, I apply my procedure to estimate a heterogeneous Cobb-Douglas production function for each of the five largest manufacturing sectors in China.
-I obtain unconditional means of the output elasticities and compare them with those derived using classic methods on the same data set.
-Furthermore, I estimate conditional means of the elasticities given the regressors.
+I obtain unconditional means of the output elasticities (APE) and compare them with those derived using classic methods on the same data set.
+Furthermore, I estimate conditional means of the elasticities given the regressors (LAR).
 Results show that there are significant across-firm variations in the output elasticities.
-In Appendix C, I conduct a simulation study motivated by production function applications to support the findings of this illustration.
+I also examine how heterogeneity in output elasticities relates to observed characteristics and obtain intuitive results.
+In Appendix B, I conduct several robustness checks—including alternative price deflators, sample trimming, alternative IV and $W_{i}$ constructions, city–year fixed effects, and comparison with OLS and naive IV estimates—and the results remain stable.
+In Appendix C, I conduct a production-function motivated simulation study to support the findings in this application.
 5.1 Data and Methodology
@@
 According to Brandt, Van Biesebroeck, Wang, and Zhang (2017), they account for 91% of the gross output, 71% of employment, 97% of exports, and 91% of total fixed assets in 2004, and thus are representative of industrial activities in China.
-Many papers on topics such as firm behavior, international trade, and growth theory have used the CASIF data (e.g., Hsieh and Klenow (2009), Brandt, Van Biesebroeck, Wang, and Zhang (2017)).
-I focus on the five largest 2-digit sectors in terms of the number of firms between 2004 and 2007.222222There are 3140 textile firms, 2186 chemical firms, 1804 nonmetallic minerals firms, 2869 general equipment firms, and 1568 transportation equipment firms in the final dataset.
-I order the industries by their 2-digit Chinese Industry Classification codes.
-[^23] The CASIF dataset spans between 1998 and 2007.
-I choose year 2004 to 2007 to (i) ensure data consistency due to the change in the Chinese Industry Classification codes in 2003, (ii) avoid the financial crisis in the early 2000s, and (iii) use the most recent data.
+Many papers on topics such as firm behavior, international trade, and growth theory have used the CASIF data (e.g., Hsieh and Klenow (2009)).
+I focus on the five largest 2-digit sectors in terms of the number of firms between 2004 and 2007.
+Note that the CASIF dataset spans between 1998 and 2007.
+I choose year 2004 to 2007 to (i) ensure data consistency due to the change in the Chinese Industry Classification codes in 2003, (ii) avoid major structural breaks in the early 2000s (e.g., China joined the WTO in 2001), and (iii) use the most recent data.
 Following Brandt, Van Biesebroeck, and Zhang (2014), appropriate price deflators for inputs and outputs are applied separately.
 I preprocess the data so that firms with strictly positive amounts of capital, employment, value-added output, real wage expense, and real interest rate are used for estimation.
-The final dataset consists of a balanced panel of 11,567 firms over four years across five sectors.
-Summary statistics of the key variables are presented in Table 1.
+The final dataset consists of a balanced panel of 11,317 firms over four years across five sectors.
+Summary statistics are presented in Table 1.
 Table 1: Summary Statistics
 Variables N mean sd min max
-[…]
-[…]
-[…]
-[…]
-[…]
+$y_{it}=\ln$$\left(\text{value-added output}\right)$ 45,268 9.55 1.34 2.24 16.96
+$k_{it}=\ln$$\left(\text{capital}\right)$ 45,268 9.17 1.56 0.98 16.84
+$l_{it}=\ln$$\left(\text{labor}\right)$ 45,268 5.08 1.08 2.08 11.97
+$Z_{it,1}=\ln\overline{r}_{cs(i),t}^{\text{real},(-i)}$ 45,268 0.73 0.45 -6.02 6.94
+$Z_{it,2}=\ln\overline{w}_{cs(i),t}^{\text{real},(-i)}$ 45,268 2.68 0.37 0.33 5.28
 Year 4 - - 2004 2007
-Firm ID 11,567 - - - -
+Firm ID 11,317 - - - -
 Industry Code 5 - - - -
-The value-added production function I consider is
-$y_{it}$ $=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
-$\beta_{it}^{K}$ $=\beta^{K}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \beta_{it}^{L}=\beta^{L}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \omega_{it}=\omega\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),$
-$k_{it}$ $=g^{K}\left(Z_{it},A_{i},\eta_{it}^{K}\right),\ l_{it}=g^{L}\left(Z_{it},A_{i},\eta_{it}^{L}\right).$ (5.1)
+  • Notes:
+    (i) Output is measured as firm-level real value added in constant 1998 RMB (10,000 yuan).
+Capital is measured as firm-level real fixed capital stock (book value net of depreciation), deflated to constant 1998 RMB (10,000 yuan).
+Labor is measured as the number of employees (persons).
+The real interest rate is defined as real interest expenditures divided by total debt (percent).
+The real wage is average real salary per worker per year in constant 1998 RMB (10,000 yuan).
+    (ii) Sample sizes by industry are: textile (3,072), chemicals (2,131), nonmetallic mineral products (1,758), general equipment (2,828), and transportation equipment (1,528).
+    (iii) $Z_{it,1}$ denotes the natural logarithm of debt-weighted average real interest rate of competitors in the same city–industry–year.
+$Z_{it,2}$ denotes the natural logarithm of employment-weighted average real wage of competitors in the same city–industry–year.
+The value-added heterogeneous production function is
+$y_{it}$ $=k_{it}\beta_{it,K}+l_{it}\beta_{it,L}+\omega_{it}+\widetilde{\epsilon}_{it},$
+$\beta_{it,K}$ $=\beta_{K}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \beta_{it,L}=\beta_{L}\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),\ \omega_{it}=\omega\left(A_{i},\varepsilon_{it},\upsilon_{it}\right),$
+$k_{it}$ $=g_{K}\left(Z_{it},A_{i},\eta_{it,K}\right),\ l_{it}=g_{L}\left(Z_{it},A_{i},\eta_{it,L}\right).$ (5.1)
 I highlight two features of model (5.1).
-First, output elasticities $\beta_{it}\coloneqq\left(\beta_{it}^{K},\beta_{it}^{L}\right)^{{}^{\prime}}$ are allowed to differ across firms and through time.
-Second, input choices $X_{it}\coloneqq\left(k_{it},l_{it}\right)^{{}^{\prime}}$ can be correlated with $\beta_{it}$ through their dependence on the pairwise correlated random variables $A_{i}$, $\eta_{it}$, and $\varepsilon_{it}$.
-It is worth noting that the output is measured by the total revenue in dollars, rather than by the physical quantity produced in units, due to the absence of individual output prices in the data.
-When firms operate in distinct imperfectly competitive output markets, this may cause problems (Klette and Griliches (1996)).
-I use input prices $Z_{it}\coloneqq\left(r_{it},w_{it}\right)^{{}^{\prime}}$ as IVs for $X_{it}$.
-[^24] Real wages are very likely to be exogenous because the labor market in the manufacturing sectors of China is close to a perfectly competitive market in which firms face a relatively flat supply curve.
-Real interest rate may be considered as a valid IV because its fluctuation is mostly driven by monetary policy set by the central bank of China.
-For estimation, I take $W_{i}$ to be the mean through time of each coordinate of $X_{it}$.
-I estimate each coordinate of $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$ by regressing $\mathbf{\mathbbm{1}}\left(k\leq k_{it}\right)$ and $\mathbf{\mathbbm{1}}\left(l\leq l_{it}\right)$ on the second-degree polynomial spline basis of $\left(Z_{it},W_{i}\right)$ with its knot at the median, respectively.
-Next, I estimate $G_{it}\coloneqq\mathbb{E}\left[\left.y_{it}\right|X_{it},V_{it},W_{i}\right]$ by regressing $y_{it}$ on the same spline basis of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$.
-Estimation of $b_{1t}\left(V_{it},W_{i}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ is then obtained by taking the partial derivative of $\widehat{G}_{it}\left(X_{it},\widehat{V}_{it}W_{i}\right)$ with respect to $X_{it}$.
-Finally, I estimate $\overline{b}\coloneqq\mathbb{E}\beta_{it}$ by averaging $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ over $i$ and $t$ and $b_{t}\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by regressing $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the second-degree polynomial spline basis of $X_{it}$ with its knot at the median.
+First, output elasticities $\beta_{it}\coloneqq(\beta_{it,K},\beta_{it,L})^{\prime}$ are allowed to differ across firms and through time.
+Second, input choices $X_{it}\coloneqq(k_{it},l_{it})^{\prime}$ can be correlated with $\beta_{it}$ through their dependence on the pairwise correlated random variables $A_{i}$, $\eta_{it}$, and $\varepsilon_{it}$.
+I use leave-one-out industry–city–year cell averages of real wages and real interest rates faced by other firms in the same industry and city, weighted by competitors’ employment (for wages) and total debt (for interest rates) as IVs.
+I write them as $Z_{it}\coloneqq(\ln\overline{r}_{cs(i),t}^{\text{real},(-i)},\,\ln\overline{w}_{cs(i),t}^{\text{real},(-i)})^{\prime}$, where $cs(i)$ represents city and sector of firm $i$.
+These instruments are relevant because firms compete locally for labor and capital, so variation in competitors’ cost conditions shifts firm $i$’s effective input prices and hence its optimal input choices.
+At the same time, they are plausibly exogenous to firm $i$’s idiosyncratic productivity shocks and fixed effects because firm $i$’s own input prices are excluded from the construction; identification therefore relies on competitors’ cost variation rather than firm-level wage or interest-rate realizations.
+In Appendix B, I examine the exogeneity issue by using a more exogenous IV constructed at the provincial level, and the results remain stable.
+For estimation, I take $W_{i}$ to be the mean through time of each coordinate of $X_{it}$ by adapting the argument of Mundlak (1978) nonparametrically.
+I also examine the effect of further including the mean through time of each coordinate of $Z_{it}$ and the results reported in Table A.3 are similar.
+For all series estimation steps, I use second-degree polynomial spline basis functions with knots at the median, and a robustness check to this choice is included in Appendix B.
+First, I estimate each coordinate of $V_{it}\coloneqq(F_{\left.k_{it}\right|Z_{it},W_{i}}(\left.k_{it}\right|Z_{it},W_{i}),F_{\left.l_{it}\right|Z_{it},W_{i}}(\left.l_{it}\right|Z_{it},W_{i}))^{\prime}$ by regressing $\mathbf{\mathbbm{1}}(k_{it}\leq k)$ and $\mathbf{\mathbbm{1}}(l_{it}\leq l)$ on the basis functions of $(Z_{it},W_{i})$, respectively.
+Next, I estimate $G_{it}\coloneqq\mathbb{E}[\left.y_{it}\right|X_{it},V_{it},W_{i}]$ by regressing $y_{it}$ on the basis functions of $(X_{it},\widehat{V}_{it},W_{i})$.
+Estimation of $b_{1t}(V_{it},W_{i})\coloneqq\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]$ is then obtained by taking the partial derivative of $\widehat{G}_{it}(X_{it},\widehat{V}_{it},W_{i})$ with respect to $X_{it}$.
+Finally, I estimate the pooled APE $\overline{b}$ by simply averaging $\widehat{b}_{1t}(\widehat{V}_{it},W_{i})$ over $i$ and $t$.
+For the LAR, as in (4.7), I regress $\widehat{b}_{1t}(\widehat{V}_{it},W_{i})$ on the basis functions of $X_{it}$ for each $t$, then average over $t$ to obtain $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ as a measure of heterogeneity in output elasticities across firms.
 5.2 Results
-First, I summarize the results for estimating the average output elasticities defined as $\widehat{\overline{b}}\coloneqq\left(\widehat{\overline{b}}^{K},\widehat{\overline{b}}^{L}\right)^{{}^{\prime}}$ in Table 2.
-I compare my estimated $\widehat{\overline{b}}$’s (TERC) with those derived by applying the methods of Olley and Pakes (1996) (OP), Levinsohn and Petrin (2003) (LP), and Ackerberg, Caves, and Frazer (2015) (ACF) to the same data set.
-Standard errors (se) and 95% confidence intervals (CI) are also included.
-[^25] I use the Stata commands prodest (Rovigatti and Mollisi (2018)) for OP and LP and acfest (Manjón and Manez (2016)) for ACF for point estimates of $\overline{b}$ and 95% CI.
-See Keiller, de Paula, and Van Reenen (2024) for a discussion of the implementation of the ACF method.
-For TERC, since $\widehat{\overline{b}}$ has slower than $n^{1/2}$ rate by Theorem 2, I follow Politis, Romano, and Wolf (1999) to randomly subsample $b=\lfloor 4n^{3/4}\rfloor$ firms from each sector without replacement for 1,000 times.
+Average Output Elasticities
+I compare my TERC estimates of average output elasticities with those obtained from OP (Olley and Pakes (1996)), LP (Levinsohn and Petrin (2003)), and ACF (Ackerberg, Caves, and Frazer (2015)) applied to the same dataset.
+Table 2 contains the estimation results.
 Table 2: Estimates of Average Output Elasticities
 Textile OP LP ACF TERC
-Capital Elasticity 0.359 0.252 0.300 0.387
-se (0.030) (0.023) (0.037) (0.031)
-95% CI [0.300, 0.417] [0.206, 0.297] [0.227, 0.373] [0.334, 0.455]
-Labor Elasticity 0.470 0.175 0.567 0.375
-se (0.014) (0.011) (0.065) (0.023)
-95% CI [0.442, 0.498] [0.154, 0.196] [0.440, 0.695] [0.327, 0.418]
+Capital Elasticity 0.359 0.252 0.300 0.415
+95% CI [0.300, 0.417] [0.206, 0.297] [0.227, 0.373] [0.393, 0.441]
+Labor Elasticity 0.470 0.175 0.567 0.452
+95% CI [0.442, 0.498] [0.154, 0.196] [0.440, 0.695] [0.415, 0.485]
 Chemical OP LP ACF TERC
-Capital Elasticity 0.294 0.288 0.344 0.380
-se (0.036) (0.031) (0.096) (0.037)
-95% CI [0.223, 0.365] [0.228, 0.348] [0.155, 0.533] [0.306, 0.451]
-Labor Elasticity 0.296 0.113 0.378 0.329
-se (0.022) (0.015) (0.213) (0.026)
-95% CI [0.253, 0.339] [0.084, 0.143] [-0.039, 0.795] [0.276, 0.377]
+Capital Elasticity 0.294 0.288 0.344 0.463
+95% CI [0.223, 0.365] [0.228, 0.348] [0.155, 0.533] [0.430, 0.497]
+Labor Elasticity 0.296 0.113 0.378 0.311
+95% CI [0.253, 0.339] [0.084, 0.143] [-0.039, 0.795] [0.263, 0.361]
 Nonmetallic Mineral OP LP ACF TERC
-Capital Elasticity 0.697 0.311 0.236 0.379
-se (0.101) (0.024) (0.101) (0.037)
-95% CI [0.499, 0.895] [0.263, 0.358] [0.038, 0.434] [0.314, 0.451]
-Labor Elasticity 0.353 0.071 0.601 0.459
-se (0.021) (0.010) (0.384) (0.027)
-95% CI [0.311, 0.394] [0.051, 0.091] [-0.151, 1.354] [0.404, 0.514]
+Capital Elasticity 0.697 0.311 0.236 0.371
+95% CI [0.499, 0.895] [0.263, 0.358] [0.038, 0.434] [0.332, 0.410]
+Labor Elasticity 0.353 0.071 0.601 0.311
+95% CI [0.311, 0.394] [0.051, 0.091] [-0.151, 1.354] [0.255, 0.365]
 General Equipment OP LP ACF TERC
-Capital Elasticity 0.416 0.246 0.176 0.504
-se (0.076) (0.023) (0.082) (0.033)
-95% CI [0.267, 0.565] [0.202, 0.291] [0.016, 0.337] [0.451, 0.578]
-Labor Elasticity 0.444 0.071 0.927 0.359
-se (0.019) (0.009) (0.149) (0.020)
-95% CI [0.406, 0.482] [0.053, 0.089] [0.635, 1.218] [0.320, 0.397]
+Capital Elasticity 0.416 0.246 0.176 0.433
+95% CI [0.267, 0.565] [0.202, 0.291] [0.016, 0.337] [0.402, 0.469]
+Labor Elasticity 0.444 0.071 0.927 0.521
+95% CI [0.406, 0.482] [0.053, 0.089] [0.635, 1.218] [0.479, 0.564]
 Transportation Equipment OP LP ACF TERC
-Capital Elasticity 0.523 0.281 0.217 0.383
-se (0.064) (0.033) (0.084) (0.035)
-95% CI [0.397, 0.649] [0.216, 0.346] [0.052, 0.381] [0.327, 0.469]
-Labor Elasticity 0.523 0.137 1.042 0.460
-se (0.025) (0.019) (0.138) (0.028)
-95% CI [0.473, 0.573] [0.100, 0.173] [0.771, 1.313] [0.407, 0.517]
-My estimated average capital elasticities are within the range of $\left[0.379,0.504\right]$ across the five sectors, which are consistent with the results obtained from applying OP, LP, and ACF to the same dataset.
-The estimated average labor elasticities present more discrepancies between methods.
-My estimates of $\widehat{\overline{b}}^{L}$’s lie between $\left[0.329,0.460\right]$ across the five sectors, similar to OP’s results.
-LP seem to derive small $\widehat{\overline{b}}^{L}$’s for all sectors, while ACF generate larger $\widehat{\overline{b}}^{L}$’s for the general equipment sector and transportation equipment sector.
-My 95% CI’s for $\widehat{\overline{b}}$’s are reasonably tight across all sectors.
-It is well documented in the literature that output elasticities in Cobb-Douglas production function estimation usually lie within $\left[0,1\right]$.
-Hsieh and Klenow (2009) also calculate that roughly half of output is distributed to capital, according to Chinese input-output tables and national accounts.
-My estimates are consistent with the empirical evidence in the literature.
-Next, I investigate the distribution of the conditional means of the output elasticities given the regressors.
-Policymakers could use the information contained in the conditional means to formulate more accurate sector specific policies, as these conditional means provide a fine approximation to the average elasticities for each subgroup of firms with a certain level of capital and labor.
-Moreover, distributional properties about the true output elasticities can be inferred from the distribution of the conditional means of the output elasticities.
-For example, by the law of total variance, the variance of the conditional means of the output elasticities provides a lower bound for the variance of the true output elasticities.
+Capital Elasticity 0.523 0.281 0.217 0.425
+95% CI [0.397, 0.649] [0.216, 0.346] [0.052, 0.381] [0.393, 0.459]
+Labor Elasticity 0.523 0.137 1.042 0.588
+95% CI [0.473, 0.573] [0.100, 0.173] [0.771, 1.313] [0.536, 0.644]
+  • Notes: For TERC, I implement the three-step series estimator as described above.
+To compute the CI, I use a subsampling method to draw $b=\lfloor 4n_{s}^{3/4}\rfloor$ firms from each sector without replacement and repeat this 1,000 times.
+For the other methods, I use the Stata commands (prodest, Rovigatti and Mollisi (2018)) for OP and LP and (acfest, Manjón and Manez (2016)) for ACF.
+The $\widehat{\overline{b}}_{K}$ estimates from TERC fall within $\left[0.371,\,0.463\right]$ across the five sectors, broadly consistent with the results obtained from applying OP, LP, and ACF to the same dataset.
+In contrast, the $\widehat{\overline{b}}_{L}$ estimates exhibit greater discrepancies across methods.
+The TERC estimates of $\widehat{\overline{b}}_{L}$ lie in $\left[0.311,0.588\right]$ across the five sectors.
+OP’s labor elasticity estimates are close to TERC’s for the first three sectors but are smaller for the last two.
+LP seems to produce smaller $\widehat{\overline{b}}_{L}$ estimates across all sectors, whereas ACF generates larger $\widehat{\overline{b}}_{L}$ estimates for the general equipment and transportation equipment sectors.
+It is worth noting that the ACF approach tends to produce wider confidence intervals in certain sectors, which may partly reflect numerical instability in the Stata implementation, an issue also discussed by Keiller, de Paula, and Van Reenen (2024).
+The 95% CIs for TERC estimates are reasonably tight across all sectors.
+I consider the estimation results from TERC reasonable based on two pieces of empirical evidence.
+First, it is well documented in the literature that output elasticities in Cobb-Douglas production function estimation usually lie within $\left[0,1\right]$.
+Second, using Chinese manufacturing data, Hsieh and Klenow (2009) show that roughly half of output accrues to capital.
+Under profit maximization with a Cobb-Douglas production function, output elasticities correspond to factor revenue shares.
+The magnitudes of the estimated capital and labor elasticities can therefore be interpreted through this lens, and my estimates are consistent with this result.
+The broad similarity of average elasticities across methods is not unexpected, as OP, LP, ACF, and the TERC approach are all designed to address simultaneity between input choices and unobserved productivity.
+While OP, LP, and ACF provide consistent estimates under a constant-coefficient production function, TERC adds value by relaxing this restriction and allowing output elasticities to vary across firms while still addressing simultaneity via a control-function approach.
+Remark 4 (Addressing Simultaneity Issue).
+OP, LP, and ACF address simultaneity via a proxy-variable strategy: they invert a strictly monotone proxy input demand—investment in OP and intermediate inputs in LP/ACF—to recover a scalar productivity term $\omega_{it}$, then impose a Markov-style law of motion for $\omega_{it}$ to form identifying moments.
+This requires (i) strict monotonicity in $\omega_{it}$ and (ii) that $\omega_{it}$ is the only econometric unobservable entering the proxy demand, which rules out random coefficients that directly affect firms’ input choices.
+In contrast, my baseline TERC model allows input choices to correlate with heterogeneous output elasticities through an arbitrary-dimensional fixed effect $A_{i}$ and a time-varying shock $\eta_{it}$, with productivity modeled flexibly as $\omega_{it}=\omega(A_{i},\varepsilon_{it})$.
+Identification proceeds via a control-function strategy using a sufficient statistic $W_{i}$ for $A_{i}$ and a period-specific control $V_{it}$ for $\eta_{it}$, so that conditional on $(V_{it},W_{i})$ the remaining input variation is driven by exogenous instruments without taking a Markov law for $\omega_{it}$ as a primitive assumption.
+Cross-Firm Heterogeneity in Output Elasticities
+Next, I examine the distribution of conditional mean output elasticities given the regressors.
+These conditional means summarize average elasticities for firm subgroups defined by their capital and labor levels, and thus provide a useful basis for designing more targeted, sector-specific policies.
+They are also informative about the underlying distribution of true elasticities.
+For instance, by the law of total variance, the variance of these conditional means is a lower bound on the variance of the true output elasticities.
 Demirer (2025) argues that the heterogeneity in output elasticities is largely explained by across-firm variation.
-To investigate into this explanation, for each firm I calculate the averages of $\widehat{b}_{t}^{K}\left(X_{it}\right)$ and $\widehat{b}_{t}^{L}\left(X_{it}\right)$ through time and denote them by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$, respectively.
-Then, I plot the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across firms for each of the five sectors.
+To study this issue, I average $\widehat{b}_{t,K}(X_{it})$ and $\widehat{b}_{t,L}(X_{it})$ for each firm through time and denote them by $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$, respectively.
+These $\widehat{\beta}_{i}$ estimates can be considered as a proxy for their output elasticities.
+Then, I plot the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ across firms for each sector.
+In Figure 1, I present the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ for the textile sector and indicate the corresponding pooled APE labeled by “Mean” in the figure.
+The left subplot of Figure 1 corresponds to the histogram of $\widehat{\beta}_{i,K}$.
+All of its probability mass lies between zero and one, with its mode at 0.4.
+Over 95% of its probability mass lies between 0.3 and 0.5.
+The distribution of $\widehat{\beta}_{i,K}$ is concentrated around the mean and symmetric, suggesting that these firms tend to be homogeneous in capital efficiency.
+The right subplot of Figure 1 is the histogram of $\widehat{\beta}_{i,L}$.
+Again, the majority of its probability mass lies between zero and one.
+The distribution of $\widehat{\beta}_{i,L}$ is more dispersed than that of $\widehat{\beta}_{i,K}$ and is slightly left-skewed, with a small number of textile firms exhibiting low labor efficiency.
 Figure 1: Distribution of Conditional Means of Output Elasticities: Textile
-In Figure 2, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the textile sector and indicate the corresponding $\widehat{\overline{b}}$’s labeled by “Mean” on the same graph.
-The left subplot of Figure 2 shows the histogram of $\widehat{\beta}_{i}^{K}$.
-All of its probability mass lies between zero and one, with its mode at 0.4.
-Over 90% of the probability mass of $\widehat{\beta}_{i}^{K}$ lies between 0.3 and 0.5.
-The right subplot of Figure 2 presents the histogram of $\widehat{\beta}_{i}^{L}$ across firms for the textile sector.
-Again, the majority of its probability mass lies between zero and one, with its mode at 0.38.
-The distributions of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ are concentrated around the mean, suggesting that firms in the textile industry tend to be homogeneous in terms of capital and labor efficiency.
+Figure 2 presents the histograms of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ for the other sectors.
+I draw the following two conclusions.
+First, for all sectors, the majority of $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ lie between zero and one, consistent with the empirical evidence in Hsieh and Klenow (2009).
+Second, there is substantial across-firm variations in both capital and labor elasticities within each sector.
+The extent of heterogeneity, however, differs across elasticities and sectors.
+For example, in the left panel of Figure 2, capital elasticity exhibits greater dispersion among chemical firms than among firms in the nonmetallic minerals sector.
 Figure 2: Distribution of Conditional Means of Output Elasticities: Other Sectors
-In Figure 3, I present the histograms of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ for the other four sectors.
-First, the majority of $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ across all sectors lie between zero and one, which is consistent with empirical evidence.
-Second, I find significant across-firm variations in both capital and labor elasticities measured by $\widehat{\beta}_{i}^{K}$ and $\widehat{\beta}_{i}^{L}$ within each of the five sectors.
-However, the degree of heterogeneity in these elasticities differs.
-For example, in the left subplot of Figure 3, there is a larger variance in $\widehat{\beta}_{i}^{K}$ in the nonmetallic mineral sector than in the chemical sector, which implies that the firms in the chemical industry may be more homogeneous in capital efficiency than those in the nonmetallic mineral sector.
+Explaining the Heterogeneity
+I examine cross-firm heterogeneity in output elasticities by regressing the estimated $\widehat{\beta}_{i,K}$ and $\widehat{\beta}_{i,L}$ on firm characteristics.
+Specifically, using the CASIF dataset, I construct five covariates: firm size, leverage, export status, industrial-park location, and state ownership.
+All specifications include industry and city fixed effects, and I report heteroskedasticity-robust standard errors.
+Table 3: Explaining the Heterogeneity
+$\widehat{\beta}^{K}$ $\widehat{\beta}^{L}$
+Firm Size 0.0328 (0.0006) 0.0002 (0.0012)
+Leverage 0.0004 (0.0024) 0.0261 (0.0044)
+Export Status 0.0098 (0.0010) 0.0023 (0.0018)
+Industrial Park -0.0014 (0.0013) -0.0043 (0.0026)
+State Ownership 0.0050 (0.0040) -0.0534 (0.0076)
+  • Notes:
+    (i) Size is defined as the natural logarithm of total average assets.
+Leverage is measured as the debt-to-asset ratio.
+Exporting firms are defined as those with positive export delivery value.
+Industrial park firms are identified based on firm address information containing keywords such as “development zone” or “industrial park.” State-owned enterprises (SOEs) are defined as firms whose state capital accounts for more than 50 percent of total paid-in capital.
+    (ii) The variables are measured as follows: levels for the elasticities; the natural logarithm of 10,000 yuan for firm size; a ratio for leverage; and binary indicators (0/1) for the last three variables in the table.
+    (iii) The numbers in parentheses in the second row of each cell represent the standard errors.
+The results summarized in Table 3 show that larger firms have significantly higher capital elasticities, while leverage is positively associated with labor elasticities.
+Exporters exhibit modestly higher capital elasticities, and state-owned enterprises have substantially lower labor elasticities.
+Conditional on these controls, the industrial-park indicator is small.
+Industry and location fixed effects are jointly significant, indicating that sectoral and regional factors account for an important share of the heterogeneity.
+The adjusted-$R^{2}$s are 0.58 and 0.57 for $\widehat{\beta}^{K}$ and $\widehat{\beta}^{L}$, respectively, suggesting that over half of the variation in estimated elasticities can be explained by observable firm characteristics and fixed effects.
+The results are generally intuitive and broadly consistent with existing empirical evidence.
+Larger firms, for instance, often operate in more capital-intensive segments or adopt technologies with stronger scale economies, automation, and standardized production processes, so output tends to be more responsive to capital at the margin (Syverson (2011)).
+Leverage may also be associated with more labor-intensive operating choices—for example, delaying or outsourcing capital investment, relying on older or rented equipment, or placing greater emphasis on variable inputs (Kalemli-Özcan, Laeven, and Moreno (2022)).
+To the extent that highly leveraged firms substitute away from capital toward labor, the estimated production relationship will place greater weight on labor.
+Finally, exporters tend to use capital more efficiently (Bernard and Jensen (1999)), whereas state-owned enterprises may employ labor in excess of the efficient level due to political considerations (Wen (2025)), which is consistent with their substantially lower labor elasticities relative to private firms.
 6 Conclusion
-In this paper, I propose a new TERC model in which the regressors are correlated with the random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock—a feature that aligns with agent optimization behavior in many empirical applications.
-I construct feasible control variables for both the fixed effect and the random shock, and subsequently utilize the residual variation in the regressors given the control variables to identify the APE and LAR.
-I provide three-step series estimators and establish their convergence rates and asymptotic normality.
-The empirical exercise of this paper reveals significant variation in the output elasticities across manufacturing firms in China.
+This paper proposes a new TERC model in which regressors are correlated with random coefficients through not only a fixed effect but also a time-varying shock—an empirically relevant feature consistent with optimizing behavior in many applications.
+I construct feasible control variables for both the fixed effect and the time-varying shock, and use the resulting residual variation in regressors to identify the APE and LAR.
+I then develop three-step series estimators and establish their convergence rates and asymptotic normality.
+In an application to Chinese manufacturing, the estimates reveal substantial cross-firm dispersion in output elasticities, part of which is explained by observable firm characteristics.
 I propose two directions for future research.
-First, beyond the first- and second-order moments analyzed in this paper, policymakers may also be interested in the density functions of the random coefficients.
-One possible, albeit laborious, approach is to identify the moments of all orders of the random coefficients by induction, which would uniquely determine their distribution (Stoyanov (2000)).
-Second, it remains an open question whether aspects of my method could be extended to dynamic linear or nonlinear panel data models, such as those discussed by Marx, Tamer, and Tang (2024) and Liu, Poirier, and Shiu (2024).
-Such an extension would likely require additional structure on how the lagged dependent variable correlates with the random coefficients.
-Appendix A Proofs
+First, beyond the low-order moments studied here, policymakers may be interested in the full distribution of random coefficients.
+One—admittedly demanding—route is to identify moments of all orders by induction and recover the distribution via moment determinacy (Stoyanov (2000)).
+Second, it remains open whether elements of the approach can be extended to dynamic linear or nonlinear panel models, such as those in Marx, Tamer, and Tang (2024) and Liu, Poirier, and Shiu (2025).
+Doing so would likely require additional structure governing how lagged outcomes (or state variables) co-move with the random coefficients.
+Appendix
+Appendix A provides proofs for all theorems and lemmas in the main text, along with additional theoretical results.
+Appendix B reports additional empirical results.
+Appendix C presents the simulation results.
+Appendix A Proofs and Supplementary Theoretical Results
+Appendix A.1 contains proofs of all the lemmas and theorems in the main text.
+Appendix A.2 introduces Proposition 1, which provides sufficient conditions for Assumption 2 by adapting the nonparametric exchangeability condition of Altonji and Matzkin (2005).
+Appendix A.3 further discusses Assumptions 2, 4, and 4.
+A.1 Proofs
 Proof of Theorem 1.
 I first prove that $V_{it}$ is a control for $\eta_{it}$ given $A_{i}$ and $W_{i}$.
-Then, I show $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ does not depend on $X_{it}$ via the LIE.
+Then, I show $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ does not depend on $X_{it}$ via the law of iterated expectations.
 Finally, I identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by leveraging the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$.
-I show how the inclusion of exogenous shocks $\upsilon_{it}$ and $\epsilon_{it}$ as well as exogenous coordinates in the regressors affects the analysis at the end of this proof.
+I present how the inclusion of exogenous shocks $\upsilon_{it}$ and $\widetilde{\epsilon}_{it}$ as well as exogenous coordinates in the regressors affects the analysis at the end of this proof.
 Without loss of generality, I assume that $d_{\eta}=d_{X}$ and that each coordinate of $\eta_{it}$ enters the corresponding coordinate of $X_{it}$.
-[^26] This is without loss of generality because under Assumption 1, I can always redefine $\widetilde{\eta}$ to be the vector that collects the coordinates of $\eta$ that enters each coordinate of $g\left(\cdot\right)$ function.
+[^3] This is without loss of generality because under Assumption 1, I can always redefine $\widetilde{\eta}$ to be the vector that collects the coordinates of $\eta$ that enters each coordinate of $g$ function.
 By Assumption 2, I have $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$, which implies $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$.
-Thus, for each $l\in\left\{1,...,d_{X}\right\}$ and any on-support $\left(x^{\left(l\right)},z,a,w\right)$, I have
-$\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.x^{\left(l\right)}\right|z,w\right)$
-$=$ $\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.x^{\left(l\right)}\right|z,a,w\right)$
-$=$ $\ \mathbb{P}\left(\left.g^{\left(l\right)}\left(z,a,\eta_{it}^{\left(l\right)}\right)\leq x^{\left(l\right)}\right|Z_{it}=z,A_{i}=a,W_{i}=w\right)$
-$=$ $\ \mathbb{P}\left(\left.\eta_{it}^{\left(l\right)}\leq g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|A_{i}=a,W_{i}=w\right)$
-$=$ $\ F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|a,w\right),$ (A.1)
-where the first equality holds by $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$, the second uses (2.2), the third holds by Assumptions 1 and 3(a), and the last holds by definition of the conditional CDF of $\eta_{it}^{\left(l\right)}$ given $\left(A_{i},W_{i}\right)$.
-By (2.2), the random variable $\eta_{it}^{\left(l\right)}=g^{\left(l\right)^{-1}}\left(X_{it}^{\left(l\right)},Z_{it},A_{i}\right)$ for each $l$, so that plugging in gives
-$V_{it}^{\left(l\right)}\coloneqq F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}\right)=F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right),$ (A.2)
-which establishes a one-to-one mapping to $\eta_{it}^{\left(l\right)}$ given $A_{i}$ and $W_{i}$ under Assumption 3(b).
-Let $V_{it}\coloneqq\left(V_{it}^{\left(1\right)},...,V_{it}^{\left(d_{X}\right)}\right)^{{}^{\prime}}$.
+Thus, for each $l\in\left\{1,\ldots,d_{X}\right\}$ and any on-support $\left(x_{l},z,a,w\right)$, I have
+$\ F_{\left.X_{it,l}\right|Z_{it},W_{i}}\left(\left.x_{l}\right|z,w\right)$
+$=$ $\ F_{\left.X_{it,l}\right|Z_{it},A_{i},W_{i}}\left(\left.x_{l}\right|z,a,w\right)$
+$=$ $\ \mathbb{P}\left(\left.g_{l}\left(z,a,\eta_{it,l}\right)\leq x_{l}\right|Z_{it}=z,A_{i}=a,W_{i}=w\right)$
+$=$ $\ \mathbb{P}\left(\left.\eta_{it,l}\leq g_{l}^{-1}\left(x_{l},z,a\right)\right|A_{i}=a,W_{i}=w\right)$
+$=$ $\ F_{\left.\eta_{it,l}\right|A_{i},W_{i}}\left(\left.g_{l}^{-1}\left(x_{l},z,a\right)\right|a,w\right),$ (A.1)
+where the first equality holds by $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$, the second uses (2.2), the third holds by Assumptions 1 and 3(a), and the last holds by definition.
+By (2.2), the random variable $\eta_{it,l}=g_{l}^{-1}\left(X_{it,l},Z_{it},A_{i}\right)$ for each $l$, so that plugging in gives
+$V_{it,l}\coloneqq F_{\left.X_{it,l}\right|Z_{it},W_{i}}\left(\left.X_{it,l}\right|Z_{it},W_{i}\right)=F_{\left.\eta_{it,l}\right|A_{i},W_{i}}\left(\left.\eta_{it,l}\right|A_{i},W_{i}\right),$ (A.2)
+which establishes a one-to-one mapping to $\eta_{it,l}$ given $A_{i}$ and $W_{i}$ under Assumption 3(b).
+Let $V_{it}\coloneqq(V_{it,1},\dots,V_{it,d_{X}})^{\prime}$.
 By (A.2), $V_{it}$ uniquely determines the vector of $\eta_{it}$ given $A_{i}$ and $W_{i}$.
@@
 I exclude $X_{it}$ from the conditioning set of $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$ and write it out as $b_{2}\left(A_{i},V_{it},W_{i}\right)$.
-By (A.3), I have
+Then, by (A.3), I have
 $\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$
@@
 $\eqqcolon$ $\ b_{1}\left(V_{it},W_{i}\right),$ (A.4)
-where the first and second equalities hold by the LIE, the third holds because $V_{it}$ is a measurable function of $X_{it},\ Z_{it},$ and $W_{i}$, all of which are also conditioned on, and the fifth equality holds by $\left.\left(X_{it},Z_{it}\right)\perp A_{i}\right|W_{i}$ by Assumption 2.
+where the first and second equalities hold by the law of iterated expectations, the third holds because $V_{it}$ is a measurable function of $X_{it},\ Z_{it},$ and $W_{i}$, all of which are also conditioned on, and the fifth equality holds by $\left.\left(X_{it},Z_{it}\right)\perp A_{i}\right|W_{i}$ by Assumption 2.
 Finally, given (A.4), I have
-$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}b_{1}\left(V_{it},W_{i}\right).$ (A.5)
+$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{\prime}b_{1}\left(V_{it},W_{i}\right).$ (A.5)
 When Assumption 4 is satisfied, I pre-multiply both sides of (A.5) by $X_{it}$ and take conditional expectation of both sides conditioning on $V_{it}$ and $W_{i}$:
-$\mathbb{E}\left[\left.\mathbb{E}\left[\left.X_{it}Y_{it}\right|X_{it},V_{it},W_{i}\right]\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]b_{1}\left(V_{it},W_{i}\right),$
+$\mathbb{E}\left[\left.\mathbb{E}\left[\left.X_{it}Y_{it}\right|X_{it},V_{it},W_{i}\right]\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]b_{1}\left(V_{it},W_{i}\right),$
 which identifies $b_{1}\left(V_{it},W_{i}\right)$
-$b_{1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]\right)^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right].$
+$b_{1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]\right)^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right].$
 When Assumption 4 is used, I take partial derivative of both sides of (A.5) with respect to $X_{it}$ and obtain
 $b_{1}\left(V_{it},W_{i}\right)=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}.$
-Given $b_{1}\left(V_{it},W_{i}\right)$, I use the LIE to identify $\overline{b}=\mathbb{E}\beta_{it}$ and $b\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by
-$\overline{b}=\mathbb{E}\left[b_{1}\left(V_{it},W_{i}\right)\right]\text{ and }b\left(X_{it}\right)=\mathbb{E}\left[\left.b_{1}\left(V_{it},W_{i}\right)\right|X_{it}\right],$ (A.6)
+Given $b_{1}\left(V_{it},W_{i}\right)$, I use the law of iterated expectations to identify $\overline{b}$ and $b(X_{it})$ by
+$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\left[b_{1}\left(V_{it},W_{i}\right)\right]\text{ and }b\left(X_{it}\right)=\mathbb{E}\left[\left.b_{1}\left(V_{it},W_{i}\right)\right|X_{it}\right],$ (A.6)
 respectively.
-When ex-post shock $\epsilon_{it}$ is included additively in (2.1) and $\mathbb{E}\epsilon_{it}=0$, equations (A.4) and (A.5) still hold, so the identification result (A.6) holds without any changes.
-When ex-post shock $\upsilon_{it}$ is also included in $\beta_{it}\coloneqq\beta\left(A_{i},\eta_{it},\upsilon_{it}\right)$, I need to change $b_{2}\left(A_{i},V_{it},W_{i}\right)$ in (A.3) to be $b_{2t}\left(A_{i},V_{it},W_{i}\right)$ because although $\upsilon_{it}$ is independent of all other variables, the PDF of $\upsilon_{it}$ itself may depend on $t$, i.e.,
-$\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$
-$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|g\left(Z_{it},A_{i},\eta_{it}\right),A_{i},V_{it},W_{i}\right]$
-$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|A_{i},V_{it},W_{i}\right]$
-$\eqqcolon$ $\ b_{1{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t}}\left(A_{i},V_{it},W_{i}\right).$ (A.7)
-As a result, I need to change $b\left(X_{it}\right)$ to be $b_{t}\left(X_{it}\right)$ in (A.6). The rest of the proof goes through as before.
-When exogenous regressors are included in (2.1), I can define $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, and rewrite model (2.1)–(2.2) to be
-$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
+When ex-post shock $\widetilde{\epsilon}_{it}$ is included additively in (2.1) and $\mathbb{E}\widetilde{\epsilon}_{it}=0$, equations (A.4) and (A.5) still hold, so the identification result (A.6) holds without any changes.
+When ex-post shock $\upsilon_{it}$ is also included in $\beta_{it}\coloneqq\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$, since $b(X_{it})$ is already a time-varying function of $X_{it}$ due to the distribution of $(\varepsilon_{it},\eta_{it})$ being possibly time-varying, adding $\upsilon_{it}$ which is independent of everything else does not affect the identification proof.
+Finally, when exogenous regressors are included in (2.1), I let $X_{it}=\left(U_{it}^{\prime},Z_{it,1}^{\prime}\right)^{\prime}$ and $Z_{it}=\left(Z_{it,1}^{\prime},Z_{it,2}^{\prime}\right)^{\prime}$, and rewrite model (2.1)–(2.2) to be
+$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$
 $U_{it}$ $=g\left(Z_{it},A_{i},\eta_{it}\right).$
@@
 Note that I keep $X_{it}$ in Assumption 4 or 4 because I need to perturb the whole $X_{it}$ vector instead of just endogenous coordinates $U_{it}$ in (A.5) to identify $b_{1}\left(V_{it},W_{i}\right)$.
-As discussed in footnote 16, the residual variation requirement of Assumption 4 or 4 is easier to be satisfied when exogenous $Z_{1,it}$ is included in $X_{it}$ as the support of $Z_{1,it}$ is unaffected by $V_{it}$ and $W_{i}$.
+As discussed in Appendix A.3, the residual variation requirement of Assumption 4 or 4 is easier to be satisfied when exogenous $Z_{it,1}$ is included in $X_{it}$ as the support of $Z_{it,2}$ is unaffected by $V_{it}$ and $W_{i}$.
 ∎
@@
 As the result for a finite-dimensional vector-valued $\beta$ can be established by proving for each of its coordinates and combining the results using the triangle inequality, I assume $\beta$ is a scalar in this proof.
-I focus on $\widehat{b}\left(x\right)$, since the result for $\widehat{\overline{b}}$ follows immediately by setting $r^{M_{3}}\left(\cdot\right)\equiv 1$.
+I focus on $\widehat{b}\left(x\right)$, since the result for $\widehat{\overline{b}}$ follows immediately by setting $r^{M_{3}}\equiv 1$.
 Let $p_{i}^{m_{2}}\coloneqq p^{m_{2}}\left(V_{i},W_{i}\right)$, $r_{i}\coloneqq r^{M_{3}}\left(X_{i}\right)$, and $\overline{p}_{i}=\overline{p}^{M_{2}}\left(S_{i}\right)$.
-Following Imbens and Newey (2009), I can normalize $\mathbb{E}p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}=I_{m_{2}}$ by Assumption 5 and $R\coloneqq\mathbb{E}r_{i}r_{i}^{{}^{\prime}}=I_{M_{3}}$ by Assumption 6, which imply $\overline{P}\coloneqq\mathbb{E}\overline{p}_{i}\overline{p}_{i}^{{}^{\prime}}=\mathbb{E}\left[I_{d_{X}}\otimes\left(p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}\right)\right]=I_{M_{2}}$.
+Following Imbens and Newey (2009), I can normalize $\mathbb{E}p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{\prime}=I_{m_{2}}$ by Assumption 5 and $R\coloneqq\mathbb{E}r_{i}r_{i}^{\prime}=I_{M_{3}}$ by Assumption 6, which imply $\overline{P}\coloneqq\mathbb{E}\overline{p}_{i}\overline{p}_{i}^{\prime}=\mathbb{E}\left[I_{d_{X}}\otimes\left(p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{\prime}\right)\right]=I_{M_{2}}$.
 Furthermore, I have $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ with probability approaching one by Newey (1997).
-Let $\widetilde{B}\coloneqq\left(b_{1}\left(\widehat{v}_{1},w_{1}\right),...,b_{1}\left(\widehat{v}_{n},w_{n}\right)\right)^{{}^{\prime}}$.
+Let $\widetilde{B}\coloneqq\left(b_{1}\left(\widehat{v}_{1},w_{1}\right),\ldots,b_{1}\left(\widehat{v}_{n},w_{n}\right)\right)^{\prime}$.
 By (4.7),
 $\ \left\|n\widehat{R}^{1/2}\left(\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right)\right\|^{2}/4$
-$\leq$ $\ \left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)+\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
-$\ +\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)+\left(B^{X}-r\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-r\rho^{M_{3}}\right),$ (A.8)
-where $B\coloneqq\left(b_{1}\left(v_{1},w_{1}\right),...,b_{1}\left(v_{n},w_{n}\right)\right)^{{}^{\prime}}$ and $B^{X}\coloneqq\left(b\left(x_{1}\right),...,b\left(x_{n}\right)\right)^{{}^{\prime}}$.
+$\leq$ $\ \left(\widehat{B}-\widetilde{B}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widehat{B}-\widetilde{B}\right)+\left(\widetilde{B}-B\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)$
+$\ +\left(B-B^{X}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B-B^{X}\right)+\left(B^{X}-R\rho^{M_{3}}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B^{X}-r\rho^{M_{3}}\right),$ (A.7)
+where $B\coloneqq\left(b_{1}\left(v_{1},w_{1}\right),\ldots,b_{1}\left(v_{n},w_{n}\right)\right)^{\prime}$ and $B^{X}\coloneqq\left(b\left(x_{1}\right),\ldots,b\left(x_{n}\right)\right)^{\prime}$.
 Hence, $\xi=B-B^{X}$.
 I analyze the RHS of (A.7) term by term.
-By Lemma S.5 of Imbens and Newey (2009), $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$. Then,
-$\ n^{-2}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)\leq Cn^{-1}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
-$=$ $\ Cn^{-1}\sum_{i}\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)\right)^{2}$
-$\leq$ $\ C\left\|\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right\|^{2}+C\sup_{s\in\mathcal{S}}\left\|\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(v,w\right)\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}\right),$ (A.9)
-where the first inequality holds because $n^{-1}r\widehat{R}^{-1}r^{{}^{\prime}}$ is idempotent, the last inequality holds by the Cauchy-Schwarz inequality (CS) and $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$, and the last equality uses Lemma 1.
+By Lemma S.5 of Imbens and Newey (2009), $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{\prime}-I\right\|=o_{p}\left(1\right)$. Then,
+[…]
+$=$ $\ Cn^{-1}\sum_{i}\left(\widehat{\overline{p}}_{i}^{\prime}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)+\left(\widehat{\overline{p}}_{i}^{\prime}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)\right)^{2}$
+$\leq$ $\ C\left\|\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right\|^{2}+C\sup_{s\in\mathcal{S}}\left\|\overline{p}^{M_{2}}\left(s\right)^{\prime}\alpha^{M_{2}}-b_{1}\left(v,w\right)\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}\right),$ (A.8)
+where the first inequality holds because $n^{-1}r\widehat{R}^{-1}r^{\prime}$ is idempotent, the last inequality holds by the Cauchy–Schwarz inequality (CS) and $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{\prime}-I\right\|=o_{p}\left(1\right)$, and the last equality uses Lemma 1.
 Next,
-$n^{-2}\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)\leq$ $\ Cn^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
+$n^{-2}\left(\widetilde{B}-B\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)\leq$ $\ Cn^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
 $\leq$ $\ Cn^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}=O_{p}\left(\Delta_{1n}^{2}\right),$ (A.9)
@@
 Finally, for the last two terms on the right-hand side of (A.7),
-$\ n^{-2}\mathbb{E}\left[\left.\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)\right|{\bf X}\right]$
-$=$ $\ n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\xi\right|{\bf X}\right]\right\}=n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\widehat{R}^{-1}r^{{}^{\prime}}\right\}$
-$\leq$ $\ n^{-2}\text{tr}\left\{CIr\widehat{R}^{-1}r^{{}^{\prime}}\right\}=Cn^{-1}\text{tr}\left\{\widehat{R}^{-1}\widehat{R}\right\}=Cn^{-1}M_{3},$
+$\ n^{-2}\mathbb{E}\left[\left.\left(B-B^{X}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B-B^{X}\right)\right|{\bf X}\right]$
+$=$ $\ n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi^{\prime}r\widehat{R}^{-1}r^{\prime}\xi\right|{\bf X}\right]\right\}=n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]r\widehat{R}^{-1}r^{\prime}\right\}$
+$\leq$ $\ n^{-2}\text{tr}\left\{CIr\widehat{R}^{-1}r^{\prime}\right\}=Cn^{-1}\text{tr}\left\{\widehat{R}^{-1}\widehat{R}\right\}=Cn^{-1}M_{3},$
 and
-$n^{-2}\left(B^{X}-R\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-R\rho^{M_{3}}\right)\leq$ $\ n^{-1}\left\|B^{X}-R\rho^{M_{3}}\right\|^{2}=O_{p}\left(M_{3}^{-2d_{3}/j_{3}}\right).$ (A.11)
-By $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ and the result that $\mathbb{E}\left[\left.\left|Y_{n}\right|\right|Z_{n}\right]=O_{p}\left(r_{n}\right)$ implies $Y_{n}=O_{p}\left(r_{n}\right)$ (Conditional Markov’s inequality, henceforth denoted as “CM”),
+$n^{-2}\left(B^{X}-R\rho^{M_{3}}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B^{X}-R\rho^{M_{3}}\right)\leq$ $\ n^{-1}\left\|B^{X}-R\rho^{M_{3}}\right\|^{2}=O_{p}\left(M_{3}^{-2d_{3}/j_{3}}\right).$ (A.10)
+By $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ and conditional Markov’s (CM) inequality (i.e., $\mathbb{E}\left[\left.\left|Y_{n}\right|\right|Z_{n}\right]=O_{p}\left(r_{n}\right)$ implies $Y_{n}=O_{p}\left(r_{n}\right)$),
 $\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}+n^{-1}M_{3}+M_{3}^{-2d_{3}/j_{3}}\right)\eqqcolon O_{p}\left(\Delta_{3n}^{2}\right),$ (A.11)
 which yields
-[…]
-$\leq$ $\ 2\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}+2\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|^{2}=O_{p}\left(\Delta_{3n}^{2}\right),$
+$\int\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|^{2}dF\left(x\right)\leq$ $\int\left(r^{M_{3}}\left(x\right)^{\prime}\left(\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right)+\left(r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}-b\left(x\right)\right)\right)^{2}dF\left(x\right)$
+$\leq$ $\ 2\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}+2\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}\right|^{2}=O_{p}\left(\Delta_{3n}^{2}\right),$
 and
-$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $\leq\sup_{x\in\mathcal{X}}\left\|r^{M_{3}}\left(x\right)\right\|\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|+\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|$
+$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $\leq\sup_{x\in\mathcal{X}}\left\|r^{M_{3}}\left(x\right)\right\|\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|+\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}\right|$
 $=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
@@
 Define
-$\Omega_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}P^{-1}\left(\Sigma+\Sigma_{1}\right)P^{-1}\overline{p}^{M_{2}}\left(v,w\right),\ \Sigma\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2},$ (A.13)
+$\Omega_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{\prime}P^{-1}\left(\Sigma+\Sigma_{1}\right)P^{-1}\overline{p}^{M_{2}}\left(v,w\right),\ \Sigma\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}u_{i}^{2},$ (A.12)
 $\Sigma_{1}$ $\coloneqq\mathbb{E}\overline{\mu}_{i}^{I}\overline{\mu}_{i}^{I^{\prime}},\ p_{i}\coloneqq p^{M_{2}}\left(S_{i}\right),\ q_{i}\coloneqq q^{M_{1}}\left(X_{i},Z_{i},W_{i}\right),$
-$\overline{\mu}_{i}^{I}$ $\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau^{{}^{\prime}}\left(V_{j}\right)p_{j}q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\ u_{i}\coloneqq Y_{i}-G\left(S_{i}\right),$
+$\overline{\mu}_{i}^{I}$ $\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau_{V}\left(V_{j}\right)p_{j}q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\ u_{i}\coloneqq Y_{i}-G\left(S_{i}\right),$
 $G_{V}\left(S_{j}\right)$ $\coloneqq\left.\partial G\left(s\right)/\partial v\right|_{s=S_{j}},\text{ and }v_{ji}\coloneqq\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-F\left(\left.x_{j}\right|z_{i},w_{i}\right).$
 and
-$\widehat{\Omega}_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}\widehat{P}^{-1}\left(\widehat{\Sigma}+\widehat{\Sigma}_{1}\right)\widehat{P}^{-1}\overline{p}^{M_{2}}\left(v,w\right),$ (A.14)
-$\widehat{\Sigma}$ $\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\left(y_{i}-\widehat{G}\left(\widehat{s}_{i}\right)\right)^{2},\ \widehat{\Sigma}_{1}\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{\overline{\mu}}_{i}^{I}\widehat{\overline{\mu}}_{i}^{I^{\prime}},$
-$\widehat{\overline{\mu}}_{i}^{I}$ $\coloneqq n^{-1}\sum_{j=1}^{n}\widehat{G}_{V}\left(\widehat{s}_{j}\right)\widehat{p}_{j}q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }\widehat{v}_{ji}\coloneqq\left(\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-\widehat{F}\left(\left.x_{j}\right|z_{i},w_{i}\right)\right).$
+$\widehat{\Omega}_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{\prime}\widehat{P}^{-1}\left(\widehat{\Sigma}+\widehat{\Sigma}_{1}\right)\widehat{P}^{-1}\overline{p}^{M_{2}}\left(v,w\right),$ (A.13)
+$\widehat{\Sigma}$ $\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{p}_{i}\widehat{p}_{i}^{\prime}\left(y_{i}-\widehat{G}\left(\widehat{s}_{i}\right)\right)^{2},\ \widehat{\Sigma}_{1}\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{\overline{\mu}}_{i}^{I}\widehat{\overline{\mu}}_{i}^{I}{}^{\prime},$
+$\widehat{\overline{\mu}}_{i}^{I}$ $\coloneqq n^{-1}\sum_{j=1}^{n}\widehat{G}_{V}\left(\widehat{s}_{j}\right)\widehat{p}_{j}q_{j}^{\prime}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }\widehat{v}_{ji}\coloneqq\left(\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-\widehat{F}\left(\left.x_{j}\right|z_{i},w_{i}\right)\right).$
 IN02 have proved asymptotic normality for known and scalar-valued functionals of $G\left(s\right)$.
-I apply their results to $c^{{}^{\prime}}\widehat{b}_{1}\left(v,w\right)$ for any constant vector $c^{{}^{\prime}}c=1$ and obtain
-[…]
-[…]
+I apply their results to $c^{\prime}\widehat{b}_{1}\left(v,w\right)$ for any constant vector $c^{\prime}c=1$ and obtain
+$c^{\prime}\sqrt{n}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$ $\overset{d}{\longrightarrow}N\left(0,1\right)\ \text{and}$
+$\left(c^{\prime}\Omega_{1}c\right)^{-1}\left[c^{\prime}\left(\widehat{\Omega}_{1}-\Omega_{1}\right)c\right]$ $\overset{p}{\longrightarrow}0.$ (A.14)
 By (A.14) and Assumption 7(f),
-[…]
+$c^{\prime}\left(c_{1n}\widehat{\Omega}_{1}-c_{1n}\Omega_{1}\right)c\overset{p}{\longrightarrow}0,$ (A.15)
 which implies
-[…]
+$c_{1n}\widehat{\Omega}_{1}\overset{p}{\longrightarrow}\overline{\Omega}_{1}.$ (A.16)
 Then,
-$\ n^{1/2}\widehat{\Omega}_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
-$=$ $\ \left(c_{1n}\widehat{\Omega}_{1}\right)^{-1/2}\left(c_{1n}\Omega_{1}\right)^{1/2}n^{1/2}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
-[…]
+$\ \sqrt{n}\widehat{\Omega}_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
+$=$ $\ \left(c_{1n}\widehat{\Omega}_{1}\right)^{-1/2}\left(c_{1n}\Omega_{1}\right)^{1/2}\sqrt{n}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
+$\overset{d}{\longrightarrow}$ $\ \overline{\Omega}_{1}^{-1/2}\overline{\Omega}_{1}^{1/2}N\left(0,I\right)=_{d}N\left(0,I\right),$ (A.17)
 where the convergence holds by (A.14), the Cramér–Wold device, (A.16), Assumption 7, and the continuous mapping theorem. ∎
 Proof of Theorem 3.
-Following the proof of Lemma 2, under regularity conditions one may use the Cramér–Wold device to extend the results from scalar-valued functionals to vector-valued functionals of $G\left(s\right)$.
-Therefore, I maintain that $b\left(x\right)$ is a scalar in this proof.
+Similarly to the proof of Lemma 2, by the Cramér–Wold device it suffices to consider the case when $b\left(x\right)$ is a scalar.
 First, I derive the influence functions for $\widehat{b}\left(x\right)$ that correctly account for the estimation errors from each step and prove its asymptotic normality.
-Then, I show consistency for the estimator of the variance of $\widehat{b}\left(x\right)$, which can be used to construct asymptotically valid confidence intervals.
+Then, I show consistency for the estimator of the variance of $\widehat{b}\left(x\right)$.
 I write $r^{M_{3}}\left(x\right)$ as $r\left(x\right)$ when there is no confusion.
@@
 Define
-$\Omega_{2}$ $\coloneqq\Omega_{21}+\Omega_{22},\ \Omega_{21}\coloneqq\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)\left(A_{1}P^{-1}p_{i}u_{i}\right)^{{}^{\prime}},$ (A.19)
-$\Omega_{22}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{{}^{\prime}},$
-$A_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}},\ \overline{\mu}_{i}^{II}\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\text{ and}$
-$b_{1,V}\left(V_{j},W_{j}\right)$ $\coloneqq\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j},w=W_{j}},$
-and $F\coloneqq\Omega_{2}^{-1/2}$, which is well-defined because $\Omega_{2}=\Omega_{21}+\Omega_{22}$ and
-$\Omega_{21}$ $=A_{1}P^{-1}\left(\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2}\right)P^{-1}A_{1}^{{}^{\prime}}$
-$=A_{1}P^{-1}\left(\mathbb{E}\left[p_{i}p_{i}^{{}^{\prime}}\mathbb{E}\left(\left.u_{i}^{2}\right|X_{i},Z_{i},W_{i}\right)\right]\right)P^{-1}A_{1}^{{}^{\prime}}$
-$\geq CA_{1}P^{-1}A_{1}^{{}^{\prime}}=Cr\left(x\right)^{{}^{\prime}}\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)P^{-1}\left(\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}\right)r\left(x\right)>0,$ (A.20)
+$\Omega_{2}$ $\coloneqq\Omega_{21}+\Omega_{22},\text{ where}$
+$\Omega_{21}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)\left(A_{1}P^{-1}p_{i}u_{i}\right)^{\prime},$
+$\Omega_{22}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r\left(x\right)^{\prime}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r\left(x\right)^{\prime}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{\prime},$
+$A_{1}$ $\coloneqq r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime},$
+$\overline{\mu}_{i}^{II}$ $\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\text{ and}$
+$b_{1,V}\left(V_{j},W_{j}\right)$ $\coloneqq\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j},w=W_{j}}.$ (A.18)
+Let $F\coloneqq\Omega_{2}^{-1/2}$, which is well-defined because $\Omega_{2}=\Omega_{21}+\Omega_{22}$ and
+$\Omega_{21}$ $=A_{1}P^{-1}\left(\mathbb{E}p_{i}p_{i}^{\prime}u_{i}^{2}\right)P^{-1}A_{1}^{\prime}$
+$=A_{1}P^{-1}\left(\mathbb{E}\left[p_{i}p_{i}^{\prime}\mathbb{E}\left(\left.u_{i}^{2}\right|X_{i},Z_{i},W_{i}\right)\right]\right)P^{-1}A_{1}^{\prime}$
+$\geq CA_{1}P^{-1}A_{1}^{\prime}=Cr\left(x\right)^{\prime}\left(\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right)P^{-1}\left(\mathbb{E}\overline{p}_{i}r_{i}^{\prime}\right)r\left(x\right)>0,$ (A.19)
 where the first inequality holds by Assumption 7(c) and the last inequality holds by Assumption 8(a).
 Define functionals:
-$\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.\widehat{b}_{1}\left(\widehat{V},W\right)\right|X=x\right]=\widehat{b}\left(x\right),$
-$\widehat{a}\left(b_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.b_{1}\left(\widehat{V},W\right)\right|X=x\right],$
+[…]
+[…]
 $\widehat{a}\left(b_{1},V\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.b_{1}\left(V,W\right)\right|X=x\right],\text{ and}$
@@
 I expand
-$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$
-$=$ $\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)+\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)+\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
+$\ \sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$
+$=$ $\ \sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)+\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)+\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
 $=$ $\ n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.20)
@@
 $\psi_{1i}$ $=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right),\ \psi_{2i}=H_{2}\overline{\mu}_{i}^{II},\text{ and }\psi_{3i}=H_{2}r_{i}\xi_{i},$ (A.21)
+where $H_{1}\coloneqq FA_{1}P^{-1}$, $A_{1}=r\left(x\right)^{\prime}R^{-1}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}=r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}$, $r_{i}\coloneqq r\left(x_{i}\right)$, $P\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}$, $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$, $u_{i}\coloneqq y_{i}-G\left(s_{i}\right)$, $\text{$\overline{\mu}_{i}^{I}$}\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau_{V}\left(V_{j}\right)p_{j}q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],$ $G_{V}\left(S_{j}\right)\coloneqq\left.\partial G\left(s\right)/\partial v\right|_{s=S_{j}},$ $v_{ji}\coloneqq\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-F\left(\left.x_{j}\right|z_{i},w_{i}\right)$, $H_{2}\coloneqq FA_{2}R^{-1}=FA_{2}$, $A_{2}\coloneqq r\left(x\right)^{\prime}$, $\overline{\mu}_{i}^{II}\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right]$, $b_{1,V}\left(V_{j},W_{j}\right)=\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j}},$ and $\xi_{i}=b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)$.
+Let $\widehat{H}_{1}\coloneqq F\widehat{A}_{1}\widehat{P}^{-1}$, $\widehat{A}_{1}\coloneqq r\left(x\right)^{\prime}\widehat{R}^{-1}n^{-1}\sum_{i=1}^{n}r_{i}\widehat{\overline{p}}_{i}^{\prime}$, $\widehat{R}\coloneqq n^{-1}\sum_{i=1}^{n}r_{i}r_{i}^{\prime}$, $\widehat{\overline{p}}_{i}\coloneqq\overline{p}\left(\widehat{s}_{i}\right)$, $\widehat{H}_{2}\coloneqq FA_{2}\widehat{R}^{-1}$, $G\coloneqq\left(G\left(s_{1}\right),\ldots,G\left(s_{n}\right)\right)^{\prime}$, and $\widetilde{G}\coloneqq\left(G\left(\widehat{s}_{1}\right),\ldots,G\left(\widehat{s}_{n}\right)\right)^{\prime}$.
+First, for $\psi_{1i}$,
 […]
-Let $\widehat{H}_{1}\coloneqq F\widehat{A}_{1}\widehat{P}^{-1}$, $\widehat{A}_{1}\coloneqq r\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}n^{-1}\sum_{i=1}^{n}r_{i}\widehat{\overline{p}}_{i}$, $\widehat{R}\coloneqq n^{-1}\sum_{i=1}^{n}r_{i}r_{i}^{{}^{\prime}}$, $\widehat{\overline{p}}_{i}\coloneqq\overline{p}\left(\widehat{s}_{i}\right)$, $\widehat{H}_{2}\coloneqq FA_{2}\widehat{R}^{-1}$, $G\coloneqq\left(G\left(s_{1}\right),...,G\left(s_{n}\right)\right)^{{}^{\prime}}$, and $\widetilde{G}\coloneqq\left(G\left(\widehat{s}_{1}\right),...,G\left(\widehat{s}_{n}\right)\right)^{{}^{\prime}}$.
-First, for $\psi_{1i}$,
-$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)\right)$
-$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
-$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}Y-\widetilde{B}\right)$
-$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left[n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}\left(Y-G+G-\widetilde{G}+\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)+\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
-$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{1}\widehat{p}_{i}\left[u_{i}-\left(G\left(\widehat{s}_{i}\right)-G\left(s_{i}\right)\right)\right]+n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)$
-$\ +n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)$ (A.23)
+[…]
+[…]
+[…]
+$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{1}\widehat{p}_{i}\left[u_{i}-\left(G\left(\widehat{s}_{i}\right)-G\left(s_{i}\right)\right)\right]+n^{-1/2}\widehat{H}_{1}\widehat{p}^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)$
+[…]
 $\eqqcolon$ $\ D_{11}+D_{12}+D_{13}.$ (A.22)
@@
 $\left\|H_{1}\right\|=O\left(1\right)\text{ and }\left\|H_{2}\right\|=O\left(1\right),$ (A.24)
-because $\left\|H_{1}\right\|^{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$ and $\left\|H_{2}\right\|^{2}=A_{2}A_{2}^{{}^{\prime}}/\Omega_{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$.
-In addition, $\left\|\widehat{P}-P\right\|=o_{p}\left(1\right)$, $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and $\left\|n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}-\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right\|=o_{p}\left(1\right)$ by Lemma A3 of IN02 and Newey (1997).
-I also know that $\lambda_{\min}\left(P\right)\geq C>0$ and $\left\|P\right\|=O\left(1\right)$ by Assumption 5.
+because $\left\|H_{1}\right\|^{2}\leq CA_{1}A_{1}^{\prime}/\Omega_{2}\leq C$ and $\left\|H_{2}\right\|^{2}=A_{2}A_{2}^{\prime}/\Omega_{2}\leq CA_{1}A_{1}^{\prime}/\Omega_{2}\leq C$.
+In addition, $\left\|\widehat{P}-P\right\|=o_{p}\left(1\right)$, $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and $\left\|n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}^{\prime}-\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right\|=o_{p}\left(1\right)$ by Lemma A3 of IN02 and Newey (1997).
+Furthermore, $\lambda_{\min}\left(P\right)\geq C>0$ and $\left\|P\right\|=O\left(1\right)$ by Assumption 5.
 By Slutsky’s theorem, $\left\|\widehat{R}^{-1}-I\right\|=o_{p}\left(1\right)$ and $\left\|\widehat{P}^{-1}-P^{-1}\right\|=o_{p}\left(1\right)$.
 By the CS inequality and Lemma A3 of IN02,
-$\left\|n^{-1}\sum_{i}r_{i}\left(\overline{p}_{i}-\widehat{\overline{p}}_{i}\right)^{{}^{\prime}}\right\|^{2}$ $\leq n^{-1}\sum_{i}\left\|r_{i}\right\|^{2}\times n^{-1}\sum_{i}\left\|\widehat{\overline{p}}_{i}-\overline{p}_{i}\right\|^{2}$
+$\left\|n^{-1}\sum_{i}r_{i}\left(\overline{p}_{i}-\widehat{\overline{p}}_{i}\right)^{\prime}\right\|^{2}$ $\leq n^{-1}\sum_{i}\left\|r_{i}\right\|^{2}\times n^{-1}\sum_{i}\left\|\widehat{\overline{p}}_{i}-\overline{p}_{i}\right\|^{2}$
 $=O_{p}\left(M_{3}\zeta_{1}\left(m_{2}\right)^{2}\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.25)
@@
 $\leq$ $\ 2\left\|F\left(\widehat{A}_{1}-A_{1}\right)\widehat{P}^{-1}\right\|^{2}+2\left\|FA_{1}\left(\widehat{P}^{-1}-P^{-1}\right)\right\|^{2}$
-$=$ $\ 2\left\|F\left(r\left(x\right)^{{}^{\prime}}\left(I+o_{p}\left(1\right)\right)\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}+o_{p}\left(1\right)\right)-r\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)\widehat{P}^{-1}\right\|^{2}$
+$=$ $\ 2\left\|F\left(r\left(x\right)^{\prime}\left(I+o_{p}\left(1\right)\right)\left(\mathbb{E}r_{i}\overline{p}_{i}^{\prime}+o_{p}\left(1\right)\right)-r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right)\widehat{P}^{-1}\right\|^{2}$
 $\ +2\left\|FA_{1}P^{-1}\left(P-\widehat{P}\right)\widehat{P}^{-1}\right\|^{2}$
@@
 Next, recall that by Assumption 7
-$n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right),\text{ and}$
-$n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right).$
-Therefore, for $D_{12}$,
-$\left|n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{1}\widehat{P}\widehat{H}_{1}^{{}^{\prime}}\right]\left[n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right]$
+$n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right)\text{ and}$
+[…]
+Therefore, for $D_{12}$, I have
+$\left|n^{-1/2}\widehat{H}_{1}\widehat{p}^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{1}\widehat{P}\widehat{H}_{1}^{\prime}\right]\left[n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right]$
 $\leq\left\|\widehat{H}_{1}\right\|^{2}O_{p}\left(nM_{2}^{-2d_{4}/j_{4}}\right)=o_{p}\left(1\right).$ (A.27)
-For $D_{13}$, similarly to (A.29),
-$\left|n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{2}\widehat{R}\widehat{H}_{2}\right]\left[n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
+For $D_{13}$, similarly to (A.27), I have
+[…]
 $\leq\left\|\widehat{H}_{2}\right\|^{2}O_{p}\left(nM_{2}^{-2d_{4}/j_{4}}\right)=o_{p}\left(1\right).$ (A.28)
@@
 $\psi_{1i}=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right).$ (A.29)
-To prove $\psi_{2i}=H_{2}\overline{\mu}_{i}^{II}$, first notice that by Taylor expansion,
-$\ n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)$
-$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
+To prove $\psi_{2i}=H_{2}\overline{\mu}_{i}^{II}$, by Taylor expansion it is true that
+$\ \sqrt{n}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)$
+$=$ $\ n^{-1/2}Fr\left(x\right)^{\prime}\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)$
 $=$ $\ \widehat{H}_{2}n^{-1/2}\sum_{i}r_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)$
@@
 where
-$\delta_{ij}=F_{\left.X_{i}\right|Z_{j},W_{j}}\left(\left.x_{i}\right|z_{j},w_{j}\right)-q_{j}^{{}^{\prime}}\gamma^{M_{1}}\left(x_{i}\right),\ \Delta_{i}^{I}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}v_{ij},$
-$\Delta_{i}^{II}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}\delta_{ij},\text{ and }\Delta_{i}^{III}=-\delta_{ii}.$ (A.34)
-Following an argument identical to the proof of Lemma B7 of IN02,
+$\delta_{ij}=F_{\left.X_{i}\right|Z_{j},W_{j}}\left(\left.x_{i}\right|z_{j},w_{j}\right)-q_{j}^{\prime}\gamma^{M_{1}}\left(x_{i}\right),\ \Delta_{i}^{I}=q_{i}^{\prime}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}v_{ij},$
+$\Delta_{i}^{II}=q_{i}^{\prime}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}\delta_{ij},\text{ and }\Delta_{i}^{III}=-\delta_{ii}.$ (A.32)
+Following an argument identical to the proof of Lemma B7 in IN02,
 $D_{211}=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.33)
 For $D_{212}$,
-$\left|D_{212}\right|^{2}$ $\leq Cn\left[\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right]$
+$\left|D_{212}\right|^{2}$ $\leq Cn\left[\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}\right]\left[n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right]$
 $=O_{p}\left\{n\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)\Delta_{1n}^{2}\right\}=o_{p}\left(1\right).$ (A.34)
 For $D_{213}$,
-$\left|D_{213}\right|^{2}\leq Cn\left[H_{2}\widehat{R}H_{2}^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\left(\Delta_{i}^{II}\right)^{2}+\left(\Delta_{i}^{III}\right)^{2}\right)\right]=O_{p}\left(nM_{1}^{1-2d_{1}/j_{1}}\right)=o_{p}\left(1\right),$ (A.37)
-where the first equality is established in the proof of Theorem 4 of IN02.
+$\left|D_{213}\right|^{2}\leq Cn\left[H_{2}\widehat{R}H_{2}^{\prime}\right]\left[n^{-1}\sum_{i}\left(\left(\Delta_{i}^{II}\right)^{2}+\left(\Delta_{i}^{III}\right)^{2}\right)\right]=O_{p}\left(nM_{1}^{1-2d_{1}/j_{1}}\right)=o_{p}\left(1\right),$ (A.35)
+where the first equality is established in the proof of Theorem 4 in IN02.
 Next, for $D_{22}$,
-$\left|D_{22}\right|$ $\leq Cn^{1/2}\left\|\widehat{H}_{2}\right\|\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|\left|n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right|$
-$=O_{p}\left(n^{1/2}\zeta\left(M_{3}\right)\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.38)
+$\left|D_{22}\right|$ $\leq C\sqrt{n}\left\|\widehat{H}_{2}\right\|\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|\left|n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right|$
+$=O_{p}\left(\sqrt{n}\zeta\left(M_{3}\right)\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.36)
 Combining the results for $D_{21}$ and $D_{22}$,
-$n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.39)
+$\sqrt{n}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.37)
 To show $\psi_{3i}=H_{2}r_{i}\xi_{i}$, I expand
-$\ n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
-$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}b_{1}\left(v_{i},w_{i}\right)-n^{1/2}Fb\left(x\right)$
+$\ \sqrt{n}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
+$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}b_{1}\left(v_{i},w_{i}\right)-\sqrt{n}Fb\left(x\right)$
 $=$ $\ n^{-1/2}\sum_{i}H_{2}r_{i}\left(b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)\right)+n^{-1/2}\sum_{i}\left(\widehat{H}_{2}-H_{2}\right)r_{i}\left(b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)\right)$
-$+n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)-n^{1/2}F\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)$
-$\eqqcolon$ $\ D_{31}+D_{32}+D_{33}+D_{34},$
+$+n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-r_{i}^{\prime}\rho^{M_{3}}\right)-\sqrt{n}F\left(b\left(x\right)-r\left(x\right)^{\prime}\rho^{M_{3}}\right)$
+$\eqqcolon$ $\ D_{31}+D_{32}+D_{33}+D_{34},$ (A.38)
 where I use the definition of $\widehat{H}_{2}$ and $\widehat{R}$ for the decomposition.
@@
 For $D_{32}$,
-$\mathbb{E}\left[\left.\left|D_{32}\right|^{2}\right|{\bf X}\right]$ $=n^{-1}\left(\widehat{H}_{2}-H_{2}\right)r^{{}^{\prime}}\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
-$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
+$\mathbb{E}\left[\left.\left|D_{32}\right|^{2}\right|{\bf X}\right]$ $=n^{-1}\left(\widehat{H}_{2}-H_{2}\right)r^{\prime}\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]r\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
+$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
 $\leq C\left\|\widehat{H}_{2}-H_{2}\right\|^{2}\left(1+\left\|\widehat{R}-I\right\|\right)$
 $=O_{p}\left\{\left\|\widehat{H}_{2}-H_{2}\right\|^{2}\right\}$
-$=O_{p}\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)=o_{p}\left(1\right),$
+$=O_{p}\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)=o_{p}\left(1\right),$ (A.39)
 where the first inequality holds by Assumption 6(c) and the fact that $\widehat{H}_{2}$ and $r$ are functions of $X$ only, the second equality holds by $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and the third equality follows similarly as in equation (A.1) and (A.6) of Newey (1997).
@@
 For $D_{33}$, by the CS inequality,
-$\left|D_{33}\right|^{2}$ $\leq n\left(\widehat{H}_{2}\widehat{R}\widehat{H}_{2}^{{}^{\prime}}\right)n^{-1}\sum_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)^{2}$
-$=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right),$
+$\left|D_{33}\right|^{2}$ $\leq n\left(\widehat{H}_{2}\widehat{R}\widehat{H}_{2}^{\prime}\right)n^{-1}\sum_{i}\left(b\left(x_{i}\right)-r_{i}^{\prime}\rho^{M_{3}}\right)^{2}$
+$=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right),$ (A.40)
 where the first equality holds by Assumption 6(a).
 Finally, for $D_{34}$,
-$\left|D_{34}\right|^{2}=nF^{2}\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)^{2}=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right).$
+$\left|D_{34}\right|^{2}=nF^{2}\left(b\left(x\right)-r\left(x\right)^{\prime}\rho^{M_{3}}\right)^{2}=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right).$ (A.41)
 Combining the results for $D_{31},$ $D_{32},$ $D_{33}$, and $D_{34}$, I have
-$n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}+o_{p}\left(1\right).$ (A.40)
+$\sqrt{n}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}+o_{p}\left(1\right).$ (A.42)
 In sum, I have proved
-$n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$ $=n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.41)
+$\sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$ $=n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.43)
 where
 $\psi_{1i}=H_{1}\left(p_{i}u_{i}-\text{$\overline{\mu}_{i}^{I}$}\right),\ \psi_{2i}=H_{2}\overline{\mu}_{i}^{II},\text{ and }\psi_{3i}=H_{2}r_{i}\xi_{i}.$ (A.44)
-Furthermore, notice that
+Furthermore, observe that
 $H_{1}p_{i}u_{i}\perp\left(H_{1}\overline{\mu}_{i}^{I},H_{2}\overline{\mu}_{i}^{II},H_{2}r_{i}\xi_{i}\right)$ (A.45)
@@
 $\leq$ $\ Cn^{-1}\mathbb{E}\left[\left(H_{1}p_{i}u_{i}\right)^{4}+\left(H_{1}\overline{\mu}_{i}^{I}\right)^{4}+\left(H_{2}\overline{\mu}_{i}^{II}\right)^{4}+\left(H_{2}r_{i}\xi_{i}\right)^{4}\right]$
-$\leq$ $\ Cn^{-1}\left(\zeta\left(M_{2}\right)^{2}M_{2}+\zeta\left(M_{2}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{2}M_{3}\right)\rightarrow 0,$
-where the last inequality holds by Lemma B5 of IN02. Then, by the Lindeberg–Feller CLT,
-[…]
-where $\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)=\widehat{b}\left(x\right)$ and $a\left(b_{1},V\right)=b\left(x\right)$.
+$\leq$ $\ Cn^{-1}\left(\zeta\left(M_{2}\right)^{2}M_{2}+\zeta\left(M_{2}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{2}M_{3}\right)\rightarrow 0,$ (A.46)
+where the last inequality holds by Lemma B5 of IN02.
+Then, by the Lindeberg–Feller CLT,
+$\sqrt{n}\Omega_{2}^{-1/2}\left(\widehat{b}\left(x\right)-b\left(x\right)\right)\overset{d}{\longrightarrow}N\left(0,1\right).$ (A.47)
 To construct a feasible confidence interval, one needs a consistent estimator of the covariance matrix $\Omega_{2}$. Define
-$\widehat{\Omega}_{2}$ $\coloneqq\widehat{\Omega}_{21}+\widehat{\Omega}_{22},\ \widehat{\Omega}_{21}\coloneqq\widehat{A}_{1}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{1}^{{}^{\prime}},$ (A.45)
-$\widehat{\Omega}_{22}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
-$\widehat{A}_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}\left(n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}\right),\ \text{$\widehat{\overline{p}}_{i}$}\coloneqq\overline{p}\left(\widehat{s}_{i}\right),$
-$\widehat{\overline{\mu}}_{i}^{II}$ $\coloneqq n^{-1}\sum_{j}r_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{and }\widehat{\xi}_{i}\coloneqq\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right).$
-[…]
+$\widehat{\Omega}_{2}$ $\coloneqq\widehat{\Omega}_{21}+\widehat{\Omega}_{22},\text{ where}$
+$\widehat{\Omega}_{21}$ $\coloneqq\widehat{A}_{1}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{\prime}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{1}^{\prime},$
+$\widehat{\Omega}_{22}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r\left(x\right)^{\prime}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r\left(x\right)^{\prime}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{\prime},$
+$\widehat{A}_{1}$ $\coloneqq r\left(x\right)^{\prime}\widehat{R}^{-1}\left(n^{-1}\sum_{i}r_{i}\overline{p}\left(\widehat{s}_{i}\right)^{\prime}\right),$
+$\widehat{\overline{\mu}}_{i}^{II}$ $\coloneqq n^{-1}\sum_{j}r_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{\prime}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }$
+$\widehat{\xi}_{i}$ $\coloneqq\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right).$ (A.48)
+I show $\widehat{\Omega}_{2}/\Omega_{2}-1\overset{p}{\longrightarrow}0$.
+Recall that
 $\Omega_{2}=\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)^{2}+\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-A_{2}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{2}=\Omega_{21}+\Omega_{22}$ (A.49)
@@
 $\widehat{\Omega}_{2}=n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{p}_{i}\widehat{u}_{i}\right)^{2}+n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\widehat{A}_{2}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{2}\eqqcolon\widehat{\Omega}_{21}+\widehat{\Omega}_{22}.$ (A.50)
-[…]
+The proof of $\widehat{\Omega}_{21}/\Omega_{2}-\Omega_{21}/\Omega_{2}\overset{p}{\longrightarrow}0$ is almost identical to the proof of Lemma B10 of IN02, except that $\widehat{A}_{1}$ instead of $A_{1}$ appears in the definition of $\widehat{H}_{1}$.
+Nonetheless, I have shown $\left\|\widehat{H}_{1}-H_{1}\right\|=o_{p}\left(1\right)$.
+Thus, the proof of Lemma B10 of IN02 directly applies.
 For $\widehat{\Omega}_{22}$, I need to show
@@
 $n^{-1}\sum_{i}D_{41i}^{2}$ $\leq\left\|\widehat{H}_{2}\right\|^{2}\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|^{2}n^{-1}\sum_{i}\left(\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}$
-$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left[\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)\right)^{2}+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}\right]$
+$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left[\left(\widehat{\overline{p}}_{i}^{\prime}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)\right)^{2}+\left(\widehat{\overline{p}}_{i}^{\prime}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}\right]$
 $=O_{p}\left(\zeta\left(M_{3}\right)^{2}\Delta_{2n}^{2}\right)=o_{p}\left(1\right),$ (A.53)
@@
 For $D_{44}$,
-$\mathbb{E}\left[\left.n^{-1}\sum_{i}D_{44i}^{2}\right|{\bf X}\right]$ $=\left(\widehat{H}_{2}-H_{2}\right)n^{-1}\sum_{i}r_{i}r_{i}^{{}^{\prime}}\mathbb{E}\left(\left.\xi_{i}^{2}\right|X_{i}\right)\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
-$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
+$\mathbb{E}\left[\left.n^{-1}\sum_{i}D_{44i}^{2}\right|{\bf X}\right]$ $=\left(\widehat{H}_{2}-H_{2}\right)n^{-1}\sum_{i}r_{i}r_{i}^{\prime}\mathbb{E}\left(\left.\xi_{i}^{2}\right|X_{i}\right)\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
+$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
 $\leq C\left\|\widehat{H}_{2}-H_{2}\right\|^{2}=o_{p}\left(1\right),$ (A.55)
@@
 $n^{-1}\sum_{i}D_{44i}^{2}=o_{p}\left(1\right).$ (A.56)
-Combining the results for $D_{41}$, $D_{42}$, $D_{43}$, and $D_{44}$, I obtain
+Combining the results for $D_{41}$, $D_{42}$, $D_{43}$, and $D_{44}$, I have
 $n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$ (A.57)
@@
 $\leq$ $\ 3n^{-1}\sum_{i}\left(\widehat{H}_{1}\widehat{\overline{\mu}}_{i}^{I}-H_{1}\overline{\mu}_{i}^{I}\right)^{2}+3n^{-1}\sum_{i}\left(\widehat{H}_{2}\widehat{\overline{\mu}}_{i}^{II}-H_{2}\overline{\mu}_{i}^{II}\right)^{2}$
-$+3n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$
+$+3n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$ (A.58)
 Since $\mathbb{E}\left(H_{1}\overline{\mu}_{i}^{I}-H_{2}\overline{\mu}_{i}^{II}-H_{2}r_{i}\xi_{i}\right)^{2}=\Omega_{22}/\Omega_{2}\leq 1$, by Markov’s inequality and Lemma B6 of IN02,
@@
 Combining the results for $\widehat{\Omega}_{21}$ and $\widehat{\Omega}_{22}$, I obtain
-[…]
-For the last result of Theorem 3, first define
-$\widehat{\Omega}_{3}$ $\coloneqq\widehat{\Omega}_{31}+\widehat{\Omega}_{32},\ \widehat{\Omega}_{31}\coloneqq\widehat{A}_{3}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{3},$ (A.58)
-$\widehat{\Omega}_{32}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
-$\widehat{A}_{3}$ $\coloneqq n^{-1}\sum_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}},\text{ and}\ \widehat{\overline{\mu}}_{i}^{III}\coloneqq n^{-1}\sum_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji}.$
-[…]
-Appendix B Sufficient Conditions for Assumption 2
-Proposition 1 (Sufficient Conditions for Assumption 2 based on Altonji and Matzkin (2005)).
+$\widehat{\Omega}_{2}/\Omega_{2}-1\overset{p}{\longrightarrow}0.$ (A.62)
+Next, for the APE estimator $\widehat{\overline{b}}$, define $b_{t}^{\text{APE}}=\mathbb{E}\beta_{it}$ and thus $\overline{b}=T^{-1}\sum_{t=1}^{T}b_{t}^{\text{APE}}$.
+I first estimate $b_{t}^{\text{APE}}$ by setting $r^{M_{3}}(x)=1$ in (4.7) for $\widehat{b}_{t}(x)$.
+Hence, the asymptotic normality proof for $\widehat{b}_{t}(x)$ directly applies to $\widehat{b}_{t}^{\text{APE}}$.
+In particular, by adapting to (A.43) with $r^{M_{3}}(x)=1$, the influence function for $\widehat{b}_{t}^{\text{APE}}$ in period $t$ is
+$\psi_{it}=A_{3t}P_{t}^{-1}(p_{it}u_{it}+\overline{\mu}_{it}^{I})-(\overline{\mu}_{it}^{III}+\xi_{it}),$ (A.63)
+where $A_{3t}=\mathbb{E}\overline{p}_{it}^{\prime}$ and $\overline{\mu}_{it}^{III}=\mathbb{E}\left[\left.b_{1t,V}\left(V_{jt},W_{j}\right)q_{jt}^{\prime}Q_{t}^{-1}q_{it}v_{ji,t}\right|\mathcal{I}_{i,t}\right]$.
+Let
+$\Omega_{3}:=\mathbb{\mathbb{V}}\left(\dfrac{1}{T}\sum_{t=1}^{T}\psi_{it}\right)=\dfrac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\text{Cov}(\psi_{it},\psi_{is}).$ (A.64)
+Following the same argument as for $\widehat{b}(x)$, I have
+$\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})=\sqrt{n}\,\Omega_{3}^{-1/2}\left[\dfrac{1}{T}\sum_{t=1}^{T}(\widehat{b}_{t}^{\text{APE}}-b_{t}^{\text{APE}})\right]=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\Omega_{3}^{-1/2}\,\dfrac{1}{T}\sum_{t=1}^{T}\psi_{it}\right)+o_{p}(1)$ (A.65)
+and the asymptotic normality follows by the Lindeberg-Feller CLT:
+$\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})\overset{d}{\longrightarrow}N(0,I).$ (A.66)
+Finally, by (A.64), an estimator $\widehat{\Omega}_{3}$ for $\Omega_{3}$ is written as
+$\widehat{\Omega}_{3}$ $=\dfrac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\left(\dfrac{1}{n}\sum_{i=1}^{n}\widehat{\psi}_{it}\widehat{\psi}_{is}^{\prime}\right),$ (A.67)
+where $\widehat{\psi}_{it}=\widehat{A}_{3t}\widehat{P}_{t}^{-1}(\widehat{p}_{it}\widehat{u}_{it}+\widehat{\overline{\mu}}_{it}^{I})-(\widehat{\overline{\mu}}_{it}^{III}+\widehat{\xi}_{it})$, $\widehat{A}_{3t}=n^{-1}\sum_{i}\widehat{\overline{p}}_{it}^{\prime}$, and $\widehat{\overline{\mu}}_{it}^{III}=n^{-1}\sum_{j}\widehat{b}_{1t,V}\left(\widehat{v}_{jt},w_{j}\right)q_{jt}^{\prime}\widehat{Q}^{-1}q_{it}\widehat{v}_{ji,t}.$ The consistency is established following the same argument as in deriving (A.62):
+$\widehat{\Omega}_{3}/\Omega_{3}-1\overset{p}{\longrightarrow}0.$ (A.68)
+Remark 5 (Degree of Basis Functions).
+The three degrees $M_{1},M_{2},M_{3}$ correspond to the series dimensions used to estimate $V_{t}(x,z,w)$, $G_{t}(x,v,w)$, and $b_{t}(x)$, respectively.
+For consistency I require $M_{i}\to\infty$ while $M_{i}/n\to 0$, so that approximation bias and sampling variance both vanish.
+Under Assumption 5, the first-step estimator satisfies
+$\text{MSE}(\widehat{V}_{t})=O\left(\underbrace{M_{1}^{1-2d_{1}/j_{1}}}_{\text{bias}}+\underbrace{M_{1}/n}_{\text{variance}}\right),$
+where the extra factor $M_{1}$ in the bias term comes from the uniform-in-$x$ analysis with $\mathbf{1}\{X_{it}\leq x\}$ as the dependent variable (as in IN02).
+For steps 2–3 the analysis is analogous, with two additional points: (i) since the model implies $G_{t}$ is linear in $x$, I use the reduced basis $p^{M_{2}}(x,v,w)=x\otimes p^{m_{2}}(v,w)$, which reduces dimension ($M_{2}=m_{2}d_{X}$); and (ii) first-step errors propagate, so the error for $\widehat{b}_{t}(x)$ aggregates across steps
+$\text{MSE}(\widehat{b}_{t}(x))=\left(M_{1}^{1-2d_{1}/j_{1}}+M_{1}/n\right)+\left(m_{2}^{-2d_{2}/j_{2}}+m_{2}/n\right)+\left(M_{3}^{-2d_{3}/j_{3}}+M_{3}/n\right).$
+Sup-norm results additionally impose standard bounds on basis functions and derivatives (Assumption 6(b)), satisfied by common choices such as splines and power series.
+Finally, Assumption 7 imposes “not too fast” growth on $M_{i}$ for asymptotic normality; for instance, for $\widehat{\overline{b}}$ (which does not involve $M_{3}$), with splines it suffices that $M_{1}=o(n^{1/7})$ and $M_{2}=o(n^{1/7})$, which is feasible when $d_{1}/j_{1}\geq 4$ and $d_{2}/j_{2}\geq 4$.
+A.2 Nonparametric Justification of the Index Exclusion Assumption
+Next, I introduce a proposition that supplies sufficient conditions for Assumption 2 by adapting the exchangeability condition of Altonji and Matzkin (2005).
+Proposition 1 (Sufficient Conditions for Assumption 2).
 Suppose the following conditions hold:
@@
 Proof of Proposition 1.
-To begin with, we use the first condition $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ to establish the exchangeability condition
-$f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}},$ (B.1)
+To begin with, I use the first condition $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ to establish the exchangeability condition
+$f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}},$ (A.69)
 where $\left(\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}\right)$ is any permutation of $\left({\bf X}_{i},{\bf Z}_{i}\right)$ in time.
-It is worth emphasizing that (B.1), which will be proved in the following, is different from Assumption 2.3 of Altonji and Matzkin (2005).
-[^27] To be specific, Altonji and Matzkin (2005) assume an exchangeability condition involving both $A$ and $\eta_{t}$, i.e., $f_{\left.A_{i},\eta_{it}\right|{\bf X}_{i}}=f_{\left.A_{i},\eta_{it}\right|{\bf\widetilde{X}}_{i}},$ which effectively rules out time-varying endogeneity through the random coefficients since it requires that the density of $\eta_{it}$ given $X_{it}=x_{it}$ is the same as that given $X_{it}=x_{is}$ for any $s\neq t$.
-Without loss of generality, we prove (B.1) for $T=2$ since any ordering of $\left\{1,...,T\right\}$ for finite $T$ can be achieved via a finite number of pairwise permutations.
-Then, we prove that one can construct $W_{i}$ such that Assumption 2 holds.
-For simplicity of notation, we assume $X_{it}$ and $Z_{it}$ are both scalars and suppress $i$ subscripts in all variables.
+It is worth emphasizing that (A.69), which will be proved in the following, is different from Assumption 2.3 of Altonji and Matzkin (2005).
+In particular, Altonji and Matzkin (2005) assume an exchangeability condition involving both $A$ and $\eta_{t}$, i.e., $f_{\left.A_{i},\eta_{it}\right|{\bf X}_{i}}=f_{\left.A_{i},\eta_{it}\right|{\bf\widetilde{X}}_{i}},$ which effectively rules out time-varying endogeneity through the random coefficients because it requires the density of $\eta_{it}$ given $X_{it}=x_{it}$ to be the same as that given $X_{it}=x_{is}$ for any $s\neq t$.
+Instead, I prove (A.69) under a more primitive assumption of $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ that is compatible with time-varying endogeneity through the random coefficients.
+Without loss of generality, I prove (A.69) for $T=2$ since any ordering of $\left\{1,\ldots,T\right\}$ for finite $T$ can be achieved via a finite number of pairwise permutations.
+Then, I prove that one can construct $W_{i}$ such that Assumption 2 holds.
+For simplicity of notation, I assume $X_{it}$ and $Z_{it}$ are both scalars and suppress $i$ subscripts in all variables.
 Thus, in this proof all subscripts of the variables denote the time period.
 By condition (a) of Proposition 1,
-$f_{A,\eta_{1},\eta_{2}}\left(a,h_{1},h_{2}\right)=f_{A,\eta_{1},\eta_{2}}\left(a,h_{2},h_{1}\right).$ (B.2)
+$f_{A,\eta_{1},\eta_{2}}\left(a,h_{1},h_{2}\right)=f_{A,\eta_{1},\eta_{2}}\left(a,h_{2},h_{1}\right).$ (A.70)
 Let $g^{-1}\left(X,Z,A\right)$ denote the inverse function of $g\left(Z,A,\eta\right)$ with respect to $\eta$, which exists by Assumption 1.
@@
 $=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right)\left|D_{2}D_{1}\right|$
-$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right),$ (B.3)
-where the first equality holds by the change of variables for $\eta_{1}$ and $\eta_{2}$, the second equality uses (B.2) and $Z\perp\left(A,\eta\right)$, the third equality holds by $X_{1}=g\left(z_{2},a,g^{-1}\left(x_{2},z_{2},a\right)\right)=x_{2}$ and $X_{2}=g\left(z_{1},a,g^{-1}\left(x_{1},z_{1},a\right)\right)=x_{1}$, and the last equality uses the fact that the product of derivatives of inverse functions equals one.
-Given (B.3), I integrate $A$ out in (B.3) to obtain
-$f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$ $=f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right),$ (B.4)
+$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right),$ (A.71)
+where the first equality holds by the change of variables for $\eta_{1}$ and $\eta_{2}$, the second equality uses (A.70) and $Z\perp\left(A,\eta\right)$, the third equality holds by $X_{1}=g\left(z_{2},a,g^{-1}\left(x_{2},z_{2},a\right)\right)=x_{2}$ and $X_{2}=g\left(z_{1},a,g^{-1}\left(x_{1},z_{1},a\right)\right)=x_{1}$, and the last equality uses the fact that the product of derivatives of inverse functions equals one.
+Given (A.71), I integrate AA out in (A.71) to obtain
+$f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$ $=f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right),$ (A.72)
 which implies
@@
 $=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right)/f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right)$
-$=$ $\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right),$ (B.5)
-where the second equality uses (B.3) and (B.4).
+$=$ $\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right),$ (A.73)
+where the second equality uses (A.71) and (A.72).
 The rest of the proof follows the same argument as in Section 2.2 of Altonji and Matzkin (2005).
-Specifically, I show that for any on-support $a$, the conditional density $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a function of the form $f_{\left.A\right|W}\left(\left.a\right|w\right)$, where $W$ is a vector-valued function symmetric in the paired elements of $\left(X,Z\right)$.
-See footnote 10 for the detailed description of $W$.
+Specifically, I show that for any on-support $a$, the conditional density $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a function of the form $f_{\left.A\right|W}\left(\left.a\right|w\right)$.
+To construct $W$, I follow Weyl (1939) and let $\varphi_{1}\left(u\right)=\sum_{t=1}^{T}u_{t}$, $\varphi_{2}\left(u,v\right)=\sum_{t\neq s}^{T}u_{t}v_{s}$, …, $\varphi_{T}\left(u,v,\ldots,w\right)=\sum_{t\neq s\neq...\neq k}^{T}u_{t}v_{s}\ldots w_{k}$, where $u,v,\ldots,w$ are generic $T\times 1$ vectors.
+Then, I substitute each column vector of ${\bf D}_{i}=\left({\bf X}_{i},{\bf Z}_{i}\right)$ for each of the arguments $u,v,\ldots,w$ (repetitions included) to construct $W_{i}$.
+See Chapter II.3 of Weyl (1939) for details of the construction and its proof.
 By condition (b) of Proposition 1, the supports of $X$ and $Z$ are compact.
@@
 Therefore, by the Stone-Weierstrass theorem, there exists a function $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}$ that is a polynomial in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$ over a compact set with the property that for any fixed $\delta$ that is arbitrarily close to 0,
-$\max_{x_{1},x_{2}\in\mathcal{X},z_{1},z_{2}\in\mathcal{Z}}\left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|\leq\delta.$ (B.6)
+$\max_{x_{1},x_{2}\in\mathcal{X},z_{1},z_{2}\in\mathcal{Z}}\left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|\leq\delta.$ (A.74)
 Let
@@
 denote the simple average of $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ over all $T!$ (here $T=2$) unique permutations of $\left(x_{t},z_{t}\right)$, and similarly for $\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$.
-By (B.5),
+By (A.73),
 $\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)=f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right).$
-By (B.5), (B.6), and the triangle inequality,
+By (A.73), (A.74), and the triangle inequality,
 $\ \left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|$
 $=$ $\ \left|\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|$
-$\leq$ $\ T!\times\left(\delta/T!\right)=\delta.$ (B.7)
-Since $f^{w}$ can be chosen to make $\delta$ arbitrarily small, equation (B.7) implies that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a polynomial $\overline{f}^{w}$ that is symmetric in $\left(x_{t},z_{t}\right)$ pairs for $t=1,2$.
-Furthermore, by the fundamental theorem of symmetric functions, for any $a$ on its support, $\overline{f}^{w}$ can be written as a polynomial function of the elementary symmetric functions $W\left(\cdot\right)$ of $\left(\left(x_{1},z_{1}\right),\left(x_{2},z_{2}\right)\right)$.
-We denote this function by $f_{\left.A\right|W}\left(\left.a\right|w\right)$ and obtain that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by $f_{\left.A\right|W}\left(\left.a\right|W\right)$.
-Let $\delta\rightarrow 0$ in (B.7).
-[^28] Note that to let $\delta\to 0$ we need the order of polynomials in the approximating function $f^{w}$ to increase.
-However, it does not affect the order of the elements of $W$ since by the fundamental theorem of symmetric functions, the elements of $W$ has a fixed order of $T$.
-The key is we have two different orders of polynomials; one is the order of the approximating functions $f^{w}$, the other is the order of the arguments of $W$.
-For example, one may let the order $P_{1}$ and $P_{2}$ of a polynomial function $h\left(x,y\right)=\sum_{i=1}^{P_{1}}\left(x+y\right)^{i}+\sum_{i=1}^{P_{2}}\left(xy\right)^{i}$ to increase while keeping the order of its symmetric arguments (in this example, it is $x+y$ and $xy$) fixed.
-See footnote 9 and 10 of Altonji and Matzkin (2005) for a more detailed discussion.
+$\leq$ $\ T!\times\left(\delta/T!\right)=\delta.$ (A.75)
+Since $f^{w}$ can be chosen to make $\delta$ arbitrarily small, equation (A.75) implies that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a polynomial $\overline{f}^{w}$ that is symmetric in $\left(x_{t},z_{t}\right)$ pairs for $t=1,2$.
+Furthermore, by the fundamental theorem of symmetric functions, for any $a$ on its support, $\overline{f}^{w}$ can be written as a polynomial function of the elementary symmetric functions $W$ of $\left(\left(x_{1},z_{1}\right),\left(x_{2},z_{2}\right)\right)$.
+I denote this function by $f_{\left.A\right|W}\left(\left.a\right|w\right)$ and obtain that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by $f_{\left.A\right|W}\left(\left.a\right|W\right)$.
+Let $\delta\rightarrow 0$ in (A.75).
+Note that to let $\delta\to 0$ the degree of polynomials in the approximating function $f^{w}$ needs to increase.
+However, it does not affect the dimension of $W$ since by the fundamental theorem of symmetric functions, the elements of $W$ have a fixed degree of $T$.
+The key is that there are two different degrees of polynomials: one is the degree of the approximating functions $f^{w}$, and the other is the degree of the arguments of WW.
+[^4] For example, one may let the degree $P_{1}$ and $P_{2}$ of a polynomial function $h\left(x,y\right)=\sum_{i=1}^{P_{1}}\left(x+y\right)^{i}+\sum_{i=1}^{P_{2}}\left(xy\right)^{i}$ to increase while keeping the degree of its symmetric arguments (in this example, $W=(x+y,xy)$) fixed.
+See also footnotes 9 and 10 of Altonji and Matzkin (2005) for a detailed discussion.
 Then, for any $t\in\left\{1,...,T\right\}$ and on-support $\left(x_{t},z_{t},a,w\right)$
 $f_{\left.A\right|X_{t},Z_{t},W}\left(\left.a\right|x_{t},z_{t},w\right)=f_{\left.A\right|W}\left(\left.a\right|w\right).$
+A.3 Further Discussion of Identification Assumptions
+When the exchangeability conditions are used to justify Assumption 2, I present a symmetry-based argument in the discussion of Assumption 4 that justifies Assumptions 2 and 4 simultaneously.
+I provide an example here to clarify this point.
+Suppose $T=2$ and $(X_{it},Z_{it})\in\mathbb{R}^{2}$.
+By Proposition 1, when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the second degree, Assumption 2 holds.
+Suppose $(X_{i1},X_{i2},Z_{i1},Z_{i2})=\left(x_{1},x_{2},z_{1},z_{2}\right)$ satisfies $\{V_{it}=v,W_{i}=w\}$ for $t=1,2$ where $x_{1},x_{2},z_{1},z_{2}\in\mathbb{R}^{2}$ and $x_{1}$ and $x_{2}$ are linearly independent.
+Then, by definition of $V_{it}$ and the symmetry of $W_{i}$ in $\left(X_{it},Z_{it}\right)$ through $t$, $(X_{i1},X_{i2},Z_{i1},Z_{i2})=\left(x_{2},x_{1},z_{2},z_{1}\right)$ must also satisfy $V_{it}=v$ and $W_{i}=w$ for $t=1,2$.
+Therefore, given $V_{it}=v$ and $W_{i}=w$ there are two linearly independent vectors of $x_{1}$ and $x_{2}$ that $X_{it}$ can take, satisfying Assumption 4.
+However, Assumption 4 requires more residual variation in $X_{it}$ given $(V_{it},W_{i})$ and can be restrictive if nonparametric justifications are used for Assumption 2.
+I summarize the proposal by Altonji and Matzkin (2005) to achieve such residual variation.
+Specifically, Altonji and Matzkin (2005) suggest to
+(i) leverage unconditional variation in the exogenous regressors that do not enter $W_{i}$, and I provide more details at the end of this section;
+(ii) restrict the conditional distribution of $A_{i}$ given $W_{i}$ to depend on not all but only a subset of coordinates of $W_{i}$, which is testable by comparing the fit of $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ with $W_{i}$ replaced by a proper subset of it;
+(iii) assume $f_{\left.A_{i}\right|W_{i}}$ to depend on $W_{i}$ only through a linear combination of the elements of $W_{i}$, i.e., $f_{\left.A_{i}\right|W_{i}}=f_{\left.A_{i}\right|\sum_{k=1}^{d_{W}}c_{k}W_{i,k}}$ where $W_{i}=(W_{i,1},\ldots,W_{i,d_{W}})^{\prime}$ and the $c_{k}$s are unknown parameters, which implies that
+$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},\sum_{k=1}^{d_{W}}c_{k}W_{i,k}\right]$
+and the $c_{k}$s can be estimated using Ichimura and Lee (1991);
+(iv) impose a priori restrictions on model (2.1)–(2.2) such that $W_{i}$ enters $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ in a parametric way (e.g.,
+$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\sum_{k=1}^{d_{W}}c_{k}W_{i,k}+\sum_{l=1}^{d_{V}}d_{l}V_{it,l}$
+when $d_{X}=1$); and,
+(v) directly impose functional restrictions on $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$.
+Finally, to see how the unconditional variations in the exogenous regressors help to satisfy the required residual variation in endogenous regressors, suppose $X_{it,1}$ is endogenous while $X_{it,-1}$ (all other coordinates of $X_{it}$ except $X_{it,1}$) are exogenous.
+Suppose the conditional support of $X_{it,1}$ given $(V_{it},W_{i})$ only contains a non-zero singleton, while the unconditional support of $X_{it,-1}$ contains a small ball of positive radius.
+Denote $b_{1}\left(V_{it},W_{i}\right)=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$.
+Then, I can identify the first-order moments of $\beta_{it}$ by taking partial derivative with respect to $X_{it,-1}$ on both sides of
+$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it,1}b_{1,1}\left(V_{it},W_{i}\right)+X_{it,-1}^{\prime}b_{1,-1}\left(V_{it},W_{i}\right)$ (A.76)
+to identify
+$b_{1,-1}\left(V_{it},W_{i}\right)=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it,-1}.$ (A.77)
+Finally, identification of $b_{1,1}\left(V_{it},W_{i}\right)$ can be obtained as
+$b_{1,1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]-X_{it,-1}^{\prime}b_{1,-1}\left(V_{it},W_{i}\right)\right)/X_{it,1}.$
+Note that there are two conditions on $X_{it,-1}$ needed for the argument to work.
+First, $X_{it,-1}$ needs to be exogenous, so that it does not enter the construction of $V_{it}$ and $W_{i}$.
+Because of this, when taking partial derivative with respect to $X_{it,-1}$ in (A.76), it directly leads to (A.77).
+Second, $X_{it,-1}$ needs to be continuous vector to enable the derivative-based argument.
+When $X_{it,-1}$ is discrete, one may use arithmetic differencing technique to identify $b_{1,-1}\left(V_{it},W_{i}\right)$.
+Appendix B Additional Empirical Results
+I conduct several robustness checks of the empirical findings in Section 5.
+Measurement error.
+To address measurement error that is typical in firm-level data, I go beyond the baseline cleaning (removing outliers, negative, and extreme values) and re-estimate the model using alternative deflators and excluding firms in the top and bottom 1% of total output.
+Table A.1 reports the results; the estimates largely remain in line with the baseline.
+Instrument validity.
+I probe instrument validity by varying the geographic aggregation (provincial rather than finer units), using unweighted competitor averages, and adding city fixed effects.
+As shown in Table A.2, the main conclusion is unchanged.
+The specifications with city fixed effects yield noisier estimates, likely because many firms have few within-city competitors, making the within-city demeaning numerically less stable.
+Control-function specification.
+I assess sensitivity to the control-function specification along two dimensions.
+First, I consider an alternative construction of $W$ that includes averages of both the instruments and the regressors.
+Second, I re-estimate the series control function using alternative bases: second-degree polynomials and third-degree cubic splines with a knot at the median.
+Table A.3 shows that the estimates are essentially unchanged.
+One systematic difference is that cubic splines tend to deliver larger point estimates and wider confidence intervals across industries—consistent with the simulation evidence in Table A.8—suggesting a curse-of-dimensionality cost when the basis becomes too rich.
+Comparison with classic methods.
+I further compare the TERC estimates with standard OLS and a naive IV approach in Table A.4.
+Overall, the TERC method yields higher capital elasticities and lower labor elasticities than OLS in three sectors; in the remaining two sectors, the estimates are broadly similar.
+These patterns echo the findings of Olley and Pakes (1996), albeit in a different context and using a different dataset.
+By contrast, the naive IV estimates are economically implausible, underscoring the importance of accounting for potentially nonlinear, nonseparable heterogeneity through time-varying random coefficients.
+Taken together, these exercises indicate that the results are robust to a wide range of alternative specifications and sample restrictions.
+Table A.1: Robustness to Measurement Error
+Textile TERC Alt Deflator Non-Extreme Output
+Capital Elasticity 0.415 0.403 0.408
+95% CI [0.393, 0.441] [0.379, 0.428] [0.385, 0.436]
+Labor Elasticity 0.452 0.472 0.441
+95% CI [0.415, 0.485] [0.436, 0.505] [0.405, 0.475]
+Chemical TERC Alt Deflator Non-Extreme Output
+Capital Elasticity 0.463 0.420 0.463
+95% CI [0.430, 0.497] [0.387, 0.453] [0.429, 0.492]
+Labor Elasticity 0.311 0.439 0.293
+95% CI [0.263, 0.361] [0.387, 0.489] [0.245, 0.339]
+Nonmetallic Mineral TERC Alt Deflator Non-Extreme Output
+Capital Elasticity 0.371 0.377 0.361
+95% CI [0.332, 0.410] [0.341, 0.417] [0.324, 0.398]
+Labor Elasticity 0.311 0.314 0.291
+95% CI [0.255, 0.365] [0.258, 0.365] [0.238, 0.341]
+General Equipment TERC Alt Deflator Non-Extreme Output
+Capital Elasticity 0.433 0.434 0.422
+95% CI [0.402, 0.469] [0.401, 0.474] [0.393, 0.457]
+Labor Elasticity 0.521 0.526 0.522
+95% CI [0.479, 0.564] [0.477, 0.569] [0.479, 0.566]
+Transportation Equipment TERC Alt Deflator Non-Extreme Output
+Capital Elasticity 0.425 0.453 0.429
+95% CI [0.393, 0.459] [0.415, 0.493] [0.396, 0.463]
+Labor Elasticity 0.588 0.566 0.571
+95% CI [0.536, 0.644] [0.508, 0.632] [0.515, 0.631]
+  • Notes: “Alt Deflator” refers to the price deflators constructed by Brandt et al.
+(2014).
+“Non-Extreme Output” stands for the subsample that excludes the top and bottom 1% firms in terms of total output.
+Table A.2: IV Validity Check
+Textile TERC Provincial IV Unweighted IV Location FE
+Capital Elasticity 0.415 0.393 0.414 0.410
+95% CI [0.393, 0.441] [0.368, 0.417] [0.390, 0.442] [0.386, 0.436]
+Labor Elasticity 0.452 0.458 0.460 0.427
+95% CI [0.415, 0.485] [0.423, 0.494] [0.425, 0.494] [0.392, 0.462]
+Chemical TERC Provincial IV Unweighted IV Location FE
+Capital Elasticity 0.463 0.444 0.459 0.449
+95% CI [0.430, 0.497] [0.411, 0.477] [0.425, 0.492] [0.409, 0.490]
+Labor Elasticity 0.311 0.321 0.303 0.360
+95% CI [0.263, 0.361] [0.274, 0.371] [0.257, 0.352] [0.300, 0.411]
+Nonmetallic Mineral TERC Provincial IV Unweighted IV Location FE
+Capital Elasticity 0.371 0.309 0.346 0.410
+95% CI [0.332, 0.410] [0.258, 0.353] [0.303, 0.387] [0.370, 0.455]
+Labor Elasticity 0.311 0.289 0.321 0.381
+95% CI [0.255, 0.365] [0.227, 0.345] [0.263, 0.377] [0.321, 0.429]
+General Equipment TERC Provincial IV Unweighted IV Location FE
+Capital Elasticity 0.433 0.433 0.434 0.414
+95% CI [0.402, 0.469] [0.400, 0.467] [0.404, 0.471] [0.378, 0.452]
+Labor Elasticity 0.521 0.535 0.520 0.572
+95% CI [0.479, 0.564] [0.492, 0.581] [0.479, 0.563] [0.523, 0.618]
+Transportation Equipment TERC Provincial IV Unweighted IV Location FE
+Capital Elasticity 0.425 0.408 0.424 0.382
+95% CI [0.393, 0.459] [0.372, 0.440] [0.393, 0.460] [0.346, 0.418]
+Labor Elasticity 0.588 0.589 0.587 0.612
+95% CI [0.536, 0.644] [0.538, 0.649] [0.538, 0.645] [0.558, 0.673]
+  • Notes: “Provincial IV” denotes the leave-one-out weighted input prices of the competitors in the same province (c.f., baseline IV constructed at city level), with the weights given by the firm’s total debt (for interest rate) and number of employees (for wages).
+“Unweighted IV” stands for the leave-one-out unweighted input prices of the competitors in the same city.
+“Location FE” includes a city fixed effect in the three-step series regression.
+Table A.3: Sensitivity to Control-Function Specification
+Textile TERC Alt $W$ Alt Basis I Alt Basis II
+Capital Elasticity 0.415 0.416 0.385 0.445
+95% CI [0.393, 0.441] [0.394, 0.443] [0.364, 0.407] [0.423, 0.478]
+Labor Elasticity 0.452 0.448 0.438 0.486
+95% CI [0.415, 0.485] [0.413, 0.482] [0.406, 0.469] [0.451, 0.528]
+Chemical TERC Alt $W$ Alt Basis I Alt Basis II
+Capital Elasticity 0.463 0.462 0.434 0.490
+95% CI [0.430, 0.497] [0.429, 0.495] [0.402, 0.464] [0.450, 0.534]
+Labor Elasticity 0.311 0.312 0.250 0.349
+95% CI [0.263, 0.361] [0.265, 0.362] [0.206, 0.291] [0.297, 0.406]
+Nonmetallic Mineral TERC Alt $W$ Alt Basis I Alt Basis II
+Capital Elasticity 0.371 0.357 0.334 0.370
+95% CI [0.332, 0.410] [0.317, 0.396] [0.298, 0.369] [0.320, 0.415]
+Labor Elasticity 0.311 0.303 0.331 0.303
+95% CI [0.255, 0.365] [0.245, 0.356] [0.280, 0.386] [0.239, 0.358]
+General Equipment TERC Alt $W$ Alt Basis I Alt Basis II
+Capital Elasticity 0.433 0.428 0.407 0.442
+95% CI [0.402, 0.469] [0.398, 0.464] [0.377, 0.441] [0.408, 0.484]
+Labor Elasticity 0.521 0.513 0.496 0.530
+95% CI [0.479, 0.564] [0.471, 0.556] [0.455, 0.540] [0.486, 0.579]
+Transportation Equipment TERC Alt $W$ Alt Basis I Alt Basis II
+Capital Elasticity 0.425 0.424 0.411 0.416
+95% CI [0.393, 0.459] [0.393, 0.458] [0.383, 0.445] [0.376, 0.451]
+Labor Elasticity 0.588 0.586 0.525 0.616
+95% CI [0.536, 0.644] [0.539, 0.644] [0.471, 0.574] [0.563, 0.678]
+  • Notes: “Alt WW” means including the time average of both the IVs and regressors into the $W$ vector.
+“Alt Basis I” represents the second-degree polynomial basis function.
+“Alt Basis II” denotes the third-degree spline basis function.
+Table A.4: Comparison with OLS and Naive IV Estimates
+Textile TERC OLS Naive IV
+Capital Elasticity 0.415 0.379 1.309
+95% CI [0.393, 0.441] [0.358, 0.400] [0.889, 1.729]
+Labor Elasticity 0.452 0.509 0.266
+95% CI [0.415, 0.485] [0.481, 0.537] [-0.281, 0.813]
+Chemical TERC OLS Naive IV
+Capital Elasticity 0.463 0.452 0.555
+95% CI [0.430, 0.497] [0.419, 0.484] [-0.094, 1.204]
+Labor Elasticity 0.311 0.339 -0.757
+95% CI [0.263, 0.361] [0.291, 0.386] [-1.302, -0.212]
+Nonmetallic Mineral TERC OLS Naive IV
+Capital Elasticity 0.371 0.382 -0.670
+95% CI [0.332, 0.410] [0.354, 0.411] [-4.605, 3.265]
+Labor Elasticity 0.311 0.384 -3.512
+95% CI [0.255, 0.365] [0.340, 0.427] [-9.599, 2.575]
+General Equipment TERC OLS Naive IV
+Capital Elasticity 0.433 0.389 0.951
+95% CI [0.402, 0.469] [0.363, 0.416] [0.765, 1.138]
+Labor Elasticity 0.521 0.501 -0.619
+95% CI [0.479, 0.564] [0.464, 0.538] [-1.220, -0.019]
+Transportation Equipment TERC OLS Naive IV
+Capital Elasticity 0.425 0.442 0.723
+95% CI [0.393, 0.459] [0.408, 0.475] [0.523, 0.924]
+Labor Elasticity 0.588 0.560 0.508
+95% CI [0.536, 0.644] [0.509, 0.612] [-0.110, 1.127]
+  • Notes: "OLS" denotes the pooled OLS method with constant coefficients to estimate the output elasticities.
+"Naive IV" represents two-stage least squares method where the IVs are the same as in TERC estimates.
+CIs are obtained using standard Stata command.
 Appendix C Simulation
@@
 A discussion of the data generating process (DGP) motivated by production function applications is first provided.
-Then, I present the baseline results of estimating the APE $\overline{b}\coloneqq\mathbb{E}\beta_{it}$.
+Then, I present the baseline results of estimating the APE $\overline{b}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$.
 Next, I show how the method performs for estimating the LAR $b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$.
-Finally, as robustness checks, I (i) vary the number of firms and periods, (ii) change the order of basis functions used for series estimation, and (iii) include ex-post shocks $\epsilon_{it}$ and $\upsilon_{it}$ to the DGP.
+Finally, as robustness checks, I (i) vary the number of firms and periods, (ii) change the degree of basis functions used for series estimation, and (iii) include ex-post shocks $\widetilde{\epsilon}_{it}$ and $\upsilon_{it}$.
 C.1 DGP
 The baseline revenue-based DGP is
-$Y_{it}=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
-[…]
-$k_{it}$ $=\dfrac{r_{it}-\left(A_{i}+\eta_{it}^{K}\right)\left(r_{it}-w_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{K}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{K}/2+1\right)}{\left(A_{i}+\eta_{it}^{K}\right)/5-1},$
-$l_{it}$ $=\dfrac{w_{it}-\left(A_{i}+\eta_{it}^{L}\right)\left(w_{it}-r_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{L}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{L}/2+1\right)}{\left(A_{i}+\eta_{it}^{L}\right)/5-1}.$
-Let $X_{it}=\left(k_{it},l_{it}\right)^{{}^{\prime}}$.
-It is clear that $X_{it}$ is correlated with $\beta_{it}$ in each period via $\left(A_{i},\eta_{it}^{K},\eta_{it}^{L}\right)$ in a nonseparable way.
+$Y_{it}=k_{it}\beta_{it,K}+l_{it}\beta_{it,L}+\omega_{it}+\widetilde{\epsilon}_{it},$
+where $\omega_{it},\ \beta_{it,K},\ \text{and }\beta_{it,L}$ are all functions of $\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$, $k_{it}$ and $l_{it}$ are the natural logs of optimal capital and labor calculated from the solution to firm $i$’s profit maximization problem, and $Y_{it}$ is the natural log of value-added output measured in dollars.
+Let $\mathcal{U}\left[a,b\right]$ denote the uniform distribution over $\left[a,b\right]$.
+I draw $A_{i}\sim_{i.i.d.}\mathcal{U}\left[1,2\right]$ and $\varepsilon_{it}\sim_{i.i.d.}\mathcal{U}\left[1,2\right]$.
+I let true $\omega_{it}=\ln\left(A_{i}+\varepsilon_{it}/2+1\right)$, $\beta_{it,K}=\left(A_{i}+\varepsilon_{it}\right)/10$, and $\beta_{it,L}=\left(A_{i}+\varepsilon_{it}\right)/10$, and write $\beta_{it}=\left(\omega_{it},\beta_{it,K},\beta_{it,L}\right)^{\prime}$.
+I compute the true $\overline{\omega}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\omega_{it}=1.1736,$ $\overline{b}_{K}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it,K}=.3$, and $\overline{b}_{L}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it,L}=.3$.
+The range of the random coefficients are set such that the second-order condition for the profit maximization problem is satisfied.
+For the baseline results, I let ex-post shocks $\widetilde{\epsilon}_{it}=\upsilon_{it}=0$ and investigate their impacts as robustness checks later.
+Suppose capital and labor choices are made separately (e.g., investment and hiring decisions made by different departments) based on ex-ante signals about $\varepsilon_{it}$ with noise $\lambda_{it,K}$ and $\lambda_{it,L}$: $\eta_{it,K}=\varepsilon_{it}+\lambda_{it,K}$ and $\eta_{it,L}=\varepsilon_{it}+\lambda_{it,L}$, where $\lambda_{it,K}\text{ and }\lambda_{it,L}\sim_{i.i.d.}\mathcal{U}\left[-.05,.05\right]$.
+Since $Y_{it}$ is measured in dollars which is the case with most real production datasets, the price of output $P_{it}$ is assumed to be 1.
+I draw each element of the IVs $Z_{it}=\left(r_{it},w_{it}\right)^{\prime}$ from $\mathcal{U}\left[0,\ln 3\right]$ independent of each other and all other variables (in particular, $Z_{it}\perp(\varepsilon_{it},\eta_{it})^{\prime}$), and solve the firm’s profit maximization problem to obtain
+$k_{it}$ $=\dfrac{r_{it}-\left(A_{i}+\eta_{it,K}\right)\left(r_{it}-w_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it,K}\right)/10\right)-\ln\left(A_{i}+\eta_{it,K}/2+1\right)}{\left(A_{i}+\eta_{it,K}\right)/5-1},$
+$l_{it}$ $=\dfrac{w_{it}-\left(A_{i}+\eta_{it,L}\right)\left(w_{it}-r_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it,L}\right)/10\right)-\ln\left(A_{i}+\eta_{it,L}/2+1\right)}{\left(A_{i}+\eta_{it,L}\right)/5-1}.$
+Let $X_{it}=\left(k_{it},l_{it}\right)^{\prime}$ denote the two endogenous choice variables.
+It is clear that $X_{it}$ is correlated with $\beta_{it}$ in each period via $\left(A_{i},\eta_{it,K},\eta_{it,L}\right)$ in a nonseparable way.
 I use $n$, $T$, and $M$ to denote the total number of firms, periods, and simulations, respectively.
-The observed data is $\left\{X_{it,m},Y_{it,m},Z_{it,m}\right\}$ for $i=1,...,n$, $t=1,...,T$, and $m=1,...,M$, which are used to construct $\widehat{\overline{b}}_{m}$ and $\widehat{b}_{t,m}\left(x\right)$ to estimate $\overline{b}$ and $b_{t}\left(x\right)$, respectively, via the estimation procedure outlined in Section 4.1.
+The simulated dataset consists of $\left\{X_{it}^{(m)},Y_{it}^{(m)},Z_{it}^{(m)}\right\}$ for $i=1,\ldots,n$, $t=1,\ldots,T$, and $m=1,\ldots,M$, which are used to construct $\widehat{\overline{b}}^{(m)}$ and $\widehat{b}_{t}^{(m)}\left(x\right)$ to estimate $\overline{b}$ and $b_{t}\left(x\right)$, respectively, via the estimation procedure outlined in Subsection 4.1.
 I evaluate the performance of the estimators by their biases, root-mean squared errors (rMSE), and mean normed deviations (MND), with the explicit mathematical definitions provided in the tables below.
 C.2 Results
-For the baseline results, I set $N=1,000$ and $T=2$, and use a second-order polynomial spline basis with its knot at the median for all series estimation steps.
+For the baseline results, I set $n=1,000$ and $T=2$, and use a second-degree polynomial spline basis with its knot at the median for all series estimation steps.
 I run $M=1,000$ simulations.
 For each $i$, I construct $W_{i}$ as the mean over time of each coordinate of $X_{it}$.
-Following the theory, I set $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$.
-The performance of $\widehat{\overline{\omega}}$, $\widehat{\overline{b}}^{K},$ and $\widehat{\overline{b}}^{L}$ is summarized in Table 3.
+Following the theory, I set $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{\prime}$.
+The performance of $\widehat{\overline{\omega}}$, $\widehat{\overline{b}}_{K},$ and $\widehat{\overline{b}}_{L}$ is summarized in Table A.5.
 For notational simplicity, I use $\sum_{m}$ for $\sum_{m=1}^{M}$ and $\sum_{i,t,m}$ for $\sum_{i=1}^{n}\sum_{t=1}^{T}\sum_{m=1}^{M}$.
-Table 3: Performance of $\widehat{\overline{b}}$
-Formula $\widehat{\overline{\omega}}$ $\widehat{\overline{b}}^{K}$ $\widehat{\overline{b}}^{L}$
-Bias $M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.22% 3.24%
-rMSE $\sqrt{M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 2.89% 3.65% 3.62%
-MND $M^{-1}\sum_{m}\left|\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.26% 3.26%
-The first row of Table 3 reports the normalized bias of each coordinate of $\widehat{\overline{b}}$.
+Table A.5: Performance of $\widehat{\overline{b}}$
+Formula $\widehat{\overline{\omega}}$ $\widehat{\overline{b}}_{K}$ $\widehat{\overline{b}}_{L}$
+Bias $M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right)/\left|\overline{b}_{d}\right|$ 2.81% 3.22% 3.24%
+rMSE $\sqrt{M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right)^{2}}/\left|\overline{b}_{d}\right|$ 2.89% 3.65% 3.62%
+MND $M^{-1}\sum_{m}\left|\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right|/\left|\overline{b}_{d}\right|$ 2.81% 3.26% 3.26%
+The first row of Table A.5 reports the normalized bias of each coordinate of $\widehat{\overline{b}}$.
 The bias is reasonably small across all three coordinates, with a magnitude between 2.81% and 3.24% of the length of the corresponding coordinate of $\overline{b}$.
 The second row shows the normalized rMSE of each coordinate of $\widehat{\overline{b}}$.
-My method achieves low normalized rMSEs between 2.89% and 3.65% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$.
+My method achieves low normalized rMSEs between 2.89% and 3.65% of the size of the corresponding coordinate $\overline{b}_{d}$ of $\overline{b}$.
 The last row presents the normalized MNDs of each coordinate of $\widehat{\overline{b}}$.
-Again, the method performs well with an MND between 2.81% and 3.26% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$.
-Next, I investigate the performance of $\widehat{b}_{t,m}\left(x\right)$, which is obtained by regressing the estimated $\widehat{b}_{1t,m}\left(\widehat{V}_{it,m},W_{i,m}\right)$ on the second-degree polynomial spline basis of $X_{it,m}$ with its knot at the median for each $\left(t,m\right)$ combination.
+Again, the method performs well with an MND between 2.81% and 3.26% of the size of the corresponding coordinate $\overline{b}_{d}$ of $\overline{b}$.
+Next, I investigate the performance of $\widehat{b}_{t}\left(x\right)$, which is obtained following the estimation procedure outlined in Subsection 4.1.
 For the true $b_{t}\left(x\right)$, due to the complex dependence structure of $X_{it}$ on $\beta_{it}$, there is no analytical solution available.
-Therefore, for each $t$ I pool the observations across all $M$ simulations ($N\times M$ observations for each $t$) and approximate $b_{t}\left(x\right)$ by regressing the true $\beta_{it,m}$ on the same spline basis functions of $X_{it,m}$.
-Table 4 presents the results.
-Note that by definition the bias of $\widehat{b}_{t}\left(x\right)$ is the same as $\widehat{\overline{b}}$ so I omit it here.
-The normalized rMSE of $\widehat{b}_{t}\left(x\right)$ is bigger than that of $\widehat{\overline{b}}$, with a magnitude between 3.36% and 5.59% of the size of the corresponding $\overline{b}^{\left(d\right)}$.
+Therefore, for each $t$ I pool the observations across all $M$ simulations ($n\times M$ observations for each $t$) and approximate $b_{t}\left(x\right)$ by regressing the true $\beta_{it}$ on the same spline basis functions of $X_{it}$.
+Table A.6 summarizes the results.
+Note that by definition the bias of $\widehat{b}_{t}\left(x\right)$ is the same as that of $\widehat{\overline{b}}$, so I omit it here.
+The normalized rMSE of $\widehat{b}_{t}\left(x\right)$ is bigger than that of $\widehat{\overline{b}}$, with a magnitude between 3.36% and 5.59% of the size of the corresponding $\overline{b}_{d}$.
 The normalized MND follows a similar pattern.
 The performance of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ is not as good as that of $\widehat{\overline{b}}$ for $\overline{b}$, which is expected because (i) $b_{t}\left(x\right)$ is a function rather than a finite-dimensional vector and (ii) there is approximation error in calculating the true $b_{t}\left(x\right)$ via simulations.
-Table 4: Performance of $\widehat{b}_{t}\left(x\right)$
-Formula $\widehat{\omega}_{t}\left(x\right)$ $\widehat{b}_{t}^{K}\left(x\right)$ $\widehat{b}_{t}^{L}\left(x\right)$
-rMSE $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left(\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 3.36% 5.57% 5.59%
-MND $\left(NTM\right)^{-1}\sum_{i,t,m}\left|\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.87% 4.36% 4.37%
-To show how robust my method is in estimating $\overline{b}$ and $b_{t}\left(x\right)$, I conduct another set of exercises.
-I evaluate the performance of my estimator using rMSE defined as $\sqrt{M^{-1}\sum_{m}\left\|\widehat{\overline{b}}_{m}-\overline{b}\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $\overline{b}$, and $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left\|\widehat{b}_{t,m}\left(x_{it,m}\right)-b_{t}\left(x_{it,m}\right)\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $b_{t}\left(x\right)$.
-First, I vary $N$ and $T$, and present the results in Table 5.
+Table A.6: Performance of $\widehat{b}_{t}\left(x\right)$
+Formula $\widehat{\omega}_{t}\left(x\right)$ $\widehat{b}_{t,K}\left(x\right)$ $\widehat{b}_{t,L}\left(x\right)$
+rMSE $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left(\widehat{b}_{t,d}^{(m)}\left(x_{it}^{(m)}\right)-b_{t,d}\left(x_{it}^{(m)}\right)\right)^{2}}/\left|\overline{b}_{d}\right|$ 3.36% 5.57% 5.59%
+MND $\left(NTM\right)^{-1}\sum_{i,t,m}\left|\widehat{b}_{t,d}^{(m)}\left(x_{it}^{(m)}\right)-b_{t,d}\left(x_{it}^{(m)}\right)\right|/\left|\overline{b}_{d}\right|$ 2.87% 4.36% 4.37%
+To show how robust my method is in estimating $\overline{b}$ and $b_{t}\left(x\right)$, I conduct another set of exercises. I evaluate the performance of my TERC estimator using rMSE defined as
+$\sqrt{M^{-1}\sum_{m}\left\|\widehat{\overline{b}}^{(m)}-\overline{b}\right\|^{2}}/\left\|\overline{b}\right\|$
+for the vector $\overline{b}$, and
+$\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left\|\widehat{b}_{t}^{(m)}\left(x_{it}^{(m)}\right)-b_{t}\left(x_{it}^{(m)}\right)\right\|^{2}}/\left\|\overline{b}\right\|$
+for the vector $b_{t}\left(x\right)$.
+First, I vary $n$ and $T$, and present the results in Table A.7.
 As expected, a larger $n$ is good for overall performance.
-However, the magnitude in the improvement of performance is mild, possibly caused by the fact that with more agents I need more data to control for the increasing degree of heterogeneity.
+However, the magnitude in the improvement of performance is mild, possibly because with more agents I need more data to control for the increasing degree of heterogeneity.
 On the other hand, I find that the method performs reasonably well even with a small sample size of $n=500$.
-Having a larger $T$ improves the performance of my method, which is again expected as I can exploit more information from repeated observations of the same individual to better control for the fixed effect $A_{i}$.
-Table 5: Performance under Varying $N$ and $T$
+Having a larger $T$ improves the performance, which is again expected as I can exploit more information from repeated observations of the same individual to better control for the fixed effect $A_{i}$.
+Table A.7: Performance under Varying $n$ and $T$
 $n$ ($T=2$) rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
-$500$ 3.25% 4.23%
+500500 3.25% 4.23%
 $1,000$ 2.99% 3.69%
@@
 4 2.66% 3.27%
-Second, I vary the degree of the spline basis functions used to construct the series estimators for all steps, and present the results in Table 6.
-The knot is placed at the median for all cases.
+Second, I vary the degree of the spline bases used to construct the series estimators for all steps.
+The knot is placed at the median for all specifications.
+Table A.8 contains the results.
 I find that increasing the degree of basis functions from one to two improves estimation accuracy significantly.
 When I increase the degree of basis functions from two to three, the performance deteriorates.
-I attribute it to the fact that there are many more regressors in each step of regression when $d=3$, leading to a possible over-fitting problem.
+When the degree becomes larger, there are more regressors in each step of regression, leading to a possible over-fitting problem.
 Based on this result, I use the second-degree splines in the empirical application.
-[^29] One may also use the AIC criterion to select the degree of basis functions.
-Table 6: Performance under Varying Degree of Basis Functions
+Note that one may also use the AIC criterion to select the degree of basis functions.
+Table A.8: Performance under Varying Degree of Basis Functions
 Degree of Basis Functions rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
@@
 3 4.82% 6.40%
-[…]
-Table 7: Performance with and without Ex Post Shocks
-Add $\epsilon_{it}$ to $Y_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
-No 2.99% 3.69%
-Yes 3.64% 5.43%
-Add $\upsilon_{it}$ to $\left(\beta_{it}^{K},\beta_{it}^{L}\right)$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
-No 2.99% 3.69%
-Yes 3.57% 5.19%
+Lastly, I examine how including the ex-post shocks $\widetilde{\epsilon}_{it}$ and $\upsilon_{it}$ into the model affects the performance of my estimators.
+I draw $\widetilde{\epsilon}_{it}\sim\mathcal{U}\left[-.25,.25\right]$, the ex-post shock to the main equation (2.1), independently of all the other variables.
+$\widetilde{\epsilon}_{it}$ can also be interpreted as an ex-post shock to $\omega_{it}$.
+For $\upsilon_{it}$, since ex-post shock to $\omega_{it}$ has been considered by $\widetilde{\epsilon}_{it}$, I draw $\upsilon_{it}^{s}\sim\mathcal{U}\left[-.1,.1\right]$ independently of all the other variables and use $\beta_{it,s}^{\text{new}}=\beta_{it,s}+\upsilon_{it,s}$ in (2.1) for $s\in\left\{K,L\right\}$.
+Results are presented in Table A.9.
+As expected, adding ex-post errors $\widetilde{\epsilon}_{it}$ or $\upsilon_{it}$ negatively affects the performance of the estimators.
+When $\widetilde{\epsilon}_{it}$ is included, the rMSE of $\widehat{\overline{b}}$ for estimating $\overline{b}$ increases from 2.99% to 3.64% and the rMSE of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ rises from 3.64% to 5.43%.
+The effect of including $\upsilon_{it}$ on the performance is similar.
+Table A.9: Performance with and without Ex Post Shocks
+Add $\epsilon_{it}$ to $Y_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$ Add $\upsilon_{it}$ to $\beta_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
+No 2.99% 3.69% No 2.99% 3.69%
+Yes 3.64% 5.43% Yes 3.57% 5.19%